Using GNU Fortran

For gcc version 13.0.1-texinfostuff (pre-release) (GCC)

The gfortran team

Published by the Free Software Foundation 51 Franklin Street, Fifth Floor Boston, MA 02110-1301, USA

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999-2023 Free Software Foundation, Inc. Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with the Invariant Sections being “Funding Free Software”, the Front-Cover Texts being (a) (see below), and with the Back-Cover Texts being (b) (see below). A copy of the license is included in the section entitled “GNU Free Documentation License”.

(a) The FSF’s Front-Cover Text is: A GNU Manual

(b) The FSF’s Back-Cover Text is: You have freedom to copy and modify this GNU Manual, like GNU software. Copies published by the Free Software Foundation raise funds for GNU development.

Short Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Invoking GNU Fortran 2 GNU Fortran Command Options . . . . . . . . . . . . . . . . . . . . . . . . . 7 3 Runtime: Influencing runtime behavior with

environment variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

Language Reference 4 Compiler Characteristics. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5 Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 6 Mixed-Language Programming. . . . . . . . . . . . . . . . . . . . . . . . . . 69 7 Coarray Programming. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 8 Intrinsic Procedures. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 9 Intrinsic Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293

Contributing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 GNU General Public License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 GNU Free Documentation License. . . . . . . . . . . . . . . . . . . . . . . . . . 315 Funding Free Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 Option Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 Keyword Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 327

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 About GNU Fortran . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 GNU Fortran and GCC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Standards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.3.1 Fortran 95 status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3.2 Fortran 2003 status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.3 Fortran 2008 status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.4 Fortran 2018 status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

Part I: Invoking GNU Fortran . . . . . . . . . . . . . . . . . . . . . 5

2 GNU Fortran Command Options . . . . . . . . . . . . . . 7

2.1 Option summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2 Options controlling Fortran dialect. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.3 Enable and customize preprocessing . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.4 Options to request or suppress errors and warnings. . . . . . . . . . . . . 17 2.5 Options for debugging your program or GNU Fortran . . . . . . . . . . 22 2.6 Options for directory search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 2.7 Influencing the linking step . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 2.8 Influencing runtime behavior. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 2.9 Options for code generation conventions . . . . . . . . . . . . . . . . . . . . . . . 25 2.10 Options for interoperability with other languages . . . . . . . . . . . . . 32 2.11 Environment variables affecting gfortran. . . . . . . . . . . . . . . . . . . . . 33

3 Runtime: Influencing runtime behavior with environment variables . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

3.1 TMPDIR—Directory for scratch files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 3.2 GFORTRAN_STDIN_UNIT—Unit number for standard input . . . . . . . 35 3.3 GFORTRAN_STDOUT_UNIT—Unit number for standard output …. 35 3.4 GFORTRAN_STDERR_UNIT—Unit number for standard error . . . . . . 35 3.5 GFORTRAN_UNBUFFERED_ALL—Do not buffer I/O on all units …. 35

3.6 GFORTRAN_UNBUFFERED_PRECONNECTED—Do not

buffer I/O on preconnected units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 3.7 GFORTRAN_SHOW_LOCUS—Show location for runtime errors . . . . . . 35 3.8 GFORTRAN_OPTIONAL_PLUS—Print leading + where permitted . . 36 3.9 GFORTRAN_LIST_SEPARATOR—Separator for list output . . . . . . . . . 36 3.10 GFORTRAN_CONVERT_UNIT—Set conversion for unformatted I/O . . 36 3.11 GFORTRAN_ERROR_BACKTRACE—Show backtrace on run-time errors . . 37

3.12 GFORTRAN_FORMATTED_BUFFER_SIZE—Set

buffer size for formatted I/O. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

3.13 GFORTRAN_UNFORMATTED_BUFFER_SIZE—Set

buffer size for unformatted I/O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

Part II: Language Reference . . . . . . . . . . . . . . . . . . . . . . 39

4 Compiler Characteristics . . . . . . . . . . . . . . . . . . . . . . 41

4.1 KIND Type Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.2 Internal representation of LOGICAL variables . . . . . . . . . . . . . . . . . 41 4.3 Evaluation of logical expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.4 MAX and MIN intrinsics with REAL NaN arguments . . . . . . . . . . 42 4.5 Thread-safety of the runtime library . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.6 Data consistency and durability. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 4.7 Files opened without an explicit ACTION= specifier . . . . . . . . . . . 44 4.8 File operations on symbolic links. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 4.9 File format of unformatted sequential files . . . . . . . . . . . . . . . . . . . . . 44 4.10 Asynchronous I/O. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 4.11 Behavior on integer overflow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

5 Extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

5.1 Extensions implemented in GNU Fortran . . . . . . . . . . . . . . . . . . . . . . 47 5.1.1 Old-style kind specifications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.1.2 Old-style variable initialization. . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.1.3 Extensions to namelist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 5.1.4 X format descriptor without count field. . . . . . . . . . . . . . . . . . . . 49 5.1.5 Commas in FORMAT specifications . . . . . . . . . . . . . . . . . . . . . . . . . 49 5.1.6 Missing period in FORMAT specifications . . . . . . . . . . . . . . . . . . . 49 5.1.7 Default widths for F, G and I format descriptors. . . . . . . . . . . 49 5.1.8 I/O item lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 5.1.9 Q exponent-letter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 5.1.10 BOZ literal constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 5.1.11 Real array indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 5.1.12 Unary operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 5.1.13 Implicitly convert LOGICAL and INTEGER values . . . . . . . . . . 50 5.1.14 Hollerith constants support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 5.1.15 Character conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 5.1.16 Cray pointers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 5.1.17 CONVERT specifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 5.1.18 OpenMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 5.1.19 OpenACC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 5.1.20 Argument list functions %VAL, %REF and %LOC . . . . . . . . . . . . 55 5.1.21 Read/Write after EOF marker . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 5.1.22 STRUCTURE and RECORD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 5.1.23 UNION and MAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 5.1.24 Type variants for integer intrinsics . . . . . . . . . . . . . . . . . . . . . . . 60 5.1.25 AUTOMATIC and STATIC attributes . . . . . . . . . . . . . . . . . . . . . . . . 61 5.1.26 Extended math intrinsics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 5.1.27 Form feed as whitespace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 5.1.28 TYPE as an alias for PRINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 5.1.29 %LOC as an rvalue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 5.1.30 .XOR. operator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 5.1.31 Bitwise logical operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 5.1.32 Extended I/O specifiers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 5.1.33 Legacy PARAMETER statements . . . . . . . . . . . . . . . . . . . . . . . 65 5.1.34 Default exponents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

5.2 Extensions not implemented in GNU Fortran . . . . . . . . . . . . . . . . . . 65 5.2.1 ENCODE and DECODE statements . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 5.2.2 Variable FORMAT expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 5.2.3 Alternate complex function syntax . . . . . . . . . . . . . . . . . . . . . . . . 67 5.2.4 Volatile COMMON blocks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 5.2.5 OPEN( ... NAME=) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 5.2.6 Q edit descriptor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

6 Mixed-Language Programming . . . . . . . . . . . . . . . 69

6.1 Interoperability with C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 6.1.1 Intrinsic Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 6.1.2 Derived Types and struct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 6.1.3 Interoperable Global Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 6.1.4 Interoperable Subroutines and Functions . . . . . . . . . . . . . . . . . . 70 6.1.5 Working with C Pointers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 6.1.6 Further Interoperability of Fortran with C . . . . . . . . . . . . . . . . 74

6.2 GNU Fortran Compiler Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 6.2.1 ATTRIBUTES directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 6.2.2 UNROLL directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 6.2.3 BUILTIN directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 6.2.4 IVDEP directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 6.2.5 VECTOR directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 6.2.6 NOVECTOR directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

6.3 Non-Fortran Main Program. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 6.3.1 _gfortran_set_args — Save command-line arguments … 77 6.3.2 _gfortran_set_options — Set library option flags . . . . . . . 77 6.3.3 _gfortran_set_convert — Set endian conversion . . . . . . . . 79

6.3.4 _gfortran_set_record_marker —

Set length of record markers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 6.3.5 _gfortran_set_fpe — Enable floating point exception traps . . 80

6.3.6 _gfortran_set_max_subrecord_

length — Set subrecord length. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 6.4 Naming and argument-passing conventions . . . . . . . . . . . . . . . . . . . . . 80

6.4.1 Naming conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 6.4.2 Argument passing conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

7 Coarray Programming . . . . . . . . . . . . . . . . . . . . . . . . . 85

7.1 Type and enum ABI Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 7.1.1 caf_token_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 7.1.2 caf_register_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 7.1.3 caf_deregister_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 7.1.4 caf_reference_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 7.1.5 caf_team_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

7.2 Function ABI Documentation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 7.2.1 _gfortran_caf_init — Initialiation function . . . . . . . . . . . . . 88 7.2.2 _gfortran_caf_finish — Finalization function . . . . . . . . . . 88 7.2.3 _gfortran_caf_this_image — Querying the image number . . 88

7.2.4 _gfortran_caf_num_images — Querying

the maximal number of images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

7.2.5 _gfortran_caf_image_status —

Query the status of an image . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

7.2.6 _gfortran_caf_failed_images — Get an array

of the indexes of the failed images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

7.2.7 _gfortran_caf_stopped_images — Get an array of

the indexes of the stopped images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 7.2.8 _gfortran_caf_register — Registering coarrays . . . . . . . . 90 7.2.9 _gfortran_caf_deregister — Deregistering coarrays …. 91 7.2.10 _gfortran_caf_is_present — Query whether an allocatable or pointer

component in a derived type coarray is allocated . . . . . . . . . . . . . . . 92 7.2.11 _gfortran_caf_send — Sending data from

a local image to a remote image . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 7.2.12 _gfortran_caf_get — Getting data from a remote image . . 93

7.2.13 _gfortran_caf_sendget — Sending

data between remote images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 7.2.14 _gfortran_caf_send_by_ref — Sending data from a local image to a

remote image with enhanced referencing options . . . . . . . . . . . . . . . 96

7.2.15 _gfortran_caf_get_by_ref — Getting data from a

remote image using enhanced references . . . . . . . . . . . . . . . . . . . . . . . 97 7.2.16 _gfortran_caf_sendget_by_ref — Sending data between remote

images using enhanced references on both sides . . . . . . . . . . . . . . . . 98 7.2.17 _gfortran_caf_lock — Locking a lock variable . . . . . . . . . 99 7.2.18 _gfortran_caf_lock — Unlocking a lock variable . . . . . . 100 7.2.19 _gfortran_caf_event_post — Post an event . . . . . . . . . . 100

7.2.20 _gfortran_caf_event_wait —

Wait that an event occurred . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 7.2.21 _gfortran_caf_event_query — Query event count …. 102 7.2.22 _gfortran_caf_sync_all — All-image barrier . . . . . . . . . 102

7.2.23 _gfortran_caf_sync_images —

Barrier for selected images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

7.2.24 _gfortran_caf_sync_memory — Wait for

completion of segment-memory operations . . . . . . . . . . . . . . . . . . . . 103

   7.2.25 _gfortran_caf_error_stop —

Error termination with exit code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

   7.2.26 _gfortran_caf_error_stop_str —

Error termination with string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

   7.2.27 _gfortran_caf_fail_image — Mark the

image failed and end its execution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

   7.2.28 _gfortran_caf_atomic_define —

Atomic variable assignment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 7.2.29 _gfortran_caf_atomic_ref — Atomic variable reference . . 104 7.2.30 _gfortran_caf_atomic_cas — Atomic compare and swap . . 105 7.2.31 _gfortran_caf_atomic_op — Atomic operation . . . . . . . . 106

   7.2.32 _gfortran_caf_co_broadcast —

Sending data to all images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 7.2.33 _gfortran_caf_co_max — Collective maximum reduction . . 107 7.2.34 _gfortran_caf_co_min — Collective minimum reduction . . 107 7.2.35 _gfortran_caf_co_sum — Collective summing reduction . . 108 7.2.36 _gfortran_caf_co_reduce — Generic collective reduction . . 108

8 Intrinsic Procedures . . . . . . . . . . . . . . . . . . . . . . . . . . 111

8.1 Introduction to intrinsic procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 8.2 ABORT — Abort the program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 8.3 ABS — Absolute value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 8.4 ACCESS — Checks file access modes . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 8.5 ACHAR — Character in ASCII collating sequence . . . . . . . . . . . . . . . 113 8.6 ACOS — Arccosine function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 8.7 ACOSD — Arccosine function, degrees . . . . . . . . . . . . . . . . . . . . . . . . . 114 8.8 ACOSH — Inverse hyperbolic cosine function . . . . . . . . . . . . . . . . . . . 115 8.9 ADJUSTL — Left adjust a string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 8.10 ADJUSTR — Right adjust a string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 8.11 AIMAG — Imaginary part of complex number. . . . . . . . . . . . . . . . . 117 8.12 AINT — Truncate to a whole number . . . . . . . . . . . . . . . . . . . . . . . . 117 8.13 ALARM — Execute a routine after a given delay. . . . . . . . . . . . . . . 118 8.14 ALL — All values in MASK along DIM are true . . . . . . . . . . . . . 119 8.15 ALLOCATED — Status of an allocatable entity. . . . . . . . . . . . . . . . . 120 8.16 AND — Bitwise logical AND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 8.17 ANINT — Nearest whole number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 8.18 ANY — Any value in MASK along DIM is true. . . . . . . . . . . . . . . 122 8.19 ASIN — Arcsine function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 8.20 ASIND — Arcsine function, degrees . . . . . . . . . . . . . . . . . . . . . . . . . . 123 8.21 ASINH — Inverse hyperbolic sine function . . . . . . . . . . . . . . . . . . . . 124 8.22 ASSOCIATED — Status of a pointer or pointer/target pair. . . . . 124 8.23 ATAN — Arctangent function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 8.24 ATAND — Arctangent function, degrees . . . . . . . . . . . . . . . . . . . . . . . 126 8.25 ATAN2 — Arctangent function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 8.26 ATAN2D — Arctangent function, degrees. . . . . . . . . . . . . . . . . . . . . . 128 8.27 ATANH — Inverse hyperbolic tangent function . . . . . . . . . . . . . . . . 129 8.28 ATOMIC_ADD — Atomic ADD operation . . . . . . . . . . . . . . . . . . . . . . 129 8.29 ATOMIC_AND — Atomic bitwise AND operation. . . . . . . . . . . . . . . 130 8.30 ATOMIC_CAS — Atomic compare and swap . . . . . . . . . . . . . . . . . . . 131 8.31 ATOMIC_DEFINE — Setting a variable atomically . . . . . . . . . . . . . 132 8.32 ATOMIC_FETCH_ADD — Atomic ADD operation with prior fetch . . 132 8.33 ATOMIC_FETCH_AND — Atomic bitwise AND

operation with prior fetch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 8.34 ATOMIC_FETCH_OR — Atomic bitwise OR

operation with prior fetch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 8.35 ATOMIC_FETCH_XOR — Atomic bitwise XOR

operation with prior fetch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 8.36 ATOMIC_OR — Atomic bitwise OR operation . . . . . . . . . . . . . . . . . 135 8.37 ATOMIC_REF — Obtaining the value of a variable atomically . . 136 8.38 ATOMIC_XOR — Atomic bitwise OR operation . . . . . . . . . . . . . . . . 137 8.39 BACKTRACE — Show a backtrace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 8.40 BESSEL_J0 — Bessel function of the first kind of order 0 . . . . . 138 8.41 BESSEL_J1 — Bessel function of the first kind of order 1 . . . . . 139 8.42 BESSEL_JN — Bessel function of the first kind . . . . . . . . . . . . . . . 139 8.43 BESSEL_Y0 — Bessel function of the second kind of order 0 . . 140 8.44 BESSEL_Y1 — Bessel function of the second kind of order 1 . . 140 8.45 BESSEL_YN — Bessel function of the second kind . . . . . . . . . . . . . 141 8.46 BGE — Bitwise greater than or equal to . . . . . . . . . . . . . . . . . . . . . . 142 8.47 BGT — Bitwise greater than . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 8.48 BIT_SIZE — Bit size inquiry function. . . . . . . . . . . . . . . . . . . . . . . . 142 8.49 BLE — Bitwise less than or equal to . . . . . . . . . . . . . . . . . . . . . . . . . 143 8.50 BLT — Bitwise less than . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 8.51 BTEST — Bit test function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 8.52 C_ASSOCIATED — Status of a C pointer . . . . . . . . . . . . . . . . . . . . . . 144 8.53 C_F_POINTER — Convert C into Fortran pointer . . . . . . . . . . . . . 145 8.54 C_F_PROCPOINTER — Convert C into Fortran procedure pointer . . 146 8.55 C_FUNLOC — Obtain the C address of a procedure . . . . . . . . . . . 146 8.56 C_LOC — Obtain the C address of an object . . . . . . . . . . . . . . . . . 147 8.57 C_SIZEOF — Size in bytes of an expression. . . . . . . . . . . . . . . . . . . 148 8.58 CEILING — Integer ceiling function . . . . . . . . . . . . . . . . . . . . . . . . . . 149 8.59 CHAR — Character conversion function. . . . . . . . . . . . . . . . . . . . . . . 149 8.60 CHDIR — Change working directory. . . . . . . . . . . . . . . . . . . . . . . . . . 150 8.61 CHMOD — Change access permissions of files . . . . . . . . . . . . . . . . . . 151 8.62 CMPLX — Complex conversion function. . . . . . . . . . . . . . . . . . . . . . . 152 8.63 CO_BROADCAST — Copy a value to all

images the current set of images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 8.64 CO_MAX — Maximal value on the current set of images . . . . . . . 153 8.65 CO_MIN — Minimal value on the current set of images. . . . . . . . 154 8.66 CO_REDUCE — Reduction of values on the current set of images . . 155 8.67 CO_SUM — Sum of values on the current set of images . . . . . . . . 156

8.68 COMMAND_ARGUMENT_COUNT — Get number

of command line arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 8.69 COMPILER_OPTIONS — Options passed to the compiler . . . . . . . 157 8.70 COMPILER_VERSION — Compiler version string . . . . . . . . . . . . . . . 158 8.71 COMPLEX — Complex conversion function . . . . . . . . . . . . . . . . . . . . 158 8.72 CONJG — Complex conjugate function . . . . . . . . . . . . . . . . . . . . . . . 159 8.73 COS — Cosine function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 8.74 COSD — Cosine function, degrees . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 8.75 COSH — Hyperbolic cosine function . . . . . . . . . . . . . . . . . . . . . . . . . . 161 8.76 COTAN — Cotangent function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 8.77 COTAND — Cotangent function, degrees . . . . . . . . . . . . . . . . . . . . . . 162 8.78 COUNT — Count function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 8.79 CPU_TIME — CPU elapsed time in seconds . . . . . . . . . . . . . . . . . . . 164 8.80 CSHIFT — Circular shift elements of an array . . . . . . . . . . . . . . . . 164 8.81 CTIME — Convert a time into a string . . . . . . . . . . . . . . . . . . . . . . . 165 8.82 DATE_AND_TIME — Date and time subroutine . . . . . . . . . . . . . . . . 166 8.83 DBLE — Double conversion function . . . . . . . . . . . . . . . . . . . . . . . . . 167 8.84 DCMPLX — Double complex conversion function . . . . . . . . . . . . . . 167 8.85 DIGITS — Significant binary digits function. . . . . . . . . . . . . . . . . . 168 8.86 DIM — Positive difference. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 8.87 DOT_PRODUCT — Dot product function . . . . . . . . . . . . . . . . . . . . . . . 169 8.88 DPROD — Double product function . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 8.89 DREAL — Double real part function . . . . . . . . . . . . . . . . . . . . . . . . . . 171 8.90 DSHIFTL — Combined left shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 8.91 DSHIFTR — Combined right shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 8.92 DTIME — Execution time subroutine (or function). . . . . . . . . . . . 172 8.93 EOSHIFT — End-off shift elements of an array. . . . . . . . . . . . . . . . 173 8.94 EPSILON — Epsilon function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 8.95 ERF — Error function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 8.96 ERFC — Error function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 8.97 ERFC_SCALED — Error function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 8.98 ETIME — Execution time subroutine (or function). . . . . . . . . . . . 176 8.99 EVENT_QUERY — Query whether a coarray event has occurred . . 177 8.100 EXECUTE_COMMAND_LINE — Execute a shell command . . . . . . . 178 8.101 EXIT — Exit the program with status.. . . . . . . . . . . . . . . . . . . . . . 179 8.102 EXP — Exponential function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 8.103 EXPONENT — Exponent function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 8.104 EXTENDS_TYPE_OF — Query dynamic type for extension. . . . . 180 8.105 FDATE — Get the current time as a string . . . . . . . . . . . . . . . . . . 181 8.106 FGET — Read a single character in stream mode from stdin . . 182 8.107 FGETC — Read a single character in stream mode . . . . . . . . . . . 183 8.108 FINDLOC — Search an array for a value . . . . . . . . . . . . . . . . . . . . . 183 8.109 FLOOR — Integer floor function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 8.110 FLUSH — Flush I/O unit(s). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 8.111 FNUM — File number function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 8.112 FPUT — Write a single character in stream mode to stdout . . 186 8.113 FPUTC — Write a single character in stream mode . . . . . . . . . . 187 8.114 FRACTION — Fractional part of the model representation …. 188 8.115 FREE — Frees memory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 8.116 FSEEK — Low level file positioning subroutine . . . . . . . . . . . . . . 189 8.117 FSTAT — Get file status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 8.118 FTELL — Current stream position . . . . . . . . . . . . . . . . . . . . . . . . . . 191 8.119 GAMMA — Gamma function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 8.120 GERROR — Get last system error message . . . . . . . . . . . . . . . . . . . 192 8.121 GETARG — Get command line arguments. . . . . . . . . . . . . . . . . . . . 192 8.122 GET_COMMAND — Get the entire command line. . . . . . . . . . . . . . . 193 8.123 GET_COMMAND_ARGUMENT — Get command line arguments … 194 8.124 GETCWD — Get current working directory . . . . . . . . . . . . . . . . . . . 195 8.125 GETENV — Get an environmental variable . . . . . . . . . . . . . . . . . . . 195 8.126 GET_ENVIRONMENT_VARIABLE — Get an environmental variable . . 196 8.127 GETGID — Group ID function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 8.128 GETLOG — Get login name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 8.129 GETPID — Process ID function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 8.130 GETUID — User ID function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 8.131 GMTIME — Convert time to GMT info . . . . . . . . . . . . . . . . . . . . . . 198 8.132 HOSTNM — Get system host name . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 8.133 HUGE — Largest number of a kind . . . . . . . . . . . . . . . . . . . . . . . . . . 200 8.134 HYPOT — Euclidean distance function. . . . . . . . . . . . . . . . . . . . . . . 200 8.135 IACHAR — Code in ASCII collating sequence . . . . . . . . . . . . . . . . 201 8.136 IALL — Bitwise AND of array elements . . . . . . . . . . . . . . . . . . . . 201 8.137 IAND — Bitwise logical and. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 8.138 IANY — Bitwise OR of array elements . . . . . . . . . . . . . . . . . . . . . . 203 8.139 IARGC — Get the number of command line arguments . . . . . . 204 8.140 IBCLR — Clear bit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 8.141 IBITS — Bit extraction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 8.142 IBSET — Set bit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 8.143 ICHAR — Character-to-integer conversion function . . . . . . . . . . 206 8.144 IDATE — Get current local time subroutine (day/month/year) . . 207 8.145 IEOR — Bitwise logical exclusive or . . . . . . . . . . . . . . . . . . . . . . . . . 208 8.146 IERRNO — Get the last system error number . . . . . . . . . . . . . . . . 209 8.147 IMAGE_INDEX — Function that converts a

cosubscript to an image index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 8.148 INDEX — Position of a substring within a string . . . . . . . . . . . . 209 8.149 INT — Convert to integer type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 8.150 INT2 — Convert to 16-bit integer type. . . . . . . . . . . . . . . . . . . . . . 211 8.151 INT8 — Convert to 64-bit integer type. . . . . . . . . . . . . . . . . . . . . . 211 8.152 IOR — Bitwise logical or . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 8.153 IPARITY — Bitwise XOR of array elements . . . . . . . . . . . . . . . . . 212 8.154 IRAND — Integer pseudo-random number . . . . . . . . . . . . . . . . . . . 213 8.155 IS_CONTIGUOUS — Test whether an array is contiguous . . . . . 214 8.156 IS_IOSTAT_END — Test for end-of-file value. . . . . . . . . . . . . . . . . 214 8.157 IS_IOSTAT_EOR — Test for end-of-record value . . . . . . . . . . . . . 215 8.158 ISATTY — Whether a unit is a terminal device. . . . . . . . . . . . . . 215 8.159 ISHFT — Shift bits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 8.160 ISHFTC — Shift bits circularly . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 8.161 ISNAN — Test for a NaN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 8.162 ITIME — Get current local time

subroutine (hour/minutes/seconds) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 8.163 KILL — Send a signal to a process. . . . . . . . . . . . . . . . . . . . . . . . . . 218 8.164 KIND — Kind of an entity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 8.165 LBOUND — Lower dimension bounds of an array . . . . . . . . . . . . . 219 8.166 LCOBOUND — Lower codimension bounds of an array . . . . . . . . 220 8.167 LEADZ — Number of leading zero bits of an integer . . . . . . . . . 220 8.168 LEN — Length of a character entity. . . . . . . . . . . . . . . . . . . . . . . . . 221 8.169 LEN_TRIM — Length of a character entity

without trailing blank characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 8.170 LGE — Lexical greater than or equal . . . . . . . . . . . . . . . . . . . . . . . . 222 8.171 LGT — Lexical greater than . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 8.172 LINK — Create a hard link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 8.173 LLE — Lexical less than or equal . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 8.174 LLT — Lexical less than . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 8.175 LNBLNK — Index of the last non-blank character in a string . . 225 8.176 LOC — Returns the address of a variable. . . . . . . . . . . . . . . . . . . . 226 8.177 LOG — Natural logarithm function. . . . . . . . . . . . . . . . . . . . . . . . . . 226 8.178 LOG10 — Base 10 logarithm function . . . . . . . . . . . . . . . . . . . . . . . 227 8.179 LOG_GAMMA — Logarithm of the Gamma function . . . . . . . . . . . 227 8.180 LOGICAL — Convert to logical type . . . . . . . . . . . . . . . . . . . . . . . . . 228 8.181 LSHIFT — Left shift bits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 8.182 LSTAT — Get file status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 8.183 LTIME — Convert time to local time info . . . . . . . . . . . . . . . . . . . 230 8.184 MALLOC — Allocate dynamic memory . . . . . . . . . . . . . . . . . . . . . . . 230 8.185 MASKL — Left justified mask. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 8.186 MASKR — Right justified mask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 8.187 MATMUL — matrix multiplication . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 8.188 MAX — Maximum value of an argument list . . . . . . . . . . . . . . . . . 232 8.189 MAXEXPONENT — Maximum exponent of a real kind . . . . . . . . . 233 8.190 MAXLOC — Location of the maximum value within an array . . 234 8.191 MAXVAL — Maximum value of an array . . . . . . . . . . . . . . . . . . . . . 235 8.192 MCLOCK — Time function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 8.193 MCLOCK8 — Time function (64-bit) . . . . . . . . . . . . . . . . . . . . . . . . . 236 8.194 MERGE — Merge variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 8.195 MERGE_BITS — Merge of bits under mask . . . . . . . . . . . . . . . . . . . 237 8.196 MIN — Minimum value of an argument list . . . . . . . . . . . . . . . . . 237 8.197 MINEXPONENT — Minimum exponent of a real kind . . . . . . . . . . 238 8.198 MINLOC — Location of the minimum value within an array . . 238 8.199 MINVAL — Minimum value of an array . . . . . . . . . . . . . . . . . . . . . . 239 8.200 MOD — Remainder function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 8.201 MODULO — Modulo function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 8.202 MOVE_ALLOC — Move allocation from one object to another . . 242 8.203 MVBITS — Move bits from one integer to another . . . . . . . . . . . 242 8.204 NEAREST — Nearest representable number . . . . . . . . . . . . . . . . . . 243 8.205 NEW_LINE — New line character . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 8.206 NINT — Nearest whole number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 8.207 NORM2 — Euclidean vector norms . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 8.208 NOT — Logical negation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 8.209 NULL — Function that returns an disassociated pointer . . . . . 246 8.210 NUM_IMAGES — Function that returns the number of images . . 246 8.211 OR — Bitwise logical OR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 8.212 PACK — Pack an array into an array of rank one . . . . . . . . . . . . 248 8.213 PARITY — Reduction with exclusive OR . . . . . . . . . . . . . . . . . . . . 249 8.214 PERROR — Print system error message . . . . . . . . . . . . . . . . . . . . . . 249 8.215 POPCNT — Number of bits set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 8.216 POPPAR — Parity of the number of bits set. . . . . . . . . . . . . . . . . . 250 8.217 PRECISION — Decimal precision of a real kind . . . . . . . . . . . . . . 251 8.218 PRESENT — Determine whether an optional

dummy argument is specified . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 8.219 PRODUCT — Product of array elements . . . . . . . . . . . . . . . . . . . . . . 252 8.220 RADIX — Base of a model number . . . . . . . . . . . . . . . . . . . . . . . . . . 252 8.221 RAN — Real pseudo-random number . . . . . . . . . . . . . . . . . . . . . . . . 253 8.222 RAND — Real pseudo-random number . . . . . . . . . . . . . . . . . . . . . . . 253 8.223 RANDOM_INIT — Initialize a pseudo-random number generator . . 254 8.224 RANDOM_NUMBER — Pseudo-random number . . . . . . . . . . . . . . . . . 255 8.225 RANDOM_SEED — Initialize a pseudo-random number sequence . . 255 8.226 RANGE — Decimal exponent range . . . . . . . . . . . . . . . . . . . . . . . . . . 256 8.227 RANK — Rank of a data object. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 8.228 REAL — Convert to real type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 8.229 RENAME — Rename a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 8.230 REPEAT — Repeated string concatenation . . . . . . . . . . . . . . . . . . . 258 8.231 RESHAPE — Function to reshape an array . . . . . . . . . . . . . . . . . . . 259 8.232 RRSPACING — Reciprocal of the relative spacing . . . . . . . . . . . . 259 8.233 RSHIFT — Right shift bits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 8.234 SAME_TYPE_AS — Query dynamic types for equality. . . . . . . . . 260 8.235 SCALE — Scale a real value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 8.236 SCAN — Scan a string for the presence of a set of characters . . 261 8.237 SECNDS — Time function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 8.238 SECOND — CPU time function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 8.239 SELECTED_CHAR_KIND — Choose character kind. . . . . . . . . . . . . 263 8.240 SELECTED_INT_KIND — Choose integer kind . . . . . . . . . . . . . . . . 264 8.241 SELECTED_REAL_KIND — Choose real kind . . . . . . . . . . . . . . . . . . 264 8.242 SET_EXPONENT — Set the exponent of the model . . . . . . . . . . . . 265 8.243 SHAPE — Determine the shape of an array . . . . . . . . . . . . . . . . . . 266 8.244 SHIFTA — Right shift with fill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 8.245 SHIFTL — Left shift. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 8.246 SHIFTR — Right shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 8.247 SIGN — Sign copying function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 8.248 SIGNAL — Signal handling subroutine (or function) . . . . . . . . . 269 8.249 SIN — Sine function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 8.250 SIND — Sine function, degrees. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 8.251 SINH — Hyperbolic sine function . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 8.252 SIZE — Determine the size of an array . . . . . . . . . . . . . . . . . . . . . 271 8.253 SIZEOF — Size in bytes of an expression . . . . . . . . . . . . . . . . . . . . 272 8.254 SLEEP — Sleep for the specified number of seconds . . . . . . . . . 273 8.255 SPACING — Smallest distance between

two numbers of a given type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 8.256 SPREAD — Add a dimension to an array . . . . . . . . . . . . . . . . . . . . 273 8.257 SQRT — Square-root function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 8.258 SRAND — Reinitialize the random number generator. . . . . . . . . 275 8.259 STAT — Get file status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 8.260 STORAGE_SIZE — Storage size in bits . . . . . . . . . . . . . . . . . . . . . . . 276 8.261 SUM — Sum of array elements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 8.262 SYMLNK — Create a symbolic link. . . . . . . . . . . . . . . . . . . . . . . . . . . 278 8.263 SYSTEM — Execute a shell command. . . . . . . . . . . . . . . . . . . . . . . . 278 8.264 SYSTEM_CLOCK — Time function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 8.265 TAN — Tangent function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 8.266 TAND — Tangent function, degrees . . . . . . . . . . . . . . . . . . . . . . . . . . 280 8.267 TANH — Hyperbolic tangent function . . . . . . . . . . . . . . . . . . . . . . . 281 8.268 THIS_IMAGE — Function that returns the

cosubscript index of this image . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 8.269 TIME — Time function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 8.270 TIME8 — Time function (64-bit) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 8.271 TINY — Smallest positive number of a real kind . . . . . . . . . . . . 284 8.272 TRAILZ — Number of trailing zero bits of an integer . . . . . . . . 284 8.273 TRANSFER — Transfer bit patterns . . . . . . . . . . . . . . . . . . . . . . . . . . 285 8.274 TRANSPOSE — Transpose an array of rank two . . . . . . . . . . . . . . 285 8.275 TRIM — Remove trailing blank characters of a string . . . . . . . . 286 8.276 TTYNAM — Get the name of a terminal device . . . . . . . . . . . . . . . 286 8.277 UBOUND — Upper dimension bounds of an array. . . . . . . . . . . . . 287 8.278 UCOBOUND — Upper codimension bounds of an array . . . . . . . . 287 8.279 UMASK — Set the file creation mask . . . . . . . . . . . . . . . . . . . . . . . . . 288 8.280 UNLINK — Remove a file from the file system . . . . . . . . . . . . . . . 288 8.281 UNPACK — Unpack an array of rank one into an array . . . . . . . 289 8.282 VERIFY — Scan a string for characters not a given set . . . . . . 289 8.283 XOR — Bitwise logical exclusive OR. . . . . . . . . . . . . . . . . . . . . . . . . 290

9 Intrinsic Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293

9.1 ISO_FORTRAN_ENV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 9.2 ISO_C_BINDING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 9.3 IEEE modules: IEEE_EXCEPTIONS,

IEEE_ARITHMETIC, and IEEE_FEATURES . . . . . . . . . . . . . . . . . . . . . . . . . . 296 9.4 OpenMP Modules OMP_LIB and OMP_LIB_KINDS . . . . . . . . . . . . . . 297 9.5 OpenACC Module OPENACC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299

Contributing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301

Contributors to GNU Fortran. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 Projects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302

GNU General Public License . . . . . . . . . . . . . . . . . . . . 303 GNU Free Documentation License . . . . . . . . . . . . . . 315

ADDENDUM: How to use this License for your documents. . . . . . . . . 322

Funding Free Software . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 Option Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 Keyword Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 327

1 Introduction

This manual documents the use of gfortran, the GNU Fortran compiler. You can find in this manual how to invoke gfortran, as well as its features and incompatibilities.

Warning: This document, and the compiler it describes, are still under development. While efforts are made to keep it up-to-date, it might not accurately reflect the status of the most recent GNU Fortran compiler.

1.1 About GNU Fortran

The GNU Fortran compiler is the successor to g77, the Fortran 77 front end included in GCC prior to version 4 (released in 2005). While it is backward-compatible with most g77 extensions and command-line options, gfortran is a completely new implemention designed to support more modern dialects of Fortran. GNU Fortran implements the Fortran 77, 90 and 95 standards completely, most of the Fortran 2003 and 2008 standards, and some features from the 2018 standard. It also implements several extensions including OpenMP and OpenACC support for parallel programming.

The GNU Fortran compiler passes the NIST Fortran 77 Test Suite (http://www.

fortran-2000.com/ArnaudRecipes/fcvs21_f95.html), and produces acceptable results
on the LAPACK Test Suite (https://www.netlib.org/lapack/faq.html). It also

provides respectable performance on the Polyhedron Fortran compiler benchmarks

(https://polyhedron.com/?page_id=175) and the Livermore Fortran Kernels test
(https://www.netlib.org/benchmark/livermore). It has been used to compile a

number of large real-world programs, including the HARMONIE and HIRLAM weather forecasting code (http://hirlam.org/) and the Tonto quantum chemistry package

(https://github.com/dylan-jayatilaka/tonto); see https://gcc.gnu.org/wiki/

GfortranApps for an extended list.

GNU Fortran provides the following functionality:

• Read a program, stored in a file and containing source code instructions written in Fortran 77.

• Translate the program into instructions a computer can carry out more quickly than it takes to translate the original Fortran instructions. The result after compilation of a program is machine code, which is efficiently translated and processed by a machine such as your computer. Humans usually are not as good writing machine code as they are at writing Fortran (or C++, Ada, or Java), because it is easy to make tiny mistakes writing machine code.

• Provide information about the reasons why the compiler may be unable to create a binary from the source code, for example if the source code is flawed. The Fortran language standards require that the compiler can point out mistakes in your code. An incorrect usage of the language causes an error message. The compiler also attempts to diagnose cases where your program contains a correct usage of the language, but instructs the computer to do something questionable. This kind of diagnostic message is called a warning message.

• Provide optional information about the translation passes from the source code to machine code. This can help you to find the cause of certain bugs which may not be obvious in the source code, but may be more easily found at a lower level compiler output. It also helps developers to find bugs in the compiler itself.

• Provide information in the generated machine code that can make it easier to find bugs in the program (using a debugging tool, called a debugger, such as the GNU Debugger

gdb).

• Locate and gather machine code already generated to perform actions requested by statements in the program. This machine code is organized into modules and is located and linked to the user program.

The GNU Fortran compiler consists of several components:

• A version of the gcc command (which also might be installed as the system’s cc command) that also understands and accepts Fortran source code. The gcc command is the driver program for all the languages in the GNU Compiler Collection (GCC); With gcc, you can compile the source code of any language for which a front end is available in GCC.

• The gfortran command itself, which also might be installed as the system’s f95 command. gfortran is just another driver program, but specifically for the Fortran compiler only. The primary difference between the gcc and gfortran commands is that the latter automatically links the correct libraries to your program.

• A collection of run-time libraries. These libraries contain the machine code needed to support capabilities of the Fortran language that are not directly provided by the machine code generated by the gfortran compilation phase, such as intrinsic functions and subroutines, and routines for interaction with files and the operating system.

• The Fortran compiler itself, (f951). This is the GNU Fortran parser and code generator, linked to and interfaced with the GCC backend library. f951 “translates” the source code to assembler code. You would typically not use this program directly; instead, the gcc or gfortran driver programs call it for you.

1.2 GNU Fortran and GCC

GNU Fortran is a part of GCC, the GNU Compiler Collection. GCC consists of a collection of front ends for various languages, which translate the source code into a languageindependent form called GENERIC. This is then processed by a common middle end which provides optimization, and then passed to one of a collection of back ends which generate code for different computer architectures and operating systems.

Functionally, this is implemented with a driver program (gcc) which provides the command-line interface for the compiler. It calls the relevant compiler front-end program (e.g., f951 for Fortran) for each file in the source code, and then calls the assembler and linker as appropriate to produce the compiled output. In a copy of GCC that has been compiled with Fortran language support enabled, gcc recognizes files with .f, .for, .ftn, .f90, .f95, .f03 and .f08 extensions as Fortran source code, and compiles it accordingly. A gfortran driver program is also provided, which is identical to gcc except that it automatically links the Fortran runtime libraries into the compiled program.

Source files with .f, .for, .fpp, .ftn, .F, .FOR, .FPP, and .FTN extensions are treated as fixed form. Source files with .f90, .f95, .f03, .f08, .F90, .F95, .F03 and .F08 extensions are treated as free form. The capitalized versions of either form are run through preprocessing. Source files with the lower case .fpp extension are also run through preprocessing.

This manual specifically documents the Fortran front end, which handles the programming language’s syntax and semantics. The aspects of GCC that relate to the optimization passes and the back-end code generation are documented in the GCC manual; see Section “Introduction” in Using the GNU Compiler Collection (GCC). The two manuals together provide a complete reference for the GNU Fortran compiler.

1.3 Standards

Fortran is developed by the Working Group 5 of Sub-Committee 22 of the Joint Technical Committee 1 of the International Organization for Standardization and the International Electrotechnical Commission (IEC). This group is known as WG5 (http://www.nag.co. uk/sc22wg5/). Official Fortran standard documents are available for purchase from ISO; a collection of free documents (typically final drafts) are also available on the wiki (https://

gcc.gnu.org/wiki/GFortranStandards).

The GNU Fortran compiler implements ISO/IEC 1539:1997 (Fortran 95). As such, it can also compile essentially all standard-compliant Fortran 90 and Fortran 77 programs. It also supports the ISO/IEC TR-15581 enhancements to allocatable arrays.

GNU Fortran also supports almost all of ISO/IEC 1539-1:2004 (Fortran 2003) and ISO/IEC 1539-1:2010 (Fortran 2008). It has partial support for features introduced in ISO/IEC 1539:2018 (Fortran 2018), the most recent version of the Fortran language standard, including full support for the Technical Specification Further Interoperability of Fortran with C (ISO/IEC TS 29113:2012). More details on support for these standards can be found in the following sections of the documentation.

Additionally, the GNU Fortran compilers supports the OpenMP specification (version 4.5 and partial support of the features of the 5.0 version, https://openmp.org/ specifications/). There also is support for the OpenACC specification (targeting version

2.6, https://www.openacc.org/). See https://gcc.gnu.org/wiki/OpenACC for more in-

formation.

1.3.1 Fortran 95 status

The Fortran 95 standard specifies in Part 2 (ISO/IEC 1539-2:2000) varying length character strings. While GNU Fortran currently does not support such strings directly, there exist two Fortran implementations for them, which work with GNU Fortran. One can be found

at http://user.astro.wisc.edu/~townsend/static.php?ref=iso-varying-string.

Deferred-length character strings of Fortran 2003 supports part of the features of ISO_ VARYING_STRING and should be considered as replacement. (Namely, allocatable or pointers

of the type character(len=:).)

Part 3 of the Fortran 95 standard (ISO/IEC 1539-3:1998) defines Conditional Compilation, which is not widely used and not directly supported by the GNU Fortran compiler. You can use the program coco to preprocess such files (http://www.daniellnagle.com/

coco.html).

1.3.2 Fortran 2003 status

GNU Fortran implements the Fortran 2003 (ISO/IEC 1539-1:2004) standard except for finalization support, which is incomplete. See the wiki page (https://gcc.gnu.org/wiki/ Fortran2003) for a full list of new features introduced by Fortran 2003 and their implementation status.

1.3.3 Fortran 2008 status

The GNU Fortran compiler supports almost all features of Fortran 2008; the wiki (https:// gcc.gnu.org/wiki/Fortran2008Status) has some information about the current implementation status. In particular, the following are not yet supported:

• DO CONCURRENT and FORALL do not recognize a type-spec in the loop header.

• The change to permit any constant expression in subscripts and nested implied-do limits in a DATA statement has not been implemented.

1.3.4 Fortran 2018 status

Fortran 2018 (ISO/IEC 1539:2018) is the most recent version of the Fortran language standard. GNU Fortran implements some of the new features of this standard:

• All Fortran 2018 features derived from ISO/IEC TS 29113:2012, “Further Interoperability of Fortran with C”, are supported by GNU Fortran. This includes assumed-type and assumed-rank objects and the SELECT RANK construct as well as the parts relating to BIND(C) functions. See also Section 6.1.6 [Further Interoperability of Fortran with C], page 74.

• GNU Fortran supports a subset of features derived from ISO/IEC TS 18508:2015,

“Additional Parallel Features in Fortran”:

• The new atomic ADD, CAS, FETCH and ADD/OR/XOR, OR and XOR intrinsics.

• The CO_MIN and CO_MAX and SUM reduction intrinsics, and the CO_BROADCAST and CO_REDUCE intrinsic, except that those do not support polymorphic types or types with allocatable, pointer or polymorphic components.

     • Events (EVENT POST, EVENT WAIT, EVENT_QUERY).
     • Failed images (FAIL IMAGE, IMAGE_STATUS, FAILED_IMAGES, STOPPED_IMAGES).

• An ERROR STOP statement is permitted in a PURE procedure.

• GNU Fortran supports the IMPLICIT NONE statement with an implicit-none-spec-

    list.

• The behavior of the INQUIRE statement with the RECL= specifier now conforms to Fortran 2018.

Chapter 1: Introduction 5

Part I: Invoking GNU Fortran

2 GNU Fortran Command Options

The gfortran command supports all the options supported by the gcc command. Only options specific to GNU Fortran are documented here.

See Section “GCC Command Options” in Using the GNU Compiler Collection (GCC), for information on the non-Fortran-specific aspects of the gcc command (and, therefore, the gfortran command).

All GCC and GNU Fortran options are accepted both by gfortran and by gcc (as well as any other drivers built at the same time, such as g++), since adding GNU Fortran to the GCC distribution enables acceptance of GNU Fortran options by all of the relevant drivers.

In some cases, options have positive and negative forms; the negative form of -ffoo would be -fno-foo. This manual documents only one of these two forms, whichever one is not the default.

2.1 Option summary

Here is a summary of all the options specific to GNU Fortran, grouped by type. Explanations are in the following sections.

Fortran Language Options

See Section 2.2 [Options controlling Fortran dialect], page 8.

-fall-intrinsics -fallow-argument-mismatch -fallow-invalid-boz
-fbackslash -fcray-pointer -fd-lines-as-code -fd-lines-as-comments
-fdec -fdec-char-conversions -fdec-structure -fdec-intrinsic-ints
-fdec-static -fdec-math -fdec-include -fdec-format-defaults
-fdec-blank-format-item -fdefault-double-8 -fdefault-integer-8
-fdefault-real-8 -fdefault-real-10 -fdefault-real-16 -fdollar-ok
-ffixed-line-length-n -ffixed-line-length-none -fpad-source
-ffree-form -ffree-line-length-n -ffree-line-length-none
-fimplicit-none -finteger-4-integer-8 -fmax-identifier-length
-fmodule-private -ffixed-form -fno-range-check -fopenacc -fopenmp
-freal-4-real-10 -freal-4-real-16 -freal-4-real-8 -freal-8-real-10
-freal-8-real-16 -freal-8-real-4 -std=std -ftest-forall-temp

Preprocessing Options

See Section 2.3 [Enable and customize preprocessing], page 14.

-A-question[=answer]
-Aquestion=answer -C -CC -Dmacro[=defn]
-H -P
-Umacro -cpp -dD -dI -dM -dN -dU -fworking-directory
-imultilib dir
-iprefix file -iquote -isysroot dir -isystem dir -nocpp
-nostdinc
-undef

Error and Warning Options

See Section 2.4 [Options to request or suppress errors and warnings], page 17.

-Waliasing -Wall -Wampersand -Warray-bounds
-Wc-binding-type -Wcharacter-truncation -Wconversion
-Wdo-subscript -Wfunction-elimination -Wimplicit-interface
-Wimplicit-procedure -Wintrinsic-shadow -Wuse-without-only

8 The GNU Fortran Compiler

-Wintrinsics-std -Wline-truncation -Wno-align-commons
-Wno-overwrite-recursive -Wno-tabs -Wreal-q-constant -Wsurprising
-Wunderflow -Wunused-parameter -Wrealloc-lhs -Wrealloc-lhs-all
-Wfrontend-loop-interchange -Wtarget-lifetime -fmax-errors=n
-fsyntax-only -pedantic
-pedantic-errors

Debugging Options See Section 2.5 [Options for debugging your program or GNU Fortran], page 22.

-fbacktrace -fdump-fortran-optimized -fdump-fortran-original
-fdebug-aux-vars -fdump-fortran-global -fdump-parse-tree -ffpe-trap=list
-ffpe-summary=list

Directory Options See Section 2.6 [Options for directory search], page 23.

-Idir  -Jdir -fintrinsic-modules-path dir

Link Options See Section 2.7 [Options for influencing the linking step], page 24.

-static-libgfortran  -static-libquadmath

Runtime Options See Section 2.8 [Options for influencing runtime behavior], page 24.

-fconvert=conversion -fmax-subrecord-length=length
-frecord-marker=length -fsign-zero

Interoperability Options See Section 2.10 [Options for interoperability], page 32.

-fc-prototypes -fc-prototypes-external

Code Generation Options See Section 2.9 [Options for code generation conventions], page 25.

-faggressive-function-elimination -fblas-matmul-limit=n
-fbounds-check -ftail-call-workaround -ftail-call-workaround=n
-fcheck-array-temporaries
-fcheck=<all|array-temps|bits|bounds|do|mem|pointer|recursion>
-fcoarray=<none|single|lib> -fexternal-blas -ff2c
-ffrontend-loop-interchange -ffrontend-optimize
-finit-character=n -finit-integer=n -finit-local-zero
-finit-derived -finit-logical=<true|false>
-finit-real=<zero|inf|-inf|nan|snan>
-finline-matmul-limit=n
-finline-arg-packing -fmax-array-constructor=n
-fmax-stack-var-size=n -fno-align-commons -fno-automatic
-fno-protect-parens -fno-underscoring -fsecond-underscore
-fpack-derived -frealloc-lhs -frecursive -frepack-arrays
-fshort-enums -fstack-arrays

2.2 Options controlling Fortran dialect

The following options control the details of the Fortran dialect accepted by the compiler:

-ffree-form
-ffixed-form

Specify the layout used by the source file. The free form layout was introduced in Fortran 90. Fixed form was traditionally used in older Fortran programs. When neither option is specified, the source form is determined by the file extension.

-fall-intrinsics

This option causes all intrinsic procedures (including the GNU-specific extensions) to be accepted. This can be useful with -std= to force standardcompliance but get access to the full range of intrinsics available with gfortran. As a consequence, -Wintrinsics-std will be ignored and no user-defined procedure with the same name as any intrinsic will be called except when it is explicitly declared EXTERNAL.

-fallow-argument-mismatch

Some code contains calls to external procedures with mismatches between the calls and the procedure definition, or with mismatches between different calls. Such code is non-conforming, and will usually be flagged with an error. This options degrades the error to a warning, which can only be disabled by disabling all warnings via -w. Only a single occurrence per argument is flagged by this

warning. -fallow-argument-mismatch is implied by -std=legacy.

Using this option is strongly discouraged. It is possible to provide standardconforming code which allows different types of arguments by using an explicit interface and TYPE(*).

-fallow-invalid-boz

A BOZ literal constant can occur in a limited number of contexts in standard conforming Fortran. This option degrades an error condition to a warning, and allows a BOZ literal constant to appear where the Fortran standard would otherwise prohibit its use.

-fd-lines-as-code
-fd-lines-as-comments

Enable special treatment for lines beginning with d or D in fixed form sources. If the -fd-lines-as-code option is given they are treated as if the first column contained a blank. If the -fd-lines-as-comments option is given, they are treated as comment lines.

-fdec DEC compatibility mode. Enables extensions and other features that mimic the default behavior of older compilers (such as DEC). These features are nonstandard and should be avoided at all costs. For details on GNU Fortran’s implementation of these extensions see the full documentation. Other flags enabled by this switch are: -fdollar-ok -fcray-pointer

-fdec-char-conversions    -fdec-structure    -fdec-intrinsic-ints
-fdec-static  -fdec-math  -fdec-include  -fdec-blank-format-item
-fdec-format-defaults
If -fd-lines-as-code/-fd-lines-as-comments are unset, then -fdec also
sets -fd-lines-as-comments.

-fdec-char-conversions

Enable the use of character literals in assignments and DATA statements for non-character variables.

-fdec-structure

Enable DEC STRUCTURE and RECORD as well as UNION, MAP, and dot (’.’) as a member separator (in addition to ’%’). This is provided for compatibility only; Fortran 90 derived types should be used instead where possible.

-fdec-intrinsic-ints

Enable B/I/J/K kind variants of existing integer functions (e.g. BIAND, IIAND, JIAND, etc...). For a complete list of intrinsics see the full documentation.

-fdec-math

Enable legacy math intrinsics such as COTAN and degree-valued trigonometric functions (e.g. TAND, ATAND, etc...) for compatability with older code.

-fdec-static

Enable DEC-style STATIC and AUTOMATIC attributes to explicitly specify the storage of variables and other objects.

-fdec-include

Enable parsing of INCLUDE as a statement in addition to parsing it as IN-CLUDE line. When parsed as INCLUDE statement, INCLUDE does not have to be on a single line and can use line continuations.

-fdec-format-defaults

Enable format specifiers F, G and I to be used without width specifiers, default widths will be used instead.

-fdec-blank-format-item

Enable a blank format item at the end of a format specification i.e. nothing following the final comma.

-fdollar-ok

Allow ‘$’ as a valid non-first character in a symbol name. Symbols that start with ‘$’ are rejected since it is unclear which rules to apply to implicit typing as different vendors implement different rules. Using ‘$’ in IMPLICIT statements is also rejected.

-fbackslash

Change the interpretation of backslashes in string literals from a single backslash character to “C-style” escape characters. The following combinations are expanded \a, \b, \f, \n, \r, \t, \v, \\, and \0 to the ASCII characters alert, backspace, form feed, newline, carriage return, horizontal tab, vertical tab, backslash, and NUL, respectively. Additionally, \xnn, \unnnn and \Unnnnnnnn (where each n is a hexadecimal digit) are translated into the Unicode characters corresponding to the specified code points. All other combinations of a character preceded by \ are unexpanded.

-fmodule-private

Set the default accessibility of module entities to PRIVATE. Use-associated entities will not be accessible unless they are explicitly declared as PUBLIC.

-ffixed-line-length-n

Set column after which characters are ignored in typical fixed-form lines in the source file, and, unless -fno-pad-source, through which spaces are assumed (as if padded to that length) after the ends of short fixed-form lines. Popular values for n include 72 (the standard and the default), 80 (card image), and 132 (corresponding to “extended-source” options in some popular compilers). n may also be ‘none’, meaning that the entire line is meaningful and that continued character constants never have implicit spaces appended to them to fill out the line. -ffixed-line-length-0 means the same thing as

-ffixed-line-length-none.
-fno-pad-source

By default fixed-form lines have spaces assumed (as if padded to that length) after the ends of short fixed-form lines. This is not done either if -ffixed-line-

length-0, -ffixed-line-length-none or if -fno-pad-source option is used.

With any of those options continued character constants never have implicit spaces appended to them to fill out the line.

-ffree-line-length-n

Set column after which characters are ignored in typical free-form lines in the source file. The default value is 132. n may be ‘none’, meaning that the entire line is meaningful. -ffree-line-length-0 means the same thing as -ffree-

line-length-none.
-fmax-identifier-length=n

Specify the maximum allowed identifier length. Typical values are 31 (Fortran 95) and 63 (Fortran 2003 and later).

-fimplicit-none

Specify that no implicit typing is allowed, unless overridden by explicit IMPLICIT statements. This is the equivalent of adding implicit none to the start of every procedure.

-fcray-pointer

Enable the Cray pointer extension, which provides C-like pointer functionality.

-fopenacc

Enable the OpenACC extensions. This includes OpenACC !$acc directives in free form and c$acc, *$acc and !$acc directives in fixed form, !$ conditional compilation sentinels in free form and c$, *$ and !$ sentinels in fixed form, and when linking arranges for the OpenACC runtime library to be linked in.

-fopenmp Enable the OpenMP extensions. This includes OpenMP !$omp directives in free form and c$omp, *$omp and !$omp directives in fixed form, !$ conditional compilation sentinels in free form and c$, *$ and !$ sentinels in fixed form, and when linking arranges for the OpenMP runtime library to be linked in. The

option -fopenmp implies -frecursive.

-fno-range-check

Disable range checking on results of simplification of constant expressions during compilation. For example, GNU Fortran will give an error at compile time when simplifying a = 1. / 0. With this option, no error will be given and a will be assigned the value +Infinity. If an expression evaluates to a value outside of the relevant range of [-HUGE():HUGE()], then the expression will be replaced by -Inf or +Inf as appropriate. Similarly, DATA i/Z'FFFFFFFF'/ will result in an integer overflow on most systems, but with -fno-range-check the value will “wrap around” and i will be initialized to −1 instead.

-fdefault-integer-8

Set the default integer and logical types to an 8 byte wide type. This option also affects the kind of integer constants like 42. Unlike -finteger-4-integer-8, it does not promote variables with explicit kind declaration.

-fdefault-real-8

Set the default real type to an 8 byte wide type. This option also affects the kind of non-double real constants like 1.0. This option promotes the default width of DOUBLE PRECISION and double real constants like 1.d0 to 16 bytes if pos-

sible. If -fdefault-double-8 is given along with fdefault-real-8, DOUBLE

PRECISION and double real constants are not promoted. Unlike -freal-4-real-8, fdefault-real-8 does not promote variables with explicit kind declarations.

-fdefault-real-10

Set the default real type to an 10 byte wide type. This option also affects the kind of non-double real constants like 1.0. This option promotes the default width of DOUBLE PRECISION and double real constants like 1.d0 to 16 bytes

if possible. If -fdefault-double-8 is given along with fdefault-real-10,

DOUBLE PRECISION and double real constants are not promoted. Unlike -freal-4-real-10, fdefault-real-10 does not promote variables with explicit kind declarations.

-fdefault-real-16

Set the default real type to an 16 byte wide type. This option also affects the kind of non-double real constants like 1.0. This option promotes the default width of DOUBLE PRECISION and double real constants like 1.d0 to 16 bytes

if possible. If -fdefault-double-8 is given along with fdefault-real-16,

DOUBLE PRECISION and double real constants are not promoted. Unlike -freal-4-real-16, fdefault-real-16 does not promote variables with explicit kind declarations.

-fdefault-double-8

Set the DOUBLE PRECISION type and double real constants like 1.d0 to an 8 byte wide type. Do nothing if this is already the default. This option prevents

-fdefault-real-8, -fdefault-real-10, and -fdefault-real-16, from pro-

moting DOUBLE PRECISION and double real constants like 1.d0 to 16 bytes.

-finteger-4-integer-8

Promote all INTEGER(KIND=4) entities to an INTEGER(KIND=8) entities. If KIND=8 is unavailable, then an error will be issued. This option should be used with care and may not be suitable for your codes. Areas of possible concern include calls to external procedures, alignment in EQUIVALENCE and/or COMMON, generic interfaces, BOZ literal constant conversion, and I/O. Inspection of the intermediate representation of the translated Fortran code, produced

by -fdump-tree-original, is suggested.
-freal-4-real-8
-freal-4-real-10
-freal-4-real-16
-freal-8-real-4
-freal-8-real-10
-freal-8-real-16
Promote all REAL(KIND=M) entities to REAL(KIND=N) entities. If REAL(KIND=N)

is unavailable, then an error will be issued. The -freal-4- flags also affect the default real kind and the -freal-8- flags also the double-precision real kind. All other real-kind types are unaffected by this option. The promotion is also applied to real literal constants of default and double-precision kind and a specified kind number of 4 or 8, respectively. However, -fdefault-

real-8, -fdefault-real-10, -fdefault-real-10, and -fdefault-double-8

take precedence for the default and double-precision real kinds, both for real literal constants and for declarations without a kind number. Note that for REAL(KIND=KIND(1.0)) the literal may get promoted and then the result may get promoted again. These options should be used with care and may not be suitable for your codes. Areas of possible concern include calls to external procedures, alignment in EQUIVALENCE and/or COMMON, generic interfaces, BOZ literal constant conversion, and I/O and calls to intrinsic procedures when passing a value to the kind= dummy argument. Inspection of the intermediate representation of the translated Fortran code, produced by -fdump-fortran-original

or -fdump-tree-original, is suggested.

-std=std Specify the standard to which the program is expected to conform, which may be one of ‘f95’, ‘f2003’, ‘f2008’, ‘f2018’, ‘gnu’, or ‘legacy’. The default value for std is ‘gnu’, which specifies a superset of the latest Fortran standard that includes all of the extensions supported by GNU Fortran, although warnings will be given for obsolete extensions not recommended for use in new code. The ‘legacy’ value is equivalent but without the warnings for obsolete extensions, and may be useful for old non-standard programs. The ‘f95’, ‘f2003’, ‘f2008’, and ‘f2018’ values specify strict conformance to the Fortran 95, Fortran 2003, Fortran 2008 and Fortran 2018 standards, respectively; errors are given for all extensions beyond the relevant language standard, and warnings are given for the Fortran 77 features that are permitted but obsolescent in later standards. The deprecated option ‘-std=f2008ts’ acts as an alias for ‘-std=f2018’. It is only present for backwards compatibility with earlier gfortran versions and should not be used any more.

-ftest-forall-temp

Enhance test coverage by forcing most forall assignments to use temporary.

2.3 Enable and customize preprocessing

Many Fortran compilers including GNU Fortran allow passing the source code through a C preprocessor (CPP; sometimes also called the Fortran preprocessor, FPP) to allow for conditional compilation. In the case of GNU Fortran, this is the GNU C Preprocessor in the traditional mode. On systems with case-preserving file names, the preprocessor is automatically invoked if the filename extension is .F, .FOR, .FTN, .fpp, .FPP, .F90, .F95, .F03 or .F08. To manually invoke the preprocessor on any file, use -cpp, to disable preprocessing on files where the preprocessor is run automatically, use -nocpp.

If a preprocessed file includes another file with the Fortran INCLUDE statement, the included file is not preprocessed. To preprocess included files, use the equivalent preprocessor statement #include.

If GNU Fortran invokes the preprocessor, __GFORTRAN__is defined. The macros __GNUC_ _, __GNUC_MINOR__and __GNUC_PATCHLEVEL__can be used to determine the version of the compiler. See Section “Overview” in The C Preprocessor for details.

GNU Fortran supports a number of INTEGER and REAL kind types in additional to the kind types required by the Fortran standard. The availability of any given kind type is architecture dependent. The following pre-defined preprocessor macros can be used to conditionally include code for these additional kind types: __GFC_INT_1__, __GFC_INT_2_

_, __GFC_INT_8__, __GFC_INT_16__, __GFC_REAL_10__, and __GFC_REAL_16__.

While CPP is the de-facto standard for preprocessing Fortran code, Part 3 of the Fortran 95 standard (ISO/IEC 1539-3:1998) defines Conditional Compilation, which is not widely used and not directly supported by the GNU Fortran compiler. You can use the program

coco to preprocess such files (http://www.daniellnagle.com/coco.html).

The following options control preprocessing of Fortran code:

-cpp

-nocpp Enable preprocessing. The preprocessor is automatically invoked if the file

          extension is .fpp, .FPP, .F, .FOR, .FTN, .F90, .F95, .F03 or .F08. Use this

option to manually enable preprocessing of any kind of Fortran file.

To disable preprocessing of files with any of the above listed extensions, use the negative form: -nocpp. The preprocessor is run in traditional mode. Any restrictions of the file-format, especially the limits on line length, apply for preprocessed output as well, so it

          might be advisable to use the -ffree-line-length-none or -ffixed-line-
          length-none options.

-dM Instead of the normal output, generate a list of '#define' directives for all the macros defined during the execution of the preprocessor, including predefined macros. This gives you a way of finding out what is predefined in your version of the preprocessor. Assuming you have no file foo.f90, the command

                    touch foo.f90; gfortran -cpp -E -dM foo.f90

will show all the predefined macros.

-dD Like -dM except in two respects: it does not include the predefined macros, and it outputs both the #define directives and the result of preprocessing. Both kinds of output go to the standard output file.

-dN Like -dD, but emit only the macro names, not their expansions.

-dU Like dD except that only macros that are expanded, or whose definedness is tested in preprocessor directives, are output; the output is delayed until the use or test of the macro; and '#undef' directives are also output for macros tested but undefined at the time.

-dI Output '#include' directives in addition to the result of preprocessing.

-fworking-directory

Enable generation of linemarkers in the preprocessor output that will let the compiler know the current working directory at the time of preprocessing. When this option is enabled, the preprocessor will emit, after the initial linemarker, a second linemarker with the current working directory followed by two slashes. GCC will use this directory, when it is present in the preprocessed input, as the directory emitted as the current working directory in some debugging information formats. This option is implicitly enabled if debugging information is enabled, but this can be inhibited with the negated form -fno-working-directory. If the -P flag is present in the command line, this option has no effect, since no #line directives are emitted whatsoever.

-idirafter dir

Search dir for include files, but do it after all directories specified with -I and the standard system directories have been exhausted. dir is treated as a system include directory. If dir begins with =, then the = will be replaced by the sysroot

prefix; see --sysroot and -isysroot.
-imultilib dir

Use dir as a subdirectory of the directory containing target-specific C++ headers.

-iprefix prefix

Specify prefix as the prefix for subsequent -iwithprefix options. If the prefix represents a directory, you should include the final '/'.

-isysroot dir

This option is like the --sysroot option, but applies only to header files. See the --sysroot option for more information.

-iquote dir

Search dir only for header files requested with #include "file"; they are not searched for #include <file>, before all directories specified by -I and before the standard system directories. If dir begins with =, then the = will be replaced by the sysroot prefix; see --sysroot and -isysroot.

-isystem dir

Search dir for header files, after all directories specified by -I but before the standard system directories. Mark it as a system directory, so that it gets the same special treatment as is applied to the standard system directories. If dir begins with =, then the = will be replaced by the sysroot prefix; see --sysroot

and -isysroot.

-nostdinc

Do not search the standard system directories for header files. Only the directories you have specified with -I options (and the directory of the current file, if appropriate) are searched.

-undef Do not predefine any system-specific or GCC-specific macros. The standard predefined macros remain defined.

-Apredicate=answer

Make an assertion with the predicate predicate and answer answer. This form is preferred to the older form -A predicate(answer), which is still supported, because it does not use shell special characters.

-A-predicate=answer

Cancel an assertion with the predicate predicate and answer answer.

-C Do not discard comments. All comments are passed through to the output file, except for comments in processed directives, which are deleted along with the directive. You should be prepared for side effects when using -C; it causes the preprocessor to treat comments as tokens in their own right. For example, comments appearing at the start of what would be a directive line have the effect of turning that line into an ordinary source line, since the first token on the line is no longer a '#'. Warning: this currently handles C-Style comments only. The preprocessor does not yet recognize Fortran-style comments.

-CC Do not discard comments, including during macro expansion. This is like -C, except that comments contained within macros are also passed through to the output file where the macro is expanded. In addition to the side-effects of the -C option, the -CC option causes all C++- style comments inside a macro to be converted to C-style comments. This is to prevent later use of that macro from inadvertently commenting out the remainder of the source line. The -CC option is generally used to support lint comments. Warning: this currently handles C- and C++-Style comments only. The preprocessor does not yet recognize Fortran-style comments.

-Dname Predefine name as a macro, with definition 1.

-Dname=definition

The contents of definition are tokenized and processed as if they appeared during translation phase three in a '#define' directive. In particular, the definition will be truncated by embedded newline characters. If you are invoking the preprocessor from a shell or shell-like program you may need to use the shell’s quoting syntax to protect characters such as spaces that have a meaning in the shell syntax. If you wish to define a function-like macro on the command line, write its argument list with surrounding parentheses before the equals sign (if any).

Parentheses are meaningful to most shells, so you will need to quote the option.

With sh and csh, -D'name(args...)=definition' works.

-D and -U options are processed in the order they are given on the command line. All -imacros file and -include file options are processed after all -D and -U options.

-H Print the name of each header file used, in addition to other normal activities. Each name is indented to show how deep in the '#include' stack it is.

-P Inhibit generation of linemarkers in the output from the preprocessor. This might be useful when running the preprocessor on something that is not C code, and will be sent to a program which might be confused by the linemarkers.

-Uname Cancel any previous definition of name, either built in or provided with a -D option.

2.4 Options to request or suppress errors and warnings

Errors are diagnostic messages that report that the GNU Fortran compiler cannot compile the relevant piece of source code. The compiler will continue to process the program in an attempt to report further errors to aid in debugging, but will not produce any compiled output.

Warnings are diagnostic messages that report constructions which are not inherently erroneous but which are risky or suggest there is likely to be a bug in the program. Unless -Werror is specified, they do not prevent compilation of the program.

You can request many specific warnings with options beginning -W, for example -Wimplicit to request warnings on implicit declarations. Each of these specific warning options also has a negative form beginning -Wno- to turn off warnings; for example, -Wno-implicit. This manual lists only one of the two forms, whichever is not the default.

These options control the amount and kinds of errors and warnings produced by GNU Fortran:

-fmax-errors=n

Limits the maximum number of error messages to n, at which point GNU Fortran bails out rather than attempting to continue processing the source code. If n is 0, there is no limit on the number of error messages produced.

-fsyntax-only

Check the code for syntax errors, but do not actually compile it. This will generate module files for each module present in the code, but no other output file.

-Wpedantic
-pedantic

Issue warnings for uses of extensions to Fortran. -pedantic also applies to C-language constructs where they occur in GNU Fortran source files, such as use of ‘\e’ in a character constant within a directive like #include. Valid Fortran programs should compile properly with or without this option. However, without this option, certain GNU extensions and traditional Fortran features are supported as well. With this option, many of them are rejected.

Some users try to use -pedantic to check programs for conformance. They soon find that it does not do quite what they want—it finds some nonstandard practices, but not all. However, improvements to GNU Fortran in this area are welcome. This should be used in conjunction with -std=f95, -std=f2003, -std=f2008

or -std=f2018.
-pedantic-errors

Like -pedantic, except that errors are produced rather than warnings.

-Wall Enables commonly used warning options pertaining to usage that we recommend avoiding and that we believe are easy to avoid. This currently

includes -Waliasing, -Wampersand, -Wconversion, -Wsurprising,
-Wc-binding-type, -Wintrinsics-std, -Wtabs, -Wintrinsic-shadow,
-Wline-truncation,  -Wtarget-lifetime,  -Winteger-division,
-Wreal-q-constant, -Wunused and -Wundefined-do-loop.
-Waliasing

Warn about possible aliasing of dummy arguments. Specifically, it warns if the same actual argument is associated with a dummy argument with INTENT(IN) and a dummy argument with INTENT(OUT) in a call with an explicit interface. The following example will trigger the warning.

        interface
          subroutine bar(a,b)
            integer, intent(in) :: a
            integer, intent(out) :: b
          end subroutine
        end interface
        integer :: a
        call bar(a,a)
-Wampersand

Warn about missing ampersand in continued character constants. The warning

is given with -Wampersand, -pedantic, -std=f95, -std=f2003, -std=f2008

and -std=f2018. Note: With no ampersand given in a continued character constant, GNU Fortran assumes continuation at the first non-comment, nonwhitespace character after the ampersand that initiated the continuation.

-Warray-temporaries

Warn about array temporaries generated by the compiler. The information generated by this warning is sometimes useful in optimization, in order to avoid such temporaries.

-Wc-binding-type

Warn if the a variable might not be C interoperable. In particular, warn if the variable has been declared using an intrinsic type with default kind instead of using a kind parameter defined for C interoperability in the intrinsic ISO_C_ Binding module. This option is implied by -Wall.

-Wcharacter-truncation

Warn when a character assignment will truncate the assigned string.

-Wline-truncation

Warn when a source code line will be truncated. This option is implied by -Wall. For free-form source code, the default is -Werror=line-truncation such that truncations are reported as error.

-Wconversion

Warn about implicit conversions that are likely to change the value of the expression after conversion. Implied by -Wall.

-Wconversion-extra

Warn about implicit conversions between different types and kinds. This option does not imply -Wconversion.

-Wextra Enables some warning options for usages of language features which may be problematic. This currently includes -Wcompare-reals, -Wunused-parameter

and -Wdo-subscript.
-Wfrontend-loop-interchange

Warn when using -ffrontend-loop-interchange for performing loop interchanges.

-Wimplicit-interface

Warn if a procedure is called without an explicit interface. Note this only checks that an explicit interface is present. It does not check that the declared interfaces are consistent across program units.

-Wimplicit-procedure

Warn if a procedure is called that has neither an explicit interface nor has been declared as EXTERNAL.

-Winteger-division

Warn if a constant integer division truncates its result. As an example, 3/5 evaluates to 0.

-Wintrinsics-std

Warn if gfortran finds a procedure named like an intrinsic not available in the currently selected standard (with -std) and treats it as EXTERNAL procedure because of this. -fall-intrinsics can be used to never trigger this behavior and always link to the intrinsic regardless of the selected standard.

-Wno-overwrite-recursive

Do not warn when -fno-automatic is used with -frecursive. Recursion will be broken if the relevant local variables do not have the attribute AUTOMATIC explicitly declared. This option can be used to suppress the warning when it is known that recursion is not broken. Useful for build environments that use

-Werror.
-Wreal-q-constant

Produce a warning if a real-literal-constant contains a q exponent-letter.

-Wsurprising

Produce a warning when “suspicious” code constructs are encountered. While technically legal these usually indicate that an error has been made.

This currently produces a warning under the following circumstances:

• An INTEGER SELECT construct has a CASE that can never be matched as its lower value is greater than its upper value.

• A LOGICAL SELECT construct has three CASE statements.

• A TRANSFER specifies a source that is shorter than the destination.

• The type of a function result is declared more than once with the same type. If -pedantic or standard-conforming mode is enabled, this is an error.

• A CHARACTER variable is declared with negative length.

• With -fopenmp, for fixed-form source code, when an omx vendor-extension sentinel is encountered. (The equivalent ompx, used in free-form source code, is diagnosed by default.)

-Wtabs By default, tabs are accepted as whitespace, but tabs are not members of the Fortran Character Set. For continuation lines, a tab followed by a digit between 1 and 9 is supported. -Wtabs will cause a warning to be issued if a tab is

encountered. Note, -Wtabs is active for -pedantic, -std=f95, -std=f2003,
-std=f2008, -std=f2018 and -Wall.
-Wundefined-do-loop

Warn if a DO loop with step either 1 or -1 yields an underflow or an overflow during iteration of an induction variable of the loop. This option is implied by

-Wall.
-Wunderflow

Produce a warning when numerical constant expressions are encountered, which yield an UNDERFLOW during compilation. Enabled by default.

-Wintrinsic-shadow

Warn if a user-defined procedure or module procedure has the same name as an intrinsic; in this case, an explicit interface or EXTERNAL or INTRINSIC declaration might be needed to get calls later resolved to the desired intrinsic/procedure. This option is implied by -Wall.

-Wuse-without-only

Warn if a USE statement has no ONLY qualifier and thus implicitly imports all public entities of the used module.

-Wunused-dummy-argument

Warn about unused dummy arguments. This option is implied by -Wall.

-Wunused-parameter

Contrary to gcc’s meaning of -Wunused-parameter, gfortran’s implementation of this option does not warn about unused dummy arguments (see

-Wunused-dummy-argument), but about unused PARAMETER values. -Wunused-
parameter is implied by -Wextra if also -Wunused or -Wall is used.
-Walign-commons

By default, gfortran warns about any occasion of variables being padded for proper alignment inside a COMMON block. This warning can be turned off via

-Wno-align-commons. See also -falign-commons.

-Wfunction-elimination

Warn if any calls to impure functions are eliminated by the optimizations enabled by the -ffrontend-optimize option. This option is implied by -Wextra.

-Wrealloc-lhs

Warn when the compiler might insert code to for allocation or reallocation of an allocatable array variable of intrinsic type in intrinsic assignments. In hot loops, the Fortran 2003 reallocation feature may reduce the performance. If the array is already allocated with the correct shape, consider using a wholearray array-spec (e.g. (:,:,:)) for the variable on the left-hand side to prevent the reallocation check. Note that in some cases the warning is shown, even if the compiler will optimize reallocation checks away. For instance, when the right-hand side contains the same variable multiplied by a scalar. See also

-frealloc-lhs.
-Wrealloc-lhs-all

Warn when the compiler inserts code to for allocation or reallocation of an allocatable variable; this includes scalars and derived types.

-Wcompare-reals

Warn when comparing real or complex types for equality or inequality. This option is implied by -Wextra.

-Wtarget-lifetime

Warn if the pointer in a pointer assignment might be longer than the its target. This option is implied by -Wall.

-Wzerotrip

Warn if a DO loop is known to execute zero times at compile time. This option is implied by -Wall.

-Wdo-subscript

Warn if an array subscript inside a DO loop could lead to an out-of-bounds access even if the compiler cannot prove that the statement is actually executed, in cases like

        real a(3)
        do i=1,4
          if (condition(i)) then
            a(i) = 1.2
          end if
        end do

This option is implied by -Wextra.

-Werror Turns all warnings into errors.

See Section “Options to Request or Suppress Errors and Warnings” in Using the GNU Compiler Collection (GCC), for information on more options offered by the GBE shared by gfortran, gcc and other GNU compilers.

Some of these have no effect when compiling programs written in Fortran.

2.5 Options for debugging your program or GNU Fortran

GNU Fortran has various special options that are used for debugging either your program or the GNU Fortran compiler.

-fdump-fortran-original

Output the internal parse tree after translating the source program into internal representation. This option is mostly useful for debugging the GNU Fortran compiler itself. The output generated by this option might change between releases. This option may also generate internal compiler errors for features which have only recently been added.

-fdump-fortran-optimized

Output the parse tree after front-end optimization. Mostly useful for debugging the GNU Fortran compiler itself. The output generated by this option might change between releases. This option may also generate internal compiler errors for features which have only recently been added.

-fdump-parse-tree

Output the internal parse tree after translating the source program into internal representation. Mostly useful for debugging the GNU Fortran compiler itself. The output generated by this option might change between releases. This option may also generate internal compiler errors for features which have only recently been added. This option is deprecated; use -fdump-fortran-original instead.

-fdebug-aux-vars

Renames internal variables created by the gfortran front end and makes them accessible to a debugger. The name of the internal variables then start with upper-case letters followed by an underscore. This option is useful for debugging the compiler’s code generation together with -fdump-tree-original and enabling debugging of the executable program by using -g or -ggdb3.

-fdump-fortran-global

Output a list of the global identifiers after translating into middle-end representation. Mostly useful for debugging the GNU Fortran compiler itself. The output generated by this option might change between releases. This option may also generate internal compiler errors for features which have only recently been added.

-ffpe-trap=list

Specify a list of floating point exception traps to enable. On most systems, if a floating point exception occurs and the trap for that exception is enabled, a SIGFPE signal will be sent and the program being aborted, producing a core file useful for debugging. list is a (possibly empty) comma-separated list of the following exceptions: ‘invalid’ (invalid floating point operation, such as SQRT(-1.0)), ‘zero’ (division by zero), ‘overflow’ (overflow in a floating point operation), ‘underflow’ (underflow in a floating point operation), ‘inexact’ (loss of precision during operation), and ‘denormal’ (operation performed on a denormal value). The first five exceptions correspond to the five IEEE 754 exceptions, whereas the last one (‘denormal’) is not part of the IEEE 754 standard but is available on some common architectures such as x86.

The first three exceptions (‘invalid’, ‘zero’, and ‘overflow’) often indicate serious errors, and unless the program has provisions for dealing with these exceptions, enabling traps for these three exceptions is probably a good idea. If the option is used more than once in the command line, the lists will be joined: ’ffpe-trap=list1 ffpe-trap=list2’ is equivalent to ffpe-trap=list1,list2. Note that once enabled an exception cannot be disabled (no negative form). Many, if not most, floating point operations incur loss of precision due to rounding, and hence the ffpe-trap=inexact is likely to be uninteresting in practice. By default no exception traps are enabled.

-ffpe-summary=list

Specify a list of floating-point exceptions, whose flag status is printed to ERROR_ UNIT when invoking STOP and ERROR STOP. list can be either ‘none’, ‘all’ or a comma-separated list of the following exceptions: ‘invalid’, ‘zero’, ‘overflow’,

‘underflow’, ‘inexact’ and ‘denormal’. (See -ffpe-trap for a description of

the exceptions.) If the option is used more than once in the command line, only the last one will be used. By default, a summary for all exceptions but ‘inexact’ is shown.

-fno-backtrace

When a serious runtime error is encountered or a deadly signal is emitted (segmentation fault, illegal instruction, bus error, floating-point exception, and the other POSIX signals that have the action ‘core’), the Fortran runtime library tries to output a backtrace of the error. -fno-backtrace disables the backtrace generation. This option only has influence for compilation of the Fortran main program.

See Section “Options for Debugging Your Program or GCC” in Using the GNU Compiler Collection (GCC), for more information on debugging options.

2.6 Options for directory search

These options affect how GNU Fortran searches for files specified by the INCLUDE directive and where it searches for previously compiled modules.

It also affects the search paths used by cpp when used to preprocess Fortran source.

-Idir These affect interpretation of the INCLUDE directive (as well as of the #include directive of the cpp preprocessor). Also note that the general behavior of -I and INCLUDE is pretty much the same as of -I with #include in the cpp preprocessor, with regard to looking for header.gcc files and other such things. This path is also used to search for .mod files when previously compiled modules are required by a USE statement. See Section “Options for Directory Search” in Using the GNU Compiler Collection (GCC), for information on the -I option.

-Jdir This option specifies where to put .mod files for compiled modules. It is also added to the list of directories to searched by an USE statement. The default is the current directory.

-fintrinsic-modules-path dir

This option specifies the location of pre-compiled intrinsic modules, if they are not in the default location expected by the compiler.

2.7 Influencing the linking step

These options come into play when the compiler links object files into an executable output file. They are meaningless if the compiler is not doing a link step.

-static-libgfortran

On systems that provide libgfortran as a shared and a static library, this option forces the use of the static version. If no shared version of libgfortran was built when the compiler was configured, this option has no effect.

-static-libquadmath

On systems that provide libquadmath as a shared and a static library, this option forces the use of the static version. If no shared version of libquadmath was built when the compiler was configured, this option has no effect. Please note that the libquadmath runtime library is licensed under the GNU Lesser General Public License (LGPL), and linking it statically introduces requirements when redistributing the resulting binaries.

2.8 Influencing runtime behavior

These options affect the runtime behavior of programs compiled with GNU Fortran.

-fconvert=conversion

Specify the representation of data for unformatted files. Valid values for conversion on most systems are: ‘native’, the default; ‘swap’, swap between bigand little-endian; ‘big-endian’, use big-endian representation for unformatted files; ‘little-endian’, use little-endian representation for unformatted files. On POWER systems which suppport -mabi=ieeelongdouble, there are additional options, which can be combined with others with commas. Those are

    -fconvert=r16_ieee Use IEEE 128-bit format for REAL(KIND=16).
    -fconvert=r16_ibm Use IBM long double format for REAL(KIND=16).

This option has an effect only when used in the main program. The CONVERT specifier and the GFORTRAN CONVERT UNIT environment variable override the default specified by -fconvert.

-frecord-marker=length

Specify the length of record markers for unformatted files. Valid values for length are 4 and 8. Default is 4. This is different from previous versions of gfortran, which specified a default record marker length of 8 on most systems. If you want to read or write files compatible with earlier versions of gfortran,

use -frecord-marker=8.

-fmax-subrecord-length=length

Specify the maximum length for a subrecord. The maximum permitted value for length is 2147483639, which is also the default. Only really useful for use by the gfortran testsuite.

-fsign-zero

When enabled, floating point numbers of value zero with the sign bit set are written as negative number in formatted output and treated as negative in the SIGN intrinsic. -fno-sign-zero does not print the negative sign of zero values (or values rounded to zero for I/O) and regards zero as positive number in the SIGN intrinsic for compatibility with Fortran 77. The default is -fsign-zero.

2.9 Options for code generation conventions

These machine-independent options control the interface conventions used in code generation.

Most of them have both positive and negative forms; the negative form of -ffoo would be -fno-foo. In the table below, only one of the forms is listed—the one which is not the default. You can figure out the other form by either removing no- or adding it.

-fno-automatic

Treat each program unit (except those marked as RECURSIVE) as if the SAVE statement were specified for every local variable and array referenced in it. Does not affect common blocks. (Some Fortran compilers provide this option under the name -static or -save.) The default, which is -fautomatic, uses the stack for local variables smaller than the value given by -fmax-stack-var-size. Use the option -frecursive to use no static memory.

Local variables or arrays having an explicit SAVE attribute are silently ignored unless the -pedantic option is added.

-ff2c Generate code designed to be compatible with code generated by g77 and f2c.

The calling conventions used by g77 (originally implemented in f2c) require functions that return type default REAL to actually return the C type double, and functions that return type COMPLEX to return the values via an extra argument in the calling sequence that points to where to store the return value. Under the default GNU calling conventions, such functions simply return their results as they would in GNU C—default REAL functions return the C type float, and COMPLEX functions return the GNU C type complex. Additionally, this option implies the -fsecond-underscore option, unless -fno-second-underscore is explicitly requested.

This does not affect the generation of code that interfaces with the libgfortran library.

Caution: It is not a good idea to mix Fortran code compiled with -ff2c with code compiled with the default -fno-f2c calling conventions as, calling COMPLEX or default REAL functions between program parts which were compiled with different calling conventions will break at execution time.

Caution: This will break code which passes intrinsic functions of type default REAL or COMPLEX as actual arguments, as the library implementations use the -fno-f2c calling conventions.

-fno-underscoring

Do not transform names of entities specified in the Fortran source file by appending underscores to them. With -funderscoring in effect, GNU Fortran appends one underscore to external names. This is done to ensure compatibility with code produced by many UNIX Fortran compilers. Caution: The default behavior of GNU Fortran is incompatible with f2c and g77, please use the -ff2c option if you want object files compiled with GNU Fortran to be compatible with object code created with these tools. Use of -fno-underscoring is not recommended unless you are experimenting with issues such as integration of GNU Fortran into existing system environments (vis-à-vis existing libraries, tools, and so on). For example, with -funderscoring, and assuming that j() and max_count() are external functions while my_var and lvar are local variables, a statement like

      I = J() + MAX_COUNT (MY_VAR, LVAR)

is implemented as something akin to:

      i = j_() + max_count_(&my_var, &lvar);

With -fno-underscoring, the same statement is implemented as:

      i = j() + max_count(&my_var, &lvar);

Use of -fno-underscoring allows direct specification of user-defined names while debugging and when interfacing GNU Fortran code with other languages. Note that just because the names match does not mean that the interface implemented by GNU Fortran for an external name matches the interface implemented by some other language for that same name. That is, getting code produced by GNU Fortran to link to code produced by some other compiler using this or any other method can be only a small part of the overall solution— getting the code generated by both compilers to agree on issues other than naming can require significant effort, and, unlike naming disagreements, linkers normally cannot detect disagreements in these other areas. Also, note that with -fno-underscoring, the lack of appended underscores introduces the very real possibility that a user-defined external name will conflict with a name in a system library, which could make finding unresolved-reference bugs quite difficult in some cases—they might occur at program run time, and show up only as buggy behavior at run time. In future versions of GNU Fortran we hope to improve naming and linking issues so that debugging always involves using the names as they appear in the source, even if the names as seen by the linker are mangled to prevent accidental linking between procedures with incompatible interfaces.

-fsecond-underscore

By default, GNU Fortran appends an underscore to external names. If this option is used GNU Fortran appends two underscores to names with underscores and one underscore to external names with no underscores. GNU Fortran also appends two underscores to internal names with underscores to avoid naming collisions with external names. This option has no effect if -fno-underscoring is in effect. It is implied by the -ff2c option. Otherwise, with this option, an external name such as MAX_COUNT is implemented as a reference to the link-time external symbol max_count__, instead of max_count_. This is required for compatibility with g77 and f2c, and is implied by use of the -ff2c option.

-fcoarray=<keyword>

‘none’ Disable coarray support; using coarray declarations and imagecontrol statements will produce a compile-time error. (Default)

‘single’ Single-image mode, i.e. num_images() is always one.

‘lib’ Library-based coarray parallelization; a suitable GNU Fortran coarray library needs to be linked.

-fcheck=<keyword>

Enable the generation of run-time checks; the argument shall be a commadelimited list of the following keywords. Prefixing a check with no- disables it if it was activated by a previous specification.

‘all’ Enable all run-time test of -fcheck.

‘array-temps’

Warns at run time when for passing an actual argument a temporary array had to be generated. The information generated by this warning is sometimes useful in optimization, in order to avoid such temporaries. Note: The warning is only printed once per location.

‘bits’ Enable generation of run-time checks for invalid arguments to the bit manipulation intrinsics.

‘bounds’ Enable generation of run-time checks for array subscripts and against the declared minimum and maximum values. It also checks array indices for assumed and deferred shape arrays against the actual allocated bounds and ensures that all string lengths are equal for character array constructors without an explicit typespec. Some checks require that -fcheck=bounds is set for the compilation of the main program. Note: In the future this may also include other forms of checking, e.g., checking substring references.

‘do’ Enable generation of run-time checks for invalid modification of loop iteration variables.

‘mem’ Enable generation of run-time checks for memory allocation. Note: This option does not affect explicit allocations using the ALLOCATE statement, which will be always checked.

‘pointer’ Enable generation of run-time checks for pointers and allocatables.

‘recursion’

Enable generation of run-time checks for recursively called subroutines and functions which are not marked as recursive. See also -frecursive. Note: This check does not work for OpenMP programs and is disabled if used together with -frecursive and

          -fopenmp.

Example: Assuming you have a file foo.f90, the command

        gfortran -fcheck=all,no-array-temps foo.f90

will compile the file with all checks enabled as specified above except warnings for generated array temporaries.

-fbounds-check

Deprecated alias for -fcheck=bounds.

-ftail-call-workaround
-ftail-call-workaround=n

Some C interfaces to Fortran codes violate the gfortran ABI by omitting the hidden character length arguments as described in See Section 6.4.2 [Argument passing conventions], page 81. This can lead to crashes because pushing arguments for tail calls can overflow the stack.

To provide a workaround for existing binary packages, this option disables tail call optimization for gfortran procedures with character arguments. With -ftail-call-workaround=2 tail call optimization is disabled in all gfortran procedures with character arguments, with -ftail-call-workaround=1 or equivalent -ftail-call-workaround only in gfortran procedures with character arguments that call implicitly prototyped procedures.

Using this option can lead to problems including crashes due to insufficient stack space.

It is very strongly recommended to fix the code in question. The -fc-prototypes-external option can be used to generate prototypes which conform to gfortran’s ABI, for inclusion in the source code.

Support for this option will likely be withdrawn in a future release of gfortran.

The negative form, -fno-tail-call-workaround or equivalent -ftail-call-

workaround=0, can be used to disable this option.

Default is currently -ftail-call-workaround, this will change in future releases.

-fcheck-array-temporaries
Deprecated alias for -fcheck=array-temps.

-fmax-array-constructor=n

This option can be used to increase the upper limit permitted in array constructors. The code below requires this option to expand the array at compile time.

      program test
      implicit none
      integer j
      integer, parameter :: n = 100000
      integer, parameter :: i(n) = (/ (2*j, j = 1, n) /)
      print '(10(I0,1X))', i
      end program test

Caution: This option can lead to long compile times and excessively large object files. The default value for n is 65535.

-fmax-stack-var-size=n

This option specifies the size in bytes of the largest array that will be put on the stack; if the size is exceeded static memory is used (except in procedures marked as RECURSIVE). Use the option -frecursive to allow for recursive procedures which do not have a RECURSIVE attribute or for parallel programs. Use -fno-automatic to never use the stack. This option currently only affects local arrays declared with constant bounds, and may not apply to all character variables. Future versions of GNU Fortran may improve this behavior. The default value for n is 65536.

-fstack-arrays

Adding this option will make the Fortran compiler put all arrays of unknown size and array temporaries onto stack memory. If your program uses very large local arrays it is possible that you will have to extend your runtime limits for stack memory on some operating systems. This flag is enabled by default at optimization level -Ofast unless -fmax-stack-var-size is specified.

-fpack-derived

This option tells GNU Fortran to pack derived type members as closely as possible. Code compiled with this option is likely to be incompatible with code compiled without this option, and may execute slower.

-frepack-arrays

In some circumstances GNU Fortran may pass assumed shape array sections via a descriptor describing a noncontiguous area of memory. This option adds code to the function prologue to repack the data into a contiguous block at runtime. This should result in faster accesses to the array. However it can introduce significant overhead to the function call, especially when the passed data is noncontiguous.

-fshort-enums

This option is provided for interoperability with C code that was compiled with the -fshort-enums option. It will make GNU Fortran choose the smallest INTEGER kind a given enumerator set will fit in, and give all its enumerators this kind.

-finline-arg-packing

When passing an assumed-shape argument of a procedure as actual argument to an assumed-size or explicit size or as argument to a procedure that does not have an explicit interface, the argument may have to be packed, that is put into contiguous memory. An example is the call to foo in

        subroutine foo(a)
           real, dimension(*) :: a
        end subroutine foo
        subroutine bar(b)
           real, dimension(:) :: b
           call foo(b)
        end subroutine bar

When -finline-arg-packing is in effect, this packing will be performed by inline code. This allows for more optimization while increasing code size.

-finline-arg-packing is implied by any of the -O options except when optimizing for size via -Os. If the code contains a very large number of argument that have to be packed, code size and also compilation time may become excessive. If that is the case, it may be better to disable this option. Instances of packing can be found by using -Warray-temporaries.

-fexternal-blas

This option will make gfortran generate calls to BLAS functions for some matrix operations like MATMUL, instead of using our own algorithms, if the size of the matrices involved is larger than a given limit (see -fblas-matmul-limit). This may be profitable if an optimized vendor BLAS library is available. The BLAS library will have to be specified at link time.

-fblas-matmul-limit=n

Only significant when -fexternal-blas is in effect. Matrix multiplication of matrices with size larger than (or equal to) n will be performed by calls to BLAS functions, while others will be handled by gfortran internal algorithms. If the matrices involved are not square, the size comparison is performed using the geometric mean of the dimensions of the argument and result matrices.

The default value for n is 30.

-finline-matmul-limit=n

When front-end optimization is active, some calls to the MATMUL intrinsic function will be inlined. This may result in code size increase if the size of the matrix cannot be determined at compile time, as code for both cases is generated. Setting -finline-matmul-limit=0 will disable inlining in all cases. Setting this option with a value of n will produce inline code for matrices with size up to n. If the matrices involved are not square, the size comparison is performed using the geometric mean of the dimensions of the argument and result matrices.

The default value for n is 30. The -fblas-matmul-limit can be used to change this value.

-frecursive

Allow indirect recursion by forcing all local arrays to be allocated on the stack. This flag cannot be used together with -fmax-stack-var-size= or

-fno-automatic.
-finit-local-zero
-finit-derived
-finit-integer=n
-finit-real=<zero|inf|-inf|nan|snan>
-finit-logical=<true|false>
-finit-character=n

The -finit-local-zero option instructs the compiler to initialize local INTEGER, REAL, and COMPLEX variables to zero, LOGICAL variables to false, and CHARACTER variables to a string of null bytes. Finer-grained initialization options are provided by the -finit-integer=n, -finit-real=<zero|inf|-inf|nan|snan> (which also initializes the real and

imaginary parts of local COMPLEX variables), -finit-logical=<true|false>,

and -finit-character=n (where n is an ASCII character value) options. With -finit-derived, components of derived type variables will be initialized according to these flags. Components whose type is not covered by an explicit -finit-* flag will be treated as described above with -finit-local-zero. These options do not initialize

• objects with the POINTER attribute

• allocatable arrays

• variables that appear in an EQUIVALENCE statement.

(These limitations may be removed in future releases). Note that the -finit-real=nan option initializes REAL and COMPLEX variables with a quiet NaN. For a signalling NaN use -finit-real=snan; note, however, that compile-time optimizations may convert them into quiet NaN and that trapping needs to be enabled (e.g. via -ffpe-trap). The -finit-integer option will parse the value into an integer of type INTEGER(kind=C_LONG) on the host. Said value is then assigned to the integer variables in the Fortran code, which might result in wraparound if the value is too large for the kind. Finally, note that enabling any of the -finit-* options will silence warnings that would have been emitted by -Wuninitialized for the affected local variables.

-falign-commons

By default, gfortran enforces proper alignment of all variables in a COMMON block by padding them as needed. On certain platforms this is mandatory, on others it increases performance. If a COMMON block is not declared with consistent data types everywhere, this padding can cause trouble, and -fno-align-commons can be used to disable automatic alignment. The same form of this option should be used for all files that share a COMMON block. To avoid potential alignment issues in COMMON blocks, it is recommended to order objects from largest to smallest.

-fno-protect-parens

By default the parentheses in expression are honored for all optimization levels such that the compiler does not do any re-association. Using -fno-protect-parens allows the compiler to reorder REAL and COMPLEX expressions to produce faster code. Note that for the re-association optimization -fno-signed-zeros and -fno-trapping-math need to be in effect. The parentheses protection is enabled by default, unless -Ofast is given.

-frealloc-lhs

An allocatable left-hand side of an intrinsic assignment is automatically (re)allocated if it is either unallocated or has a different shape. The option is enabled by default except when -std=f95 is given. See also -Wrealloc-lhs.

-faggressive-function-elimination

Functions with identical argument lists are eliminated within statements, regardless of whether these functions are marked PURE or not. For example, in

        a = f(b,c) + f(b,c)

there will only be a single call to f. This option only works if -ffrontend-optimize is in effect.

-ffrontend-optimize

This option performs front-end optimization, based on manipulating parts the Fortran parse tree. Enabled by default by any -O option except -O0 and -Og. Optimizations enabled by this option include:

• inlining calls to MATMUL,

• elimination of identical function calls within expressions,

• removing unnecessary calls to TRIM in comparisons and assignments,

  • replacing TRIM(a) with a(1:LEN_TRIM(a)) and

• short-circuiting of logical operators (.AND. and .OR.).

It can be deselected by specifying -fno-frontend-optimize.

-ffrontend-loop-interchange

Attempt to interchange loops in the Fortran front end where profitable. Enabled by default by any -O option. At the moment, this option only affects FORALL and DO CONCURRENT statements with several forall triplets.

See Section “Options for Code Generation Conventions” in Using the GNU Compiler Collection (GCC), for information on more options offered by the GBE shared by gfortran, gcc, and other GNU compilers.

2.10 Options for interoperability with other languages

-fc-prototypes This option will generate C prototypes from BIND(C) variable declarations, types and procedure interfaces and writes them to standard output. ENUM is not yet supported.

The generated prototypes may need inclusion of an appropriate header, such as <stdint.h> or <stdlib.h>. For types which are not specified using the appropriate kind from the iso_c_binding module, a warning is added as a comment to the code. For function pointers, a pointer to a function returning int without an explicit argument list is generated. Example of use:

      $ gfortran -fc-prototypes -fsyntax-only foo.f90 > foo.h

where the C code intended for interoperating with the Fortran code then uses

#include "foo.h".

-fc-prototypes-external This option will generate C prototypes from external functions and subroutines and write them to standard output. This may be useful for making sure that C bindings to Fortran code are correct. This option does not generate prototypes for BIND(C) procedures, use -fc-prototypes for that. The generated prototypes may need inclusion of an appropriate header, such

as <stdint.h> or <stdlib.h>.

This is primarily meant for legacy code to ensure that existing C bindings match what gfortran emits. The generated C prototypes should be correct for the current version of the compiler, but may not match what other compilers or earlier versions of gfortran need. For new developments, use of the BIND(C) features is recommended. Example of use:

      $ gfortran -fc-prototypes-external -fsyntax-only foo.f > foo.h

where the C code intended for interoperating with the Fortran code then uses

#include "foo.h".

2.11 Environment variables affecting gfortran

The gfortran compiler currently does not make use of any environment variables to control its operation above and beyond those that affect the operation of gcc.

See Section “Environment Variables Affecting GCC” in Using the GNU Compiler Collection (GCC), for information on environment variables.

See Chapter 3 [Runtime], page 35, for environment variables that affect the run-time behavior of programs compiled with GNU Fortran.

3 Runtime: Influencing runtime behavior with environment variables

The behavior of the gfortran can be influenced by environment variables.

Malformed environment variables are silently ignored.

3.1 TMPDIR—Directory for scratch files

When opening a file with STATUS='SCRATCH', GNU Fortran tries to create the file in one of the potential directories by testing each directory in the order below.

1. The environment variable TMPDIR, if it exists.

2. On the MinGW target, the directory returned by the GetTempPath function. Alternatively, on the Cygwin target, the TMP and TEMP environment variables, if they exist, in that order.

3. The P_tmpdir macro if it is defined, otherwise the directory /tmp.

3.2 GFORTRAN_STDIN_UNIT—Unit number for standard input

This environment variable can be used to select the unit number preconnected to standard input. This must be a positive integer. The default value is 5.

3.3 GFORTRAN_STDOUT_UNIT—Unit number for standard output

This environment variable can be used to select the unit number preconnected to standard output. This must be a positive integer. The default value is 6.

3.4 GFORTRAN_STDERR_UNIT—Unit number for standard error

This environment variable can be used to select the unit number preconnected to standard error. This must be a positive integer. The default value is 0.

3.5 GFORTRAN_UNBUFFERED_ALL—Do not buffer I/O on all units

This environment variable controls whether all I/O is unbuffered. If the first letter is ‘y’, ‘Y’ or ‘1’, all I/O is unbuffered. This will slow down small sequential reads and writes. If the first letter is ‘n’, ‘N’ or ‘0’, I/O is buffered. This is the default.

3.6 GFORTRAN_UNBUFFERED_PRECONNECTED—Do not buffer I/O on

preconnected units

The environment variable named GFORTRAN_UNBUFFERED_PRECONNECTED controls whether I/O on a preconnected unit (i.e. STDOUT or STDERR) is unbuffered. If the first letter is ‘y’, ‘Y’ or ‘1’, I/O is unbuffered. This will slow down small sequential reads and writes. If the first letter is ‘n’, ‘N’ or ‘0’, I/O is buffered. This is the default.

3.7 GFORTRAN_SHOW_LOCUS—Show location for runtime errors

If the first letter is ‘y’, ‘Y’ or ‘1’, filename and line numbers for runtime errors are printed. If the first letter is ‘n’, ‘N’ or ‘0’, do not print filename and line numbers for runtime errors. The default is to print the location.

3.8 GFORTRAN_OPTIONAL_PLUS—Print leading + where

permitted

If the first letter is ‘y’, ‘Y’ or ‘1’, a plus sign is printed where permitted by the Fortran standard. If the first letter is ‘n’, ‘N’ or ‘0’, a plus sign is not printed in most cases. Default is not to print plus signs.

3.9 GFORTRAN_LIST_SEPARATOR—Separator for list output

This environment variable specifies the separator when writing list-directed output. It may contain any number of spaces and at most one comma. If you specify this on the command line, be sure to quote spaces, as in

      $ GFORTRAN_LIST_SEPARATOR='  ,  ' ./a.out

when a.out is the compiled Fortran program that you want to run. Default is a single space.

3.10 GFORTRAN_CONVERT_UNIT—Set conversion for unformatted I/O

By setting the GFORTRAN_CONVERT_UNIT variable, it is possible to change the representation of data for unformatted files. The syntax for the GFORTRAN_CONVERT_UNIT variable for most systems is:

      GFORTRAN_CONVERT_UNIT: mode | mode ';' exception | exception ;
      mode: 'native' | 'swap' | 'big_endian' | 'little_endian' ;
      exception: mode ':' unit_list | unit_list ;
      unit_list: unit_spec | unit_list unit_spec ;
      unit_spec: INTEGER | INTEGER '-' INTEGER ;

The variable consists of an optional default mode, followed by a list of optional exceptions, which are separated by semicolons from the preceding default and each other. Each exception consists of a format and a comma-separated list of units. Valid values for the modes are the same as for the CONVERT specifier:

NATIVE Use the native format. This is the default.

SWAP Swap between little- and big-endian.

LITTLE_ENDIAN Use the little-endian format for unformatted files. BIG_ENDIAN Use the big-endian format for unformatted files.

For POWER systems which support -mabi=ieeelongdouble, there are additional options, which can be combined with the others with commas. Those are

R16_IEEE Use IEEE 128-bit format for REAL(KIND=16).

    R16_IBM Use IBM long double format for REAL(KIND=16).

A missing mode for an exception is taken to mean BIG_ENDIAN. Examples of values for

GFORTRAN_CONVERT_UNIT are:

'big_endian' Do all unformatted I/O in big endian mode.

'little_endian;native:10-20,25' Do all unformatted I/O in little endian mode, except for units 10 to 20 and 25, which are in native format. '10-20' Units 10 to 20 are big-endian, the rest is native. 'big_endian,r16_ibm' Do all unformatted I/O in big-endian mode and use IBM long double for output of REAL(KIND=16) values.

Setting the environment variables should be done on the command line or via the export command for sh-compatible shells and via setenv for csh-compatible shells.

Example for sh:

      $ gfortran foo.f90
      $ GFORTRAN_CONVERT_UNIT='big_endian;native:10-20' ./a.out

Example code for csh:

      % gfortran foo.f90
      % setenv GFORTRAN_CONVERT_UNIT 'big_endian;native:10-20'
      % ./a.out

Using anything but the native representation for unformatted data carries a significant speed overhead. If speed in this area matters to you, it is best if you use this only for data that needs to be portable.

See Section 5.1.17 [CONVERT specifier], page 53, for an alternative way to specify the data representation for unformatted files. See Section 2.8 [Runtime Options], page 24, for setting a default data representation for the whole program. The CONVERT specifier overrides the -fconvert compile options.

Note that the values specified via the GFORTRAN CONVERT UNIT environment variable will override the CONVERT specifier in the open statement. This is to give control over data formats to users who do not have the source code of their program available.

3.11 GFORTRAN_ERROR_BACKTRACE—Show backtrace on run-time errors

If the GFORTRAN_ERROR_BACKTRACE variable is set to ‘y’, ‘Y’ or ‘1’ (only the first letter is relevant) then a backtrace is printed when a serious run-time error occurs. To disable the backtracing, set the variable to ‘n’, ‘N’, ‘0’. Default is to print a backtrace unless the -fno-backtrace compile option was used.

3.12 GFORTRAN_FORMATTED_BUFFER_SIZE—Set buffer size for

formatted I/O

The GFORTRAN_FORMATTED_BUFFER_SIZE environment variable specifies buffer size in bytes to be used for formatted output. The default value is 8192.

3.13 GFORTRAN_UNFORMATTED_BUFFER_SIZE—Set buffer size for

unformatted I/O

The GFORTRAN_UNFORMATTED_BUFFER_SIZE environment variable specifies buffer size in bytes to be used for unformatted output. The default value is 131072.

Chapter 3: Runtime: Influencing runtime behavior with environment variables 39

Part II: Language Reference

4 Compiler Characteristics

This chapter describes certain characteristics of the GNU Fortran compiler, that are not specified by the Fortran standard, but which might in some way or another become visible to the programmer.

4.1 KIND Type Parameters

The KIND type parameters supported by GNU Fortran for the primitive data types are:

INTEGER 1, 2, 4, 8*, 16*, default: 4**

LOGICAL 1, 2, 4, 8*, 16*, default: 4**

REAL 4, 8, 10*, 16*, default: 4***

COMPLEX 4, 8, 10*, 16*, default: 4***

DOUBLE PRECISION

4, 8, 10*, 16*, default: 8***

CHARACTER

1, 4, default: 1

* not available on all systems

** unless -fdefault-integer-8 is used

*** unless -fdefault-real-8 is used (see Section 2.2 [Fortran Dialect Options], page 8)

The KIND value matches the storage size in bytes, except for COMPLEX where the storage size is twice as much (or both real and imaginary part are a real value of the given size). It is recommended to use the Section 8.239 [SELECTED CHAR KIND], page 263, Section 8.240 [SELECTED INT KIND], page 264, and Section 8.241 [SELECTED REAL KIND], page 264, intrinsics or the INT8, INT16, INT32, INT64, REAL32, REAL64, and REAL128 parameters of the ISO_FORTRAN_ENV module instead of the concrete values. The available kind parameters can be found in the constant arrays

CHARACTER_KINDS, INTEGER_KINDS, LOGICAL_KINDS and REAL_KINDS in the Section 9.1

[ISO FORTRAN ENV], page 293, module. For C interoperability, the kind parameters of the Section 9.2 [ISO C BINDING], page 295, module should be used.

4.2 Internal representation of LOGICAL variables

The Fortran standard does not specify how variables of LOGICAL type are represented, beyond requiring that LOGICAL variables of default kind have the same storage size as default INTEGER and REAL variables. The GNU Fortran internal representation is as follows.

A LOGICAL(KIND=N) variable is represented as an INTEGER(KIND=N) variable, however, with only two permissible values: 1 for .TRUE. and 0 for .FALSE.. Any other integer value results in undefined behavior.

See also Section 6.4.2 [Argument passing conventions], page 81, and Section 6.1 [Interoperability with C], page 69.

4.3 Evaluation of logical expressions

The Fortran standard does not require the compiler to evaluate all parts of an expression, if they do not contribute to the final result. For logical expressions with .AND. or .OR. operators, in particular, GNU Fortran will optimize out function calls (even to impure functions) if the result of the expression can be established without them. However, since not all compilers do that, and such an optimization can potentially modify the program flow and subsequent results, GNU Fortran throws warnings for such situations with the

-Wfunction-elimination flag.

4.4 MAX and MIN intrinsics with REAL NaN arguments

The Fortran standard does not specify what the result of the MAX and MIN intrinsics are if one of the arguments is a NaN. Accordingly, the GNU Fortran compiler does not specify that either, as this allows for faster and more compact code to be generated. If the programmer wishes to take some specific action in case one of the arguments is a NaN, it is necessary to explicitly test the arguments before calling MAX or MIN, e.g. with the IEEE_IS_NAN function from the intrinsic module IEEE_ARITHMETIC.

4.5 Thread-safety of the runtime library

GNU Fortran can be used in programs with multiple threads, e.g. by using OpenMP, by calling OS thread handling functions via the ISO_C_BINDING facility, or by GNU Fortran compiled library code being called from a multi-threaded program.

The GNU Fortran runtime library, (libgfortran), supports being called concurrently from multiple threads with the following exceptions.

During library initialization, the C getenv function is used, which need not be threadsafe. Similarly, the getenv function is used to implement the GET_ENVIRONMENT_VARIABLE and GETENV intrinsics. It is the responsibility of the user to ensure that the environment is not being updated concurrently when any of these actions are taking place.

The EXECUTE_COMMAND_LINE and SYSTEM intrinsics are implemented with the system function, which need not be thread-safe. It is the responsibility of the user to ensure that system is not called concurrently.

For platforms not supporting thread-safe POSIX functions, further functionality might not be thread-safe. For details, please consult the documentation for your operating system.

The GNU Fortran runtime library uses various C library functions that depend on the locale, such as strtod and snprintf. In order to work correctly in locale-aware programs that set the locale using setlocale, the locale is reset to the default “C” locale while executing a formatted READ or WRITE statement. On targets supporting the POSIX 2008 per-thread locale functions (e.g. newlocale, uselocale, freelocale), these are used and thus the global locale set using setlocale or the per-thread locales in other threads are not affected. However, on targets lacking this functionality, the global LC NUMERIC locale is set to “C” during the formatted I/O. Thus, on such targets it’s not safe to call setlocale concurrently from another thread while a Fortran formatted I/O operation is in progress. Also, other threads doing something dependent on the LC NUMERIC locale might not work correctly if a formatted I/O operation is in progress in another thread.

4.6 Data consistency and durability

This section contains a brief overview of data and metadata consistency and durability issues when doing I/O.

With respect to durability, GNU Fortran makes no effort to ensure that data is committed to stable storage. If this is required, the GNU Fortran programmer can use the intrinsic FNUM to retrieve the low level file descriptor corresponding to an open Fortran unit. Then, using e.g. the ISO_C_BINDING feature, one can call the underlying system call to flush dirty data to stable storage, such as fsync on POSIX, _commit on MingW, or fcntl(fd, F_FULLSYNC, 0) on Mac OS X. The following example shows how to call fsync:

        ! Declare the interface for POSIX fsync function
        interface
          function fsync (fd) bind(c,name="fsync")
          use iso_c_binding, only: c_int
            integer(c_int), value :: fd
            integer(c_int) :: fsync
          end function fsync
        end interface
        ! Variable declaration
        integer :: ret
        ! Opening unit 10
        open (10,file="foo")
        ! ...
        ! Perform I/O on unit 10
        ! ...
        ! Flush and sync
        flush(10)
        ret = fsync(fnum(10))
        ! Handle possible error
        if (ret /= 0) stop "Error calling FSYNC"

With respect to consistency, for regular files GNU Fortran uses buffered I/O in order to improve performance. This buffer is flushed automatically when full and in some other situations, e.g. when closing a unit. It can also be explicitly flushed with the FLUSH statement. Also, the buffering can be turned off with the GFORTRAN_UNBUFFERED_ALL and GFORTRAN_UNBUFFERED_PRECONNECTED environment variables. Special files, such as terminals and pipes, are always unbuffered. Sometimes, however, further things may need to be done in order to allow other processes to see data that GNU Fortran has written, as follows.

The Windows platform supports a relaxed metadata consistency model, where file metadata is written to the directory lazily. This means that, for instance, the dir command can show a stale size for a file. One can force a directory metadata update by closing the unit, or by calling _commit on the file descriptor. Note, though, that _commit will force all dirty data to stable storage, which is often a very slow operation.

The Network File System (NFS) implements a relaxed consistency model called open-to-close consistency. Closing a file forces dirty data and metadata to be flushed to the server, and opening a file forces the client to contact the server in order to revalidate cached data. fsync will also force a flush of dirty data and metadata to the server. Similar to open and close, acquiring and releasing fcntl file locks, if the server supports them, will also force cache validation and flushing dirty data and metadata.

4.7 Files opened without an explicit ACTION= specifier

The Fortran standard says that if an OPEN statement is executed without an explicit ACTION= specifier, the default value is processor dependent. GNU Fortran behaves as follows:

1. Attempt to open the file with ACTION='READWRITE'

2. If that fails, try to open with ACTION='READ'

3. If that fails, try to open with ACTION='WRITE'

4. If that fails, generate an error

4.8 File operations on symbolic links

This section documents the behavior of GNU Fortran for file operations on symbolic links, on systems that support them.

• Results of INQUIRE statements of the “inquire by file” form will relate to the target of the symbolic link. For example, INQUIRE(FILE="foo",EXIST=ex) will set ex to .true. if foo is a symbolic link pointing to an existing file, and .false. if foo points to an non-existing file (“dangling” symbolic link).

• Using the OPEN statement with a STATUS="NEW" specifier on a symbolic link will result in an error condition, whether the symbolic link points to an existing target or is dangling.

• If a symbolic link was connected, using the CLOSE statement with a STATUS="DELETE" specifier will cause the symbolic link itself to be deleted, not its target.

4.9 File format of unformatted sequential files

Unformatted sequential files are stored as logical records using record markers. Each logical record consists of one of more subrecords.

Each subrecord consists of a leading record marker, the data written by the user program, and a trailing record marker. The record markers are four-byte integers by default, and eight-byte integers if the -fmax-subrecord-length=8 option (which exists for backwards compability only) is in effect.

The representation of the record markers is that of unformatted files given with the -fconvert option, the Section 5.1.17 [CONVERT specifier], page 53, in an open statement or the Section 3.10 [GFORTRAN CONVERT UNIT], page 36, environment variable.

The maximum number of bytes of user data in a subrecord is 2147483639 (2 GiB - 9) for a four-byte record marker. This limit can be lowered with the -fmax-subrecord-length option, although this is rarely useful. If the length of a logical record exceeds this limit, the data is distributed among several subrecords.

The absolute of the number stored in the record markers is the number of bytes of user data in the corresponding subrecord. If the leading record marker of a subrecord contains a negative number, another subrecord follows the current one. If the trailing record marker contains a negative number, then there is a preceding subrecord.

In the most simple case, with only one subrecord per logical record, both record markers contain the number of bytes of user data in the record.

The format for unformatted sequential data can be duplicated using unformatted stream, as shown in the example program for an unformatted record containing a single subrecord:

      program main
        use iso_fortran_env, only: int32
        implicit none
        integer(int32) :: i
        real, dimension(10) :: a, b
        call random_number(a)
        open (10,file='test.dat',form='unformatted',access='stream')
        inquire (iolength=i) a
        write (10) i, a, i
        close (10)
        open (10,file='test.dat',form='unformatted')
        read (10) b
        if (all (a == b)) print *,'success!'
      end program main

4.10 Asynchronous I/O

Asynchronous I/O is supported if the program is linked against the POSIX thread library. If that is not the case, all I/O is performed as synchronous. On systems which do not support pthread condition variables, such as AIX, I/O is also performed as synchronous.

On some systems, such as Darwin or Solaris, the POSIX thread library is always linked in, so asynchronous I/O is always performed. On other sytems, such as Linux, it is necessary to specify -pthread, -lpthread or -fopenmp during the linking step.

4.11 Behavior on integer overflow

Integer overflow is prohibited by the Fortran standard. The behavior of gfortran on integer overflow is undefined by default. Traditional code, like linear congruential pseudo-random number generators in old programs that rely on specific, non-standard behavior may generate unexpected results. The -fsanitize=undefined option can be used to detect such code at runtime.

It is recommended to use the intrinsic subroutine RANDOM_NUMBER for random number generators or, if the old behavior is desired, to use the -fwrapv option. Note that this option can impact performance.

5 Extensions

The two sections below detail the extensions to standard Fortran that are implemented in GNU Fortran, as well as some of the popular or historically important extensions that are not (or not yet) implemented. For the latter case, we explain the alternatives available to GNU Fortran users, including replacement by standard-conforming code or GNU extensions.

5.1 Extensions implemented in GNU Fortran

GNU Fortran implements a number of extensions over standard Fortran. This chapter contains information on their syntax and meaning. There are currently two categories of GNU Fortran extensions, those that provide functionality beyond that provided by any standard, and those that are supported by GNU Fortran purely for backward compatibility with legacy compilers. By default, -std=gnu allows the compiler to accept both types of extensions, but to warn about the use of the latter. Specifying either -std=f95, -std=f2003, -std=f2008, or -std=f2018 disables both types of extensions, and -std=legacy allows both without warning. The special compile flag -fdec enables additional compatibility extensions along with those enabled by -std=legacy.

5.1.1 Old-style kind specifications

GNU Fortran allows old-style kind specifications in declarations. These look like:

            TYPESPEC*size x,y,z

where TYPESPEC is a basic type (INTEGER, REAL, etc.), and where size is a byte count corresponding to the storage size of a valid kind for that type. (For COMPLEX variables, size is the total size of the real and imaginary parts.) The statement then declares x, y and z to be of type TYPESPEC with the appropriate kind. This is equivalent to the standardconforming declaration

            TYPESPEC(k) x,y,z

where k is the kind parameter suitable for the intended precision. As kind parameters are

implementation-dependent, use the KIND, SELECTED_INT_KIND and SELECTED_REAL_KIND

intrinsics to retrieve the correct value, for instance REAL*8 x can be replaced by:

      INTEGER, PARAMETER :: dbl = KIND(1.0d0)
      REAL(KIND=dbl) :: x

5.1.2 Old-style variable initialization

GNU Fortran allows old-style initialization of variables of the form:

            INTEGER i/1/,j/2/
            REAL x(2,2) /3*0.,1./

The syntax for the initializers is as for the DATA statement, but unlike in a DATA statement, an initializer only applies to the variable immediately preceding the initialization. In other words, something like INTEGER I,J/2,3/ is not valid. This style of initialization is only allowed in declarations without double colons (::); the double colons were introduced in Fortran 90, which also introduced a standard syntax for initializing variables in type declarations.

Examples of standard-conforming code equivalent to the above example are:

      ! Fortran 90

            INTEGER :: i = 1, j = 2
            REAL :: x(2,2) = RESHAPE((/0.,0.,0.,1./),SHAPE(x))
      ! Fortran 77
            INTEGER i, j
            REAL x(2,2)
            DATA i/1/, j/2/, x/3*0.,1./

Note that variables which are explicitly initialized in declarations or in DATA statements automatically acquire the SAVE attribute.

5.1.3 Extensions to namelist

GNU Fortran fully supports the Fortran 95 standard for namelist I/O including array qualifiers, substrings and fully qualified derived types. The output from a namelist write is compatible with namelist read. The output has all names in upper case and indentation to column 1 after the namelist name. Two extensions are permitted:

Old-style use of ‘$’ instead of ‘&’

      $MYNML
       X(:)%Y(2) = 1.0 2.0 3.0
       CH(1:4) = "abcd"
      $END

It should be noted that the default terminator is ‘/’ rather than ‘&END’.

Querying of the namelist when inputting from stdin. After at least one space, entering ‘?’ sends to stdout the namelist name and the names of the variables in the namelist:

       ?
      &mynml
       x
       x%y
       ch
      &end

Entering ‘=?’ outputs the namelist to stdout, as if WRITE(*,NML = mynml) had been called:

      =?
      &MYNML
       X(1)%Y=  0.000000    ,  1.000000    ,  0.000000    ,
       X(2)%Y=  0.000000    ,  2.000000    ,  0.000000    ,
       X(3)%Y=  0.000000    ,  3.000000    ,  0.000000    ,
       CH=abcd,  /

To aid this dialog, when input is from stdin, errors send their messages to stderr and execution continues, even if IOSTAT is set.

PRINT namelist is permitted. This causes an error if -std=f95 is used.

      PROGRAM test_print
        REAL, dimension (4)  ::  x = (/1.0, 2.0, 3.0, 4.0/)
        NAMELIST /mynml/ x
        PRINT mynml
      END PROGRAM test_print

Expanded namelist reads are permitted. This causes an error if -std=f95 is used. In the following example, the first element of the array will be given the value 0.00 and the two succeeding elements will be given the values 1.00 and 2.00.

      &MYNML

        X(1,1) = 0.00 , 1.00 , 2.00
      /

When writing a namelist, if no DELIM= is specified, by default a double quote is used to delimit character strings. If -std=F95, F2003, or F2008, etc, the delim status is set to ’none’. Defaulting to quotes ensures that namelists with character strings can be subsequently read back in accurately.

5.1.4 X format descriptor without count field

To support legacy codes, GNU Fortran permits the count field of the X edit descriptor in FORMAT statements to be omitted. When omitted, the count is implicitly assumed to be one.

             PRINT 10, 2, 3
      10     FORMAT (I1, X, I1)

5.1.5 Commas in FORMAT specifications

To support legacy codes, GNU Fortran allows the comma separator to be omitted immediately before and after character string edit descriptors in FORMAT statements. A comma with no following format decriptor is permited if the -fdec-blank-format-item is given on the command line. This is considered non-conforming code and is discouraged.

             PRINT 10, 2, 3
      10     FORMAT ('FOO='I1' BAR='I2)
             print 20, 5, 6
      20     FORMAT (I3, I3,)

5.1.6 Missing period in FORMAT specifications

To support legacy codes, GNU Fortran allows missing periods in format specifications if and only if -std=legacy is given on the command line. This is considered non-conforming code and is discouraged.

             REAL :: value
             READ(*,10) value
      10     FORMAT ('F4')

5.1.7 Default widths for F, G and I format descriptors

To support legacy codes, GNU Fortran allows width to be omitted from format specifications if and only if -fdec-format-defaults is given on the command line. Default widths will be used. This is considered non-conforming code and is discouraged.

             REAL :: value1
             INTEGER :: value2
             WRITE(*,10) value1, value1, value2
      10     FORMAT ('F, G, I')

5.1.8 I/O item lists

To support legacy codes, GNU Fortran allows the input item list of the READ statement, and the output item lists of the WRITE and PRINT statements, to start with a comma.

5.1.9 Q exponent-letter

GNU Fortran accepts real literal constants with an exponent-letter of Q, for example, 1.23Q45. The constant is interpreted as a REAL(16) entity on targets that support this type. If the target does not support REAL(16) but has a REAL(10) type, then the realliteral-constant will be interpreted as a REAL(10) entity. In the absence of REAL(16) and REAL(10), an error will occur.

5.1.10 BOZ literal constants

Besides decimal constants, Fortran also supports binary (b), octal (o) and hexadecimal (z) integer constants. The syntax is: ‘prefix quote digits quote’, where the prefix is either b, o or z, quote is either ' or " and the digits are 0 or 1 for binary, between 0 and 7 for octal, and between 0 and F for hexadecimal. (Example: b'01011101'.)

Up to Fortran 95, BOZ literal constants were only allowed to initialize integer variables in DATA statements. Since Fortran 2003 BOZ literal constants are also allowed as actual arguments to the REAL, DBLE, INT and CMPLX intrinsic functions. The BOZ literal constant is simply a string of bits, which is padded or truncated as needed, during conversion to a numeric type. The Fortran standard states that the treatment of the sign bit is processor dependent. Gfortran interprets the sign bit as a user would expect.

As a deprecated extension, GNU Fortran allows hexadecimal BOZ literal constants to be specified using the X prefix. That the BOZ literal constant can also be specified by adding a suffix to the string, for example, Z'ABC' and 'ABC'X are equivalent. Additionally, as extension, BOZ literals are permitted in some contexts outside of DATA and the intrinsic functions listed in the Fortran standard. Use -fallow-invalid-boz to enable the extension.

5.1.11 Real array indices

As an extension, GNU Fortran allows the use of REAL expressions or variables as array indices.

5.1.12 Unary operators

As an extension, GNU Fortran allows unary plus and unary minus operators to appear as the second operand of binary arithmetic operators without the need for parenthesis.

             X = Y * -Z

5.1.13 Implicitly convert LOGICAL and INTEGER values

As an extension for backwards compatibility with other compilers, GNU Fortran allows the implicit conversion of LOGICAL values to INTEGER values and vice versa. When converting from a LOGICAL to an INTEGER, .FALSE. is interpreted as zero, and .TRUE. is interpreted as one. When converting from INTEGER to LOGICAL, the value zero is interpreted as .FALSE. and any nonzero value is interpreted as .TRUE..

              LOGICAL :: l
              l = 1
              INTEGER :: i
              i = .TRUE.

However, there is no implicit conversion of INTEGER values in if-statements, nor of LOGICAL or INTEGER values in I/O operations.

5.1.14 Hollerith constants support

GNU Fortran supports Hollerith constants in assignments, DATA statements, function and subroutine arguments. A Hollerith constant is written as a string of characters preceded by an integer constant indicating the character count, and the letter H or h, and stored in bytewise fashion in a numeric (INTEGER, REAL, or COMPLEX), LOGICAL or CHARACTER variable. The constant will be padded with spaces or truncated to fit the size of the variable in which it is stored.

Examples of valid uses of Hollerith constants:

            complex*16 x(2)
            data x /16Habcdefghijklmnop, 16Hqrstuvwxyz012345/
            x(1) = 16HABCDEFGHIJKLMNOP
            call foo (4h abc)

Examples of Hollerith constants:

            integer*4 a
            a = 0H         ! Invalid, at least one character is needed.
            a = 4HAB12     ! Valid
            a = 8H12345678 ! Valid, but the Hollerith constant will be truncated.
            a = 3Hxyz      ! Valid, but the Hollerith constant will be padded.

In general, Hollerith constants were used to provide a rudimentary facility for handling character strings in early Fortran compilers, prior to the introduction of CHARACTER variables in Fortran 77; in those cases, the standard-compliant equivalent is to convert the program to use proper character strings. On occasion, there may be a case where the intent is specifically to initialize a numeric variable with a given byte sequence. In these cases, the same result can be obtained by using the TRANSFER statement, as in this example.

            integer(kind=4) :: a
            a = transfer ("abcd", a)     ! equivalent to: a = 4Habcd

The use of the -fdec option extends support of Hollerith constants to comparisons:

            integer*4 a
            a = 4hABCD
            if (a .ne. 4habcd) then
              write(*,*) "no match"
            end if

Supported types are numeric (INTEGER, REAL, or COMPLEX), and CHARACTER.

5.1.15 Character conversion

Allowing character literals to be used in a similar way to Hollerith constants is a nonstandard extension. This feature is enabled using -fdec-char-conversions and only applies to character literals of kind=1.

Character literals can be used in DATA statements and assignments with numeric (INTEGER, REAL, or COMPLEX) or LOGICAL variables. Like Hollerith constants they are copied byte-wise fashion. The constant will be padded with spaces or truncated to fit the size of the variable in which it is stored.

Examples:

            integer*4 x
            data x / 'abcd' /
            x = 'A'       ! Will be padded.
            x = 'ab1234'  ! Will be truncated.

5.1.16 Cray pointers

Cray pointers are part of a non-standard extension that provides a C-like pointer in Fortran. This is accomplished through a pair of variables: an integer "pointer" that holds a memory address, and a "pointee" that is used to dereference the pointer.

Pointer/pointee pairs are declared in statements of the form:

              pointer ( <pointer> , <pointee> )

or,

              pointer ( <pointer1> , <pointee1> ), ( <pointer2> , <pointee2> ), ...

The pointer is an integer that is intended to hold a memory address. The pointee may be an array or scalar. If an assumed-size array is permitted within the scoping unit, a pointee can be an assumed-size array. That is, the last dimension may be left unspecified by using a * in place of a value. A pointee cannot be an assumed shape array. No space is allocated for the pointee.

The pointee may have its type declared before or after the pointer statement, and its array specification (if any) may be declared before, during, or after the pointer statement. The pointer may be declared as an integer prior to the pointer statement. However, some machines have default integer sizes that are different than the size of a pointer, and so the following code is not portable:

              integer ipt
              pointer (ipt, iarr)

If a pointer is declared with a kind that is too small, the compiler will issue a warning; the resulting binary will probably not work correctly, because the memory addresses stored in the pointers may be truncated. It is safer to omit the first line of the above example; if explicit declaration of ipt’s type is omitted, then the compiler will ensure that ipt is an integer variable large enough to hold a pointer.

Pointer arithmetic is valid with Cray pointers, but it is not the same as C pointer arithmetic. Cray pointers are just ordinary integers, so the user is responsible for determining how many bytes to add to a pointer in order to increment it. Consider the following example:

              real target(10)
              real pointee(10)
              pointer (ipt, pointee)
              ipt = loc (target)
              ipt = ipt + 1

The last statement does not set ipt to the address of target(1), as it would in C pointer arithmetic. Adding 1 to ipt just adds one byte to the address stored in ipt.

Any expression involving the pointee will be translated to use the value stored in the pointer as the base address.

To get the address of elements, this extension provides an intrinsic function LOC(). The LOC() function is equivalent to the & operator in C, except the address is cast to an integer type:

              real ar(10)
              pointer(ipt, arpte(10))
              real arpte
              ipt = loc(ar)  ! Makes arpte is an alias for ar
              arpte(1) = 1.0 ! Sets ar(1) to 1.0

The pointer can also be set by a call to the MALLOC intrinsic (see Section 8.184 [MALLOC], page 230).

Cray pointees often are used to alias an existing variable. For example:

              integer target(10)
              integer iarr(10)
              pointer (ipt, iarr)
              ipt = loc(target)

As long as ipt remains unchanged, iarr is now an alias for target. The optimizer, however, will not detect this aliasing, so it is unsafe to use iarr and target simultaneously. Using a pointee in any way that violates the Fortran aliasing rules or assumptions is illegal. It is the user’s responsibility to avoid doing this; the compiler works under the assumption that no such aliasing occurs.

Cray pointers will work correctly when there is no aliasing (i.e., when they are used to access a dynamically allocated block of memory), and also in any routine where a pointee is used, but any variable with which it shares storage is not used. Code that violates these rules may not run as the user intends. This is not a bug in the optimizer; any code that violates the aliasing rules is illegal. (Note that this is not unique to GNU Fortran; any Fortran compiler that supports Cray pointers will “incorrectly” optimize code with illegal aliasing.)

There are a number of restrictions on the attributes that can be applied to Cray pointers and pointees. Pointees may not have the ALLOCATABLE, INTENT, OPTIONAL, DUMMY,

TARGET, INTRINSIC, or POINTER attributes. Pointers may not have the DIMENSION, POINTER,

TARGET, ALLOCATABLE, EXTERNAL, or INTRINSIC attributes, nor may they be function results. Pointees may not occur in more than one pointer statement. A pointee cannot be a pointer. Pointees cannot occur in equivalence, common, or data statements.

A Cray pointer may also point to a function or a subroutine. For example, the following excerpt is valid:

        implicit none
        external sub
        pointer (subptr,subpte)
        external subpte
        subptr = loc(sub)
        call subpte()
        [...]
        subroutine sub
        [...]
        end subroutine sub

A pointer may be modified during the course of a program, and this will change the location to which the pointee refers. However, when pointees are passed as arguments, they are treated as ordinary variables in the invoked function. Subsequent changes to the pointer will not change the base address of the array that was passed.

5.1.17 CONVERT specifier

GNU Fortran allows the conversion of unformatted data between little- and big-endian representation to facilitate moving of data between different systems. The conversion can be indicated with the CONVERT specifier on the OPEN statement. See Section 3.10 [GFOR-TRAN CONVERT UNIT], page 36, for an alternative way of specifying the data format via an environment variable.

Valid values for CONVERT on most systems are:

CONVERT='NATIVE' Use the native format. This is the default.

CONVERT='SWAP' Swap between little- and big-endian.

CONVERT='LITTLE_ENDIAN' Use the little-endian representation for unformatted files. CONVERT='BIG_ENDIAN' Use the big-endian representation for unformatted files.

On POWER systems which support -mabi=ieeelongdouble, there are additional options, which can be combined with the others with commas. Those are

    CONVERT='R16_IEEE' Use IEEE 128-bit format for REAL(KIND=16).
    CONVERT='R16_IBM' Use IBM long double format for realREAL(KIND=16).

Using the option could look like this:

        open(file='big.dat',form='unformatted',access='sequential', &
             convert='big_endian')

The value of the conversion can be queried by using INQUIRE(CONVERT=ch). The values

returned are 'BIG_ENDIAN' and 'LITTLE_ENDIAN'.

CONVERT works between big- and little-endian for INTEGER values of all supported kinds and for REAL on IEEE systems of kinds 4 and 8. Conversion between different “extended double” types on different architectures such as m68k and x86 64, which GNU Fortran supports as REAL(KIND=10) and REAL(KIND=16), will probably not work.

Note that the values specified via the GFORTRAN CONVERT UNIT environment variable will override the CONVERT specifier in the open statement. This is to give control over data formats to users who do not have the source code of their program available.

Using anything but the native representation for unformatted data carries a significant speed overhead. If speed in this area matters to you, it is best if you use this only for data that needs to be portable.

5.1.18 OpenMP

OpenMP (Open Multi-Processing) is an application programming interface (API) that supports multi-platform shared memory multiprocessing programming in C/C++ and Fortran on many architectures, including Unix and Microsoft Windows platforms. It consists of a set of compiler directives, library routines, and environment variables that influence run-time behavior.

GNU Fortran strives to be compatible to the OpenMP Application Program Interface

v4.5 (https://openmp.org/specifications/).

To enable the processing of the OpenMP directive !$omp in free-form source code; the c$omp, *$omp and !$omp directives in fixed form; the !$ conditional compilation sentinels in free form; and the c$, *$ and !$ sentinels in fixed form, gfortran needs to be invoked with the -fopenmp. This also arranges for automatic linking of the GNU Offloading and Multi Processing Runtime Library Section “libgomp” in GNU Offloading and Multi Processing Runtime Library.

The OpenMP Fortran runtime library routines are provided both in a form of a Fortran 90 module named omp_lib and in a form of a Fortran include file named omp_lib.h.

An example of a parallelized loop taken from Appendix A.1 of the OpenMP Application Program Interface v2.5:

      SUBROUTINE A1(N, A, B)

        INTEGER I, N
        REAL B(N), A(N)
      !$OMP PARALLEL DO !I is private by default
        DO I=2,N
          B(I) = (A(I) + A(I-1)) / 2.0
        ENDDO
      !$OMP END PARALLEL DO
      END SUBROUTINE A1

Please note:

• -fopenmp implies -frecursive, i.e., all local arrays will be allocated on the stack. When porting existing code to OpenMP, this may lead to surprising results, especially to segmentation faults if the stacksize is limited.

• On glibc-based systems, OpenMP enabled applications cannot be statically linked due to limitations of the underlying pthreads-implementation. It might be possible to get

    a working solution if -Wl,--whole-archive -lpthread -Wl,--no-whole-archive is

added to the command line. However, this is not supported by gcc and thus not recommended.

5.1.19 OpenACC

OpenACC is an application programming interface (API) that supports offloading of code to accelerator devices. It consists of a set of compiler directives, library routines, and environment variables that influence run-time behavior.

GNU Fortran strives to be compatible to the OpenACC Application Programming In-

terface v2.6 (https://www.openacc.org/).

To enable the processing of the OpenACC directive !$acc in free-form source code; the c$acc, *$acc and !$acc directives in fixed form; the !$ conditional compilation sentinels in free form; and the c$, *$ and !$ sentinels in fixed form, gfortran needs to be invoked with the -fopenacc. This also arranges for automatic linking of the GNU Offloading and Multi Processing Runtime Library Section “libgomp” in GNU Offloading and Multi Processing Runtime Library.

The OpenACC Fortran runtime library routines are provided both in a form of a Fortran 90 module named openacc and in a form of a Fortran include file named openacc_lib.h.

5.1.20 Argument list functions %VAL, %REF and %LOC

GNU Fortran supports argument list functions %VAL, %REF and %LOC statements, for backward compatibility with g77. It is recommended that these should be used only for code that is accessing facilities outside of GNU Fortran, such as operating system or windowing facilities. It is best to constrain such uses to isolated portions of a program–portions that deal specifically and exclusively with low-level, system-dependent facilities. Such portions might well provide a portable interface for use by the program as a whole, but are themselves not portable, and should be thoroughly tested each time they are rebuilt using a new compiler or version of a compiler.

%VAL passes a scalar argument by value, %REF passes it by reference and %LOC passes its memory location. Since gfortran already passes scalar arguments by reference, %REF is in effect a do-nothing. %LOC has the same effect as a Fortran pointer.

An example of passing an argument by value to a C subroutine foo.:

      C prototype      void foo_(float x);
      C
            external foo
            real*4 x
            x = 3.14159
            call foo (%VAL (x))
            end
   For details refer to the g77 manual https://gcc.gnu.org/onlinedocs/gcc-3.4.6/
g77/index.html#Top.

Also, c_by_val.f and its partner c_by_val.c of the GNU Fortran testsuite are worth a look.

5.1.21 Read/Write after EOF marker

Some legacy codes rely on allowing READ or WRITE after the EOF file marker in order to find the end of a file. GNU Fortran normally rejects these codes with a run-time error message and suggests the user consider BACKSPACE or REWIND to properly position the file before the EOF marker. As an extension, the run-time error may be disabled using -std=legacy.

5.1.22 STRUCTURE and RECORD

Record structures are a pre-Fortran-90 vendor extension to create user-defined aggregate data types. Support for record structures in GNU Fortran can be enabled with the -fdec-structure compile flag. If you have a choice, you should instead use Fortran 90’s “derived types”, which have a different syntax.

In many cases, record structures can easily be converted to derived types. To convert, replace STRUCTURE /structure-name/ by TYPE type-name. Additionally, replace RECORD /structure-name/ by TYPE(type-name). Finally, in the component access, replace the period (.) by the percent sign (%).

Here is an example of code using the non portable record structure syntax:

     ! Declaring a structure named `  ` item'' and containing three fields:
     ! an integer ID, an description string and a floating-point price.
     STRUCTURE /item/
       INTEGER id
       CHARACTER(LEN=200) description
       REAL price
     END STRUCTURE
     ! Define two variables, an single record of type `  ` item''
     ! named `  ` pear'', and an array of items named `  ` store_catalog''
     RECORD /item/ pear, store_catalog(100)
     ! We can directly access the fields of both variables
     pear.id = 92316
     pear.description = "juicy D'Anjou pear"
     pear.price = 0.15
     store_catalog(7).id = 7831
     store_catalog(7).description = "milk bottle"
     store_catalog(7).price = 1.2

! We can also manipulate the whole structure
store_catalog(12) = pear
print *, store_catalog(12)

This code can easily be rewritten in the Fortran 90 syntax as following:

! `  ` STRUCTURE /name/ ... END STRUCTURE'' becomes
! `  ` TYPE name ... END TYPE''
TYPE item
  INTEGER id
  CHARACTER(LEN=200) description
  REAL price
END TYPE
! `  ` RECORD /name/ variable'' becomes `  ` TYPE(name) variable''
TYPE(item) pear, store_catalog(100)
! Instead of using a dot (.) to access fields of a record, the
! standard syntax uses a percent sign (%)
pear%id = 92316
pear%description = "juicy D'Anjou pear"
pear%price = 0.15
store_catalog(7)%id = 7831
store_catalog(7)%description = "milk bottle"
store_catalog(7)%price = 1.2
! Assignments of a whole variable do not change
store_catalog(12) = pear
print *, store_catalog(12)

GNU Fortran implements STRUCTURES like derived types with the following rules and exceptions:

• Structures act like derived types with the SEQUENCE attribute. Otherwise they may contain no specifiers.

• Structures may contain a special field with the name %FILL. This will create an anonymous component which cannot be accessed but occupies space just as if a component of the same type was declared in its place, useful for alignment purposes. As an example, the following structure will consist of at least sixteen bytes:

     structure /padded/
       character(4) start
       character(8) %FILL
       character(4) end
     end structure

• Structures may share names with other symbols. For example, the following is invalid

for derived types, but valid for structures:

     structure /header/
       ! ...
     end structure
     record /header/ header

• Structure types may be declared nested within another parent structure. The syntax

is:

      structure /type-name/
          ...
          structure [/<type-name>/] <field-list>
      ...

The type name may be ommitted, in which case the structure type itself is anonymous, and other structures of the same type cannot be instantiated. The following shows some examples:

     structure /appointment/
       ! nested structure definition: app_time is an array of two 'time'
       structure /time/ app_time (2)
         integer(1) hour, minute
       end structure
       character(10) memo
     end structure
     ! The 'time' structure is still usable
     record /time/ now
     now = time(5, 30)
     ...
     structure /appointment/
       ! anonymous nested structure definition
       structure start, end
         integer(1) hour, minute
       end structure
       character(10) memo
     end structure

• Structures may contain UNION blocks. For more detail see the section on Section 5.1.23 [UNION and MAP], page 58.

• Structures support old-style initialization of components, like those described in Section 5.1.2 [Old-style variable initialization], page 47. For array initializers, an initializer may contain a repeat specification of the form <literal-integer> * <constantinitializer>. The value of the integer indicates the number of times to repeat the constant initializer when expanding the initializer list.

5.1.23 UNION and MAP

Unions are an old vendor extension which were commonly used with the non-standard Section 5.1.22 [STRUCTURE and RECORD], page 56, extensions. Use of UNION and MAP is automatically enabled with -fdec-structure.

A UNION declaration occurs within a structure; within the definition of each union is a number of MAP blocks. Each MAP shares storage with its sibling maps (in the same union), and the size of the union is the size of the largest map within it, just as with unions in C.

The major difference is that component references do not indicate which union or map the component is in (the compiler gets to figure that out).

Here is a small example:

structure /myunion/
union
  map
    character(2) w0, w1, w2
  end map
  map
    character(6) long
  end map
end union
end structure
record /myunion/ rec
! After this assignment...
rec.long = 'hello!'
! The following is true:
! rec.w0 === 'he'
! rec.w1 === 'll'
! rec.w2 === 'o!'

The two maps share memory, and the size of the union is ultimately six bytes:

0    1    2    3    4   5   6     Byte offset
-------------------------------
|    |    |    |    |    |    |
-------------------------------
^W0   ^W1   ^W2   ^
 \-------/ \-------/ \-------/
^LONG            ^
 \---------------------------/

Following is an example mirroring the layout of an Intel x86 64 register:

structure /reg/
  union ! U0                ! rax
    map
      character(16) rx
    end map
    map
      character(8) rh         ! rah
      union ! U1
        map
          character(8) rl     ! ral
        end map
        map
          character(8) ex     ! eax
        end map
        map

          character(4) eh     ! eah
          union ! U2
            map
              character(4) el ! eal
            end map
            map
              character(4) x  ! ax
            end map
            map
              character(2) h  ! ah
              character(2) l  ! al
            end map
          end union
        end map
      end union
    end map
  end union
end structure
record /reg/ a
! After this assignment...
a.rx     =     'AAAAAAAA.BBB.C.D'
! The following is true:
a.rx === 'AAAAAAAA.BBB.C.D'
a.rh === 'AAAAAAAA'
a.rl ===         '.BBB.C.D'
a.ex ===         '.BBB.C.D'
a.eh ===         '.BBB'
a.el ===             '.C.D'
a.x  ===             '.C.D'
a.h  ===             '.C'
a.l  ===               '.D'

5.1.24 Type variants for integer intrinsics

Similar to the D/C prefixes to real functions to specify the input/output types, GNU Fortran offers B/I/J/K prefixes to integer functions for compatibility with DEC programs. The types implied by each are:

B - INTEGER(kind=1)
I - INTEGER(kind=2)
J - INTEGER(kind=4)
K - INTEGER(kind=8)

GNU Fortran supports these with the flag -fdec-intrinsic-ints. Intrinsics for which prefixed versions are available and in what form are noted in Chapter 8 [Intrinsic Procedures], page 111. The complete list of supported intrinsics is here:

Intrinsic  B  I  J  K

Section 8.3
[ABS], page 112

                 BABS

                                IIABS

                                               JIABS

                                                              KIABS
Section 8.51
[BTEST],
page 144

                 BBTEST

                                BITEST

                                               BJTEST

                                                              BKTEST

Section 8.137
[IAND],
page 202

                 BIAND

                                IIAND

                                               JIAND

                                                              KIAND

Section 8.140
[IBCLR],
page 204

                 BBCLR

                                IIBCLR

                                               JIBCLR

                                                              KIBCLR

Section 8.141
[IBITS],
page 205

                 BBITS

                                IIBITS

                                               JIBITS

                                                              KIBITS

Section 8.142
[IBSET],
page 206

                 BBSET

                                IIBSET

                                               JIBSET

                                                              KIBSET

Section 8.145
[IEOR],
page 208

                 BIEOR

                                IIEOR

                                               JIEOR

                                                              KIEOR

Section 8.152
[IOR], page 212

                 BIOR

                                IIOR

                                               JIOR

                                                              KIOR
Section 8.159
[ISHFT],
page 216

                 BSHFT

                                IISHFT

                                               JISHFT

                                                              KISHFT

Section 8.160
[ISHFTC],
page 217

                 BSHFTC

                                IISHFTC

                                               JISHFTC

                                                              KISHFTC

Section 8.200
[MOD], page 240

                 BMOD

                                IMOD

                                               JMOD

                                                              KMOD
Section 8.208
[NOT], page 245

                 BNOT

                                INOT

                                               JNOT

                                                              KNOT
Section 8.228
[REAL],
page 257

                 --

                                FLOATI

                                               FLOATJ

                                                              FLOATK

5.1.25 AUTOMATIC and STATIC attributes

With -fdec-static GNU Fortran supports the DEC extended attributes STATIC and AUTOMATIC to provide explicit specification of entity storage. These follow the syntax of the Fortran standard SAVE attribute.

STATIC is exactly equivalent to SAVE, and specifies that an entity should be allocated in static memory. As an example, STATIC local variables will retain their values across multiple calls to a function.

Entities marked AUTOMATIC will be stack automatic whenever possible. AUTOMATIC is the default for local variables smaller than -fmax-stack-var-size, unless -fno-automatic is given. This attribute overrides -fno-automatic, -fmax-stack-var-size, and blanket SAVE statements.

Examples:

     subroutine f
       integer, automatic :: i  ! automatic variable
       integer x, y             ! static variables
       save
       ...
     endsubroutine
     subroutine f
       integer a, b, c, x, y, z
       static :: x
       save y
       automatic z, c
       ! a, b, c, and z are automatic
       ! x and y are static
     endsubroutine
     ! Compiled with -fno-automatic
     subroutine f
       integer a, b, c, d
       automatic :: a
       ! a is automatic; b, c, and d are static
     endsubroutine

5.1.26 Extended math intrinsics

GNU Fortran supports an extended list of mathematical intrinsics with the compile flag -fdec-math for compatability with legacy code. These intrinsics are described fully in Chapter 8 [Intrinsic Procedures], page 111, where it is noted that they are extensions and should be avoided whenever possible.

Specifically, -fdec-math enables the Section 8.76 [COTAN], page 162, intrinsic, and trigonometric intrinsics which accept or produce values in degrees instead of radians. Here is a summary of the new intrinsics:

Radians  Degrees

Section 8.6 [ACOS], page 114             Section 8.7 [ACOSD], page 114*
Section 8.19 [ASIN], page 122            Section 8.20 [ASIND], page 123*
Section 8.23 [ATAN], page 126            Section 8.24 [ATAND], page 126*
Section 8.25 [ATAN2], page 127           Section 8.26 [ATAN2D], page 128*
Section 8.73 [COS], page 160             Section 8.74 [COSD], page 160*
Section 8.76 [COTAN], page 162*  Section 8.77 [COTAND], page 162*
Section 8.249 [SIN], page 269            Section 8.250 [SIND], page 270*
Section 8.265 [TAN], page 280            Section 8.266 [TAND], page 280*

* Enabled with -fdec-math.

For advanced users, it may be important to know the implementation of these functions. They are simply wrappers around the standard radian functions, which have more accurate builtin versions. These functions convert their arguments (or results) to degrees (or radians) by taking the value modulus 360 (or 2*pi) and then multiplying it by a constant radian-to-degree (or degree-to-radian) factor, as appropriate. The factor is computed at compile-time as 180/pi (or pi/180).

5.1.27 Form feed as whitespace

Historically, legacy compilers allowed insertion of form feed characters (’\f’, ASCII 0xC) at the beginning of lines for formatted output to line printers, though the Fortran standard does not mention this. GNU Fortran supports the interpretation of form feed characters in source as whitespace for compatibility.

5.1.28 TYPE as an alias for PRINT

For compatibility, GNU Fortran will interpret TYPE statements as PRINT statements with the flag -fdec. With this flag asserted, the following two examples are equivalent:

      TYPE *, 'hello world'
      PRINT *, 'hello world'

5.1.29 %LOC as an rvalue

Normally %LOC is allowed only in parameter lists. However the intrinsic function LOC does the same thing, and is usable as the right-hand-side of assignments. For compatibility, GNU Fortran supports the use of %LOC as an alias for the builtin LOC with -std=legacy. With this feature enabled the following two examples are equivalent:

      integer :: i, l
      l = %loc(i)
      call sub(l)
      integer :: i
      call sub(%loc(i))

5.1.30 .XOR. operator

GNU Fortran supports .XOR. as a logical operator with -std=legacy for compatibility with legacy code. .XOR. is equivalent to .NEQV.. That is, the output is true if and only if the inputs differ.

5.1.31 Bitwise logical operators

With -fdec, GNU Fortran relaxes the type constraints on logical operators to allow integer operands, and performs the corresponding bitwise operation instead. This flag is for compatibility only, and should be avoided in new code. Consider:

        INTEGER :: i, j
        i = z'33'
        j = z'cc'
        print *, i .AND. j

In this example, compiled with -fdec, GNU Fortran will replace the .AND. operation with a call to the intrinsic Section 8.137 [IAND], page 202 function, yielding the bitwiseand of i and j.

Note that this conversion will occur if at least one operand is of integral type. As a result, a logical operand will be converted to an integer when the other operand is an integer in a logical operation. In this case, .TRUE. is converted to 1 and .FALSE. to 0.

Here is the mapping of logical operator to bitwise intrinsic used with -fdec:

Operator  Intrinsic  Bitwise operation

.NOT.

           Section 8.208
           [NOT], page 245

complement

.AND.

           Section 8.137
           [IAND], page 202

intersection

.OR.

           Section 8.152
           [IOR], page 212

union

.NEQV.

           Section 8.145
           [IEOR], page 208

exclusive or

.EQV.

           Section 8.208
           [NOT],
           page 245(Section 8.145
           [IEOR], page 208)

complement of exclusive or

5.1.32 Extended I/O specifiers

GNU Fortran supports the additional legacy I/O specifiers CARRIAGECONTROL, READONLY, and SHARE with the compile flag -fdec, for compatibility.

CARRIAGECONTROL

The CARRIAGECONTROL specifier allows a user to control line termination settings between output records for an I/O unit. The specifier has no meaning for readonly files. When CARRAIGECONTROL is specified upon opening a unit for formatted writing, the exact CARRIAGECONTROL setting determines what characters to write between output records. The syntax is:

      OPEN(..., CARRIAGECONTROL=cc)

Where cc is a character expression that evaluates to one of the following values:

'LIST'  One line feed between records (default)
'FORTRAN'  Legacy interpretation of the first character (see below)
'NONE'  No separator between records

With CARRIAGECONTROL='FORTRAN', when a record is written, the first character of the input record is not written, and instead determines the output record separator as follows: Leading character Meaning Output separating character(s)

'+'  Overprinting  Carriage return only
'-'  New line  Line feed and carriage return

'0'

Skip line Two line feeds and carriage return

'1'

New page Form feed and carriage return

'$'  Prompting  Line feed (no carriage return)

CHAR(0)

Overprinting (no advance)

None

READONLY The READONLY specifier may be given upon opening a unit, and is equivalent to specifying ACTION='READ', except that the file may not be deleted on close (i.e. CLOSE with STATUS="DELETE"). The syntax is:

      OPEN(..., READONLY)

SHARE The SHARE specifier allows system-level locking on a unit upon opening it for controlled access from multiple processes/threads. The SHARE specifier has several forms:

      OPEN(..., SHARE=sh)
      OPEN(..., SHARED)
      OPEN(..., NOSHARED)

Where sh in the first form is a character expression that evaluates to a value as seen in the table below. The latter two forms are aliases for particular values of sh:

Explicit form  Short form  Meaning

SHARE='DENYRW'        NOSHARED  Exclusive (write) lock
SHARE='DENYNONE'      SHARED  Shared (read) lock

In general only one process may hold an exclusive (write) lock for a given file at a time, whereas many processes may hold shared (read) locks for the same file. The behavior of locking may vary with your operating system. On POSIX systems, locking is implemented with fcntl. Consult your corresponding operating system’s manual pages for further details. Locking via SHARE= is not supported on other systems.

5.1.33 Legacy PARAMETER statements

For compatibility, GNU Fortran supports legacy PARAMETER statements without parentheses with -std=legacy. A warning is emitted if used with -std=gnu, and an error is acknowledged with a real Fortran standard flag (-std=f95, etc...). These statements take the following form:

implicit real (E)
parameter e = 2.718282
real c
parameter c = 3.0e8

5.1.34 Default exponents

For compatibility, GNU Fortran supports a default exponent of zero in real constants with -fdec. For example, 9e would be interpreted as 9e0, rather than an error.

5.2 Extensions not implemented in GNU Fortran

The long history of the Fortran language, its wide use and broad userbase, the large number of different compiler vendors and the lack of some features crucial to users in the first standards have lead to the existence of a number of important extensions to the language. While some of the most useful or popular extensions are supported by the GNU Fortran compiler, not all existing extensions are supported. This section aims at listing these extensions and offering advice on how best make code that uses them running with the GNU Fortran compiler.

5.2.1 ENCODE and DECODE statements

GNU Fortran does not support the ENCODE and DECODE statements. These statements are best replaced by READ and WRITE statements involving internal files (CHARACTER variables and arrays), which have been part of the Fortran standard since Fortran 77. For example, replace a code fragment like

            INTEGER*1 LINE(80)
            REAL A, B, C
      c     ... Code that sets LINE
            DECODE (80, 9000, LINE) A, B, C
       9000 FORMAT (1X, 3(F10.5))

with the following:

            CHARACTER(LEN=80) LINE
            REAL A, B, C
      c     ... Code that sets LINE
            READ (UNIT=LINE, FMT=9000) A, B, C
       9000 FORMAT (1X, 3(F10.5))

Similarly, replace a code fragment like

            INTEGER*1 LINE(80)
            REAL A, B, C
      c     ... Code that sets A, B and C
            ENCODE (80, 9000, LINE) A, B, C
       9000 FORMAT (1X, 'OUTPUT IS ', 3(F10.5))

with the following:

            CHARACTER(LEN=80) LINE
            REAL A, B, C
      c     ... Code that sets A, B and C
            WRITE (UNIT=LINE, FMT=9000) A, B, C
       9000 FORMAT (1X, 'OUTPUT IS ', 3(F10.5))

5.2.2 Variable FORMAT expressions

A variable FORMAT expression is format statement which includes angle brackets enclosing a Fortran expression: FORMAT(I<N>). GNU Fortran does not support this legacy extension. The effect of variable format expressions can be reproduced by using the more powerful (and standard) combination of internal output and string formats. For example, replace a code fragment like this:

            WRITE(6,20) INT1
       20   FORMAT(I<N+1>)

with the following:

      c     Variable declaration
            CHARACTER(LEN=20) FMT
      c
      c     Other code here...
      c
            WRITE(FMT,'("(I", I0, ")")') N+1
            WRITE(6,FMT) INT1

or with:

      c     Variable declaration
            CHARACTER(LEN=20) FMT
      c
      c     Other code here...
      c
            WRITE(FMT,*) N+1
            WRITE(6,"(I" // ADJUSTL(FMT) // ")") INT1

5.2.3 Alternate complex function syntax

Some Fortran compilers, including g77, let the user declare complex functions with the

syntax COMPLEX FUNCTION name*16(), as well as COMPLEX*16 FUNCTION name(). Both are

non-standard, legacy extensions. gfortran accepts the latter form, which is more common, but not the former.

5.2.4 Volatile COMMON blocks

Some Fortran compilers, including g77, let the user declare COMMON with the VOLATILE attribute. This is invalid standard Fortran syntax and is not supported by gfortran. Note that gfortran accepts VOLATILE variables in COMMON blocks since revision 4.3.

5.2.5 OPEN( ... NAME=)

Some Fortran compilers, including g77, let the user declare OPEN( ... NAME=). This is invalid standard Fortran syntax and is not supported by gfortran. OPEN( ... NAME=) should be replaced with OPEN( ... FILE=).

5.2.6 Q edit descriptor

Some Fortran compilers provide the Q edit descriptor, which transfers the number of characters left within an input record into an integer variable.

A direct replacement of the Q edit descriptor is not available in gfortran. How to replicate its functionality using standard-conforming code depends on what the intent of the original code is.

Options to replace Q may be to read the whole line into a character variable and then counting the number of non-blank characters left using LEN_TRIM. Another method may be to use formatted stream, read the data up to the position where the Q descriptor occurred, use INQUIRE to get the file position, count the characters up to the next NEW_LINE and then start reading from the position marked previously.

6 Mixed-Language Programming

This chapter is about mixed-language interoperability, but also applies if you link Fortran code compiled by different compilers. In most cases, use of the C Binding features of the Fortran 2003 and later standards is sufficient.

For example, it is possible to mix Fortran code with C++ code as well as C, if you declare the interface functions as extern "C" on the C++ side and BIND(C) on the Fortran side, and follow the rules for interoperability with C. Note that you cannot manipulate C++ class objects in Fortran or vice versa except as opaque pointers.

You can use the gfortran command to link both Fortran and non-Fortran code into the same program, or you can use gcc or g++ if you also add an explicit -lgfortran option to link with the Fortran library. If your main program is written in C or some other language instead of Fortran, see Section 6.3 [Non-Fortran Main Program], page 76, below.

6.1 Interoperability with C

Since Fortran 2003 (ISO/IEC 1539-1:2004(E)) there is a standardized way to generate procedure and derived-type declarations and global variables that are interoperable with C (ISO/IEC 9899:1999). The BIND(C) attribute has been added to inform the compiler that a symbol shall be interoperable with C; also, some constraints are added. Note, however, that not all C features have a Fortran equivalent or vice versa. For instance, neither C’s unsigned integers nor C’s functions with variable number of arguments have an equivalent in Fortran.

Note that array dimensions are reversely ordered in C and that arrays in C always start with index 0 while in Fortran they start by default with 1. Thus, an array declaration A(n,m) in Fortran matches A[m][n] in C and accessing the element A(i,j) matches A[j-1][i-1]. The element following A(i,j) (C: A[j-1][i-1]; assuming i < n) in memory is

A(i+1,j) (C: A[j-1][i]).

6.1.1 Intrinsic Types

In order to ensure that exactly the same variable type and kind is used in C and Fortran, you should use the named constants for kind parameters that are defined in the ISO_C_BINDING intrinsic module. That module contains named constants of character type representing the escaped special characters in C, such as newline. For a list of the constants, see Section 9.2 [ISO C BINDING], page 295.

For logical types, please note that the Fortran standard only guarantees interoperability between C99’s _Bool and Fortran’s C_Bool-kind logicals and C99 defines that true has the value 1 and false the value 0. Using any other integer value with GNU Fortran’s LOGICAL (with any kind parameter) gives an undefined result. (Passing other integer values than 0 and 1 to GCC’s _Bool is also undefined, unless the integer is explicitly or implicitly casted

to _Bool.)

6.1.2 Derived Types and struct

For compatibility of derived types with struct, use the BIND(C) attribute in the type declaration. For instance, the following type declaration

       USE ISO_C_BINDING

       TYPE, BIND(C) :: myType
         INTEGER(C_INT) :: i1, i2
         INTEGER(C_SIGNED_CHAR) :: i3
         REAL(C_DOUBLE) :: d1
         COMPLEX(C_FLOAT_COMPLEX) :: c1
         CHARACTER(KIND=C_CHAR) :: str(5)
       END TYPE

matches the following struct declaration in C

       struct {
         int i1, i2;
         /* Note: "char" might be signed or unsigned.  */
         signed char i3;
         double d1;
         float _Complex c1;
         char str[5];
       } myType;

Derived types with the C binding attribute shall not have the sequence attribute, type parameters, the extends attribute, nor type-bound procedures. Every component must be of interoperable type and kind and may not have the pointer or allocatable attribute. The names of the components are irrelevant for interoperability.

As there exist no direct Fortran equivalents, neither unions nor structs with bit field or variable-length array members are interoperable.

6.1.3 Interoperable Global Variables

Variables can be made accessible from C using the C binding attribute, optionally together with specifying a binding name. Those variables have to be declared in the declaration part of a MODULE, be of interoperable type, and have neither the pointer nor the allocatable attribute.

        MODULE m
          USE myType_module
          USE ISO_C_BINDING
          integer(C_INT), bind(C, name="_MyProject_flags") :: global_flag
          type(myType), bind(C) :: tp
        END MODULE

Here, _MyProject_flags is the case-sensitive name of the variable as seen from C programs while global_flag is the case-insensitive name as seen from Fortran. If no binding name is specified, as for tp, the C binding name is the (lowercase) Fortran binding name. If a binding name is specified, only a single variable may be after the double colon. Note of warning: You cannot use a global variable to access errno of the C library as the C standard allows it to be a macro. Use the IERRNO intrinsic (GNU extension) instead.

6.1.4 Interoperable Subroutines and Functions

Subroutines and functions have to have the BIND(C) attribute to be compatible with C. The dummy argument declaration is relatively straightforward. However, one needs to be careful because C uses call-by-value by default while Fortran behaves usually similar to call-by-reference. Furthermore, strings and pointers are handled differently.

To pass a variable by value, use the VALUE attribute. Thus, the following C prototype

      int func(int i, int *j)

matches the Fortran declaration

        integer(c_int) function func(i,j)
          use iso_c_binding, only: c_int
          integer(c_int), VALUE :: i
          integer(c_int) :: j

Note that pointer arguments also frequently need the VALUE attribute, see Section 6.1.5 [Working with C Pointers], page 72.

Strings are handled quite differently in C and Fortran. In C a string is a NUL-terminated array of characters while in Fortran each string has a length associated with it and is thus not terminated (by e.g. NUL). For example, if you want to use the following C function,

        #include <stdio.h>
        void print_C(char *string) /* equivalent: char string[]  */
        {
           printf("%s\n", string);
        }

to print “Hello World” from Fortran, you can call it using

        use iso_c_binding, only: C_CHAR, C_NULL_CHAR
        interface
          subroutine print_c(string) bind(C, name="print_C")
            use iso_c_binding, only: c_char
            character(kind=c_char) :: string(*)
          end subroutine print_c
        end interface
        call print_c(C_CHAR_"Hello World"//C_NULL_CHAR)

As the example shows, you need to ensure that the string is NUL terminated. Additionally, the dummy argument string of print_C is a length-one assumed-size array; using character(len=*) is not allowed. The example above uses c_char_"Hello World" to ensure the string literal has the right type; typically the default character kind and c_char are the same and thus "Hello World" is equivalent. However, the standard does not guarantee this.

The use of strings is now further illustrated using the C library function strncpy, whose prototype is

        char *strncpy(char *restrict s1, const char *restrict s2, size_t n);

The function strncpy copies at most n characters from string s2 to s1 and returns s1. In the following example, we ignore the return value:

        use iso_c_binding
        implicit none
        character(len=30) :: str,str2
        interface
          ! Ignore the return value of strncpy -> subroutine
          ! "restrict" is always assumed if we do not pass a pointer
          subroutine strncpy(dest, src, n) bind(C)
            import
            character(kind=c_char),  intent(out) :: dest(*)
            character(kind=c_char),  intent(in)  :: src(*)
            integer(c_size_t), value, intent(in) :: n
          end subroutine strncpy
        end interface
        str = repeat('X',30) ! Initialize whole string with 'X'
        call strncpy(str, c_char_"Hello World"//C_NULL_CHAR, &
                     len(c_char_"Hello World",kind=c_size_t))

        print '(a)', str ! prints: "Hello WorldXXXXXXXXXXXXXXXXXXX"
        end

The intrinsic procedures are described in Chapter 8 [Intrinsic Procedures], page 111.

6.1.5 Working with C Pointers

C pointers are represented in Fortran via the special opaque derived type type(c_ptr) (with private components). C pointers are distinct from Fortran objects with the POINTER attribute. Thus one needs to use intrinsic conversion procedures to convert from or to C pointers. For some applications, using an assumed type (TYPE(*)) can be an alternative to a C pointer, and you can also use library routines to access Fortran pointers from C. See Section 6.1.6 [Further Interoperability of Fortran with C], page 74.

Here is an example of using C pointers in Fortran:

        use iso_c_binding
        type(c_ptr) :: cptr1, cptr2
        integer, target :: array(7), scalar
        integer, pointer :: pa(:), ps
        cptr1 = c_loc(array(1)) ! The programmer needs to ensure that the
                                ! array is contiguous if required by the C
                                ! procedure
        cptr2 = c_loc(scalar)
        call c_f_pointer(cptr2, ps)
        call c_f_pointer(cptr2, pa, shape=[7])

When converting C to Fortran arrays, the one-dimensional SHAPE argument has to be passed.

If a pointer is a dummy argument of an interoperable procedure, it usually has to be declared using the VALUE attribute. void* matches TYPE(C_PTR), VALUE, while TYPE(C_ PTR) alone matches void**.

Procedure pointers are handled analogously to pointers; the C type is TYPE(C_FUNPTR) and the intrinsic conversion procedures are C_F_PROCPOINTER and C_FUNLOC.

Let us consider two examples of actually passing a procedure pointer from C to Fortran and vice versa. Note that these examples are also very similar to passing ordinary pointers between both languages. First, consider this code in C:

      /* Procedure implemented in Fortran.  */
      void get_values (void (*)(double));
      /* Call-back routine we want called from Fortran.  */
      void
      print_it (double x)
      {
        printf ("Number is %f.\n", x);
      }
      /* Call Fortran routine and pass call-back to it.  */
      void
      foobar ()
      {
        get_values (&print_it);
      }

A matching implementation for get_values in Fortran, that correctly receives the procedure pointer from C and is able to call it, is given in the following MODULE:

      MODULE m

Chapter 6: Mixed-Language Programming 73

IMPLICIT NONE
! Define interface of call-back routine.
ABSTRACT INTERFACE
  SUBROUTINE callback (x)
    USE, INTRINSIC :: ISO_C_BINDING
    REAL(KIND=C_DOUBLE), INTENT(IN), VALUE :: x
  END SUBROUTINE callback
END INTERFACE
CONTAINS
! Define C-bound procedure.
SUBROUTINE get_values (cproc) BIND(C)
  USE, INTRINSIC :: ISO_C_BINDING
  TYPE(C_FUNPTR), INTENT(IN), VALUE :: cproc
  PROCEDURE(callback), POINTER :: proc
  ! Convert C to Fortran procedure pointer.
  CALL C_F_PROCPOINTER (cproc, proc)
  ! Call it.
  CALL proc (1.0_C_DOUBLE)
  CALL proc (-42.0_C_DOUBLE)
  CALL proc (18.12_C_DOUBLE)
END SUBROUTINE get_values
END MODULE m

Next, we want to call a C routine that expects a procedure pointer argument and pass it a Fortran procedure (which clearly must be interoperable!). Again, the C function may be:

int
call_it (int (*func)(int), int arg)
{
return func (arg);
}

It can be used as in the following Fortran code:

MODULE m
USE, INTRINSIC :: ISO_C_BINDING
IMPLICIT NONE
! Define interface of C function.
INTERFACE
  INTEGER(KIND=C_INT) FUNCTION call_it (func, arg) BIND(C)
    USE, INTRINSIC :: ISO_C_BINDING
    TYPE(C_FUNPTR), INTENT(IN), VALUE :: func
    INTEGER(KIND=C_INT), INTENT(IN), VALUE :: arg
  END FUNCTION call_it
END INTERFACE
CONTAINS
! Define procedure passed to C function.
! It must be interoperable!
INTEGER(KIND=C_INT) FUNCTION double_it (arg) BIND(C)

          INTEGER(KIND=C_INT), INTENT(IN), VALUE :: arg
          double_it = arg + arg
        END FUNCTION double_it
        ! Call C function.
        SUBROUTINE foobar ()
          TYPE(C_FUNPTR) :: cproc
          INTEGER(KIND=C_INT) :: i
          ! Get C procedure pointer.
          cproc = C_FUNLOC (double_it)
          ! Use it.
          DO i = 1_C_INT, 10_C_INT
            PRINT *, call_it (cproc, i)
          END DO
        END SUBROUTINE foobar
      END MODULE m

6.1.6 Further Interoperability of Fortran with C

GNU Fortran implements the Technical Specification ISO/IEC TS 29113:2012, which extends the interoperability support of Fortran 2003 and Fortran 2008 and is now part of the 2018 Fortran standard. Besides removing some restrictions and constraints, the Technical Specification adds assumed-type (TYPE(*)) and assumed-rank (DIMENSION(..)) variables and allows for interoperability of assumed-shape, assumed-rank, and deferred-shape arrays, as well as allocatables and pointers. Objects of these types are passed to BIND(C) functions as descriptors with a standard interface, declared in the header file <ISO_Fortran_

binding.h>.

Note: Currently, GNU Fortran does not use internally the array descriptor (dope vector) as specified in the Technical Specification, but uses an array descriptor with different fields in functions without the BIND(C) attribute. Arguments to functions marked BIND(C) are converted to the specified form. If you need to access GNU Fortran’s internal array descriptor, you can use the Chasm Language Interoperability Tools, http://chasm-interop.

sourceforge.net/.

6.2 GNU Fortran Compiler Directives

6.2.1 ATTRIBUTES directive

The Fortran standard describes how a conforming program shall behave; however, the exact implementation is not standardized. In order to allow the user to choose specific implementation details, compiler directives can be used to set attributes of variables and procedures which are not part of the standard. Whether a given attribute is supported and its exact effects depend on both the operating system and on the processor; see Section “C Extensions” in Using the GNU Compiler Collection (GCC) for details.

For procedures and procedure pointers, the following attributes can be used to change the calling convention:

• CDECL – standard C calling convention

• STDCALL – convention where the called procedure pops the stack • FASTCALL – part of the arguments are passed via registers instead using the stack

Besides changing the calling convention, the attributes also influence the decoration of the symbol name, e.g., by a leading underscore or by a trailing at-sign followed by the number of bytes on the stack. When assigning a procedure to a procedure pointer, both should use the same calling convention.

On some systems, procedures and global variables (module variables and COMMON blocks) need special handling to be accessible when they are in a shared library. The following attributes are available:

• DLLEXPORT – provide a global pointer to a pointer in the DLL

• DLLIMPORT – reference the function or variable using a global pointer

For dummy arguments, the NO_ARG_CHECK attribute can be used; in other compilers, it is also known as IGNORE_TKR. For dummy arguments with this attribute actual arguments of any type and kind (similar to TYPE(*)), scalars and arrays of any rank (no equivalent in Fortran standard) are accepted. As with TYPE(*), the argument is unlimited polymorphic and no type information is available. Additionally, the argument may only be passed to dummy arguments with the NO_ARG_CHECK attribute and as argument to the PRESENT intrinsic function and to C_LOC of the ISO_C_BINDING module.

Variables with NO_ARG_CHECK attribute shall be of assumed-type (TYPE(*); recommended) or of type INTEGER, LOGICAL, REAL or COMPLEX. They shall not have the

ALLOCATE, CODIMENSION, INTENT(OUT), POINTER or VALUE attribute; furthermore, they

shall be either scalar or of assumed-size (dimension(*)). As TYPE(*), the NO_ARG_CHECK attribute requires an explicit interface.

• NO_ARG_CHECK – disable the type, kind and rank checking

• DEPRECATED – print a warning when using a such-tagged deprecated procedure, variable or parameter; the warning can be suppressed with -Wno-deprecated-declarations.

• NOINLINE – prevent inlining given function.

• NORETURN – add a hint that a given function cannot return.

• WEAK – emit the declaration of an external symbol as a weak symbol rather than a global. This is primarily useful in defining library functions that can be overridden in user code, though it can also be used with non-function declarations. The overriding symbol must have the same type as the weak symbol.

The attributes are specified using the syntax

!GCC$ ATTRIBUTES attribute-list :: variable-list

where in free-form source code only whitespace is allowed before !GCC$ and in fixed-form source code !GCC$, cGCC$ or *GCC$ shall start in the first column.

For procedures, the compiler directives shall be placed into the body of the procedure; for variables and procedure pointers, they shall be in the same declaration part as the variable or procedure pointer.

6.2.2 UNROLL directive

The syntax of the directive is

   !GCC$ unroll N

You can use this directive to control how many times a loop should be unrolled. It must be placed immediately before a DO loop and applies only to the loop that follows. N is an integer constant specifying the unrolling factor. The values of 0 and 1 block any unrolling of the loop.

6.2.3 BUILTIN directive

The syntax of the directive is

   !GCC$ BUILTIN (B) attributes simd FLAGS IF('target')

You can use this directive to define which middle-end built-ins provide vector implementations. B is name of the middle-end built-in. FLAGS are optional and must be either "(inbranch)" or "(notinbranch)". IF statement is optional and is used to filter multilib ABIs for the built-in that should be vectorized. Example usage:

      !GCC$ builtin (sinf) attributes simd (notinbranch) if('x86_64')

The purpose of the directive is to provide an API among the GCC compiler and the GNU C Library which would define vector implementations of math routines.

6.2.4 IVDEP directive

The syntax of the directive is

   !GCC$ ivdep

This directive tells the compiler to ignore vector dependencies in the following loop. It must be placed immediately before a DO loop and applies only to the loop that follows.

Sometimes the compiler may not have sufficient information to decide whether a particular loop is vectorizable due to potential dependencies between iterations. The purpose of the directive is to tell the compiler that vectorization is safe.

This directive is intended for annotation of existing code. For new code it is recommended to consider OpenMP SIMD directives as potential alternative.

6.2.5 VECTOR directive

The syntax of the directive is

   !GCC$ vector

This directive tells the compiler to vectorize the following loop. It must be placed immediately before a DO loop and applies only to the loop that follows.

6.2.6 NOVECTOR directive

The syntax of the directive is

   !GCC$ novector

This directive tells the compiler to not vectorize the following loop. It must be placed immediately before a DO loop and applies only to the loop that follows.

6.3 Non-Fortran Main Program

Even if you are doing mixed-language programming, it is very likely that you do not need to know or use the information in this section. Since it is about the internal structure of GNU Fortran, it may also change in GCC minor releases.

When you compile a PROGRAM with GNU Fortran, a function with the name main (in the symbol table of the object file) is generated, which initializes the libgfortran library and then calls the actual program which uses the name MAIN__, for historic reasons. If you link GNU Fortran compiled procedures to, e.g., a C or C++ program or to a Fortran program compiled by a different compiler, the libgfortran library is not initialized and thus a few intrinsic procedures do not work properly, e.g. those for obtaining the command-line arguments.

Therefore, if your PROGRAM is not compiled with GNU Fortran and the GNU Fortran compiled procedures require intrinsics relying on the library initialization, you need to initialize the library yourself. Using the default options, gfortran calls _gfortran_set_ args and _gfortran_set_options. The initialization of the former is needed if the called procedures access the command line (and for backtracing); the latter sets some flags based on the standard chosen or to enable backtracing. In typical programs, it is not necessary to call any initialization function.

If your PROGRAM is compiled with GNU Fortran, you shall not call any of the following functions. The libgfortran initialization functions are shown in C syntax but using C bindings they are also accessible from Fortran.

6.3.1 _gfortran_set_args — Save command-line arguments

Description:

_gfortran_set_args saves the command-line arguments; this initialization is required if any of the command-line intrinsics is called. Additionally, it shall be called if backtracing is enabled (see _gfortran_set_options).

Syntax: void _gfortran_set_args (int argc, char *argv[])

Arguments:

argc  number of command line argument strings

argv the command-line argument strings; argv[0] is the pathname of the executable itself.

Example:

                  int main (int argc, char *argv[])
                  {
                    /* Initialize libgfortran.  */
                    _gfortran_set_args (argc, argv);
                    return 0;
                  }

6.3.2 _gfortran_set_options — Set library option flags

Description:

_gfortran_set_options sets several flags related to the Fortran standard to be used, whether backtracing should be enabled and whether range checks should be performed. The syntax allows for upward compatibility since the number of passed flags is specified; for non-passed flags, the default value is used. See also see Section 2.9 [Code Gen Options], page 25. Please note that not all flags are actually used.

Syntax: void _gfortran_set_options (int num, int options[])

Arguments:

num  number of options passed
argv  The list of flag values

option flag list:

option[0] Allowed standard; can give run-time errors if e.g. an input-output edit descriptor is invalid in a given standard. Possible values are (bitwise or-ed)

GFC_STD_F77 (1), GFC_STD_F95_OBS (2), GFC_STD_
F95_DEL (4), GFC_STD_F95 (8), GFC_STD_F2003
(16), GFC_STD_GNU (32), GFC_STD_LEGACY (64),
GFC_STD_F2008 (128), GFC_STD_F2008_OBS (256),
GFC_STD_F2008_TS (512), GFC_STD_F2018 (1024),
GFC_STD_F2018_OBS (2048), and GFC_STD=F2018_
DEL (4096).  Default:  GFC_STD_F95_OBS |
GFC_STD_F95_DEL | GFC_STD_F95 | GFC_STD_F2003
| GFC_STD_F2008 | GFC_STD_F2008_TS | GFC_
STD_F2008_OBS | GFC_STD_F77 | GFC_STD_F2018
| GFC_STD_F2018_OBS | GFC_STD_F2018_DEL |
GFC_STD_GNU | GFC_STD_LEGACY.

option[1] Standard-warning flag; prints a warning to stan-

dard error. Default: GFC_STD_F95_DEL | GFC_STD_
LEGACY.

option[2]  If non zero, enable pedantic checking. Default: off.
option[3]  Unused.

option[4] If non zero, enable backtracing on run-time errors. Default: off. (Default in the compiler: on.) Note: Installs a signal handler and requires command-line initialization using _gfortran_set_args.

option[5]  If non zero, supports signed zeros. Default: enabled.

option[6] Enables run-time checking. Possible values are (bitwise or-ed): GFC RTCHECK BOUNDS (1), GFC RTCHECK ARRAY TEMPS (2), GFC RTCHECK RECURSION (4), GFC RTCHECK DO (8), GFC RTCHECK POINTER (16), GFC RTCHECK MEM (32), GFC RTCHECK BITS (64). Default: disabled.

option[7]  Unused.

option[8] Show a warning when invoking STOP and ERROR STOP if a floating-point exception occurred. Possible values are (bitwise or-ed) GFC_FPE_INVALID

(1), GFC_FPE_DENORMAL (2), GFC_FPE_ZERO (4),
GFC_FPE_OVERFLOW (8), GFC_FPE_UNDERFLOW (16),

GFC_FPE_INEXACT (32). Default: None (0). (Default

in the compiler: GFC_FPE_INVALID | GFC_FPE_
DENORMAL | GFC_FPE_ZERO | GFC_FPE_OVERFLOW |
GFC_FPE_UNDERFLOW.)

Example:

/* Use gfortran 4.9 default options.  */
static int options[] = {68, 511, 0, 0, 1, 1, 0, 0, 31};
_gfortran_set_options (9, &options);

6.3.3 _gfortran_set_convert — Set endian conversion

Description:

_gfortran_set_convert set the representation of data for unformatted files.

Syntax: void _gfortran_set_convert (int conv)

Arguments:

conv

Endian  conversion,  possible  values:
GFC CONVERT NATIVE  (0,  default),

GFC CONVERT SWAP (1), GFC CONVERT BIG (2), GFC CONVERT LITTLE (3).

Example:

int main (int argc, char *argv[])
{
/* Initialize libgfortran.  */
_gfortran_set_args (argc, argv);
_gfortran_set_convert (1);
return 0;
}

6.3.4 _gfortran_set_record_marker — Set length of record markers

Description:

_gfortran_set_record_marker sets the length of record markers for unformatted files.

Syntax: void _gfortran_set_record_marker (int val)

Arguments:

val Length of the record marker; valid values are 4 and 8. Default is 4.

Example:

int main (int argc, char *argv[])
{
/* Initialize libgfortran.  */
_gfortran_set_args (argc, argv);
_gfortran_set_record_marker (8);

                    return 0;
                  }

6.3.5 _gfortran_set_fpe — Enable floating point exception traps

Description:

_gfortran_set_fpe enables floating point exception traps for the specified exceptions. On most systems, this will result in a SIGFPE signal being sent and the program being aborted.

Syntax: void _gfortran_set_fpe (int val)

Arguments:

option[0] IEEE exceptions. Possible values are (bitwise ored) zero (0, default) no trapping, GFC_FPE_INVALID

                      (1), GFC_FPE_DENORMAL (2), GFC_FPE_ZERO (4), GFC_
                      FPE_OVERFLOW (8), GFC_FPE_UNDERFLOW (16), and
                      GFC_FPE_INEXACT (32).

Example:

                  int main (int argc, char *argv[])
                  {
                    /* Initialize libgfortran.  */
                    _gfortran_set_args (argc, argv);
                    /* FPE for invalid operations such as SQRT(-1.0).  */
                    _gfortran_set_fpe (1);
                    return 0;
                  }
6.3.6 _gfortran_set_max_subrecord_length — Set subrecord length

Description:

          _gfortran_set_max_subrecord_length set the maximum length for a sub-

record. This option only makes sense for testing and debugging of unformatted I/O.

Syntax: void _gfortran_set_max_subrecord_length (int val)

Arguments:

val the maximum length for a subrecord; the maximum permitted value is 2147483639, which is also the default.

Example:

                  int main (int argc, char *argv[])
                  {
                    /* Initialize libgfortran.  */
                    _gfortran_set_args (argc, argv);
                    _gfortran_set_max_subrecord_length (8);
                    return 0;
                  }

6.4 Naming and argument-passing conventions

This section gives an overview about the naming convention of procedures and global variables and about the argument passing conventions used by GNU Fortran. If a C binding has been specified, the naming convention and some of the argument-passing conventions change. If possible, mixed-language and mixed-compiler projects should use the better defined C binding for interoperability. See see Section 6.1 [Interoperability with C], page 69.

6.4.1 Naming conventions

According the Fortran standard, valid Fortran names consist of a letter between A to Z, a to z, digits 0, 1 to 9 and underscores (_) with the restriction that names may only start with a letter. As vendor extension, the dollar sign ($) is additionally permitted with the option -fdollar-ok, but not as first character and only if the target system supports it.

By default, the procedure name is the lower-cased Fortran name with an appended underscore (_); using -fno-underscoring no underscore is appended while -fsecond-underscore appends two underscores. Depending on the target system and the calling convention, the procedure might be additionally dressed; for instance, on 32bit Windows with stdcall, an at-sign @ followed by an integer number is appended. For the changing the calling convention, see see Section 6.2 [GNU Fortran Compiler Directives], page 74.

For common blocks, the same convention is used, i.e. by default an underscore is appended to the lower-cased Fortran name. Blank commons have the name __BLNK__.

For procedures and variables declared in the specification space of a module, the name is formed by __, followed by the lower-cased module name, _MOD_, and the lower-cased Fortran name. Note that no underscore is appended.

6.4.2 Argument passing conventions

Subroutines do not return a value (matching C99’s void) while functions either return a value as specified in the platform ABI or the result variable is passed as hidden argument to the function and no result is returned. A hidden result variable is used when the result variable is an array or of type CHARACTER.

Arguments are passed according to the platform ABI. In particular, complex arguments might not be compatible to a struct with two real components for the real and imaginary part. The argument passing matches the one of C99’s _Complex. Functions with scalar complex result variables return their value and do not use a by-reference argument. Note that with the -ff2c option, the argument passing is modified and no longer completely matches the platform ABI. Some other Fortran compilers use f2c semantic by default; this might cause problems with interoperablility.

GNU Fortran passes most arguments by reference, i.e. by passing a pointer to the data. Note that the compiler might use a temporary variable into which the actual argument has been copied, if required semantically (copy-in/copy-out).

For arguments with ALLOCATABLE and POINTER attribute (including procedure pointers), a pointer to the pointer is passed such that the pointer address can be modified in the procedure.

For dummy arguments with the VALUE attribute: Scalar arguments of the type INTEGER, LOGICAL, REAL and COMPLEX are passed by value according to the platform ABI. (As vendor extension and not recommended, using %VAL() in the call to a procedure has the same effect.) For TYPE(C_PTR) and procedure pointers, the pointer itself is passed such that it can be modified without affecting the caller.

For Boolean (LOGICAL) arguments, please note that GCC expects only the integer value 0 and 1. If a GNU Fortran LOGICAL variable contains another integer value, the result is undefined. As some other Fortran compilers use −1 for .TRUE., extra care has to be taken – such as passing the value as INTEGER. (The same value restriction also applies to other front ends of GCC, e.g. to GCC’s C99 compiler for _Bool or GCC’s Ada compiler for Boolean.)

For arguments of CHARACTER type, the character length is passed as a hidden argument at the end of the argument list. For deferred-length strings, the value is passed by reference, otherwise by value. The character length has the C type size_t (or INTEGER(kind=C_SIZE_ T) in Fortran). Note that this is different to older versions of the GNU Fortran compiler, where the type of the hidden character length argument was a C int. In order to retain compatibility with older versions, one can e.g. for the following Fortran procedure

      subroutine fstrlen (s, a)
         character(len=*) :: s
         integer :: a
         print*, len(s)
      end subroutine fstrlen

define the corresponding C prototype as follows:

      #if __GNUC__> 7
      typedef size_t fortran_charlen_t;
      #else
      typedef int fortran_charlen_t;
      #endif
      void fstrlen_(char*, int*, fortran_charlen_t);

In order to avoid such compiler-specific details, for new code it is instead recommended to use the ISO C BINDING feature.

Note with C binding, CHARACTER(len=1) result variables are returned according to the platform ABI and no hidden length argument is used for dummy arguments; with VALUE, those variables are passed by value.

For OPTIONAL dummy arguments, an absent argument is denoted by a NULL pointer, except for scalar dummy arguments of intrinsic type which have the VALUE attribute. For those, a hidden Boolean argument (logical(kind=C_bool),value) is used to indicate whether the argument is present.

Arguments which are assumed-shape, assumed-rank or deferred-rank arrays or, with -fcoarray=lib, allocatable scalar coarrays use an array descriptor. All other arrays pass the address of the first element of the array. With -fcoarray=lib, the token and the offset belonging to nonallocatable coarrays dummy arguments are passed as hidden argument along the character length hidden arguments. The token is an opaque pointer identifying the coarray and the offset is a passed-by-value integer of kind C_PTRDIFF_T, denoting the byte offset between the base address of the coarray and the passed scalar or first element of the passed array.

The arguments are passed in the following order

• Result variable, when the function result is passed by reference

• Character length of the function result, if it is a of type CHARACTER and no C binding is used

• The arguments in the order in which they appear in the Fortran declaration • The present status for optional arguments with value attribute, which are internally

passed by value

• The character length and/or coarray token and offset for the first argument which is a CHARACTER or a nonallocatable coarray dummy argument, followed by the hidden arguments of the next dummy argument of such a type

7 Coarray Programming

7.1 Type and enum ABI Documentation

7.1.1 caf_token_t

Typedef of type void * on the compiler side. Can be any data type on the library side.

7.1.2 caf_register_t

Indicates which kind of coarray variable should be registered.

typedef enum caf_register_t {
  CAF_REGTYPE_COARRAY_STATIC,
  CAF_REGTYPE_COARRAY_ALLOC,
  CAF_REGTYPE_LOCK_STATIC,
  CAF_REGTYPE_LOCK_ALLOC,
  CAF_REGTYPE_CRITICAL,
  CAF_REGTYPE_EVENT_STATIC,
  CAF_REGTYPE_EVENT_ALLOC,
  CAF_REGTYPE_COARRAY_ALLOC_REGISTER_ONLY,
  CAF_REGTYPE_COARRAY_ALLOC_ALLOCATE_ONLY
}
caf_register_t;
   The values CAF_REGTYPE_COARRAY_ALLOC_REGISTER_ONLY and CAF_REGTYPE_COARRAY_

ALLOC_ALLOCATE_ONLY are for allocatable components in derived type coarrays only. The first one sets up the token without allocating memory for allocatable component. The latter one only allocates the memory for an allocatable component in a derived type coarray. The token needs to be setup previously by the REGISTER ONLY. This allows to have allocatable components un-allocated on some images. The status whether an allocatable component is allocated on a remote image can be queried by _caf_is_present which used internally by the ALLOCATED intrinsic.

7.1.3 caf_deregister_t
typedef enum caf_deregister_t {
  CAF_DEREGTYPE_COARRAY_DEREGISTER,
  CAF_DEREGTYPE_COARRAY_DEALLOCATE_ONLY
}
caf_deregister_t;

Allows to specifiy the type of deregistration of a coarray object. The CAF_DEREGTYPE_ COARRAY_DEALLOCATE_ONLY flag is only allowed for allocatable components in derived type coarrays.

7.1.4 caf_reference_t

The structure used for implementing arbitrary reference chains. A CAF_REFERENCE_T allows to specify a component reference or any kind of array reference of any rank supported by gfortran. For array references all kinds as known by the compiler/Fortran standard are supported indicated by a MODE.

typedef enum caf_ref_type_t {
/* Reference a component of a derived type, either regular one or an
   allocatable or pointer type.  For regular ones idx in caf_reference_t is
   set to -1.  */
CAF_REF_COMPONENT,
/* Reference an allocatable array.  */
CAF_REF_ARRAY,
/* Reference a non-allocatable/non-pointer array.  I.e., the coarray object
   has no array descriptor associated and the addressing is done
   completely using the ref.  */
CAF_REF_STATIC_ARRAY
} caf_ref_type_t;
typedef enum caf_array_ref_t {
/* No array ref.  This terminates the array ref.  */
CAF_ARR_REF_NONE = 0,
/* Reference array elements given by a vector.  Only for this mode
   caf_reference_t.u.a.dim[i].v is valid.  */
CAF_ARR_REF_VECTOR,
/* A full array ref (:).  */
CAF_ARR_REF_FULL,
/* Reference a range on elements given by start, end and stride.  */
CAF_ARR_REF_RANGE,
/* Only a single item is referenced given in the start member.  */
CAF_ARR_REF_SINGLE,
/* An array ref of the kind (i:), where i is an arbitrary valid index in the
   array.  The index i is given in the start member.  */
CAF_ARR_REF_OPEN_END,
/* An array ref of the kind (:i), where the lower bound of the array ref
   is given by the remote side.  The index i is given in the end member.  */
CAF_ARR_REF_OPEN_START
} caf_array_ref_t;
/* References to remote components of a derived type.  */
typedef struct caf_reference_t {
/* A pointer to the next ref or NULL.  */
struct caf_reference_t *next;
/* The type of the reference.  */
/* caf_ref_type_t, replaced by int to allow specification in fortran FE.  */
int type;
/* The size of an item referenced in bytes.  I.e. in an array ref this is
   the factor to advance the array pointer with to get to the next item.
   For component refs this gives just the size of the element referenced.  */
size_t item_size;
union {
  struct {

/* The offset (in bytes) of the component in the derived type.
   Unused for allocatable or pointer components.  */
ptrdiff_t offset;
/* The offset (in bytes) to the caf_token associated with this
   component.  NULL, when not allocatable/pointer ref.  */
ptrdiff_t caf_token_offset;
} c;
struct {
/* The mode of the array ref.  See CAF_ARR_REF_*.  */
/* caf_array_ref_t, replaced by unsigend char to allow specification in
   fortran FE.  */
unsigned char mode[GFC_MAX_DIMENSIONS];
/* The type of a static array.  Unset for array's with descriptors.  */
int static_array_type;
/* Subscript refs (s) or vector refs (v).  */
union {
  struct {
    /* The start and end boundary of the ref and the stride.  */
    index_type start, end, stride;
  } s;
  struct {
    /* nvec entries of kind giving the elements to reference.  */
    void *vector;
    /* The number of entries in vector.  */
    size_t nvec;
    /* The integer kind used for the elements in vector.  */
    int kind;
  } v;
} dim[GFC_MAX_DIMENSIONS];
} a;
} u;
} caf_reference_t;

The references make up a single linked list of reference operations. The NEXT member links to the next reference or NULL to indicate the end of the chain. Component and array refs can be arbitrarily mixed as long as they comply to the Fortran standard.

NOTES The member STATIC_ARRAY_TYPE is used only when the TYPE is CAF_REF_ STATIC_ARRAY. The member gives the type of the data referenced. Because no array descriptor is available for a descriptor-less array and type conversion still needs to take place the type is transported here.

At the moment CAF_ARR_REF_VECTOR is not implemented in the front end for descriptorless arrays. The library caf single has untested support for it.

7.1.5 caf_team_t

Opaque pointer to represent a team-handle. This type is a stand-in for the future implementation of teams. It is about to change without further notice.

7.2 Function ABI Documentation

7.2.1 _gfortran_caf_init — Initialiation function

Description:

This function is called at startup of the program before the Fortran main program, if the latter has been compiled with -fcoarray=lib. It takes as arguments the command-line arguments of the program. It is permitted to pass two NULL pointers as argument; if non-NULL, the library is permitted to modify the arguments.

Syntax: void _gfortran_caf_init (int *argc, char ***argv)

Arguments:

argc intent(inout) An integer pointer with the number of arguments passed to the program or NULL.

argv intent(inout) A pointer to an array of strings with the command-line arguments or NULL.

NOTES The function is modelled after the initialization function of the Message Passing Interface (MPI) specification. Due to the way coarray registration works, it might not be the first call to the library. If the main program is not written in Fortran and only a library uses coarrays, it can happen that this function is never called. Therefore, it is recommended that the library does not rely on the passed arguments and whether the call has been done.

7.2.2 _gfortran_caf_finish — Finalization function

Description:

This function is called at the end of the Fortran main program, if it has been compiled with the -fcoarray=lib option.

Syntax: void _gfortran_caf_finish (void)

NOTES For non-Fortran programs, it is recommended to call the function at the end of the main program. To ensure that the shutdown is also performed for programs where this function is not explicitly invoked, for instance non-Fortran programs or calls to the system’s exit() function, the library can use a destructor function. Note that programs can also be terminated using the STOP and ERROR STOP statements; those use different library calls.

7.2.3 _gfortran_caf_this_image — Querying the image number

Description:

This function returns the current image number, which is a positive number.

Syntax: int _gfortran_caf_this_image (int distance)

Arguments:

distance As specified for the this_image intrinsic in TS18508. Shall be a non-negative number.

NOTES If the Fortran intrinsic this_image is invoked without an argument, which is the only permitted form in Fortran 2008, GCC passes 0 as first argument.

7.2.4 _gfortran_caf_num_images — Querying the maximal number of images

Description:

This function returns the number of images in the current team, if distance is 0 or the number of images in the parent team at the specified distance. If failed is -1, the function returns the number of all images at the specified distance; if it is 0, the function returns the number of nonfailed images, and if it is 1, it returns the number of failed images.

Syntax: int _gfortran_caf_num_images(int distance, int failed)

Arguments:

distance the distance from this image to the ancestor. Shall be positive.

failed  shall be -1, 0, or 1

NOTES This function follows TS18508. If the num image intrinsic has no arguments, then the compiler passes distance=0 and failed=-1 to the function.

7.2.5 _gfortran_caf_image_status — Query the status of an image

Description:

Get the status of the image given by the id image of the team given by team. Valid results are zero, for image is ok, STAT_STOPPED_IMAGE from the ISO FORTRAN ENV module to indicate that the image has been stopped and STAT_FAILED_IMAGE also from ISO FORTRAN ENV to indicate that the image has executed a FAIL IMAGE statement.

Syntax: int _gfortran_caf_image_status (int image, caf_team_t * team)

Arguments:

image the positive scalar id of the image in the current TEAM.

team optional; team on the which the inquiry is to be performed.

NOTES This function follows TS18508. Because team-functionality is not yet implemented a null-pointer is passed for the team argument at the moment.

7.2.6 _gfortran_caf_failed_images — Get an array of the indexes of the failed images

Description:

Get an array of image indexes in the current team that have failed. The array is sorted ascendingly. When team is not provided the current team is to be used. When kind is provided then the resulting array is of that integer kind else it is of default integer kind. The returns an unallocated size zero array when no images have failed.

Syntax: int _gfortran_caf_failed_images (caf_team_t * team, int * kind)

Arguments:

team optional; team on the which the inquiry is to be performed.

image  optional; the kind of the resulting integer array.

NOTES This function follows TS18508. Because team-functionality is not yet implemented a null-pointer is passed for the team argument at the moment.

7.2.7 _gfortran_caf_stopped_images — Get an array of the indexes

of the stopped images

Description:

Get an array of image indexes in the current team that have stopped. The array is sorted ascendingly. When team is not provided the current team is to be used. When kind is provided then the resulting array is of that integer kind else it is of default integer kind. The returns an unallocated size zero array when no images have failed.

Syntax: int _gfortran_caf_stopped_images (caf_team_t * team, int * kind)

Arguments:

team optional; team on the which the inquiry is to be performed.

image  optional; the kind of the resulting integer array.

NOTES This function follows TS18508. Because team-functionality is not yet implemented a null-pointer is passed for the team argument at the moment.

7.2.8 _gfortran_caf_register — Registering coarrays

Description:

Registers memory for a coarray and creates a token to identify the coarray. The routine is called for both coarrays with SAVE attribute and using an explicit ALLOCATE statement. If an error occurs and STAT is a NULL pointer, the function shall abort with printing an error message and starting the error termination. If no error occurs and STAT is present, it shall be set to zero. Otherwise, it shall be set to a positive value and, if not-NULL, ERRMSG shall be set to a string describing the failure. The routine shall register the memory provided in the DATA-component of the array descriptor DESC, when that component is non-NULL, else it shall allocate sufficient memory and provide a pointer to it in the DATA-component of DESC. The array descriptor has rank zero, when a scalar object is to be registered and the array descriptor may be invalid after the call to _gfortran_caf_register. When an array is to be allocated the descriptor persists.

For CAF_REGTYPE_COARRAY_STATIC and CAF_REGTYPE_COARRAY_ALLOC, the

passed size is the byte size requested. For CAF_REGTYPE_LOCK_STATIC,

CAF_REGTYPE_LOCK_ALLOC and CAF_REGTYPE_CRITICAL it is the array size or

one for a scalar.

When CAF_REGTYPE_COARRAY_ALLOC_REGISTER_ONLY is used, then only a token

for an allocatable or pointer component is created. The SIZE parameter is not

used then. On the contrary when CAF_REGTYPE_COARRAY_ALLOC_ALLOCATE_

ONLY is specified, then the token needs to be registered by a previous call with

regtype CAF_REGTYPE_COARRAY_ALLOC_REGISTER_ONLY and either the memory

specified in the DESC’s data-ptr is registered or allocate when the data-ptr is

NULL.
Syntax: void caf_register (size_t size, caf_register_t type, caf_token_t
*token, gfc_descriptor_t *desc, int *stat, char *errmsg, size_t
errmsg_len)

Arguments:

size For normal coarrays, the byte size of the coarray to be allocated; for lock types and event types, the number of elements.

type  one of the caf register t types.

token intent(out) An opaque pointer identifying the coarray.

desc  intent(inout) The (pseudo) array descriptor.

stat intent(out) For allocatable coarrays, stores the STAT=; may be NULL

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL

errmsg len the buffer size of errmsg.

NOTES Nonallocatable coarrays have to be registered prior use from remote images. In order to guarantee this, they have to be registered before the main program. This can be achieved by creating constructor functions. That is what GCC does such that also for nonallocatable coarrays the memory is allocated and no static memory is used. The token permits to identify the coarray; to the processor, the token is a nonaliasing pointer. The library can, for instance, store the base address of the coarray in the token, some handle or a more complicated struct. The library may also store the array descriptor DESC when its rank is non-zero.

For lock types, the value shall only be used for checking the allocation status. Note that for critical blocks, the locking is only required on one image; in the locking statement, the processor shall always pass an image index of one for critical-block lock variables (CAF_REGTYPE_CRITICAL). For lock types and critical-block variables, the initial value shall be unlocked (or, respectively, not in critical section) such as the value false; for event types, the initial state should be no event, e.g. zero.

7.2.9 _gfortran_caf_deregister — Deregistering coarrays

Description:

Called to free or deregister the memory of a coarray; the processor calls this function for automatic and explicit deallocation. In case of an error, this function shall fail with an error message, unless the STAT variable is not null. The library is only expected to free memory it allocated itself during a call to

_gfortran_caf_register.
Syntax: void caf_deregister (caf_token_t *token, caf_deregister_t type, int
*stat, char *errmsg, size_t errmsg_len)

Arguments:

token  the token to free.

type the type of action to take for the coarray. A CAF_

            DEREGTYPE_COARRAY_DEALLOCATE_ONLY is allowed

only for allocatable or pointer components of derived type coarrays. The action only deallocates the local memory without deleting the token.

stat  intent(out) Stores the STAT=; may be NULL

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL

errmsg len the buffer size of errmsg.

NOTES For nonalloatable coarrays this function is never called. If a cleanup is required, it has to be handled via the finish, stop and error stop functions, and via destructors.

7.2.10 _gfortran_caf_is_present — Query whether an allocatable or pointer component in a derived type coarray is allocated

Description:

Used to query the coarray library whether an allocatable component in a derived type coarray is allocated on a remote image.

Syntax: void _gfortran_caf_is_present (caf_token_t token, int image_index,
gfc_reference_t *ref)

Arguments:

token  An opaque pointer identifying the coarray.

image index The ID of the remote image; must be a positive number.

ref A chain of references to address the allocatable or pointer component in the derived type coarray. The object reference needs to be a scalar or a full array reference, respectively.

7.2.11 _gfortran_caf_send — Sending data from a local image to a remote image

Description:

Called to send a scalar, an array section or a whole array from a local to a remote image identified by the image index.

Syntax: void _gfortran_caf_send (caf_token_t token, size_t offset,
int image_index, gfc_descriptor_t *dest, caf_vector_t *dst_
vector, gfc_descriptor_t *src, int dst_kind, int src_kind, bool
may_require_tmp, int *stat)

Arguments:

token  intent(in) An opaque pointer identifying the coarray.

offset intent(in) By which amount of bytes the actual data is shifted compared to the base address of the coarray. image index intent(in) The ID of the remote image; must be a positive number.

dest intent(in) Array descriptor for the remote image for the bounds and the size. The base_addr shall not be accessed.

dst vector intent(in) If not NULL, it contains the vector subscript of the destination array; the values are relative to the dimension triplet of the dest argument.

src intent(in) Array descriptor of the local array to be transferred to the remote image

dst kind  intent(in) Kind of the destination argument
src kind  intent(in) Kind of the source argument

may require tmp intent(in) The variable is false when it is known at compile time that the dest and src either cannot overlap or overlap (fully or partially) such that walking src and dest in element wise element order (honoring the stride value) will not lead to wrong results. Otherwise, the value is true.

stat intent(out) when non-NULL give the result of the operation, i.e., zero on success and non-zero on error. When NULL and an error occurs, then an error message is printed and the program is terminated.

NOTES It is permitted to have image index equal the current image; the memory of the send-to and the send-from might (partially) overlap in that case. The implementation has to take care that it handles this case, e.g. using memmove which handles (partially) overlapping memory. If may require tmp is true, the library might additionally create a temporary variable, unless additional checks show that this is not required (e.g. because walking backward is possible or because both arrays are contiguous and memmove takes care of overlap issues). Note that the assignment of a scalar to an array is permitted. In addition, the library has to handle numeric-type conversion and for strings, padding and different character kinds.

7.2.12 _gfortran_caf_get — Getting data from a remote image

Description:

Called to get an array section or a whole array from a remote, image identified by the image index.

Syntax: void _gfortran_caf_get (caf_token_t token, size_t offset, int
image_index, gfc_descriptor_t *src, caf_vector_t *src_vector,
gfc_descriptor_t *dest, int src_kind, int dst_kind, bool may_
require_tmp, int *stat)

Arguments:

token  intent(in) An opaque pointer identifying the coarray.

offset intent(in) By which amount of bytes the actual data is shifted compared to the base address of the coarray. image index intent(in) The ID of the remote image; must be a positive number.

dest intent(out) Array descriptor of the local array to store the data retrieved from the remote image

src intent(in) Array descriptor for the remote image for the bounds and the size. The base_addr shall not be accessed.

src vector intent(in) If not NULL, it contains the vector subscript of the source array; the values are relative to the dimension triplet of the src argument.

dst kind  intent(in) Kind of the destination argument
src kind  intent(in) Kind of the source argument

may require tmp intent(in) The variable is false when it is known at compile time that the dest and src either cannot overlap or overlap (fully or partially) such that walking src and dest in element wise element order (honoring the stride value) will not lead to wrong results. Otherwise, the value is true.

stat intent(out) When non-NULL give the result of the operation, i.e., zero on success and non-zero on error. When NULL and an error occurs, then an error message is printed and the program is terminated.

NOTES It is permitted to have image index equal the current image; the memory of the send-to and the send-from might (partially) overlap in that case. The implementation has to take care that it handles this case, e.g. using memmove which handles (partially) overlapping memory. If may require tmp is true, the library might additionally create a temporary variable, unless additional checks show that this is not required (e.g. because walking backward is possible or because both arrays are contiguous and memmove takes care of overlap issues).

Note that the library has to handle numeric-type conversion and for strings, padding and different character kinds.

7.2.13 _gfortran_caf_sendget — Sending data between remote images

Description:

Called to send a scalar, an array section or a whole array from a remote image identified by the src image index to a remote image identified by the dst image index.

Syntax: void _gfortran_caf_sendget (caf_token_t dst_token, size_t dst_
offset, int dst_image_index, gfc_descriptor_t *dest, caf_vector_t
*dst_vector, caf_token_t src_token, size_t src_offset, int src_
image_index, gfc_descriptor_t *src, caf_vector_t *src_vector, int
dst_kind, int src_kind, bool may_require_tmp, int *stat)

Arguments:

dst token intent(in) An opaque pointer identifying the destination coarray.

dst offset intent(in) By which amount of bytes the actual data is shifted compared to the base address of the destination coarray.

dst image index intent(in) The ID of the destination remote image; must be a positive number.

dest intent(in) Array descriptor for the destination remote image for the bounds and the size. The base_addr shall not be accessed.

dst vector intent(int) If not NULL, it contains the vector subscript of the destination array; the values are relative to the dimension triplet of the dest argument.

src token intent(in) An opaque pointer identifying the source coarray.

src offset intent(in) By which amount of bytes the actual data is shifted compared to the base address of the source coarray.

src image index intent(in) The ID of the source remote image; must be a positive number.

src intent(in) Array descriptor of the local array to be transferred to the remote image.

src vector intent(in) Array descriptor of the local array to be transferred to the remote image

dst kind  intent(in) Kind of the destination argument
src kind  intent(in) Kind of the source argument

may require tmp intent(in) The variable is false when it is known at compile time that the dest and src either cannot overlap or overlap (fully or partially) such that walking src and dest in element wise element order (honoring the stride value) will not lead to wrong results. Otherwise, the value is true.

stat intent(out) when non-NULL give the result of the operation, i.e., zero on success and non-zero on error. When NULL and an error occurs, then an error message is printed and the program is terminated.

NOTES It is permitted to have the same image index for both src image index and dst image index; the memory of the send-to and the send-from might (partially) overlap in that case. The implementation has to take care that it handles this case, e.g. using memmove which handles (partially) overlapping memory. If may require tmp is true, the library might additionally create a temporary variable, unless additional checks show that this is not required (e.g. because walking backward is possible or because both arrays are contiguous and memmove takes care of overlap issues).

Note that the assignment of a scalar to an array is permitted. In addition, the library has to handle numeric-type conversion and for strings, padding and different character kinds.

7.2.14 _gfortran_caf_send_by_ref — Sending data from a local image to a remote image with enhanced referencing options

Description:

Called to send a scalar, an array section or a whole array from a local to a remote image identified by the image index.

Syntax: void _gfortran_caf_send_by_ref (caf_token_t token, int image_index,
gfc_descriptor_t *src, caf_reference_t *refs, int dst_kind, int
src_kind, bool may_require_tmp, bool dst_reallocatable, int *stat,
int dst_type)

Arguments:

token  intent(in) An opaque pointer identifying the coarray.

image index intent(in) The ID of the remote image; must be a positive number.

src intent(in) Array descriptor of the local array to be transferred to the remote image

refs intent(in) The references on the remote array to store the data given by src. Guaranteed to have at least one entry.

dst kind  intent(in) Kind of the destination argument
src kind  intent(in) Kind of the source argument

may require tmp intent(in) The variable is false when it is known at compile time that the dest and src either cannot overlap or overlap (fully or partially) such that walking src and dest in element wise element order (honoring the stride value) will not lead to wrong results. Otherwise, the value is true.

dst reallocatable intent(in) Set when the destination is of allocatable or pointer type and the refs will allow reallocation, i.e., the ref is a full array or component ref.

stat intent(out) When non-NULL give the result of the operation, i.e., zero on success and non-zero on error. When NULL and an error occurs, then an error message is printed and the program is terminated.

dst type intent(in) Give the type of the destination. When the destination is not an array, than the precise type, e.g. of a component in a derived type, is not known, but provided here.

NOTES It is permitted to have image index equal the current image; the memory of the send-to and the send-from might (partially) overlap in that case. The implementation has to take care that it handles this case, e.g. using memmove which handles (partially) overlapping memory. If may require tmp is true, the library might additionally create a temporary variable, unless additional checks show that this is not required (e.g. because walking backward is possible or because both arrays are contiguous and memmove takes care of overlap issues).

Note that the assignment of a scalar to an array is permitted. In addition, the library has to handle numeric-type conversion and for strings, padding and different character kinds.

Because of the more complicated references possible some operations may be unsupported by certain libraries. The library is expected to issue a precise error message why the operation is not permitted.

7.2.15 _gfortran_caf_get_by_ref — Getting data from a remote image using enhanced references

Description:

Called to get a scalar, an array section or a whole array from a remote image identified by the image index.

Syntax: void _gfortran_caf_get_by_ref (caf_token_t token, int image_index,
caf_reference_t *refs, gfc_descriptor_t *dst, int dst_kind, int
src_kind, bool may_require_tmp, bool dst_reallocatable, int *stat,
int src_type)

Arguments:

token  intent(in) An opaque pointer identifying the coarray.

image index intent(in) The ID of the remote image; must be a positive number.

refs intent(in) The references to apply to the remote structure to get the data.

dst intent(in) Array descriptor of the local array to store the data transferred from the remote image. May be reallocated where needed and when DST REALLOCATABLE allows it.

dst kind  intent(in) Kind of the destination argument
src kind  intent(in) Kind of the source argument

may require tmp intent(in) The variable is false when it is known at compile time that the dest and src either cannot overlap or overlap (fully or partially) such that walking src and dest in element wise element order (honoring the stride value) will not lead to wrong results. Otherwise, the value is true.

dst reallocatable intent(in) Set when DST is of allocatable or pointer type and its refs allow reallocation, i.e., the full array or a component is referenced.

stat intent(out) When non-NULL give the result of the operation, i.e., zero on success and non-zero on error. When NULL and an error occurs, then an error message is printed and the program is terminated. src type intent(in) Give the type of the source. When the source is not an array, than the precise type, e.g. of a component in a derived type, is not known, but provided here.

NOTES It is permitted to have image_index equal the current image; the memory of the send-to and the send-from might (partially) overlap in that case. The implementation has to take care that it handles this case, e.g. using memmove which handles (partially) overlapping memory. If may require tmp is true, the library might additionally create a temporary variable, unless additional checks show that this is not required (e.g. because walking backward is possible or because both arrays are contiguous and memmove takes care of overlap issues). Note that the library has to handle numeric-type conversion and for strings, padding and different character kinds. Because of the more complicated references possible some operations may be unsupported by certain libraries. The library is expected to issue a precise error message why the operation is not permitted.

7.2.16 _gfortran_caf_sendget_by_ref — Sending data between

remote images using enhanced references on both sides

Description:

Called to send a scalar, an array section or a whole array from a remote image identified by the src image index to a remote image identified by the dst image index.

Syntax: void _gfortran_caf_sendget_by_ref (caf_token_t dst_token, int
dst_image_index, caf_reference_t *dst_refs, caf_token_t src_token,
int src_image_index, caf_reference_t *src_refs, int dst_kind, int
src_kind, bool may_require_tmp, int *dst_stat, int *src_stat, int
dst_type, int src_type)

Arguments:

dst token intent(in) An opaque pointer identifying the destination coarray.

dst image index intent(in) The ID of the destination remote image; must be a positive number.

dst refs intent(in) The references on the remote array to store the data given by the source. Guaranteed to have at least one entry.

src token intent(in) An opaque pointer identifying the source coarray.

src image index intent(in) The ID of the source remote image; must be a positive number.

src refs intent(in) The references to apply to the remote structure to get the data.

dst kind  intent(in) Kind of the destination argument
src kind  intent(in) Kind of the source argument

may require tmp intent(in) The variable is false when it is known at compile time that the dest and src either cannot overlap or overlap (fully or partially) such that walking src and dest in element wise element order (honoring the stride value) will not lead to wrong results. Otherwise, the value is true.

dst stat intent(out) when non-NULL give the result of the sendoperation, i.e., zero on success and non-zero on error. When NULL and an error occurs, then an error message is printed and the program is terminated.

src stat intent(out) When non-NULL give the result of the getoperation, i.e., zero on success and non-zero on error. When NULL and an error occurs, then an error message is printed and the program is terminated.

dst type intent(in) Give the type of the destination. When the destination is not an array, than the precise type, e.g. of a component in a derived type, is not known, but provided here.

src type intent(in) Give the type of the source. When the source is not an array, than the precise type, e.g. of a component in a derived type, is not known, but provided here.

NOTES It is permitted to have the same image index for both src image index and dst image index; the memory of the send-to and the send-from might (partially) overlap in that case. The implementation has to take care that it handles this case, e.g. using memmove which handles (partially) overlapping memory. If may require tmp is true, the library might additionally create a temporary variable, unless additional checks show that this is not required (e.g. because walking backward is possible or because both arrays are contiguous and memmove takes care of overlap issues). Note that the assignment of a scalar to an array is permitted. In addition, the library has to handle numeric-type conversion and for strings, padding and different character kinds. Because of the more complicated references possible some operations may be unsupported by certain libraries. The library is expected to issue a precise error message why the operation is not permitted.

7.2.17 _gfortran_caf_lock — Locking a lock variable

Description:

Acquire a lock on the given image on a scalar locking variable or for the given array element for an array-valued variable. If the acquired lock is NULL, the function returns after having obtained the lock. If it is non-NULL, then acquired lock is assigned the value true (one) when the lock could be obtained and false (zero) otherwise. Locking a lock variable which has already been locked by the same image is an error.

Syntax: void _gfortran_caf_lock (caf_token_t token, size_t index, int
image_index, int *acquired_lock, int *stat, char *errmsg, size_t
errmsg_len)

Arguments:

token  intent(in) An opaque pointer identifying the coarray.

index intent(in) Array index; first array index is 0. For scalars, it is always 0.

image index intent(in) The ID of the remote image; must be a positive number.

acquired lock intent(out) If not NULL, it returns whether lock could be obtained.

stat  intent(out) Stores the STAT=; may be NULL.

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL.

errmsg len intent(in) the buffer size of errmsg

NOTES This function is also called for critical blocks; for those, the array index is always zero and the image index is one. Libraries are permitted to use other images for critical-block locking variables.

7.2.18 _gfortran_caf_lock — Unlocking a lock variable

Description:

Release a lock on the given image on a scalar locking variable or for the given array element for an array-valued variable. Unlocking a lock variable which is unlocked or has been locked by a different image is an error.

Syntax: void _gfortran_caf_unlock (caf_token_t token, size_t index, int
image_index, int *stat, char *errmsg, size_t errmsg_len)

Arguments:

token  intent(in) An opaque pointer identifying the coarray.

index intent(in) Array index; first array index is 0. For scalars, it is always 0.

image index intent(in) The ID of the remote image; must be a positive number.

stat intent(out) For allocatable coarrays, stores the STAT=; may be NULL.

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL.

errmsg len intent(in) the buffer size of errmsg

NOTES This function is also called for critical block; for those, the array index is always zero and the image index is one. Libraries are permitted to use other images for critical-block locking variables.

7.2.19 _gfortran_caf_event_post — Post an event

Description:

Increment the event count of the specified event variable.

Syntax: void _gfortran_caf_event_post (caf_token_t token, size_t index, int
image_index, int *stat, char *errmsg, size_t errmsg_len)

Arguments:

token  intent(in) An opaque pointer identifying the coarray.

index intent(in) Array index; first array index is 0. For scalars, it is always 0.

image index intent(in) The ID of the remote image; must be a positive number; zero indicates the current image, when accessed noncoindexed.

stat  intent(out) Stores the STAT=; may be NULL.

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL.

errmsg len intent(in) the buffer size of errmsg

NOTES This acts like an atomic add of one to the remote image’s event variable. The statement is an image-control statement but does not imply sync memory. Still, all preceeding push communications of this image to the specified remote image have to be completed before event_wait on the remote image returns.

7.2.20 _gfortran_caf_event_wait — Wait that an event occurred

Description:

Wait until the event count has reached at least the specified until count; if so, atomically decrement the event variable by this amount and return.

Syntax: void _gfortran_caf_event_wait (caf_token_t token, size_t index, int
until_count, int *stat, char *errmsg, size_t errmsg_len)

Arguments:

token  intent(in) An opaque pointer identifying the coarray.

index intent(in) Array index; first array index is 0. For scalars, it is always 0.

until count intent(in) The number of events which have to be available before the function returns.

stat  intent(out) Stores the STAT=; may be NULL.

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL.

errmsg len intent(in) the buffer size of errmsg

NOTES This function only operates on a local coarray. It acts like a loop checking atomically the value of the event variable, breaking if the value is greater or equal the requested number of counts. Before the function returns, the event variable has to be decremented by the requested until count value. A possible implementation would be a busy loop for a certain number of spins (possibly depending on the number of threads relative to the number of available cores) followed by another waiting strategy such as a sleeping wait (possibly with an increasing number of sleep time) or, if possible, a futex wait.

The statement is an image-control statement but does not imply sync memory. Still, all preceeding push communications of this image to the specified remote image have to be completed before event_wait on the remote image returns.

7.2.21 _gfortran_caf_event_query — Query event count

Description:

Return the event count of the specified event variable.

Syntax: void _gfortran_caf_event_query (caf_token_t token, size_t index,
int image_index, int *count, int *stat)

Arguments:

token  intent(in) An opaque pointer identifying the coarray.

index intent(in) Array index; first array index is 0. For scalars, it is always 0.

image index intent(in) The ID of the remote image; must be a positive number; zero indicates the current image when accessed noncoindexed.

count intent(out) The number of events currently posted to the event variable.

stat  intent(out) Stores the STAT=; may be NULL.

NOTES The typical use is to check the local event variable to only call event_wait when the data is available. However, a coindexed variable is permitted; there is no ordering or synchronization implied. It acts like an atomic fetch of the value of the event variable.

7.2.22 _gfortran_caf_sync_all — All-image barrier

Description:

Synchronization of all images in the current team; the program only continues on a given image after this function has been called on all images of the current team. Additionally, it ensures that all pending data transfers of previous segment have completed.

Syntax: void _gfortran_caf_sync_all (int *stat, char *errmsg, size_t
errmsg_len)

Arguments:

stat intent(out) Stores the status STAT= and may be NULL.

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL.

errmsg len intent(in) the buffer size of errmsg

7.2.23 _gfortran_caf_sync_images — Barrier for selected images

Description:

Synchronization between the specified images; the program only continues on a given image after this function has been called on all images specified for that image. Note that one image can wait for all other images in the current team (e.g. via sync images(*)) while those only wait for that specific image. Additionally, sync images ensures that all pending data transfers of previous segments have completed.

Syntax: void _gfortran_caf_sync_images (int count, int images[], int *stat,
char *errmsg, size_t errmsg_len)

Arguments:

count intent(in) The number of images which are provided in the next argument. For a zero-sized array, the value is zero. For sync images (*), the value is −1.

images intent(in) An array with the images provided by the user. If count is zero, a NULL pointer is passed.

stat intent(out) Stores the status STAT= and may be NULL.

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL.

errmsg len intent(in) the buffer size of errmsg

7.2.24 _gfortran_caf_sync_memory — Wait for completion of segment-memory operations

Description:

Acts as optimization barrier between different segments. It also ensures that all pending memory operations of this image have been completed.

Syntax: void _gfortran_caf_sync_memory (int *stat, char *errmsg, size_t
errmsg_len)

Arguments:

stat intent(out) Stores the status STAT= and may be NULL.

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL.

errmsg len intent(in) the buffer size of errmsg

NOTE A simple implementation could be

__asm____volatile__("":::"memory") to prevent code movements.

7.2.25 _gfortran_caf_error_stop — Error termination with exit code

Description:

Invoked for an ERROR STOP statement which has an integer argument. The function should terminate the program with the specified exit code.

Syntax: void _gfortran_caf_error_stop (int error)

Arguments:

error  intent(in) The exit status to be used.

7.2.26 _gfortran_caf_error_stop_str — Error termination with

string

Description:

Invoked for an ERROR STOP statement which has a string as argument. The function should terminate the program with a nonzero-exit code.

Syntax: void _gfortran_caf_error_stop (const char *string, size_t len)

Arguments:

string  intent(in) the error message (not zero terminated)
len  intent(in) the length of the string

7.2.27 _gfortran_caf_fail_image — Mark the image failed and end its execution

Description:

Invoked for an FAIL IMAGE statement. The function should terminate the current image.

Syntax: void _gfortran_caf_fail_image ()

NOTES This function follows TS18508.

7.2.28 _gfortran_caf_atomic_define — Atomic variable assignment

Description:

Assign atomically a value to an integer or logical variable.

Syntax: void _gfortran_caf_atomic_define (caf_token_t token, size_t offset,
int image_index, void *value, int *stat, int type, int kind)

Arguments:

token  intent(in) An opaque pointer identifying the coarray.

offset intent(in) By which amount of bytes the actual data is shifted compared to the base address of the coarray.

image index intent(in) The ID of the remote image; must be a positive number; zero indicates the current image when used noncoindexed.

value intent(in) the value to be assigned, passed by reference

stat intent(out) Stores the status STAT= and may be NULL.

type intent(in) The data type, i.e. BT_INTEGER (1) or BT_

            LOGICAL (2).

kind  intent(in) The kind value (only 4; always int)

7.2.29 _gfortran_caf_atomic_ref — Atomic variable reference

Description:

Reference atomically a value of a kind-4 integer or logical variable.

Syntax: void _gfortran_caf_atomic_ref (caf_token_t token, size_t offset,
int image_index, void *value, int *stat, int type, int kind)

Arguments:

token  intent(in) An opaque pointer identifying the coarray.

offset intent(in) By which amount of bytes the actual data is shifted compared to the base address of the coarray.

image index intent(in) The ID of the remote image; must be a positive number; zero indicates the current image when used noncoindexed.

value intent(out) The variable assigned the atomically referenced variable.

stat intent(out) Stores the status STAT= and may be NULL.

type the data type, i.e. BT_INTEGER (1) or BT_LOGICAL (2).

kind  The kind value (only 4; always int)

7.2.30 _gfortran_caf_atomic_cas — Atomic compare and swap

Description:

Atomic compare and swap of a kind-4 integer or logical variable. Assigns atomically the specified value to the atomic variable, if the latter has the value specified by the passed condition value.

Syntax: void _gfortran_caf_atomic_cas (caf_token_t token, size_t offset,
int image_index, void *old, void *compare, void *new_val, int *stat,
int type, int kind)

Arguments:

token  intent(in) An opaque pointer identifying the coarray.

offset intent(in) By which amount of bytes the actual data is shifted compared to the base address of the coarray.

image index intent(in) The ID of the remote image; must be a positive number; zero indicates the current image when used noncoindexed.

old intent(out) The value which the atomic variable had just before the cas operation.

compare  intent(in) The value used for comparision.

new val intent(in) The new value for the atomic variable, assigned to the atomic variable, if compare equals the value of the atomic variable.

stat intent(out) Stores the status STAT= and may be NULL.

type intent(in) the data type, i.e. BT_INTEGER (1) or BT_

LOGICAL (2).

kind  intent(in) The kind value (only 4; always int)

7.2.31 _gfortran_caf_atomic_op — Atomic operation

Description:

Apply an operation atomically to an atomic integer or logical variable. After the operation, old contains the value just before the operation, which, respectively, adds (GFC CAF ATOMIC ADD) atomically the value to the atomic integer variable or does a bitwise AND, OR or exclusive OR between the atomic variable and value; the result is then stored in the atomic variable.

Syntax: void _gfortran_caf_atomic_op (int op, caf_token_t token, size_t
offset, int image_index, void *value, void *old, int *stat, int type,
int kind)

Arguments:

op intent(in) the operation to be performed; possible

values GFC_CAF_ATOMIC_ADD (1), GFC_CAF_ATOMIC_
AND (2), GFC_CAF_ATOMIC_OR (3), GFC_CAF_ATOMIC_

XOR (4).

token  intent(in) An opaque pointer identifying the coarray.

offset intent(in) By which amount of bytes the actual data is shifted compared to the base address of the coarray.

image index intent(in) The ID of the remote image; must be a positive number; zero indicates the current image when used noncoindexed.

old intent(out) The value which the atomic variable had just before the atomic operation.

val intent(in) The new value for the atomic variable, assigned to the atomic variable, if compare equals the value of the atomic variable.

stat intent(out) Stores the status STAT= and may be NULL.

type intent(in) the data type, i.e. BT_INTEGER (1) or BT_

LOGICAL (2)

kind  intent(in) the kind value (only 4; always int)

7.2.32 _gfortran_caf_co_broadcast — Sending data to all images

Description:

Distribute a value from a given image to all other images in the team. Has to be called collectively.

Syntax: void _gfortran_caf_co_broadcast (gfc_descriptor_t *a, int
source_image, int *stat, char *errmsg, size_t errmsg_len)

Arguments:

a intent(inout) An array descriptor with the data to be broadcasted (on source image) or to be received (other images).

source image intent(in) The ID of the image from which the data should be broadcasted. stat intent(out) Stores the status STAT= and may be NULL.

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL.

errmsg len intent(in) the buffer size of errmsg.

7.2.33 _gfortran_caf_co_max — Collective maximum reduction

Description:

Calculates for each array element of the variable a the maximum value for that element in the current team; if result image has the value 0, the result shall be stored on all images, otherwise, only on the specified image. This function operates on numeric values and character strings.

Syntax: void _gfortran_caf_co_max (gfc_descriptor_t *a, int result_image,
int *stat, char *errmsg, int a_len, size_t errmsg_len)

Arguments:

a intent(inout) An array descriptor for the data to be processed. On the destination image(s) the result overwrites the old content.

result image intent(in) The ID of the image to which the reduced value should be copied to; if zero, it has to be copied to all images.

stat intent(out) Stores the status STAT= and may be NULL.

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL.

a len  intent(in) the string length of argument a
errmsg len intent(in) the buffer size of errmsg

NOTES If result image is nonzero, the data in the array descriptor a on all images except of the specified one become undefined; hence, the library may make use of this.

7.2.34 _gfortran_caf_co_min — Collective minimum reduction

Description:

Calculates for each array element of the variable a the minimum value for that element in the current team; if result image has the value 0, the result shall be stored on all images, otherwise, only on the specified image. This function operates on numeric values and character strings.

Syntax: void _gfortran_caf_co_min (gfc_descriptor_t *a, int result_image,
int *stat, char *errmsg, int a_len, size_t errmsg_len)

Arguments:

a intent(inout) An array descriptor for the data to be processed. On the destination image(s) the result overwrites the old content. result image intent(in) The ID of the image to which the reduced value should be copied to; if zero, it has to be copied to all images.

stat intent(out) Stores the status STAT= and may be NULL.

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL.

a len  intent(in) the string length of argument a
errmsg len intent(in) the buffer size of errmsg

NOTES If result image is nonzero, the data in the array descriptor a on all images except of the specified one become undefined; hence, the library may make use of this.

7.2.35 _gfortran_caf_co_sum — Collective summing reduction

Description:

Calculates for each array element of the variable a the sum of all values for that element in the current team; if result image has the value 0, the result shall be stored on all images, otherwise, only on the specified image. This function operates on numeric values only.

Syntax: void _gfortran_caf_co_sum (gfc_descriptor_t *a, int result_image,
int *stat, char *errmsg, size_t errmsg_len)

Arguments:

a intent(inout) An array descriptor with the data to be processed. On the destination image(s) the result overwrites the old content.

result image intent(in) The ID of the image to which the reduced value should be copied to; if zero, it has to be copied to all images.

stat intent(out) Stores the status STAT= and may be NULL.

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL.

errmsg len intent(in) the buffer size of errmsg

NOTES If result image is nonzero, the data in the array descriptor a on all images except of the specified one become undefined; hence, the library may make use of this.

7.2.36 _gfortran_caf_co_reduce — Generic collective reduction

Description:

Calculates for each array element of the variable a the reduction value for that element in the current team; if result image has the value 0, the result shall be stored on all images, otherwise, only on the specified image. The opr is a pure function doing a mathematically commutative and associative operation.

The opr flags denote the following; the values are bitwise ored. GFC_CAF_ BYREF (1) if the result should be returned by reference; GFC_CAF_HIDDENLEN (2) whether the result and argument string lengths shall be specified as hidden arguments; GFC_CAF_ARG_VALUE (4) whether the arguments shall be passed by value, GFC_CAF_ARG_DESC (8) whether the arguments shall be passed by descriptor.

Syntax: void _gfortran_caf_co_reduce (gfc_descriptor_t *a, void * (*opr)
(void *, void *), int opr_flags, int result_image, int *stat, char
*errmsg, int a_len, size_t errmsg_len)

Arguments:

a intent(inout) An array descriptor with the data to be processed. On the destination image(s) the result overwrites the old content.

opr  intent(in) Function pointer to the reduction function
opr flags  intent(in) Flags regarding the reduction function

result image intent(in) The ID of the image to which the reduced value should be copied to; if zero, it has to be copied to all images.

stat intent(out) Stores the status STAT= and may be NULL.

errmsg intent(out) When an error occurs, this will be set to an error message; may be NULL.

a len  intent(in) the string length of argument a
errmsg len intent(in) the buffer size of errmsg

NOTES If result image is nonzero, the data in the array descriptor a on all images except of the specified one become undefined; hence, the library may make use of this. For character arguments, the result is passed as first argument, followed by the result string length, next come the two string arguments, followed by the two hidden string length arguments. With C binding, there are no hidden arguments and by-reference passing and either only a single character is passed or an array descriptor.

8 Intrinsic Procedures

8.1 Introduction to intrinsic procedures

The intrinsic procedures provided by GNU Fortran include procedures required by the Fortran 95 and later supported standards, and a set of intrinsic procedures for backwards compatibility with G77. Any conflict between a description here and a description in the Fortran standards is unintentional, and the standard(s) should be considered authoritative.

The enumeration of the KIND type parameter is processor defined in the Fortran 95 standard. GNU Fortran defines the default integer type and default real type by INTEGER(KIND=4) and REAL(KIND=4), respectively. The standard mandates that both data types shall have another kind, which have more precision. On typical target architectures supported by gfortran, this kind type parameter is KIND=8. Hence, REAL(KIND=8) and DOUBLE PRECISION are equivalent. In the description of generic intrinsic procedures, the kind type parameter will be specified by KIND=*, and in the description of specific names for an intrinsic procedure the kind type parameter will be explicitly given (e.g., REAL(KIND=4) or REAL(KIND=8)). Finally, for brevity the optional KIND= syntax will be omitted.

Many of the intrinsic procedures take one or more optional arguments. This document follows the convention used in the Fortran 95 standard, and denotes such arguments by square brackets.

GNU Fortran offers the -std= command-line option, which can be used to restrict the set of intrinsic procedures to a given standard. By default, gfortran sets the -std=gnu option, and so all intrinsic procedures described here are accepted. There is one caveat. For a select group of intrinsic procedures, g77 implemented both a function and a subroutine. Both classes have been implemented in gfortran for backwards compatibility with g77. It is noted here that these functions and subroutines cannot be intermixed in a given subprogram. In the descriptions that follow, the applicable standard for each intrinsic procedure is noted.

8.2 ABORT — Abort the program

Description:

ABORT causes immediate termination of the program. On operating systems that support a core dump, ABORT will produce a core dump. It will also print a backtrace, unless -fno-backtrace is given.

Standard: GNU extension

Class: Subroutine

Syntax: CALL ABORT

Return value:

Does not return.

Example:

                  program test_abort
                    integer :: i = 1, j = 2
                    if (i /= j) call abort
                  end program test_abort

See also: Section 8.101 [EXIT], page 179, Section 8.163 [KILL], page 218, Section 8.39 [BACKTRACE], page 138,

8.3 ABS — Absolute value

Description:

ABS(A) computes the absolute value of A.

Standard: Fortran 77 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = ABS(A)

Arguments:

A The type of the argument shall be an INTEGER, REAL,

            or COMPLEX.

Return value:

The return value is of the same type and kind as the argument except the return value is REAL for a COMPLEX argument.

Example:

      program test_abs
        integer :: i = -1
        real :: x = -1.e0
        complex :: z = (-1.e0,0.e0)
        i = abs(i)
        x = abs(x)
        x = abs(z)
      end program test_abs

Specific names:

Name  Argument  Return type Standard

ABS(A)         REAL(4) A      REAL(4)  Fortran 77 and later
CABS(A)        COMPLEX(4) A   REAL(4)  Fortran 77 and later
DABS(A)        REAL(8) A      REAL(8)  Fortran 77 and later
IABS(A)        INTEGER(4) A   INTEGER(4) Fortran 77 and later
BABS(A)        INTEGER(1) A   INTEGER(1) GNU extension
IIABS(A)       INTEGER(2) A   INTEGER(2) GNU extension
JIABS(A)       INTEGER(4) A   INTEGER(4) GNU extension
KIABS(A)       INTEGER(8) A   INTEGER(8) GNU extension
ZABS(A)        COMPLEX(8) A   REAL(8)  GNU extension
CDABS(A)       COMPLEX(8) A   REAL(8)  GNU extension

8.4 ACCESS — Checks file access modes

Description:

ACCESS(NAME, MODE) checks whether the file NAME exists, is readable, writable or executable. Except for the executable check, ACCESS can be replaced by Fortran 95’s INQUIRE.

Standard: GNU extension

Class: Inquiry function

Syntax: RESULT = ACCESS(NAME, MODE)

Arguments:

NAME Scalar CHARACTER of default kind with the file name. Trailing blank are ignored unless the character achar(0) is present, then all characters up to and excluding achar(0) are used as file name.

MODE Scalar CHARACTER of default kind with the file access mode, may be any concatenation of "r" (readable), "w" (writable) and "x" (executable), or " " to check for existence.

Return value:

Returns a scalar INTEGER, which is 0 if the file is accessible in the given mode; otherwise or if an invalid argument has been given for MODE the value 1 is returned.

Example:

                  program access_test
                    implicit none
                    character(len=*), parameter :: file  = 'test.dat'
                    character(len=*), parameter :: file2 = 'test.dat  '//achar(0)
                    if(access(file,' ') == 0) print *, trim(file),' is exists'
                    if(access(file,'r') == 0) print *, trim(file),' is readable'
                    if(access(file,'w') == 0) print *, trim(file),' is writable'
                    if(access(file,'x') == 0) print *, trim(file),' is executable'
                    if(access(file2,'rwx') == 0) &
                      print *, trim(file2),' is readable, writable and executable'
                  end program access_test

8.5 ACHAR — Character in ASCII collating sequence

Description:

ACHAR(I) returns the character located at position I in the ASCII collating sequence.

Standard: Fortran 77 and later, with KIND argument Fortran 2003 and later

Class: Elemental function

Syntax: RESULT = ACHAR(I [, KIND])

Arguments:

I  The type shall be INTEGER.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type CHARACTER with a length of one. If the KIND argument is present, the return value is of the specified kind and of the default kind otherwise.

Example:

      program test_achar
        character c
        c = achar(32)
      end program test_achar

Note: See Section 8.143 [ICHAR], page 206, for a discussion of converting between numerical values and formatted string representations.

See also: Section 8.59 [CHAR], page 149, Section 8.135 [IACHAR], page 201, Section 8.143 [ICHAR], page 206,

8.6 ACOS — Arccosine function

Description:

ACOS(X) computes the arccosine of X (inverse of COS(X)).

Standard: Fortran 77 and later, for a complex argument Fortran 2008 or later

Class: Elemental function

Syntax: RESULT = ACOS(X)

Arguments:

X The type shall either be REAL with a magnitude that is less than or equal to one - or the type shall be

            COMPLEX.

Return value:

The return value is of the same type and kind as X. The real part of the result is in radians and lies in the range 0 ≤ ℜ acos(x) ≤ π.

Example:

      program test_acos
        real(8) :: x = 0.866_8
        x = acos(x)
      end program test_acos

Specific names:

Name  Argument  Return type Standard

ACOS(X)        REAL(4) X      REAL(4)  Fortran 77 and later
DACOS(X)       REAL(8) X      REAL(8)  Fortran 77 and later

See also: Inverse function:

Section 8.73 [COS], page 160, Degrees function: Section 8.7 [ACOSD], page 114,

8.7 ACOSD — Arccosine function, degrees

Description:

ACOSD(X) computes the arccosine of X in degrees (inverse of COSD(X)). This function is for compatibility only and should be avoided in favor of standard constructs wherever possible.

Standard: GNU extension, enabled with -fdec-math

Class: Elemental function

Syntax: RESULT = ACOSD(X)

Arguments:

X The type shall either be REAL with a magnitude that is less than or equal to one - or the type shall be

                      COMPLEX.

Return value:

The return value is of the same type and kind as X. The real part of the result is in degrees and lies in the range 0 ≤ ℜ acos(x) ≤ 180.

Example:

                  program test_acosd
                    real(8) :: x = 0.866_8
                    x = acosd(x)
                  end program test_acosd

Specific names:

Name  Argument  Return type Standard

          ACOSD(X)       REAL(4) X      REAL(4)  GNU extension
          DACOSD(X)      REAL(8) X      REAL(8)  GNU extension

See also: Inverse function:

Section 8.74 [COSD], page 160,

Radians function:

Section 8.6 [ACOS], page 114,

8.8 ACOSH — Inverse hyperbolic cosine function

Description:

ACOSH(X) computes the inverse hyperbolic cosine of X.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = ACOSH(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value has the same type and kind as X. If X is complex, the imaginary part of the result is in radians and lies between 0 ≤ ℑ acosh(x) ≤ π.

Example:

                  PROGRAM test_acosh
                    REAL(8), DIMENSION(3) :: x = (/ 1.0, 2.0, 3.0 /)
                    WRITE (*,*) ACOSH(x)
                  END PROGRAM

Specific names:

Name  Argument  Return type Standard

          DACOSH(X)      REAL(8) X      REAL(8)  GNU extension

See also: Inverse function:

Section 8.75 [COSH], page 161,

8.9 ADJUSTL — Left adjust a string

Description:

ADJUSTL(STRING) will left adjust a string by removing leading spaces. Spaces are inserted at the end of the string as needed.

Standard: Fortran 90 and later

Class: Elemental function

Syntax: RESULT = ADJUSTL(STRING)

Arguments:

STRING  The type shall be CHARACTER.

Return value:

The return value is of type CHARACTER and of the same kind as STRING where leading spaces are removed and the same number of spaces are inserted on the end of STRING.

Example:

                  program test_adjustl
                    character(len=20) :: str = '   gfortran'
                    str = adjustl(str)
                    print *, str
                  end program test_adjustl

See also: Section 8.10 [ADJUSTR], page 116, Section 8.275 [TRIM], page 286,

8.10 ADJUSTR — Right adjust a string

Description:

ADJUSTR(STRING) will right adjust a string by removing trailing spaces. Spaces are inserted at the start of the string as needed.

Standard: Fortran 90 and later

Class: Elemental function

Syntax: RESULT = ADJUSTR(STRING)

Arguments:

STR  The type shall be CHARACTER.

Return value:

The return value is of type CHARACTER and of the same kind as STRING where trailing spaces are removed and the same number of spaces are inserted at the start of STRING.

Example:

                  program test_adjustr
                    character(len=20) :: str = 'gfortran'
                    str = adjustr(str)
                    print *, str
                  end program test_adjustr

See also: Section 8.9 [ADJUSTL], page 116, Section 8.275 [TRIM], page 286,

8.11 AIMAG — Imaginary part of complex number

Description:

AIMAG(Z) yields the imaginary part of complex argument Z. The IMAG(Z) and IMAGPART(Z) intrinsic functions are provided for compatibility with g77, and their use in new code is strongly discouraged.

Standard: Fortran 77 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = AIMAG(Z)

Arguments:

Z  The type of the argument shall be COMPLEX.

Return value:

The return value is of type REAL with the kind type parameter of the argument.

Example:

                  program test_aimag
                    complex(4) z4
                    complex(8) z8
                    z4 = cmplx(1.e0_4, 0.e0_4)
                    z8 = cmplx(0.e0_8, 1.e0_8)
                    print *, aimag(z4), dimag(z8)
                  end program test_aimag

Specific names:

Name  Argument  Return type Standard

          AIMAG(Z)       COMPLEX Z      REAL  Fortran 77 and later
          DIMAG(Z)       COMPLEX(8) Z   REAL(8)  GNU extension
          IMAG(Z)        COMPLEX Z      REAL  GNU extension
          IMAGPART(Z)    COMPLEX Z      REAL  GNU extension

8.12 AINT — Truncate to a whole number

Description:

AINT(A [, KIND]) truncates its argument to a whole number.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = AINT(A [, KIND])

Arguments:

A  The type of the argument shall be REAL.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type REAL with the kind type parameter of the argument if the optional KIND is absent; otherwise, the kind type parameter will be given by KIND. If the magnitude of X is less than one, AINT(X) returns zero. If the magnitude is equal to or greater than one then it returns the largest whole number that does not exceed its magnitude. The sign is the same as the sign of X.

Example:

      program test_aint
        real(4) x4
        real(8) x8
        x4 = 1.234E0_4
        x8 = 4.321_8
        print *, aint(x4), dint(x8)
        x8 = aint(x4,8)
      end program test_aint

Specific names:

Name  Argument  Return type Standard

AINT(A)        REAL(4) A      REAL(4)  Fortran 77 and later
DINT(A)        REAL(8) A      REAL(8)  Fortran 77 and later

8.13 ALARM — Execute a routine after a given delay

Description:

ALARM(SECONDS, HANDLER [, STATUS]) causes external subroutine HANDLER

to be executed after a delay of SECONDS by using alarm(2) to set up a signal and signal(2) to catch it. If STATUS is supplied, it will be returned with the number of seconds remaining until any previously scheduled alarm was due to be delivered, or zero if there was no previously scheduled alarm.

Standard: GNU extension

Class: Subroutine

Syntax: CALL ALARM(SECONDS, HANDLER [, STATUS])

Arguments:

SECONDS The type of the argument shall be a scalar INTEGER.

            It is INTENT(IN).

HANDLER

            Signal handler (INTEGER FUNCTION or SUBROUTINE)

or dummy/global INTEGER scalar. The scalar values may be either SIG_IGN=1 to ignore the alarm generated or SIG_DFL=0 to set the default action. It is

            INTENT(IN).

STATUS (Optional) STATUS shall be a scalar variable of the

            default INTEGER kind. It is INTENT(OUT).

Example:

      program test_alarm
        external handler_print
        integer i
        call alarm (3, handler_print, i)
        print *, i
        call sleep(10)
      end program test_alarm

This will cause the external routine handler print to be called after 3 seconds.

8.14 ALL — All values in MASK along DIM are true

Description:

ALL(MASK [, DIM]) determines if all the values are true in MASK in the array along dimension DIM.

Standard: Fortran 90 and later

Class: Transformational function

Syntax: RESULT = ALL(MASK [, DIM])

Arguments:

MASK The type of the argument shall be LOGICAL and it shall not be scalar.

DIM (Optional) DIM shall be a scalar integer with a value that lies between one and the rank of MASK.

Return value:

ALL(MASK) returns a scalar value of type LOGICAL where the kind type parameter is the same as the kind type parameter of MASK. If DIM is present, then ALL(MASK, DIM) returns an array with the rank of MASK minus 1. The shape is determined from the shape of MASK where the DIM dimension is elided.

(A) ALL(MASK) is true if all elements of MASK are true. It also is true if MASK has zero size; otherwise, it is false.

(B) If the rank of MASK is one, then ALL(MASK,DIM) is equivalent to ALL(MASK). If the rank is greater than one, then ALL(MASK,DIM) is determined by applying ALL to the array sections.

Example:

      program test_all
        logical l
        l = all((/.true., .true., .true./))
        print *, l
        call section
        contains
          subroutine section
            integer a(2,3), b(2,3)
            a = 1
            b = 1
            b(2,2) = 2
            print *, all(a .eq. b, 1)
            print *, all(a .eq. b, 2)

          end subroutine section
      end program test_all

8.15 ALLOCATED — Status of an allocatable entity

Description:

ALLOCATED(ARRAY) and ALLOCATED(SCALAR) check the allocation status of AR-

RAY and SCALAR, respectively.

Standard: Fortran 90 and later. Note, the SCALAR= keyword and allocatable scalar entities are available in Fortran 2003 and later.

Class: Inquiry function

Syntax:

RESULT = ALLOCATED(ARRAY)
RESULT = ALLOCATED(SCALAR)

Arguments:

ARRAY  The argument shall be an ALLOCATABLE array.
SCALAR  The argument shall be an ALLOCATABLE scalar.

Return value:

The return value is a scalar LOGICAL with the default logical kind type parameter. If the argument is allocated, then the result is .TRUE.; otherwise, it returns

.FALSE.

Example:

      program test_allocated
        integer :: i = 4
        real(4), allocatable :: x(:)
        if (.not. allocated(x)) allocate(x(i))
      end program test_allocated

8.16 AND — Bitwise logical AND

Description:

Bitwise logical AND. This intrinsic routine is provided for backwards compatibility with GNU Fortran 77. For integer arguments, programmers should consider the use of the Section 8.137 [IAND], page 202, intrinsic defined by the Fortran standard.

Standard: GNU extension

Class: Function

Syntax: RESULT = AND(I, J)

Arguments:

I The type shall be either a scalar INTEGER type or a scalar LOGICAL type or a boz-literal-constant.

J The type shall be the same as the type of I or a bozliteral-constant. I and J shall not both be boz-literal-constants. If either I or J is a boz-literal-constant, then the other argument must be a scalar INTEGER.

Return value:

The return type is either a scalar INTEGER or a scalar LOGICAL. If the kind type parameters differ, then the smaller kind type is implicitly converted to larger kind, and the return has the larger kind. A boz-literal-constant is converted to an INTEGER with the kind type parameter of the other argument as-if a call to Section 8.149 [INT], page 210, occurred.

Example:

      PROGRAM test_and
        LOGICAL :: T = .TRUE., F = .FALSE.
        INTEGER :: a, b
        DATA a / Z'F' /, b / Z'3' /
        WRITE (*,*) AND(T, T), AND(T, F), AND(F, T), AND(F, F)
        WRITE (*,*) AND(a, b)
      END PROGRAM

See also: Fortran 95 elemental function:

Section 8.137 [IAND], page 202,

8.17 ANINT — Nearest whole number

Description:

ANINT(A [, KIND]) rounds its argument to the nearest whole number.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = ANINT(A [, KIND])

Arguments:

A  The type of the argument shall be REAL.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type real with the kind type parameter of the argument if the optional KIND is absent; otherwise, the kind type parameter will be given by KIND. If A is greater than zero, ANINT(A) returns AINT(X+0.5). If A is less than or equal to zero then it returns AINT(X-0.5).

Example:

      program test_anint
        real(4) x4
        real(8) x8
        x4 = 1.234E0_4
        x8 = 4.321_8
        print *, anint(x4), dnint(x8)
        x8 = anint(x4,8)
      end program test_anint

Specific names:

Name  Argument  Return type Standard

ANINT(A)       REAL(4) A      REAL(4)  Fortran 77 and later
DNINT(A)       REAL(8) A      REAL(8)  Fortran 77 and later

8.18 ANY — Any value in MASK along DIM is true

Description:

ANY(MASK [, DIM]) determines if any of the values in the logical array MASK along dimension DIM are .TRUE..

Standard: Fortran 90 and later

Class: Transformational function

Syntax: RESULT = ANY(MASK [, DIM])

Arguments:

MASK The type of the argument shall be LOGICAL and it shall not be scalar.

DIM (Optional) DIM shall be a scalar integer with a value that lies between one and the rank of MASK.

Return value:

ANY(MASK) returns a scalar value of type LOGICAL where the kind type parameter is the same as the kind type parameter of MASK. If DIM is present, then ANY(MASK, DIM) returns an array with the rank of MASK minus 1. The shape is determined from the shape of MASK where the DIM dimension is elided.

(A) ANY(MASK) is true if any element of MASK is true; otherwise, it is false. It also is false if MASK has zero size.

(B) If the rank of MASK is one, then ANY(MASK,DIM) is equivalent to ANY(MASK). If the rank is greater than one, then ANY(MASK,DIM) is determined by applying ANY to the array sections.

Example:

                  program test_any
                    logical l
                    l = any((/.true., .true., .true./))
                    print *, l
                    call section
                    contains
                      subroutine section
                        integer a(2,3), b(2,3)
                        a = 1
                        b = 1
                        b(2,2) = 2
                        print *, any(a .eq. b, 1)
                        print *, any(a .eq. b, 2)
                      end subroutine section
                  end program test_any

8.19 ASIN — Arcsine function

Description:

ASIN(X) computes the arcsine of its X (inverse of SIN(X)).

Standard: Fortran 77 and later, for a complex argument Fortran 2008 or later

Class: Elemental function

Syntax: RESULT = ASIN(X)

Arguments:

X The type shall be either REAL and a magnitude that is less than or equal to one - or be COMPLEX.

Return value:

The return value is of the same type and kind as X. The real part of the result is in radians and lies in the range −π/2 ≤ ℜ asin(x) ≤ π/2.

Example:

      program test_asin
        real(8) :: x = 0.866_8
        x = asin(x)
      end program test_asin

Specific names:

Name  Argument  Return type Standard

ASIN(X)        REAL(4) X      REAL(4)  Fortran 77 and later
DASIN(X)       REAL(8) X      REAL(8)  Fortran 77 and later

See also: Inverse function:

Section 8.249 [SIN], page 269, Degrees function: Section 8.20 [ASIND], page 123,

8.20 ASIND — Arcsine function, degrees

Description:

ASIND(X) computes the arcsine of its X in degrees (inverse of SIND(X)). This function is for compatibility only and should be avoided in favor of standard constructs wherever possible.

Standard: GNU extension, enabled with -fdec-math.

Class: Elemental function

Syntax: RESULT = ASIND(X)

Arguments:

X The type shall be either REAL and a magnitude that is less than or equal to one - or be COMPLEX.

Return value:

The return value is of the same type and kind as X. The real part of the result is in degrees and lies in the range −90 ≤ ℜ asin(x) ≤ 90.

Example:

      program test_asind
        real(8) :: x = 0.866_8
        x = asind(x)
      end program test_asind

Specific names:

Name  Argument  Return type Standard

ASIND(X)       REAL(4) X      REAL(4)  GNU extension

          DASIND(X)      REAL(8) X      REAL(8)  GNU extension

See also: Inverse function:

Section 8.250 [SIND], page 270,

Radians function:

Section 8.19 [ASIN], page 122,

8.21 ASINH — Inverse hyperbolic sine function

Description:

ASINH(X) computes the inverse hyperbolic sine of X.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = ASINH(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value is of the same type and kind as X. If X is complex, the imaginary part of the result is in radians and lies between −π/2 ≤ ℑ asinh(x) ≤ π/2.

Example:

                  PROGRAM test_asinh
                    REAL(8), DIMENSION(3) :: x = (/ -1.0, 0.0, 1.0 /)
                    WRITE (*,*) ASINH(x)
                  END PROGRAM

Specific names:

Name  Argument  Return type Standard

          DASINH(X)      REAL(8) X      REAL(8)  GNU extension.

See also: Inverse function:

Section 8.251 [SINH], page 271,

8.22 ASSOCIATED — Status of a pointer or pointer/target pair

Description:

ASSOCIATED(POINTER [, TARGET]) determines the status of the pointer POINTER or if POINTER is associated with the target TARGET.

Standard: Fortran 90 and later

Class: Inquiry function

Syntax: RESULT = ASSOCIATED(POINTER [, TARGET])

Arguments:

POINTER POINTER shall have the POINTER attribute and it can be of any type.

TARGET (Optional) TARGET shall be a pointer or a target. It must have the same type, kind type parameter, and array rank as POINTER.

The association status of neither POINTER nor TARGET shall be undefined.

Return value:

ASSOCIATED(POINTER) returns a scalar value of type LOGICAL(4). There are several cases:

(A) When the optional TARGET is not present then ASSOCIATED(POINTER) is true if POINTER is associated with a target; otherwise, it returns false.

(B) If TARGET is present and a scalar target, the result is true if TARGET is not a zero-sized storage sequence and the target associated with POINTER occupies the same storage units. If POINTER is disassociated, the result is false.

(C) If TARGET is present and an array target, the result is true if TARGET and POINTER have the same shape, are not zero-sized arrays, are arrays whose elements are not zero-sized storage sequences, and TARGET and POINTER occupy the same storage units in array element order. As in case(B), the result is false, if POINTER is disassociated.

(D) If TARGET is present and an scalar pointer, the result is true if TARGET is associated with POINTER, the target associated with TARGET are not zero-sized storage sequences and occupy the same storage units. The result is false, if either TARGET or POINTER is disassociated.

(E) If TARGET is present and an array pointer, the result is true if target associated with POINTER and the target associated with TARGET have the same shape, are not zero-sized arrays, are arrays whose elements are not zero-sized storage sequences, and TAR-GET and POINTER occupy the same storage units in array element order. The result is false, if either TARGET or POINTER is disassociated.

Example:

program test_associated
implicit none
real, target  :: tgt(2) = (/1., 2./)
real, pointer :: ptr(:)
ptr => tgt
if (associated(ptr)     .eqv. .false.) call abort
if (associated(ptr,tgt) .eqv. .false.) call abort
end program test_associated

See also: Section 8.209 [NULL], page 246,

8.23 ATAN — Arctangent function

Description:

ATAN(X) computes the arctangent of X.

Standard: Fortran 77 and later, for a complex argument and for two arguments Fortran 2008 or later

Class: Elemental function

Syntax:

RESULT = ATAN(X)
RESULT = ATAN(Y, X)

Arguments:

X The type shall be REAL or COMPLEX; if Y is present, X shall be REAL.

Y The type and kind type parameter shall be the same as X.

Return value:

The return value is of the same type and kind as X. If Y is present, the result is identical to ATAN2(Y,X). Otherwise, it the arcus tangent of X, where the real part of the result is in radians and lies in the range −π/2 ≤ ℜ atan(x) ≤ π/2.

Example:

      program test_atan
        real(8) :: x = 2.866_8
        x = atan(x)
      end program test_atan

Specific names:

Name  Argument  Return type Standard

ATAN(X)        REAL(4) X      REAL(4)  Fortran 77 and later
DATAN(X)       REAL(8) X      REAL(8)  Fortran 77 and later

See also: Inverse function:

Section 8.265 [TAN], page 280, Degrees function: Section 8.24 [ATAND], page 126,

8.24 ATAND — Arctangent function, degrees

Description:

ATAND(X) computes the arctangent of X in degrees (inverse of Section 8.266 [TAND], page 280). This function is for compatibility only and should be avoided in favor of standard constructs wherever possible.

Standard: GNU extension, enabled with -fdec-math.

Class: Elemental function Syntax:

RESULT = ATAND(X)
RESULT = ATAND(Y, X)

Arguments:

X The type shall be REAL or COMPLEX; if Y is present, X shall be REAL.

Y The type and kind type parameter shall be the same as X.

Return value:

The return value is of the same type and kind as X. If Y is present, the result is identical to ATAND2(Y,X). Otherwise, it is the arcus tangent of X, where the real part of the result is in degrees and lies in the range −90 ≤ ℜ atand(x) ≤ 90.

Example:

      program test_atand
        real(8) :: x = 2.866_8
        x = atand(x)
      end program test_atand

Specific names:

Name  Argument  Return type Standard

ATAND(X)         REAL(4) X      REAL(4)  GNU extension
DATAND(X)        REAL(8) X      REAL(8)  GNU extension

See also: Inverse function:

Section 8.266 [TAND], page 280, Radians function: Section 8.23 [ATAN], page 126,

8.25 ATAN2 — Arctangent function

Description:

ATAN2(Y, X) computes the principal value of the argument function of the complex number X+iY. This function can be used to transform from Cartesian into polar coordinates and allows to determine the angle in the correct quadrant.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = ATAN2(Y, X)

Arguments:

Y  The type shall be REAL.

X The type and kind type parameter shall be the same as Y. If Y is zero, then X must be nonzero.

Return value:

The return value has the same type and kind type parameter as Y. It is the principal value of the complex number X + iY. If X is nonzero, then it lies in the range −π ≤ atan(x) ≤ π. The sign is positive if Y is positive. If Y is zero, then the return value is zero if X is strictly positive, π if X is negative and Y is positive zero (or the processor does not handle signed zeros), and −π if X is negative and Y is negative zero. Finally, if X is zero, then the magnitude of the result is π/2.

Example:

      program test_atan2
        real(4) :: x = 1.e0_4, y = 0.5e0_4
        x = atan2(y,x)
      end program test_atan2

Specific names:

Name  Argument  Return type Standard

ATAN2(X, Y)      REAL(4) X, Y   REAL(4)  Fortran 77 and later
DATAN2(X, Y)     REAL(8) X, Y   REAL(8)  Fortran 77 and later

See also: Alias:

Section 8.23 [ATAN], page 126, Degrees function: Section 8.26 [ATAN2D], page 128,

8.26 ATAN2D — Arctangent function, degrees

Description:

ATAN2D(Y, X) computes the principal value of the argument function of the complex number X + iY in degrees. This function can be used to transform from Cartesian into polar coordinates and allows to determine the angle in the correct quadrant. This function is for compatibility only and should be avoided in favor of standard constructs wherever possible.

Standard: GNU extension, enabled with -fdec-math.

Class: Elemental function

Syntax: RESULT = ATAN2D(Y, X)

Arguments:

Y  The type shall be REAL.

X The type and kind type parameter shall be the same as Y. If Y is zero, then X must be nonzero.

Return value:

The return value has the same type and kind type parameter as Y. It is the principal value of the complex number X + iY. If X is nonzero, then it lies in the range −180 ≤ atan(x) ≤ 180. The sign is positive if Y is positive. If Y is zero, then the return value is zero if X is strictly positive, 180 if X is negative and Y is positive zero (or the processor does not handle signed zeros), and −180 if X is negative and Y is negative zero. Finally, if X is zero, then the magnitude of the result is 90.

Example:

      program test_atan2d

        real(4) :: x = 1.e0_4, y = 0.5e0_4
        x = atan2d(y,x)
      end program test_atan2d

Specific names:

Name  Argument  Return type Standard

ATAN2D(X, Y)     REAL(4) X, Y    REAL(4)  GNU extension
DATAN2D(X, Y)    REAL(8) X, Y    REAL(8)  GNU extension

See also: Alias:

Section 8.24 [ATAND], page 126, Radians function: Section 8.25 [ATAN2], page 127,

8.27 ATANH — Inverse hyperbolic tangent function

Description:

ATANH(X) computes the inverse hyperbolic tangent of X.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = ATANH(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value has same type and kind as X. If X is complex, the imaginary part of the result is in radians and lies between −π/2 ≤ ℑ atanh(x) ≤ π/2.

Example:

      PROGRAM test_atanh
        REAL, DIMENSION(3) :: x = (/ -1.0, 0.0, 1.0 /)
        WRITE (*,*) ATANH(x)
      END PROGRAM

Specific names:

Name  Argument  Return type Standard

DATANH(X)      REAL(8) X      REAL(8)  GNU extension

See also: Inverse function:

Section 8.267 [TANH], page 281,

8.28 ATOMIC_ADD — Atomic ADD operation

Description:

ATOMIC_ADD(ATOM, VALUE) atomically adds the value of VALUE to the variable ATOM. When STAT is present and the invocation was successful, it is assigned the value 0. If it is present and the invocation has failed, it is assigned a positive value; in particular, for a coindexed ATOM, if the remote image has stopped,

it is assigned the value of ISO_FORTRAN_ENV’s STAT_STOPPED_IMAGE and if the

remote image has failed, the value STAT_FAILED_IMAGE.

Standard: TS 18508 or later

Class: Atomic subroutine

Syntax: CALL ATOMIC_ADD (ATOM, VALUE [, STAT])

Arguments:

ATOM Scalar coarray or coindexed variable of integer type

            with ATOMIC_INT_KIND kind.

VALUE Scalar of the same type as ATOM. If the kind is different, the value is converted to the kind of ATOM.

STAT  (optional) Scalar default-kind integer variable.

Example:

      program atomic
        use iso_fortran_env
        integer(atomic_int_kind) :: atom[*]
        call atomic_add (atom[1], this_image())
      end program atomic

See also: Section 8.31 [ATOMIC DEFINE], page 132, Section 8.32 [ATOMIC FETCH ADD], page 132, Section 9.1 [ISO FORTRAN ENV], page 293, Section 8.29 [ATOMIC AND], page 130, Section 8.36 [ATOMIC OR], page 135, Section 8.38 [ATOMIC XOR], page 137,

8.29 ATOMIC_AND — Atomic bitwise AND operation

Description:

ATOMIC_AND(ATOM, VALUE) atomically defines ATOM with the bitwise AND between the values of ATOM and VALUE. When STAT is present and the invocation was successful, it is assigned the value 0. If it is present and the invocation has failed, it is assigned a positive value; in particular, for a coindexed ATOM, if the remote image has stopped, it is assigned the value of

ISO_FORTRAN_ENV’s STAT_STOPPED_IMAGE and if the remote image has failed,
the value STAT_FAILED_IMAGE.

Standard: TS 18508 or later

Class: Atomic subroutine

Syntax: CALL ATOMIC_AND (ATOM, VALUE [, STAT])

Arguments:

ATOM Scalar coarray or coindexed variable of integer type

            with ATOMIC_INT_KIND kind.

VALUE Scalar of the same type as ATOM. If the kind is different, the value is converted to the kind of ATOM.

STAT  (optional) Scalar default-kind integer variable.

Example:

      program atomic

        use iso_fortran_env
        integer(atomic_int_kind) :: atom[*]
        call atomic_and (atom[1], int(b'10100011101'))
      end program atomic

See also: Section 8.31 [ATOMIC DEFINE], page 132, Section 8.33 [ATOMIC FETCH AND], page 133, Section 9.1 [ISO FORTRAN ENV], page 293, Section 8.28 [ATOMIC ADD], page 129, Section 8.36 [ATOMIC OR], page 135, Section 8.38 [ATOMIC XOR], page 137,

8.30 ATOMIC_CAS — Atomic compare and swap

Description:

ATOMIC_CAS compares the variable ATOM with the value of COMPARE; if the value is the same, ATOM is set to the value of NEW. Additionally, OLD is set to the value of ATOM that was used for the comparison. When STAT is present and the invocation was successful, it is assigned the value 0. If it is present and the invocation has failed, it is assigned a positive value; in particular, for a coindexed ATOM, if the remote image has stopped, it is assigned the value of

ISO_FORTRAN_ENV’s STAT_STOPPED_IMAGE and if the remote image has failed,
the value STAT_FAILED_IMAGE.

Standard: TS 18508 or later

Class: Atomic subroutine

Syntax: CALL ATOMIC_CAS (ATOM, OLD, COMPARE, NEW [, STAT])

Arguments:

ATOM Scalar coarray or coindexed variable of either integer type with ATOMIC_INT_KIND kind or logical type with

            ATOMIC_LOGICAL_KIND kind.

OLD  Scalar of the same type and kind as ATOM.
COMPARE Scalar variable of the same type and kind as ATOM.

NEW Scalar variable of the same type as ATOM. If kind is different, the value is converted to the kind of ATOM.

STAT  (optional) Scalar default-kind integer variable.

Example:

      program atomic
        use iso_fortran_env
        logical(atomic_logical_kind) :: atom[*], prev
        call atomic_cas (atom[1], prev, .false., .true.))
      end program atomic

See also: Section 8.31 [ATOMIC DEFINE], page 132, Section 8.37 [ATOMIC REF], page 136, Section 9.1 [ISO FORTRAN ENV], page 293,

8.31 ATOMIC_DEFINE — Setting a variable atomically

Description:

ATOMIC_DEFINE(ATOM, VALUE) defines the variable ATOM with the value VALUE atomically. When STAT is present and the invocation was successful, it is assigned the value 0. If it is present and the invocation has failed, it is assigned a positive value; in particular, for a coindexed ATOM, if the remote image has stopped, it is assigned the value of ISO_FORTRAN_ENV’s STAT_STOPPED_IMAGE and if the remote image has failed, the value

STAT_FAILED_IMAGE.

Standard: Fortran 2008 and later; with STAT, TS 18508 or later

Class: Atomic subroutine

Syntax: CALL ATOMIC_DEFINE (ATOM, VALUE [, STAT])

Arguments:

ATOM Scalar coarray or coindexed variable of either integer type with ATOMIC_INT_KIND kind or logical type with

            ATOMIC_LOGICAL_KIND kind.

VALUE Scalar of the same type as ATOM. If the kind is different, the value is converted to the kind of ATOM.

STAT  (optional) Scalar default-kind integer variable.

Example:

      program atomic
        use iso_fortran_env
        integer(atomic_int_kind) :: atom[*]
        call atomic_define (atom[1], this_image())
      end program atomic

See also: Section 8.37 [ATOMIC REF], page 136, Section 8.30 [ATOMIC CAS], page 131, Section 9.1 [ISO FORTRAN ENV], page 293, Section 8.28 [ATOMIC ADD], page 129, Section 8.29 [ATOMIC AND], page 130, Section 8.36 [ATOMIC OR], page 135, Section 8.38 [ATOMIC XOR], page 137,

8.32 ATOMIC_FETCH_ADD — Atomic ADD operation with prior fetch

Description:

ATOMIC_FETCH_ADD(ATOM, VALUE, OLD) atomically stores the value of ATOM

in OLD and adds the value of VALUE to the variable ATOM. When STAT is present and the invocation was successful, it is assigned the value 0. If it is present and the invocation has failed, it is assigned a positive value; in particular, for a coindexed ATOM, if the remote image has stopped, it is assigned the

value of ISO_FORTRAN_ENV’s STAT_STOPPED_IMAGE and if the remote image has
failed, the value STAT_FAILED_IMAGE.

Standard: TS 18508 or later

Class: Atomic subroutine

Syntax: CALL ATOMIC_FETCH_ADD (ATOM, VALUE, old [, STAT])

Arguments:

ATOM Scalar coarray or coindexed variable of integer type

            with ATOMIC_INT_KIND kind. ATOMIC_LOGICAL_KIND

kind.

VALUE Scalar of the same type as ATOM. If the kind is different, the value is converted to the kind of ATOM.

OLD  Scalar of the same type and kind as ATOM.
STAT  (optional) Scalar default-kind integer variable.

Example:

      program atomic
        use iso_fortran_env
        integer(atomic_int_kind) :: atom[*], old
        call atomic_add (atom[1], this_image(), old)
      end program atomic

See also: Section 8.31 [ATOMIC DEFINE], page 132, Section 8.28 [ATOMIC ADD], page 129, Section 9.1 [ISO FORTRAN ENV], page 293, Section 8.33 [ATOMIC FETCH AND], page 133, Section 8.34 [ATOMIC FETCH OR], page 134, Section 8.35 [ATOMIC FETCH XOR], page 135,

8.33 ATOMIC_FETCH_AND — Atomic bitwise AND operation with prior fetch

Description:

ATOMIC_AND(ATOM, VALUE) atomically stores the value of ATOM in OLD and defines ATOM with the bitwise AND between the values of ATOM and VALUE. When STAT is present and the invocation was successful, it is assigned the value 0. If it is present and the invocation has failed, it is assigned a positive value; in particular, for a coindexed ATOM, if the remote image has

stopped, it is assigned the value of ISO_FORTRAN_ENV’s STAT_STOPPED_IMAGE

and if the remote image has failed, the value STAT_FAILED_IMAGE.

Standard: TS 18508 or later

Class: Atomic subroutine

Syntax: CALL ATOMIC_FETCH_AND (ATOM, VALUE, OLD [, STAT])

Arguments:

ATOM Scalar coarray or coindexed variable of integer type

            with ATOMIC_INT_KIND kind.

VALUE Scalar of the same type as ATOM. If the kind is different, the value is converted to the kind of ATOM.

OLD  Scalar of the same type and kind as ATOM.
STAT  (optional) Scalar default-kind integer variable.

Example:

      program atomic
        use iso_fortran_env
        integer(atomic_int_kind) :: atom[*], old
        call atomic_fetch_and (atom[1], int(b'10100011101'), old)
      end program atomic

See also: Section 8.31 [ATOMIC DEFINE], page 132, Section 8.29 [ATOMIC AND], page 130, Section 9.1 [ISO FORTRAN ENV], page 293, Section 8.32 [ATOMIC FETCH ADD], page 132, Section 8.34 [ATOMIC FETCH OR], page 134, Section 8.35 [ATOMIC FETCH XOR], page 135,

8.34 ATOMIC_FETCH_OR — Atomic bitwise OR operation with prior fetch

Description:

ATOMIC_OR(ATOM, VALUE) atomically stores the value of ATOM in OLD and defines ATOM with the bitwise OR between the values of ATOM and VALUE. When STAT is present and the invocation was successful, it is assigned the value 0. If it is present and the invocation has failed, it is assigned a positive value; in particular, for a coindexed ATOM, if the remote image has stopped,

it is assigned the value of ISO_FORTRAN_ENV’s STAT_STOPPED_IMAGE and if the

remote image has failed, the value STAT_FAILED_IMAGE.

Standard: TS 18508 or later

Class: Atomic subroutine

Syntax: CALL ATOMIC_FETCH_OR (ATOM, VALUE, OLD [, STAT])

Arguments:

ATOM Scalar coarray or coindexed variable of integer type

            with ATOMIC_INT_KIND kind.

VALUE Scalar of the same type as ATOM. If the kind is different, the value is converted to the kind of ATOM.

OLD  Scalar of the same type and kind as ATOM.
STAT  (optional) Scalar default-kind integer variable.

Example:

      program atomic
        use iso_fortran_env
        integer(atomic_int_kind) :: atom[*], old
        call atomic_fetch_or (atom[1], int(b'10100011101'), old)
      end program atomic

See also: Section 8.31 [ATOMIC DEFINE], page 132, Section 8.36 [ATOMIC OR], page 135, Section 9.1 [ISO FORTRAN ENV], page 293, Section 8.32 [ATOMIC FETCH ADD], page 132, Section 8.33 [ATOMIC FETCH AND], page 133, Section 8.35 [ATOMIC FETCH XOR], page 135,

8.35 ATOMIC_FETCH_XOR — Atomic bitwise XOR operation with prior fetch

Description:

ATOMIC_XOR(ATOM, VALUE) atomically stores the value of ATOM in OLD and defines ATOM with the bitwise XOR between the values of ATOM and VALUE. When STAT is present and the invocation was successful, it is assigned the value 0. If it is present and the invocation has failed, it is assigned a positive value; in particular, for a coindexed ATOM, if the remote image has stopped,

it is assigned the value of ISO_FORTRAN_ENV’s STAT_STOPPED_IMAGE and if the

remote image has failed, the value STAT_FAILED_IMAGE.

Standard: TS 18508 or later

Class: Atomic subroutine

Syntax: CALL ATOMIC_FETCH_XOR (ATOM, VALUE, OLD [, STAT])

Arguments:

ATOM Scalar coarray or coindexed variable of integer type

            with ATOMIC_INT_KIND kind.

VALUE Scalar of the same type as ATOM. If the kind is different, the value is converted to the kind of ATOM.

OLD  Scalar of the same type and kind as ATOM.
STAT  (optional) Scalar default-kind integer variable.

Example:

      program atomic
        use iso_fortran_env
        integer(atomic_int_kind) :: atom[*], old
        call atomic_fetch_xor (atom[1], int(b'10100011101'), old)
      end program atomic

See also: Section 8.31 [ATOMIC DEFINE], page 132, Section 8.38 [ATOMIC XOR], page 137, Section 9.1 [ISO FORTRAN ENV], page 293, Section 8.32 [ATOMIC FETCH ADD], page 132, Section 8.33 [ATOMIC FETCH AND], page 133, Section 8.34 [ATOMIC FETCH OR], page 134,

8.36 ATOMIC_OR — Atomic bitwise OR operation

Description:

ATOMIC_OR(ATOM, VALUE) atomically defines ATOM with the bitwise AND between the values of ATOM and VALUE. When STAT is present and the invocation was successful, it is assigned the value 0. If it is present and the invocation has failed, it is assigned a positive value; in particular, for a coindexed ATOM, if the remote image has stopped, it is assigned the value of

ISO_FORTRAN_ENV’s STAT_STOPPED_IMAGE and if the remote image has failed,
the value STAT_FAILED_IMAGE.

Standard: TS 18508 or later

Class: Atomic subroutine

Syntax: CALL ATOMIC_OR (ATOM, VALUE [, STAT])

Arguments:

ATOM Scalar coarray or coindexed variable of integer type

            with ATOMIC_INT_KIND kind.

VALUE Scalar of the same type as ATOM. If the kind is different, the value is converted to the kind of ATOM.

STAT  (optional) Scalar default-kind integer variable.

Example:

      program atomic
        use iso_fortran_env
        integer(atomic_int_kind) :: atom[*]
        call atomic_or (atom[1], int(b'10100011101'))
      end program atomic

See also: Section 8.31 [ATOMIC DEFINE], page 132, Section 8.34 [ATOMIC FETCH OR], page 134, Section 9.1 [ISO FORTRAN ENV], page 293, Section 8.28 [ATOMIC ADD], page 129, Section 8.36 [ATOMIC OR], page 135, Section 8.38 [ATOMIC XOR], page 137,

8.37 ATOMIC_REF — Obtaining the value of a variable atomically

Description:

ATOMIC_DEFINE(ATOM, VALUE) atomically assigns the value of the variable ATOM to VALUE. When STAT is present and the invocation was successful, it is assigned the value 0. If it is present and the invocation has failed, it is assigned a positive value; in particular, for a coindexed ATOM, if the remote image has stopped, it is assigned the value of ISO_FORTRAN_ENV’s STAT_STOPPED_IMAGE and if the remote image has failed, the value

STAT_FAILED_IMAGE.

Standard: Fortran 2008 and later; with STAT, TS 18508 or later

Class: Atomic subroutine

Syntax: CALL ATOMIC_REF(VALUE, ATOM [, STAT])

Arguments:

VALUE Scalar of the same type as ATOM. If the kind is different, the value is converted to the kind of ATOM.

ATOM Scalar coarray or coindexed variable of either integer type with ATOMIC_INT_KIND kind or logical type with

            ATOMIC_LOGICAL_KIND kind.

STAT  (optional) Scalar default-kind integer variable.

Example:

      program atomic
        use iso_fortran_env
        logical(atomic_logical_kind) :: atom[*]
        logical :: val
        call atomic_ref (atom, .false.)
        ! ...
        call atomic_ref (atom, val)
        if (val) then
          print *, "Obtained"
        end if
      end program atomic

See also: Section 8.31 [ATOMIC DEFINE], page 132, Section 8.30 [ATOMIC CAS], page 131, Section 9.1 [ISO FORTRAN ENV], page 293, Section 8.32 [ATOMIC FETCH ADD], page 132, Section 8.33 [ATOMIC FETCH AND], page 133, Section 8.34 [ATOMIC FETCH OR], page 134, Section 8.35 [ATOMIC FETCH XOR], page 135,

8.38 ATOMIC_XOR — Atomic bitwise OR operation

Description:

ATOMIC_AND(ATOM, VALUE) atomically defines ATOM with the bitwise XOR between the values of ATOM and VALUE. When STAT is present and the invocation was successful, it is assigned the value 0. If it is present and the invocation has failed, it is assigned a positive value; in particular, for a coindexed ATOM, if the remote image has stopped, it is assigned the value of

ISO_FORTRAN_ENV’s STAT_STOPPED_IMAGE and if the remote image has failed,
the value STAT_FAILED_IMAGE.

Standard: TS 18508 or later

Class: Atomic subroutine

Syntax: CALL ATOMIC_XOR (ATOM, VALUE [, STAT])

Arguments:

ATOM Scalar coarray or coindexed variable of integer type

            with ATOMIC_INT_KIND kind.

VALUE Scalar of the same type as ATOM. If the kind is different, the value is converted to the kind of ATOM.

STAT  (optional) Scalar default-kind integer variable.

Example:

      program atomic
        use iso_fortran_env
        integer(atomic_int_kind) :: atom[*]
        call atomic_xor (atom[1], int(b'10100011101'))
      end program atomic

See also: Section 8.31 [ATOMIC DEFINE], page 132, Section 8.35 [ATOMIC FETCH XOR], page 135, Section 9.1 [ISO FORTRAN ENV], page 293, Section 8.28 [ATOMIC ADD], page 129, Section 8.36 [ATOMIC OR], page 135, Section 8.38 [ATOMIC XOR], page 137,

8.39 BACKTRACE — Show a backtrace

Description:

BACKTRACE shows a backtrace at an arbitrary place in user code. Program execution continues normally afterwards. The backtrace information is printed to the unit corresponding to ERROR_UNIT in ISO_FORTRAN_ENV.

Standard: GNU extension

Class: Subroutine

Syntax: CALL BACKTRACE

Arguments:

None

See also: Section 8.2 [ABORT], page 111,

8.40 BESSEL_J0 — Bessel function of the first kind of order 0

Description:

BESSEL_J0(X) computes the Bessel function of the first kind of order 0 of X. This function is available under the name BESJ0 as a GNU extension.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = BESSEL_J0(X)

Arguments:

X  The type shall be REAL.

Return value:

The return value is of type REAL and lies in the range −0.4027... ≤ Bessel(0, x) ≤ 1. It has the same kind as X.

Example:

                  program test_besj0
                    real(8) :: x = 0.0_8
                    x = bessel_j0(x)
                  end program test_besj0

Specific names:

Name  Argument  Return type Standard

          DBESJ0(X)       REAL(8) X      REAL(8)  GNU extension

8.41 BESSEL_J1 — Bessel function of the first kind of order 1

Description:

BESSEL_J1(X) computes the Bessel function of the first kind of order 1 of X. This function is available under the name BESJ1 as a GNU extension.

Standard: Fortran 2008

Class: Elemental function

Syntax: RESULT = BESSEL_J1(X)

Arguments:

X  The type shall be REAL.

Return value:

The return value is of type REAL and lies in the range −0.5818... ≤ Bessel(0, x) ≤ 0.5818. It has the same kind as X.

Example:

      program test_besj1
        real(8) :: x = 1.0_8
        x = bessel_j1(x)
      end program test_besj1

Specific names:

Name  Argument  Return type Standard

DBESJ1(X)      REAL(8) X      REAL(8)  GNU extension

8.42 BESSEL_JN — Bessel function of the first kind

Description:

BESSEL_JN(N, X) computes the Bessel function of the first kind of order N of X. This function is available under the name BESJN as a GNU extension. If N and X are arrays, their ranks and shapes shall conform. BESSEL_JN(N1, N2, X) returns an array with the Bessel functions of the first kind of the orders N1 to N2.

Standard: Fortran 2008 and later, negative N is allowed as GNU extension

Class: Elemental function, except for the transformational function BESSEL_JN(N1,

N2, X)

Syntax:

RESULT = BESSEL_JN(N, X)
RESULT = BESSEL_JN(N1, N2, X)

Arguments:

N  Shall be a scalar or an array of type INTEGER.
N1  Shall be a non-negative scalar of type INTEGER.
N2  Shall be a non-negative scalar of type INTEGER.

X Shall be a scalar or an array of type REAL; for BESSEL_ JN(N1, N2, X) it shall be scalar.

Return value:

The return value is a scalar of type REAL. It has the same kind as X.

Note: The transformational function uses a recurrence algorithm which might, for

some values of X, lead to different results than calls to the elemental function.

Example:

                  program test_besjn
                    real(8) :: x = 1.0_8
                    x = bessel_jn(5,x)
                  end program test_besjn

Specific names:

Name  Argument  Return type Standard

          DBESJN(N, X)     INTEGER N      REAL(8)  GNU extension

                           REAL(8) X

8.43 BESSEL_Y0 — Bessel function of the second kind of

order 0

Description:

BESSEL_Y0(X) computes the Bessel function of the second kind of order 0 of X. This function is available under the name BESY0 as a GNU extension.

Standard: Fortran 2008 and later Class: Elemental function

Syntax: RESULT = BESSEL_Y0(X)

Arguments:

X  The type shall be REAL.

Return value:

The return value is of type REAL. It has the same kind as X.

Example:

                  program test_besy0
                    real(8) :: x = 0.0_8
                    x = bessel_y0(x)
                  end program test_besy0

Specific names:

Name  Argument  Return type Standard

          DBESY0(X)      REAL(8) X      REAL(8)  GNU extension

8.44 BESSEL_Y1 — Bessel function of the second kind of

order 1

Description:

BESSEL_Y1(X) computes the Bessel function of the second kind of order 1 of X. This function is available under the name BESY1 as a GNU extension.

Standard: Fortran 2008 and later Class: Elemental function

Syntax: RESULT = BESSEL_Y1(X)

Arguments:

X  The type shall be REAL.

Return value:

The return value is of type REAL. It has the same kind as X.

Example:

      program test_besy1
        real(8) :: x = 1.0_8
        x = bessel_y1(x)
      end program test_besy1

Specific names:

Name  Argument  Return type Standard

DBESY1(X)      REAL(8) X      REAL(8)  GNU extension

8.45 BESSEL_YN — Bessel function of the second kind

Description:

BESSEL_YN(N, X) computes the Bessel function of the second kind of order N of X. This function is available under the name BESYN as a GNU extension. If N and X are arrays, their ranks and shapes shall conform. BESSEL_YN(N1, N2, X) returns an array with the Bessel functions of the first kind of the orders N1 to N2.

Standard: Fortran 2008 and later, negative N is allowed as GNU extension Class: Elemental function, except for the transformational function BESSEL_YN(N1,

N2, X)

Syntax:

RESULT = BESSEL_YN(N, X)
RESULT = BESSEL_YN(N1, N2, X)

Arguments:

N  Shall be a scalar or an array of type INTEGER .
N1  Shall be a non-negative scalar of type INTEGER.
N2  Shall be a non-negative scalar of type INTEGER.

X Shall be a scalar or an array of type REAL; for BESSEL_ YN(N1, N2, X) it shall be scalar.

Return value:

The return value is a scalar of type REAL. It has the same kind as X.

Note: The transformational function uses a recurrence algorithm which might, for

some values of X, lead to different results than calls to the elemental function.

Example:

      program test_besyn
        real(8) :: x = 1.0_8
        x = bessel_yn(5,x)
      end program test_besyn

Specific names:

Name  Argument  Return type Standard

          DBESYN(N,X)    INTEGER N      REAL(8)  GNU extension

                         REAL(8) X

8.46 BGE — Bitwise greater than or equal to

Description:

Determines whether an integral is a bitwise greater than or equal to another.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = BGE(I, J)

Arguments:

I  Shall be of INTEGER type.
J  Shall be of INTEGER type, and of the same kind as I.

Return value:

The return value is of type LOGICAL and of the default kind.

See also: Section 8.47 [BGT], page 142, Section 8.49 [BLE], page 143, Section 8.50 [BLT], page 143,

8.47 BGT — Bitwise greater than

Description:

Determines whether an integral is a bitwise greater than another.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = BGT(I, J)

Arguments:

I  Shall be of INTEGER type.
J  Shall be of INTEGER type, and of the same kind as I.

Return value:

The return value is of type LOGICAL and of the default kind.

See also: Section 8.46 [BGE], page 142, Section 8.49 [BLE], page 143, Section 8.50 [BLT], page 143,

8.48 BIT_SIZE — Bit size inquiry function

Description:

BIT_SIZE(I) returns the number of bits (integer precision plus sign bit) represented by the type of I. The result of BIT_SIZE(I) is independent of the actual value of I.

Standard: Fortran 90 and later Class: Inquiry function

Syntax: RESULT = BIT_SIZE(I)

Arguments:

I  The type shall be INTEGER.

Return value:

The return value is of type INTEGER

Example:

                  program test_bit_size
                      integer :: i = 123
                      integer :: size
                      size = bit_size(i)
                      print *, size
                  end program test_bit_size

8.49 BLE — Bitwise less than or equal to

Description:

Determines whether an integral is a bitwise less than or equal to another.

Standard: Fortran 2008 and later Class: Elemental function

Syntax: RESULT = BLE(I, J)

Arguments:

I  Shall be of INTEGER type.
J  Shall be of INTEGER type, and of the same kind as I.

Return value:

The return value is of type LOGICAL and of the default kind.

See also: Section 8.47 [BGT], page 142, Section 8.46 [BGE], page 142, Section 8.50 [BLT], page 143,

8.50 BLT — Bitwise less than

Description:

Determines whether an integral is a bitwise less than another.

Standard: Fortran 2008 and later Class: Elemental function

Syntax: RESULT = BLT(I, J)

Arguments:

I  Shall be of INTEGER type.
J  Shall be of INTEGER type, and of the same kind as I.

Return value:

The return value is of type LOGICAL and of the default kind.

See also: Section 8.46 [BGE], page 142, Section 8.47 [BGT], page 142, Section 8.49 [BLE], page 143,

8.51 BTEST — Bit test function

Description:

BTEST(I,POS) returns logical .TRUE. if the bit at POS in I is set. The counting of the bits starts at 0.

Standard: Fortran 90 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = BTEST(I, POS)

Arguments:

I  The type shall be INTEGER.
POS  The type shall be INTEGER.

Return value:

The return value is of type LOGICAL

Example:

      program test_btest
          integer :: i = 32768 + 1024 + 64
          integer :: pos
          logical :: bool
          do pos=0,16
              bool = btest(i, pos)
              print *, pos, bool
          end do
      end program test_btest

Specific names:

Name  Argument  Return type  Standard

BTEST(I,POS)    INTEGER I,POS      LOGICAL  Fortran 95 and later
BBTEST(I,POS)   INTEGER(1) I,POS   LOGICAL(1) GNU extension
BITEST(I,POS)   INTEGER(2) I,POS   LOGICAL(2) GNU extension
BJTEST(I,POS)   INTEGER(4) I,POS   LOGICAL(4) GNU extension
BKTEST(I,POS)   INTEGER(8) I,POS   LOGICAL(8) GNU extension

8.52 C_ASSOCIATED — Status of a C pointer

Description:

C_ASSOCIATED(c_ptr_1[, c_ptr_2]) determines the status of the C pointer c ptr 1 or if c ptr 1 is associated with the target c ptr 2.

Standard: Fortran 2003 and later Class: Inquiry function

Syntax: RESULT = C_ASSOCIATED(c_ptr_1[, c_ptr_2])

Arguments:

c ptr 1  Scalar of the type C_PTR or C_FUNPTR.
c ptr 2  (Optional) Scalar of the same type as c ptr 1.

Return value:

The return value is of type LOGICAL; it is .false. if either c ptr 1 is a C NULL pointer or if c ptr1 and c ptr 2 point to different addresses.

Example:

                  subroutine association_test(a,b)
                    use iso_c_binding, only: c_associated, c_loc, c_ptr
                    implicit none
                    real, pointer :: a
                    type(c_ptr) :: b
                    if(c_associated(b, c_loc(a))) &
                       stop 'b and a do not point to same target'
                  end subroutine association_test

See also: Section 8.56 [C LOC], page 147, Section 8.55 [C FUNLOC], page 146,

8.53 C_F_POINTER — Convert C into Fortran pointer

Description:

C_F_POINTER(CPTR, FPTR[, SHAPE]) assigns the target of the C pointer CPTR to the Fortran pointer FPTR and specifies its shape.

Standard: Fortran 2003 and later

Class: Subroutine

Syntax: CALL C_F_POINTER(CPTR, FPTR[, SHAPE])

Arguments:

CPTR  scalar of the type C_PTR. It is INTENT(IN).
FPTR  pointer interoperable with cptr. It is INTENT(OUT).

SHAPE (Optional) Rank-one array of type INTEGER with INTENT(IN). It shall be present if and only if fptr is an array. The size must be equal to the rank of fptr.

Example:

                  program main
                    use iso_c_binding
                    implicit none
                    interface
                      subroutine my_routine(p) bind(c,name='myC_func')
                        import :: c_ptr
                        type(c_ptr), intent(out) :: p
                      end subroutine
                    end interface
                    type(c_ptr) :: cptr

                    real,pointer :: a(:)
                    call my_routine(cptr)
                    call c_f_pointer(cptr, a, [12])
                  end program main

See also: Section 8.56 [C LOC], page 147, Section 8.54 [C F PROCPOINTER], page 146,

8.54 C_F_PROCPOINTER — Convert C into Fortran procedure pointer

Description:

C_F_PROCPOINTER(CPTR, FPTR) Assign the target of the C function pointer CPTR to the Fortran procedure pointer FPTR.

Standard: Fortran 2003 and later

Class: Subroutine

Syntax: CALL C_F_PROCPOINTER(cptr, fptr)

Arguments:

CPTR  scalar of the type C_FUNPTR. It is INTENT(IN).

FPTR procedure pointer interoperable with cptr. It is

                      INTENT(OUT).

Example:

                  program main
                    use iso_c_binding
                    implicit none
                    abstract interface
                      function func(a)
                        import :: c_float
                        real(c_float), intent(in) :: a
                        real(c_float) :: func
                      end function
                    end interface
                    interface
                       function getIterFunc() bind(c,name="getIterFunc")
                         import :: c_funptr
                         type(c_funptr) :: getIterFunc
                       end function
                    end interface
                    type(c_funptr) :: cfunptr
                    procedure(func), pointer :: myFunc
                    cfunptr = getIterFunc()
                    call c_f_procpointer(cfunptr, myFunc)
                  end program main

See also: Section 8.56 [C LOC], page 147, Section 8.53 [C F POINTER], page 145,

8.55 C_FUNLOC — Obtain the C address of a procedure

Description:

C_FUNLOC(x) determines the C address of the argument.

Standard: Fortran 2003 and later

Class: Inquiry function

Syntax: RESULT = C_FUNLOC(x)

Arguments:

x  Interoperable function or pointer to such function.

Return value:

The return value is of type C_FUNPTR and contains the C address of the argument.

Example:

                  module x
                    use iso_c_binding
                    implicit none
                  contains
                    subroutine sub(a) bind(c)
                      real(c_float) :: a
                      a = sqrt(a)+5.0
                    end subroutine sub
                  end module x
                  program main
                    use iso_c_binding
                    use x
                    implicit none
                    interface
                      subroutine my_routine(p) bind(c,name='myC_func')
                        import :: c_funptr
                        type(c_funptr), intent(in) :: p
                      end subroutine
                    end interface
                    call my_routine(c_funloc(sub))
                  end program main

See also: Section 8.52 [C ASSOCIATED], page 144, Section 8.56 [C LOC], page 147, Section 8.53 [C F POINTER], page 145, Section 8.54 [C F PROCPOINTER], page 146,

8.56 C_LOC — Obtain the C address of an object

Description:

C_LOC(X) determines the C address of the argument.

Standard: Fortran 2003 and later

Class: Inquiry function

Syntax: RESULT = C_LOC(X)

Arguments:

X Shall have either the POINTER or TARGET attribute. It shall not be a coindexed object. It shall either be a variable with interoperable type and kind type parameters, or be a scalar, nonpolymorphic variable with no length type parameters.

Return value:

The return value is of type C_PTR and contains the C address of the argument.

Example:

      subroutine association_test(a,b)
        use iso_c_binding, only: c_associated, c_loc, c_ptr
        implicit none
        real, pointer :: a
        type(c_ptr) :: b
        if(c_associated(b, c_loc(a))) &
           stop 'b and a do not point to same target'
      end subroutine association_test

See also: Section 8.52 [C ASSOCIATED], page 144, Section 8.55 [C FUNLOC], page 146, Section 8.53 [C F POINTER], page 145, Section 8.54 [C F PROCPOINTER], page 146,

8.57 C_SIZEOF — Size in bytes of an expression

Description:

C_SIZEOF(X) calculates the number of bytes of storage the expression X occupies.

Standard: Fortran 2008

Class: Inquiry function of the module ISO_C_BINDING

Syntax: N = C_SIZEOF(X)

Arguments:

X  The argument shall be an interoperable data entity.

Return value:

The return value is of type integer and of the system-dependent kind C_SIZE_T (from the ISO_C_BINDING module). Its value is the number of bytes occupied by the argument. If the argument has the POINTER attribute, the number of bytes of the storage area pointed to is returned. If the argument is of a derived type with POINTER or ALLOCATABLE components, the return value does not account for the sizes of the data pointed to by these components.

Example:

         use iso_c_binding
         integer(c_int) :: i
         real(c_float) :: r, s(5)
         print *, (c_sizeof(s)/c_sizeof(r) == 5)

                     end

The example will print T unless you are using a platform where default REAL variables are unusually padded.

See also: Section 8.253 [SIZEOF], page 272, Section 8.260 [STORAGE SIZE], page 276,

8.58 CEILING — Integer ceiling function

Description:

CEILING(A) returns the least integer greater than or equal to A.

Standard: Fortran 95 and later

Class: Elemental function

Syntax: RESULT = CEILING(A [, KIND])

Arguments:

A  The type shall be REAL.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER(KIND) if KIND is present and a defaultkind INTEGER otherwise.

Example:

                  program test_ceiling
                      real :: x = 63.29
                      real :: y = -63.59
                      print *, ceiling(x) ! returns 64
                      print *, ceiling(y) ! returns -63
                  end program test_ceiling

See also: Section 8.109 [FLOOR], page 184, Section 8.206 [NINT], page 244,

8.59 CHAR — Character conversion function

Description:

CHAR(I [, KIND]) returns the character represented by the integer I.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = CHAR(I [, KIND])

Arguments:

I  The type shall be INTEGER.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type CHARACTER(1)

Example:

                  program test_char
                      integer :: i = 74
                      character(1) :: c
                      c = char(i)
                      print *, i, c ! returns 'J'
                  end program test_char

Specific names:

Name  Argument  Return type  Standard

          CHAR(I)        INTEGER I    CHARACTER(LEN=1) Fortran 77 and later

Note: See Section 8.143 [ICHAR], page 206, for a discussion of converting between numerical values and formatted string representations.

See also: Section 8.5 [ACHAR], page 113, Section 8.135 [IACHAR], page 201, Section 8.143 [ICHAR], page 206,

8.60 CHDIR — Change working directory

Description:

Change current working directory to a specified path.

This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

          CALL CHDIR(NAME [, STATUS])
          STATUS = CHDIR(NAME)

Arguments:

NAME The type shall be CHARACTER of default kind and shall specify a valid path within the file system.

STATUS (Optional) INTEGER status flag of the default kind. Returns 0 on success, and a system specific and nonzero error code otherwise.

Example:

                  PROGRAM test_chdir
                    CHARACTER(len=255) :: path
                    CALL getcwd(path)
                    WRITE(*,*) TRIM(path)
                    CALL chdir("/tmp")
                    CALL getcwd(path)
                    WRITE(*,*) TRIM(path)
                  END PROGRAM

See also: Section 8.124 [GETCWD], page 195,

8.61 CHMOD — Change access permissions of files

Description:

CHMOD changes the permissions of a file.

This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL CHMOD(NAME, MODE[, STATUS])
STATUS = CHMOD(NAME, MODE)

Arguments:

NAME Scalar CHARACTER of default kind with the file name. Trailing blanks are ignored unless the character achar(0) is present, then all characters up to and excluding achar(0) are used as the file name.

MODE Scalar CHARACTER of default kind giving the file permission. MODE uses the same syntax as the chmod utility as defined by the POSIX standard. The argument shall either be a string of a nonnegative octal number or a symbolic mode.

STATUS (optional) scalar INTEGER, which is 0 on success and nonzero otherwise.

Return value:

In either syntax, STATUS is set to 0 on success and nonzero otherwise.

Example: CHMOD as subroutine

program chmod_test
implicit none
integer :: status
call chmod('test.dat','u+x',status)
print *, 'Status: ', status
end program chmod_test

CHMOD as function:

program chmod_test
implicit none
integer :: status
status = chmod('test.dat','u+x')
print *, 'Status: ', status
end program chmod_test

8.62 CMPLX — Complex conversion function

Description:

CMPLX(X [, Y [, KIND]]) returns a complex number where X is converted to the real component. If Y is present it is converted to the imaginary component. If Y is not present then the imaginary component is set to 0.0. If X is complex then Y must not be present.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = CMPLX(X [, Y [, KIND]])

Arguments:

X  The type may be INTEGER, REAL, or COMPLEX.

Y (Optional; only allowed if X is not COMPLEX.) May

            be INTEGER or REAL.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of COMPLEX type, with a kind equal to KIND if it is specified. If KIND is not specified, the result is of the default COMPLEX kind, regardless of the kinds of X and Y.

Example:

      program test_cmplx
          integer :: i = 42
          real :: x = 3.14
          complex :: z
          z = cmplx(i, x)
          print *, z, cmplx(x)
      end program test_cmplx

See also: Section 8.71 [COMPLEX], page 158,

8.63 CO_BROADCAST — Copy a value to all images the current set of images

Description:

CO_BROADCAST copies the value of argument A on the image with image index SOURCE_IMAGE to all images in the current team. A becomes defined as if by intrinsic assignment. If the execution was successful and STAT is present, it is assigned the value zero. If the execution failed, STAT gets assigned a nonzero value and, if present, ERRMSG gets assigned a value describing the occurred error.

Standard: Technical Specification (TS) 18508 or later

Class: Collective subroutine

Syntax: CALL CO_BROADCAST(A, SOURCE_IMAGE [, STAT, ERRMSG])

Arguments:

A INTENT(INOUT) argument; shall have the same dynamic type and type parameters on all images of the current team. If it is an array, it shall have the same shape on all images.

SOURCE IMAGE a scalar integer expression. It shall have the same value on all images and refer to an image of the current team.

STAT  (optional) a scalar integer variable
ERRMSG  (optional) a scalar character variable

Example:

      program test
        integer :: val(3)
        if (this_image() == 1) then
          val = [1, 5, 3]
        end if
        call co_broadcast (val, source_image=1)
        print *, this_image, ":", val
      end program test

See also: Section 8.64 [CO MAX], page 153, Section 8.65 [CO MIN], page 154, Section 8.67 [CO SUM], page 156, Section 8.66 [CO REDUCE], page 155,

8.64 CO_MAX — Maximal value on the current set of images

Description:

CO_MAX determines element-wise the maximal value of A on all images of the current team. If RESULT IMAGE is present, the maximum values are returned in A on the specified image only and the value of A on the other images become undefined. If RESULT IMAGE is not present, the value is returned on all images. If the execution was successful and STAT is present, it is assigned the value zero. If the execution failed, STAT gets assigned a nonzero value and, if present, ERRMSG gets assigned a value describing the occurred error.

Standard: Technical Specification (TS) 18508 or later

Class: Collective subroutine

Syntax: CALL CO_MAX(A [, RESULT_IMAGE, STAT, ERRMSG])

Arguments:

A shall be an integer, real or character variable, which has the same type and type parameters on all images of the team.

RESULT IMAGE (optional) a scalar integer expression; if present, it shall have the same value on all images and refer to an image of the current team.

STAT  (optional) a scalar integer variable
ERRMSG  (optional) a scalar character variable

Example:

      program test
        integer :: val
        val = this_image ()
        call co_max (val, result_image=1)
        if (this_image() == 1) then
          write(*,*) "Maximal value", val  ! prints num_images()
        end if
      end program test

See also: Section 8.65 [CO MIN], page 154, Section 8.67 [CO SUM], page 156, Section 8.66 [CO REDUCE], page 155, Section 8.63 [CO BROADCAST], page 152,

8.65 CO_MIN — Minimal value on the current set of images

Description:

CO_MIN determines element-wise the minimal value of A on all images of the current team. If RESULT IMAGE is present, the minimal values are returned in A on the specified image only and the value of A on the other images become undefined. If RESULT IMAGE is not present, the value is returned on all images. If the execution was successful and STAT is present, it is assigned the value zero. If the execution failed, STAT gets assigned a nonzero value and, if present, ERRMSG gets assigned a value describing the occurred error.

Standard: Technical Specification (TS) 18508 or later

Class: Collective subroutine

Syntax: CALL CO_MIN(A [, RESULT_IMAGE, STAT, ERRMSG])

Arguments:

A shall be an integer, real or character variable, which has the same type and type parameters on all images of the team.

RESULT IMAGE (optional) a scalar integer expression; if present, it shall have the same value on all images and refer to an image of the current team.

STAT  (optional) a scalar integer variable
ERRMSG  (optional) a scalar character variable

Example:

      program test
        integer :: val
        val = this_image ()
        call co_min (val, result_image=1)
        if (this_image() == 1) then
          write(*,*) "Minimal value", val  ! prints 1
        end if
      end program test

See also: Section 8.64 [CO MAX], page 153, Section 8.67 [CO SUM], page 156, Section 8.66 [CO REDUCE], page 155, Section 8.63 [CO BROADCAST], page 152,

8.66 CO_REDUCE — Reduction of values on the current set of images

Description:

CO_REDUCE determines element-wise the reduction of the value of A on all images of the current team. The pure function passed as OPERATION is used to pairwise reduce the values of A by passing either the value of A of different images or the result values of such a reduction as argument. If A is an array, the deduction is done element wise. If RESULT IMAGE is present, the result values are returned in A on the specified image only and the value of A on the other images become undefined. If RESULT IMAGE is not present, the value is returned on all images. If the execution was successful and STAT is present, it is assigned the value zero. If the execution failed, STAT gets assigned a nonzero value and, if present, ERRMSG gets assigned a value describing the occurred error.

Standard: Technical Specification (TS) 18508 or later

Class: Collective subroutine

Syntax: CALL CO_REDUCE(A, OPERATION, [, RESULT_IMAGE, STAT, ERRMSG])

Arguments:

A is an INTENT(INOUT) argument and shall be nonpolymorphic. If it is allocatable, it shall be allocated; if it is a pointer, it shall be associated. A shall have the same type and type parameters on all images of the team; if it is an array, it shall have the same shape on all images.

OPERATION pure function with two scalar nonallocatable arguments, which shall be nonpolymorphic and have the same type and type parameters as A. The function shall return a nonallocatable scalar of the same type and type parameters as A. The function shall be the same on all images and with regards to the arguments mathematically commutative and associative. Note that OPERA-TION may not be an elemental function, unless it is an intrisic function.

RESULT IMAGE (optional) a scalar integer expression; if present, it shall have the same value on all images and refer to an image of the current team.

STAT  (optional) a scalar integer variable
ERRMSG  (optional) a scalar character variable

Example:

program test

        integer :: val
        val = this_image ()
        call co_reduce (val, result_image=1, operation=myprod)
        if (this_image() == 1) then
          write(*,*) "Product value", val  ! prints num_images() factorial
        end if
      contains
        pure function myprod(a, b)
          integer, value :: a, b
          integer :: myprod
          myprod = a * b
        end function myprod
      end program test

Note: While the rules permit in principle an intrinsic function, none of the intrinsics in the standard fulfill the criteria of having a specific function, which takes two arguments of the same type and returning that type as result.

See also: Section 8.65 [CO MIN], page 154, Section 8.64 [CO MAX], page 153, Section 8.67 [CO SUM], page 156, Section 8.63 [CO BROADCAST], page 152,

8.67 CO_SUM — Sum of values on the current set of images

Description:

CO_SUM sums up the values of each element of A on all images of the current team. If RESULT IMAGE is present, the summed-up values are returned in A on the specified image only and the value of A on the other images become undefined. If RESULT IMAGE is not present, the value is returned on all images. If the execution was successful and STAT is present, it is assigned the value zero. If the execution failed, STAT gets assigned a nonzero value and, if present, ERRMSG gets assigned a value describing the occurred error.

Standard: Technical Specification (TS) 18508 or later

Class: Collective subroutine

Syntax: CALL CO_SUM(A [, RESULT_IMAGE, STAT, ERRMSG])

Arguments:

A shall be an integer, real or complex variable, which has the same type and type parameters on all images of the team.

RESULT IMAGE (optional) a scalar integer expression; if present, it shall have the same value on all images and refer to an image of the current team.

STAT  (optional) a scalar integer variable
ERRMSG  (optional) a scalar character variable

Example:

      program test
        integer :: val
        val = this_image ()

                    call co_sum (val, result_image=1)
                    if (this_image() == 1) then
                      write(*,*) "The sum is ", val ! prints (n**2 + n)/2,
                                                    ! with n = num_images()
                    end if
                  end program test

See also: Section 8.64 [CO MAX], page 153, Section 8.65 [CO MIN], page 154, Section 8.66 [CO REDUCE], page 155, Section 8.63 [CO BROADCAST], page 152,

8.68 COMMAND_ARGUMENT_COUNT — Get number of command

line arguments

Description:

COMMAND_ARGUMENT_COUNT returns the number of arguments passed on the command line when the containing program was invoked.

Standard: Fortran 2003 and later Class: Inquiry function

Syntax: RESULT = COMMAND_ARGUMENT_COUNT()

Arguments:

None

Return value:

The return value is an INTEGER of default kind.

Example:

                  program test_command_argument_count
                      integer :: count
                      count = command_argument_count()
                      print *, count
                  end program test_command_argument_count

See also: Section 8.122 [GET COMMAND], page 193, Section 8.123 [GET COMMAND ARGUMENT], page 194,

8.69 COMPILER_OPTIONS — Options passed to the compiler

Description:

COMPILER_OPTIONS returns a string with the options used for compiling.

Standard: Fortran 2008 Class: Inquiry function of the module ISO_FORTRAN_ENV

Syntax: STR = COMPILER_OPTIONS()

Arguments:

None Return value:

The return value is a default-kind string with system-dependent length. It contains the compiler flags used to compile the file, which called the COMPILER_ OPTIONS intrinsic.

Example:

                     use iso_fortran_env
                     print '(4a)', 'This file was compiled by ', &
                                   compiler_version(), ' using the options ', &
                                   compiler_options()
                     end

See also: Section 8.70 [COMPILER VERSION], page 158, Section 9.1 [ISO FORTRAN ENV], page 293,

8.70 COMPILER_VERSION — Compiler version string

Description:

COMPILER_VERSION returns a string with the name and the version of the compiler.

Standard: Fortran 2008

Class: Inquiry function of the module ISO_FORTRAN_ENV

Syntax: STR = COMPILER_VERSION()

Arguments:

None

Return value:

The return value is a default-kind string with system-dependent length. It contains the name of the compiler and its version number.

Example:

                     use iso_fortran_env
                     print '(4a)', 'This file was compiled by ', &
                                   compiler_version(), ' using the options ', &
                                   compiler_options()
                     end

See also: Section 8.69 [COMPILER OPTIONS], page 157, Section 9.1 [ISO FORTRAN ENV], page 293,

8.71 COMPLEX — Complex conversion function

Description:

COMPLEX(X, Y) returns a complex number where X is converted to the real component and Y is converted to the imaginary component.

Standard: GNU extension

Class: Elemental function

Syntax: RESULT = COMPLEX(X, Y)

Arguments:

X  The type may be INTEGER or REAL.
Y  The type may be INTEGER or REAL.

Return value:

If X and Y are both of INTEGER type, then the return value is of default COMPLEX type.

If X and Y are of REAL type, or one is of REAL type and one is of INTEGER type, then the return value is of COMPLEX type with a kind equal to that of the REAL argument with the highest precision.

Example:

                  program test_complex
                      integer :: i = 42
                      real :: x = 3.14
                      print *, complex(i, x)
                  end program test_complex

See also: Section 8.62 [CMPLX], page 152,

8.72 CONJG — Complex conjugate function

Description:

CONJG(Z) returns the conjugate of Z. If Z is (x, y) then the result is (x, -y)

Standard: Fortran 77 and later, has an overload that is a GNU extension

Class: Elemental function

Syntax: Z = CONJG(Z)

Arguments:

Z  The type shall be COMPLEX.

Return value:

The return value is of type COMPLEX.

Example:

                  program test_conjg
                      complex :: z = (2.0, 3.0)
                      complex(8) :: dz = (2.71_8, -3.14_8)
                      z= conjg(z)
                      print *, z
                      dz = dconjg(dz)
                      print *, dz
                  end program test_conjg

Specific names:

Name  Argument  Return type Standard

          DCONJG(Z)      COMPLEX(8) Z   COMPLEX(8) GNU extension

8.73 COS — Cosine function

Description:

COS(X) computes the cosine of X.

Standard: Fortran 77 and later, has overloads that are GNU extensions Class: Elemental function

Syntax: RESULT = COS(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value is of the same type and kind as X. The real part of the result is in radians. If X is of the type REAL, the return value lies in the range −1 ≤ cos(x) ≤ 1.

Example:

      program test_cos
        real :: x = 0.0
        x = cos(x)
      end program test_cos

Specific names:

Name  Argument  Return type Standard

COS(X)         REAL(4) X      REAL(4)  Fortran 77 and later
DCOS(X)        REAL(8) X      REAL(8)  Fortran 77 and later
CCOS(X)        COMPLEX(4) X   COMPLEX(4) Fortran 77 and later
ZCOS(X)        COMPLEX(8) X   COMPLEX(8) GNU extension
CDCOS(X)       COMPLEX(8) X   COMPLEX(8) GNU extension

See also: Inverse function:

Section 8.6 [ACOS], page 114, Degrees function: Section 8.74 [COSD], page 160,

8.74 COSD — Cosine function, degrees

Description:

COSD(X) computes the cosine of X in degrees. This function is for compatibility only and should be avoided in favor of standard constructs wherever possible.

Standard: GNU extension, enabled with -fdec-math. Class: Elemental function

Syntax: RESULT = COSD(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value is of the same type and kind as X. The real part of the result is in degrees. If X is of the type REAL, the return value lies in the range −1 ≤ cosd(x) ≤ 1.

Example:

      program test_cosd
        real :: x = 0.0
        x = cosd(x)
      end program test_cosd

Specific names:

Name  Argument  Return type Standard

COSD(X)        REAL(4) X      REAL(4)  GNU extension
DCOSD(X)       REAL(8) X      REAL(8)  GNU extension
CCOSD(X)       COMPLEX(4) X   COMPLEX(4) GNU extension
ZCOSD(X)       COMPLEX(8) X   COMPLEX(8) GNU extension
CDCOSD(X)      COMPLEX(8) X   COMPLEX(8) GNU extension

See also: Inverse function:

Section 8.7 [ACOSD], page 114, Radians function: Section 8.73 [COS], page 160,

8.75 COSH — Hyperbolic cosine function

Description:

COSH(X) computes the hyperbolic cosine of X.

Standard: Fortran 77 and later, for a complex argument Fortran 2008 or later

Class: Elemental function

Syntax: X = COSH(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value has same type and kind as X. If X is complex, the imaginary part of the result is in radians. If X is REAL, the return value has a lower bound of one, cosh(x) ≥ 1.

Example:

      program test_cosh
        real(8) :: x = 1.0_8
        x = cosh(x)
      end program test_cosh

Specific names:

Name  Argument  Return type Standard

COSH(X)        REAL(4) X      REAL(4)  Fortran 77 and later
DCOSH(X)       REAL(8) X      REAL(8)  Fortran 77 and later

See also: Inverse function:

Section 8.8 [ACOSH], page 115,

8.76 COTAN — Cotangent function

Description:

COTAN(X) computes the cotangent of X. Equivalent to COS(x) divided by

SIN(x), or 1 / TAN(x).

This function is for compatibility only and should be avoided in favor of standard constructs wherever possible.

Standard: GNU extension, enabled with -fdec-math.

Class: Elemental function

Syntax: RESULT = COTAN(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value has same type and kind as X, and its value is in radians.

Example:

      program test_cotan
        real(8) :: x = 0.165_8
        x = cotan(x)
      end program test_cotan

Specific names:

Name  Argument  Return type Standard

COTAN(X)       REAL(4) X      REAL(4)  GNU extension
DCOTAN(X)      REAL(8) X      REAL(8)  GNU extension

See also: Converse function:

Section 8.265 [TAN], page 280, Degrees function: Section 8.77 [COTAND], page 162,

8.77 COTAND — Cotangent function, degrees

Description:

COTAND(X) computes the cotangent of X in degrees. Equivalent to COSD(x)

divided by SIND(x), or 1 / TAND(x).

Standard: GNU extension, enabled with -fdec-math.

This function is for compatibility only and should be avoided in favor of standard constructs wherever possible.

Class: Elemental function

Syntax: RESULT = COTAND(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value has same type and kind as X, and its value is in degrees.

Example:

      program test_cotand
        real(8) :: x = 0.165_8
        x = cotand(x)
      end program test_cotand

Specific names:

Name  Argument  Return type Standard

COTAND(X)      REAL(4) X      REAL(4)  GNU extension
DCOTAND(X)     REAL(8) X      REAL(8)  GNU extension

See also: Converse function:

Section 8.266 [TAND], page 280, Radians function: Section 8.76 [COTAN], page 162,

8.78 COUNT — Count function

Description:

Counts the number of .TRUE. elements in a logical MASK, or, if the DIM argument is supplied, counts the number of elements along each row of the array in the DIM direction. If the array has zero size, or all of the elements of MASK are .FALSE., then the result is 0.

Standard: Fortran 90 and later, with KIND argument Fortran 2003 and later

Class: Transformational function

Syntax: RESULT = COUNT(MASK [, DIM, KIND])

Arguments:

MASK  The type shall be LOGICAL.
DIM  (Optional) The type shall be INTEGER.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind. If DIM is present, the result is an array with a rank one less than the rank of ARRAY, and a size corresponding to the shape of ARRAY with the DIM dimension removed.

Example:

      program test_count
          integer, dimension(2,3) :: a, b
          logical, dimension(2,3) :: mask
          a = reshape( (/ 1, 2, 3, 4, 5, 6 /), (/ 2, 3 /))
          b = reshape( (/ 0, 7, 3, 4, 5, 8 /), (/ 2, 3 /))
          print '(3i3)', a(1,:)
          print '(3i3)', a(2,:)
          print *
          print '(3i3)', b(1,:)
          print '(3i3)', b(2,:)
          print *

          mask = a.ne.b
          print '(3l3)', mask(1,:)
          print '(3l3)', mask(2,:)
          print *
          print '(3i3)', count(mask)
          print *
          print '(3i3)', count(mask, 1)
          print *
          print '(3i3)', count(mask, 2)
      end program test_count

8.79 CPU_TIME — CPU elapsed time in seconds

Description:

Returns a REAL value representing the elapsed CPU time in seconds. This is useful for testing segments of code to determine execution time. If a time source is available, time will be reported with microsecond resolution. If no time source is available, TIME is set to -1.0. Note that TIME may contain a, system dependent, arbitrary offset and may not start with 0.0. For CPU_TIME, the absolute value is meaningless, only differences between subsequent calls to this subroutine, as shown in the example below, should be used.

Standard: Fortran 95 and later

Class: Subroutine

Syntax: CALL CPU_TIME(TIME)

Arguments:

TIME  The type shall be REAL with INTENT(OUT).

Return value:

None

Example:

      program test_cpu_time
          real :: start, finish
          call cpu_time(start)
              ! put code to test here
          call cpu_time(finish)
          print '("Time = ",f6.3," seconds.")',finish-start
      end program test_cpu_time

See also: Section 8.264 [SYSTEM CLOCK], page 279, Section 8.82 [DATE AND TIME], page 166,

8.80 CSHIFT — Circular shift elements of an array

Description:

CSHIFT(ARRAY, SHIFT [, DIM]) performs a circular shift on elements of AR-RAY along the dimension of DIM. If DIM is omitted it is taken to be 1. DIM is a scalar of type INTEGER in the range of 1 ≤ DIM ≤ n) where n is the rank of ARRAY. If the rank of ARRAY is one, then all elements of ARRAY are shifted by SHIFT places. If rank is greater than one, then all complete rank one sections of ARRAY along the given dimension are shifted. Elements shifted out one end of each rank one section are shifted back in the other end.

Standard: Fortran 90 and later

Class: Transformational function

Syntax: RESULT = CSHIFT(ARRAY, SHIFT [, DIM])

Arguments:

ARRAY  Shall be an array of any type.
SHIFT  The type shall be INTEGER.
DIM  The type shall be INTEGER.

Return value:

Returns an array of same type and rank as the ARRAY argument.

Example:

      program test_cshift
          integer, dimension(3,3) :: a
          a = reshape( (/ 1, 2, 3, 4, 5, 6, 7, 8, 9 /), (/ 3, 3 /))
          print '(3i3)', a(1,:)
          print '(3i3)', a(2,:)
          print '(3i3)', a(3,:)
          a = cshift(a, SHIFT=(/1, 2, -1/), DIM=2)
          print *
          print '(3i3)', a(1,:)
          print '(3i3)', a(2,:)
          print '(3i3)', a(3,:)
      end program test_cshift

8.81 CTIME — Convert a time into a string

Description:

CTIME converts a system time value, such as returned by Section 8.270 [TIME8], page 283, to a string. The output will be of the form ‘Sat Aug 19 18:13:14

1995’.

This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL CTIME(TIME, RESULT).
RESULT = CTIME(TIME).

Arguments:

TIME  The type shall be of type INTEGER.

RESULT The type shall be of type CHARACTER and of default kind. It is an INTENT(OUT) argument. If the length of this variable is too short for the time and date string to fit completely, it will be blank on procedure return.

Return value:

The converted date and time as a string.

Example:

      program test_ctime
          integer(8) :: i
          character(len=30) :: date
          i = time8()
          ! Do something, main part of the program
          call ctime(i,date)
          print *, 'Program was started on ', date
      end program test_ctime

See Also: Section 8.82 [DATE AND TIME], page 166, Section 8.131 [GMTIME], page 198, Section 8.183 [LTIME], page 230, Section 8.269 [TIME], page 282, Section 8.270 [TIME8], page 283,

8.82 DATE_AND_TIME — Date and time subroutine

Description:

DATE_AND_TIME(DATE, TIME, ZONE, VALUES) gets the corresponding date and

time information from the real-time system clock. DATE is INTENT(OUT) and has form ccyymmdd. TIME is INTENT(OUT) and has form hhmmss.sss. ZONE is INTENT(OUT) and has form (+-)hhmm, representing the difference with respect to Coordinated Universal Time (UTC). Unavailable time and date parameters return blanks. VALUES is INTENT(OUT) and provides the following:

VALUE(1): The year

VALUE(2): The month
VALUE(3): The day of the month
VALUE(4): Time difference with UTC in minutes
VALUE(5): The hour of the day
VALUE(6): The minutes of the hour
VALUE(7): The seconds of the minute
VALUE(8): The milliseconds of the second

Standard: Fortran 90 and later

Class: Subroutine

Syntax: CALL DATE_AND_TIME([DATE, TIME, ZONE, VALUES])

Arguments:

DATE (Optional) The type shall be CHARACTER(LEN=8) or larger, and of default kind.

TIME (Optional) The type shall be CHARACTER(LEN=10) or larger, and of default kind.

ZONE (Optional) The type shall be CHARACTER(LEN=5) or larger, and of default kind.

VALUES  (Optional) The type shall be INTEGER(8).

Return value:

None

Example:

                  program test_time_and_date
                      character(8)  :: date
                      character(10) :: time
                      character(5)  :: zone
                      integer,dimension(8) :: values
                      ! using keyword arguments
                      call date_and_time(date,time,zone,values)
                      call date_and_time(DATE=date,ZONE=zone)
                      call date_and_time(TIME=time)
                      call date_and_time(VALUES=values)
                      print '(a,2x,a,2x,a)', date, time, zone
                      print '(8i5)', values
                  end program test_time_and_date

See also: Section 8.79 [CPU TIME], page 164, Section 8.264 [SYSTEM CLOCK], page 279,

8.83 DBLE — Double conversion function

Description:

DBLE(A) Converts A to double precision real type.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = DBLE(A)

Arguments:

A  The type shall be INTEGER, REAL, or COMPLEX.

Return value:

The return value is of type double precision real.

Example:

                  program test_dble
                      real    :: x = 2.18
                      integer :: i = 5
                      complex :: z = (2.3,1.14)
                      print *, dble(x), dble(i), dble(z)
                  end program test_dble

See also: Section 8.228 [REAL], page 257,

8.84 DCMPLX — Double complex conversion function

Description:

DCMPLX(X [,Y]) returns a double complex number where X is converted to the real component. If Y is present it is converted to the imaginary component. If Y is not present then the imaginary component is set to 0.0. If X is complex then Y must not be present.

Standard: GNU extension

Class: Elemental function

Syntax: RESULT = DCMPLX(X [, Y])

Arguments:

          X  The type may be INTEGER, REAL, or COMPLEX.

Y (Optional if X is not COMPLEX.) May be INTEGER or

                      REAL.

Return value:

The return value is of type COMPLEX(8)

Example:

                  program test_dcmplx
                      integer :: i = 42
                      real :: x = 3.14
                      complex :: z
                      z = cmplx(i, x)
                      print *, dcmplx(i)
                      print *, dcmplx(x)
                      print *, dcmplx(z)
                      print *, dcmplx(x,i)
                  end program test_dcmplx

8.85 DIGITS — Significant binary digits function

Description:

DIGITS(X) returns the number of significant binary digits of the internal model representation of X. For example, on a system using a 32-bit floating point representation, a default real number would likely return 24.

Standard: Fortran 90 and later

Class: Inquiry function

Syntax: RESULT = DIGITS(X)

Arguments:

X  The type may be INTEGER or REAL.

Return value:

The return value is of type INTEGER.

Example:

                  program test_digits
                      integer :: i = 12345
                      real :: x = 3.143
                      real(8) :: y = 2.33
                      print *, digits(i)
                      print *, digits(x)
                      print *, digits(y)
                  end program test_digits

8.86 DIM — Positive difference

Description:

DIM(X,Y) returns the difference X-Y if the result is positive; otherwise returns zero.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = DIM(X, Y)

Arguments:

X  The type shall be INTEGER or REAL

Y The type shall be the same type and kind as X. (As a GNU extension, arguments of different kinds are permitted.)

Return value:

The return value is of type INTEGER or REAL. (As a GNU extension, kind is the largest kind of the actual arguments.)

Example:

      program test_dim
          integer :: i
          real(8) :: x
          i = dim(4, 15)
          x = dim(4.345_8, 2.111_8)
          print *, i
          print *, x
      end program test_dim

Specific names:

Name  Argument  Return type Standard

DIM(X,Y)       REAL(4) X, Y      REAL(4)  Fortran 77 and later
IDIM(X,Y)      INTEGER(4) X, Y   INTEGER(4) Fortran 77 and later
DDIM(X,Y)      REAL(8) X, Y      REAL(8)  Fortran 77 and later

8.87 DOT_PRODUCT — Dot product function

Description:

DOT_PRODUCT(VECTOR_A, VECTOR_B) computes the dot product multiplication of two vectors VECTOR A and VECTOR B. The two vectors may be either numeric or logical and must be arrays of rank one and of equal size. If the

vectors are INTEGER or REAL, the result is SUM(VECTOR_A*VECTOR_B). If the
vectors are COMPLEX, the result is SUM(CONJG(VECTOR_A)*VECTOR_B). If the
vectors are LOGICAL, the result is ANY(VECTOR_A .AND. VECTOR_B).

Standard: Fortran 90 and later

Class: Transformational function

Syntax: RESULT = DOT_PRODUCT(VECTOR_A, VECTOR_B)

Arguments:

VECTOR A The type shall be numeric or LOGICAL, rank 1.

VECTOR B The type shall be numeric if VECTOR A is of numeric type or LOGICAL if VECTOR A is of type LOGICAL. VECTOR B shall be a rank-one array.

Return value:

If the arguments are numeric, the return value is a scalar of numeric type, INTEGER, REAL, or COMPLEX. If the arguments are LOGICAL, the return value is

          .TRUE. or .FALSE..

Example:

                  program test_dot_prod
                      integer, dimension(3) :: a, b
                      a = (/ 1, 2, 3 /)
                      b = (/ 4, 5, 6 /)
                      print '(3i3)', a
                      print *
                      print '(3i3)', b
                      print *
                      print *, dot_product(a,b)
                  end program test_dot_prod

8.88 DPROD — Double product function

Description:

DPROD(X,Y) returns the product X*Y.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = DPROD(X, Y)

Arguments:

X  The type shall be REAL.
Y  The type shall be REAL.

Return value:

The return value is of type REAL(8).

Example:

                  program test_dprod
                      real :: x = 5.2
                      real :: y = 2.3
                      real(8) :: d
                      d = dprod(x,y)
                      print *, d
                  end program test_dprod

Specific names:

Name  Argument  Return type Standard

          DPROD(X,Y)     REAL(4) X, Y    REAL(8)  Fortran 77 and later

8.89 DREAL — Double real part function

Description:

DREAL(Z) returns the real part of complex variable Z.

Standard: GNU extension

Class: Elemental function

Syntax: RESULT = DREAL(A)

Arguments:

A  The type shall be COMPLEX(8).

Return value:

The return value is of type REAL(8).

Example:

                  program test_dreal
                      complex(8) :: z = (1.3_8,7.2_8)
                      print *, dreal(z)
                  end program test_dreal

See also: Section 8.11 [AIMAG], page 117,

8.90 DSHIFTL — Combined left shift

Description:

DSHIFTL(I, J, SHIFT) combines bits of I and J. The rightmost SHIFT bits of the result are the leftmost SHIFT bits of J, and the remaining bits are the rightmost bits of I.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = DSHIFTL(I, J, SHIFT)

Arguments:

I  Shall be of type INTEGER or a BOZ constant.

J Shall be of type INTEGER or a BOZ constant. If both I and J have integer type, then they shall have the same kind type parameter. I and J shall not both be BOZ constants.

SHIFT Shall be of type INTEGER. It shall be nonnegative. If I is not a BOZ constant, then SHIFT shall be less than or equal to BIT_SIZE(I); otherwise, SHIFT shall be less than or equal to BIT_SIZE(J).

Return value:

If either I or J is a BOZ constant, it is first converted as if by the intrinsic function INT to an integer type with the kind type parameter of the other.

See also: Section 8.91 [DSHIFTR], page 172,

8.91 DSHIFTR — Combined right shift

Description:

DSHIFTR(I, J, SHIFT) combines bits of I and J. The leftmost SHIFT bits of the result are the rightmost SHIFT bits of I, and the remaining bits are the leftmost bits of J.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = DSHIFTR(I, J, SHIFT)

Arguments:

I  Shall be of type INTEGER or a BOZ constant.

J Shall be of type INTEGER or a BOZ constant. If both I and J have integer type, then they shall have the same kind type parameter. I and J shall not both be BOZ constants.

SHIFT Shall be of type INTEGER. It shall be nonnegative. If I is not a BOZ constant, then SHIFT shall be less than or equal to BIT_SIZE(I); otherwise, SHIFT shall be less than or equal to BIT_SIZE(J).

Return value:

If either I or J is a BOZ constant, it is first converted as if by the intrinsic function INT to an integer type with the kind type parameter of the other.

See also: Section 8.90 [DSHIFTL], page 171,

8.92 DTIME — Execution time subroutine (or function)

Description:

DTIME(VALUES, TIME) initially returns the number of seconds of runtime since the start of the process’s execution in TIME. VALUES returns the user and system components of this time in VALUES(1) and VALUES(2) respectively. TIME

is equal to VALUES(1) + VALUES(2).

Subsequent invocations of DTIME return values accumulated since the previous invocation. On some systems, the underlying timings are represented using types with sufficiently small limits that overflows (wrap around) are possible, such as 32bit types. Therefore, the values returned by this intrinsic might be, or become, negative, or numerically less than previous values, during a single run of the compiled program. Please note, that this implementation is thread safe if used within OpenMP directives, i.e., its state will be consistent while called from multiple threads. However, if DTIME is called from multiple threads, the result is still the time since the last invocation. This may not give the intended results. If possible, use CPU_TIME instead.

This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit. VALUES and TIME are INTENT(OUT) and provide the following:

VALUES(1): User time in seconds.
VALUES(2): System time in seconds.
TIME:  Run time since start in seconds.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL DTIME(VALUES, TIME).
TIME = DTIME(VALUES), (not recommended).

Arguments:

VALUES  The type shall be REAL(4), DIMENSION(2).
TIME  The type shall be REAL(4).

Return value:

Elapsed time in seconds since the last invocation or since the start of program execution if not called before.

Example:

      program test_dtime
          integer(8) :: i, j
          real, dimension(2) :: tarray
          real :: result
          call dtime(tarray, result)
          print *, result
          print *, tarray(1)
          print *, tarray(2)
          do i=1,100000000    ! Just a delay
              j = i * i - i
          end do
          call dtime(tarray, result)
          print *, result
          print *, tarray(1)
          print *, tarray(2)
      end program test_dtime

See also: Section 8.79 [CPU TIME], page 164,

8.93 EOSHIFT — End-off shift elements of an array

Description:

EOSHIFT(ARRAY, SHIFT[, BOUNDARY, DIM]) performs an end-off shift on ele-

ments of ARRAY along the dimension of DIM. If DIM is omitted it is taken to be 1. DIM is a scalar of type INTEGER in the range of 1 ≤ DIM ≤ n) where n is the rank of ARRAY. If the rank of ARRAY is one, then all elements of ARRAY are shifted by SHIFT places. If rank is greater than one, then all complete rank one sections of ARRAY along the given dimension are shifted. Elements shifted out one end of each rank one section are dropped. If BOUND-ARY is present then the corresponding value of from BOUNDARY is copied back in the other end. If BOUNDARY is not present then the following are copied in depending on the type of ARRAY.

Array Type Boundary Value
Numeric  0 of the type and kind of ARRAY.
Logical  .FALSE..

Character(len)

len blanks.

Standard: Fortran 90 and later

Class: Transformational function

Syntax: RESULT = EOSHIFT(ARRAY, SHIFT [, BOUNDARY, DIM])

Arguments:

ARRAY  May be any type, not scalar.
SHIFT  The type shall be INTEGER.
BOUNDARY Same type as ARRAY.
DIM  The type shall be INTEGER.

Return value:

Returns an array of same type and rank as the ARRAY argument.

Example:

      program test_eoshift
          integer, dimension(3,3) :: a
          a = reshape( (/ 1, 2, 3, 4, 5, 6, 7, 8, 9 /), (/ 3, 3 /))
          print '(3i3)', a(1,:)
          print '(3i3)', a(2,:)
          print '(3i3)', a(3,:)
          a = EOSHIFT(a, SHIFT=(/1, 2, 1/), BOUNDARY=-5, DIM=2)
          print *
          print '(3i3)', a(1,:)
          print '(3i3)', a(2,:)
          print '(3i3)', a(3,:)
      end program test_eoshift

8.94 EPSILON — Epsilon function

Description:

EPSILON(X) returns the smallest number E of the same kind as X such that 1 + E > 1.

Standard: Fortran 90 and later

Class: Inquiry function

Syntax: RESULT = EPSILON(X)

Arguments:

X  The type shall be REAL.

Return value:

The return value is of same type as the argument.

Example:

      program test_epsilon

                      real :: x = 3.143
                      real(8) :: y = 2.33
                      print *, EPSILON(x)
                      print *, EPSILON(y)
                  end program test_epsilon

8.95 ERF — Error function

Description:

ERF(X) computes the error function of X.

Standard: Fortran 2008 and later Class: Elemental function

Syntax: RESULT = ERF(X)

Arguments:

X  The type shall be REAL.

Return value:

The return value is of type REAL, of the same kind as X and lies in the range −1 ≤ erf(x) ≤ 1.

Example:

                  program test_erf
                    real(8) :: x = 0.17_8
                    x = erf(x)
                  end program test_erf

Specific names:

Name  Argument  Return type Standard

          DERF(X)        REAL(8) X      REAL(8)  GNU extension

8.96 ERFC — Error function

Description:

ERFC(X) computes the complementary error function of X.

Standard: Fortran 2008 and later Class: Elemental function

Syntax: RESULT = ERFC(X)

Arguments:

X  The type shall be REAL.

Return value:

The return value is of type REAL and of the same kind as X. It lies in the range 0 ≤ erfc(x) ≤ 2.

Example:

                  program test_erfc
                    real(8) :: x = 0.17_8
                    x = erfc(x)
                  end program test_erfc

Specific names:

Name  Argument  Return type Standard

DERFC(X)       REAL(8) X      REAL(8)  GNU extension

8.97 ERFC_SCALED — Error function

Description:

ERFC_SCALED(X) computes the exponentially-scaled complementary error function of X.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = ERFC_SCALED(X)

Arguments:

X  The type shall be REAL.

Return value:

The return value is of type REAL and of the same kind as X.

Example:

      program test_erfc_scaled
        real(8) :: x = 0.17_8
        x = erfc_scaled(x)
      end program test_erfc_scaled

8.98 ETIME — Execution time subroutine (or function)

Description:

ETIME(VALUES, TIME) returns the number of seconds of runtime since the start of the process’s execution in TIME. VALUES returns the user and system components of this time in VALUES(1) and VALUES(2) respectively. TIME is

equal to VALUES(1) + VALUES(2).

On some systems, the underlying timings are represented using types with sufficiently small limits that overflows (wrap around) are possible, such as 32bit types. Therefore, the values returned by this intrinsic might be, or become, negative, or numerically less than previous values, during a single run of the compiled program. This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit. VALUES and TIME are INTENT(OUT) and provide the following:

VALUES(1): User time in seconds.
VALUES(2): System time in seconds.
TIME:  Run time since start in seconds.

Standard: GNU extension

Class: Subroutine, function Syntax:

CALL ETIME(VALUES, TIME).
TIME = ETIME(VALUES), (not recommended).

Arguments:

VALUES  The type shall be REAL(4), DIMENSION(2).
TIME  The type shall be REAL(4).

Return value:

Elapsed time in seconds since the start of program execution.

Example:

      program test_etime
          integer(8) :: i, j
          real, dimension(2) :: tarray
          real :: result
          call ETIME(tarray, result)
          print *, result
          print *, tarray(1)
          print *, tarray(2)
          do i=1,100000000    ! Just a delay
              j = i * i - i
          end do
          call ETIME(tarray, result)
          print *, result
          print *, tarray(1)
          print *, tarray(2)
      end program test_etime

See also: Section 8.79 [CPU TIME], page 164,

8.99 EVENT_QUERY — Query whether a coarray event has occurred

Description:

EVENT_QUERY assignes the number of events to COUNT which have been posted to the EVENT variable and not yet been removed by calling EVENT WAIT. When STAT is present and the invocation was successful, it is assigned the value 0. If it is present and the invocation has failed, it is assigned a positive value and COUNT is assigned the value −1.

Standard: TS 18508 or later

Class: subroutine

Syntax: CALL EVENT_QUERY (EVENT, COUNT [, STAT])

Arguments:

EVENT (intent(IN)) Scalar of type EVENT_TYPE, defined in ISO_FORTRAN_ENV; shall not be coindexed.

COUNT (intent(out))Scalar integer with at least the precision of default integer.

STAT  (optional) Scalar default-kind integer variable.

Example:

      program atomic
        use iso_fortran_env
        implicit none
        type(event_type) :: event_value_has_been_set[*]
        integer :: cnt
        if (this_image() == 1) then
          call event_query (event_value_has_been_set, cnt)
          if (cnt > 0) write(*,*) "Value has been set"
        elseif (this_image() == 2) then
          event post (event_value_has_been_set[1])
        end if
      end program atomic

8.100 EXECUTE_COMMAND_LINE — Execute a shell command

Description:

EXECUTE_COMMAND_LINE runs a shell command, synchronously or asynchronously. The COMMAND argument is passed to the shell and executed (The shell is sh on Unix systems, and cmd.exe on Windows.). If WAIT is present and has the value false, the execution of the command is asynchronous if the system supports it; otherwise, the command is executed synchronously using the C library’s system call. The three last arguments allow the user to get status information. After synchronous execution, EXITSTAT contains the integer exit code of the command, as returned by system. CMDSTAT is set to zero if the command line was executed (whatever its exit status was). CMDMSG is assigned an error message if an error has occurred. Note that the system function need not be thread-safe. It is the responsibility of the user to ensure that system is not called concurrently. For asynchronous execution on supported targets, the POSIX posix_spawn or fork functions are used. Also, a signal handler for the SIGCHLD signal is installed.

Standard: Fortran 2008 and later

Class: Subroutine

Syntax: CALL EXECUTE_COMMAND_LINE(COMMAND [, WAIT, EXITSTAT, CMDSTAT,
CMDMSG ])

Arguments:

COMMAND Shall be a default CHARACTER scalar.
WAIT  (Optional) Shall be a default LOGICAL scalar.
EXITSTAT (Optional) Shall be an INTEGER of the default kind.
CMDSTAT (Optional) Shall be an INTEGER of the default kind.

CMDMSG (Optional) Shall be an CHARACTER scalar of the default kind.

Example:

      program test_exec

                    integer :: i
                    call execute_command_line ("external_prog.exe", exitstat=i)
                    print *, "Exit status of external_prog.exe was ", i
                    call execute_command_line ("reindex_files.exe", wait=.false.)
                    print *, "Now reindexing files in the background"
                  end program test_exec

Note:

Because this intrinsic is implemented in terms of the system function call, its behavior with respect to signaling is processor dependent. In particular, on POSIX-compliant systems, the SIGINT and SIGQUIT signals will be ignored, and the SIGCHLD will be blocked. As such, if the parent process is terminated, the child process might not be terminated alongside.

See also: Section 8.263 [SYSTEM], page 278,

8.101 EXIT — Exit the program with status.

Description:

EXIT causes immediate termination of the program with status. If status is omitted it returns the canonical success for the system. All Fortran I/O units are closed.

Standard: GNU extension

Class: Subroutine

Syntax: CALL EXIT([STATUS])

Arguments:

STATUS  Shall be an INTEGER of the default kind.

Return value:

STATUS is passed to the parent process on exit.

Example:

                  program test_exit
                    integer :: STATUS = 0
                    print *, 'This program is going to exit.'
                    call EXIT(STATUS)
                  end program test_exit

See also: Section 8.2 [ABORT], page 111, Section 8.163 [KILL], page 218,

8.102 EXP — Exponential function

Description:

EXP(X) computes the base e exponential of X.

Standard: Fortran 77 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = EXP(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value has same type and kind as X.

Example:

                  program test_exp
                    real :: x = 1.0
                    x = exp(x)
                  end program test_exp

Specific names:

Name  Argument  Return type Standard

          EXP(X)         REAL(4) X      REAL(4)  Fortran 77 and later
          DEXP(X)        REAL(8) X      REAL(8)  Fortran 77 and later
          CEXP(X)        COMPLEX(4) X   COMPLEX(4) Fortran 77 and later
          ZEXP(X)        COMPLEX(8) X   COMPLEX(8) GNU extension
          CDEXP(X)       COMPLEX(8) X   COMPLEX(8) GNU extension

8.103 EXPONENT — Exponent function

Description:

EXPONENT(X) returns the value of the exponent part of X. If X is zero the value returned is zero.

Standard: Fortran 90 and later Class: Elemental function

Syntax: RESULT = EXPONENT(X)

Arguments:

X  The type shall be REAL.

Return value:

The return value is of type default INTEGER.

Example:

                  program test_exponent
                    real :: x = 1.0
                    integer :: i
                    i = exponent(x)
                    print *, i
                    print *, exponent(0.0)
                  end program test_exponent

8.104 EXTENDS_TYPE_OF — Query dynamic type for extension

Description:

Query dynamic type for extension.

Standard: Fortran 2003 and later Class: Inquiry function

Syntax: RESULT = EXTENDS_TYPE_OF(A, MOLD)

Arguments:

A Shall be an object of extensible declared type or unlimited polymorphic.

MOLD Shall be an object of extensible declared type or unlimited polymorphic.

Return value:

The return value is a scalar of type default logical. It is true if and only if the dynamic type of A is an extension type of the dynamic type of MOLD.

See also: Section 8.234 [SAME TYPE AS], page 260,

8.105 FDATE — Get the current time as a string

Description:

FDATE(DATE) returns the current date (using the same format as Section 8.81 [CTIME], page 165) in DATE. It is equivalent to CALL CTIME(DATE, TIME()). This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

          CALL FDATE(DATE).
          DATE = FDATE().

Arguments:

DATE The type shall be of type CHARACTER of the default kind. It is an INTENT(OUT) argument. If the length of this variable is too short for the date and time string to fit completely, it will be blank on procedure return.

Return value:

The current date and time as a string.

Example:

                  program test_fdate
                      integer(8) :: i, j
                      character(len=30) :: date
                      call fdate(date)
                      print *, 'Program started on ', date
                      do i = 1, 100000000 ! Just a delay
                          j = i * i - i
                      end do
                      call fdate(date)
                      print *, 'Program ended on ', date
                  end program test_fdate

See also: Section 8.82 [DATE AND TIME], page 166, Section 8.81 [CTIME], page 165,

8.106 FGET — Read a single character in stream mode from stdin

Description:

Read a single character in stream mode from stdin by bypassing normal formatted output. Stream I/O should not be mixed with normal record-oriented (formatted or unformatted) I/O on the same unit; the results are unpredictable.

This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Note that the FGET intrinsic is provided for backwards compatibility with g77. GNU Fortran provides the Fortran 2003 Stream facility. Programmers should consider the use of new stream IO feature in new code for future portability. See also Section 1.3.2 [Fortran 2003 status], page 4.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL FGET(C [, STATUS])
STATUS = FGET(C)

Arguments:

C  The type shall be CHARACTER and of default kind.

STATUS (Optional) status flag of type INTEGER. Returns 0 on success, -1 on end-of-file, and a system specific positive error code otherwise.

Example:

      PROGRAM test_fget
        INTEGER, PARAMETER :: strlen = 100
        INTEGER :: status, i = 1
        CHARACTER(len=strlen) :: str = ""
        WRITE (*,*) 'Enter text:'
        DO
          CALL fget(str(i:i), status)
          if (status /= 0 .OR. i > strlen) exit
          i = i + 1
        END DO
        WRITE (*,*) TRIM(str)
      END PROGRAM

See also: Section 8.107 [FGETC], page 183, Section 8.112 [FPUT], page 186, Section 8.113 [FPUTC], page 187,

8.107 FGETC — Read a single character in stream mode

Description:

Read a single character in stream mode by bypassing normal formatted output. Stream I/O should not be mixed with normal record-oriented (formatted or unformatted) I/O on the same unit; the results are unpredictable. This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit. Note that the FGET intrinsic is provided for backwards compatibility with g77. GNU Fortran provides the Fortran 2003 Stream facility. Programmers should consider the use of new stream IO feature in new code for future portability. See also Section 1.3.2 [Fortran 2003 status], page 4.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL FGETC(UNIT, C [, STATUS])
STATUS = FGETC(UNIT, C)

Arguments:

UNIT  The type shall be INTEGER.
C  The type shall be CHARACTER and of default kind.

STATUS (Optional) status flag of type INTEGER. Returns 0 on success, -1 on end-of-file and a system specific positive error code otherwise.

Example:

      PROGRAM test_fgetc
        INTEGER :: fd = 42, status
        CHARACTER :: c
        OPEN(UNIT=fd, FILE="/etc/passwd", ACTION="READ", STATUS = "OLD")
        DO
          CALL fgetc(fd, c, status)
          IF (status /= 0) EXIT
          call fput(c)
        END DO
        CLOSE(UNIT=fd)
      END PROGRAM

See also: Section 8.106 [FGET], page 182, Section 8.112 [FPUT], page 186, Section 8.113 [FPUTC], page 187,

8.108 FINDLOC — Search an array for a value

Description:

Determines the location of the element in the array with the value given in the VALUE argument, or, if the DIM argument is supplied, determines the locations of the elements equal to the VALUE argument element along each row of the array in the DIM direction. If MASK is present, only the elements for which MASK is .TRUE. are considered. If more than one element in the array has the value VALUE, the location returned is that of the first such element in array element order if the BACK is not present or if it is .FALSE.. If BACK is true, the location returned is that of the last such element. If the array has zero size, or all of the elements of MASK are .FALSE., then the result is an array of zeroes. Similarly, if DIM is supplied and all of the elements of MASK along a given row are zero, the result value for that row is zero.

Standard: Fortran 2008 and later.

Class: Transformational function

Syntax:

RESULT = FINDLOC(ARRAY, VALUE, DIM [, MASK] [,KIND]
[,BACK])
RESULT = FINDLOC(ARRAY, VALUE, [, MASK] [,KIND]
[,BACK])

Arguments:

ARRAY  Shall be an array of intrinsic type.

VALUE A scalar of intrinsic type which is in type conformance with ARRAY.

DIM (Optional) Shall be a scalar of type INTEGER, with a value between one and the rank of ARRAY, inclusive. It may not be an optional dummy argument.

MASK (Optional) Shall be of type LOGICAL, and conformable with ARRAY.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

BACK  (Optional) A scalar of type LOGICAL.

Return value:

If DIM is absent, the result is a rank-one array with a length equal to the rank of ARRAY. If DIM is present, the result is an array with a rank one less than the rank of ARRAY, and a size corresponding to the size of ARRAY with the DIM dimension removed. If DIM is present and ARRAY has a rank of one, the result is a scalar. If the optional argument KIND is present, the result is an integer of kind KIND, otherwise it is of default kind.

See also: Section 8.190 [MAXLOC], page 234, Section 8.198 [MINLOC], page 238,

8.109 FLOOR — Integer floor function

Description:

FLOOR(A) returns the greatest integer less than or equal to A.

Standard: Fortran 95 and later

Class: Elemental function

Syntax: RESULT = FLOOR(A [, KIND])

Arguments:

A  The type shall be REAL.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER(KIND) if KIND is present and of defaultkind INTEGER otherwise.

Example:

      program test_floor
          real :: x = 63.29
          real :: y = -63.59
          print *, floor(x) ! returns 63
          print *, floor(y) ! returns -64
      end program test_floor

See also: Section 8.58 [CEILING], page 149, Section 8.206 [NINT], page 244,

8.110 FLUSH — Flush I/O unit(s)

Description:

Flushes Fortran unit(s) currently open for output. Without the optional argument, all units are flushed, otherwise just the unit specified.

Standard: GNU extension

Class: Subroutine

Syntax: CALL FLUSH(UNIT)

Arguments:

UNIT  (Optional) The type shall be INTEGER.

Note: Beginning with the Fortran 2003 standard, there is a FLUSH statement that should be preferred over the FLUSH intrinsic. The FLUSH intrinsic and the Fortran 2003 FLUSH statement have identical effect: they flush the runtime library’s I/O buffer so that the data becomes visible to other processes. This does not guarantee that the data is committed to disk. On POSIX systems, you can request that all data is transferred to the storage device by calling the fsync function, with the POSIX file descriptor of the I/O unit as argument (retrieved with GNU intrinsic FNUM). The following example shows how:

        ! Declare the interface for POSIX fsync function
        interface
          function fsync (fd) bind(c,name="fsync")
          use iso_c_binding, only: c_int
            integer(c_int), value :: fd
            integer(c_int) :: fsync
          end function fsync
        end interface

                    ! Variable declaration
                    integer :: ret
                    ! Opening unit 10
                    open (10,file="foo")
                    ! ...
                    ! Perform I/O on unit 10
                    ! ...
                    ! Flush and sync
                    flush(10)
                    ret = fsync(fnum(10))
                    ! Handle possible error
                    if (ret /= 0) stop "Error calling FSYNC"

8.111 FNUM — File number function

Description:

FNUM(UNIT) returns the POSIX file descriptor number corresponding to the open Fortran I/O unit UNIT.

Standard: GNU extension

Class: Function

Syntax: RESULT = FNUM(UNIT)

Arguments:

UNIT  The type shall be INTEGER.

Return value:

The return value is of type INTEGER

Example:

                  program test_fnum
                    integer :: i
                    open (unit=10, status = "scratch")
                    i = fnum(10)
                    print *, i
                    close (10)
                  end program test_fnum

8.112 FPUT — Write a single character in stream mode to stdout

Description:

Write a single character in stream mode to stdout by bypassing normal formatted output. Stream I/O should not be mixed with normal record-oriented (formatted or unformatted) I/O on the same unit; the results are unpredictable. This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Note that the FGET intrinsic is provided for backwards compatibility with g77. GNU Fortran provides the Fortran 2003 Stream facility. Programmers should consider the use of new stream IO feature in new code for future portability. See also Section 1.3.2 [Fortran 2003 status], page 4.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL FPUT(C [, STATUS])
STATUS = FPUT(C)

Arguments:

C  The type shall be CHARACTER and of default kind.

STATUS (Optional) status flag of type INTEGER. Returns 0 on success, -1 on end-of-file and a system specific positive error code otherwise.

Example:

      PROGRAM test_fput
        CHARACTER(len=10) :: str = "gfortran"
        INTEGER :: i
        DO i = 1, len_trim(str)
          CALL fput(str(i:i))
        END DO
      END PROGRAM

See also: Section 8.113 [FPUTC], page 187, Section 8.106 [FGET], page 182, Section 8.107 [FGETC], page 183,

8.113 FPUTC — Write a single character in stream mode

Description:

Write a single character in stream mode by bypassing normal formatted output. Stream I/O should not be mixed with normal record-oriented (formatted or unformatted) I/O on the same unit; the results are unpredictable. This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit. Note that the FGET intrinsic is provided for backwards compatibility with g77. GNU Fortran provides the Fortran 2003 Stream facility. Programmers should consider the use of new stream IO feature in new code for future portability. See also Section 1.3.2 [Fortran 2003 status], page 4.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL FPUTC(UNIT, C [, STATUS])
STATUS = FPUTC(UNIT, C)

Arguments:

UNIT  The type shall be INTEGER.
C  The type shall be CHARACTER and of default kind.

STATUS (Optional) status flag of type INTEGER. Returns 0 on success, -1 on end-of-file and a system specific positive error code otherwise.

Example:

                  PROGRAM test_fputc
                    CHARACTER(len=10) :: str = "gfortran"
                    INTEGER :: fd = 42, i
                    OPEN(UNIT = fd, FILE = "out", ACTION = "WRITE", STATUS="NEW")
                    DO i = 1, len_trim(str)
                      CALL fputc(fd, str(i:i))
                    END DO
                    CLOSE(fd)
                  END PROGRAM

See also: Section 8.112 [FPUT], page 186, Section 8.106 [FGET], page 182, Section 8.107 [FGETC], page 183,

8.114 FRACTION — Fractional part of the model representation

Description:

FRACTION(X) returns the fractional part of the model representation of X.

Standard: Fortran 90 and later

Class: Elemental function

Syntax: Y = FRACTION(X)

Arguments:

X  The type of the argument shall be a REAL.

Return value:

The return value is of the same type and kind as the argument. The fractional part of the model representation of X is returned; it is X * RADIX(X)**(-

          EXPONENT(X)).

Example:

                  program test_fraction
                    real :: x
                    x = 178.1387e-4
                    print *, fraction(x), x * radix(x)**(-exponent(x))
                  end program test_fraction

8.115 FREE — Frees memory

Description:

Frees memory previously allocated by MALLOC. The FREE intrinsic is an extension intended to be used with Cray pointers, and is provided in GNU Fortran to allow user to compile legacy code. For new code using Fortran 95 pointers, the memory de-allocation intrinsic is DEALLOCATE.

Standard: GNU extension

Class: Subroutine

Syntax: CALL FREE(PTR)

Arguments:

PTR The type shall be INTEGER. It represents the location of the memory that should be de-allocated.

Return value:

None

Example: See MALLOC for an example.

See also: Section 8.184 [MALLOC], page 230,

8.116 FSEEK — Low level file positioning subroutine

Description:

Moves UNIT to the specified OFFSET. If WHENCE is set to 0, the OFFSET is taken as an absolute value SEEK_SET, if set to 1, OFFSET is taken to be relative to the current position SEEK_CUR, and if set to 2 relative to the end of the file SEEK_END. On error, STATUS is set to a nonzero value. If STATUS the seek fails silently. This intrinsic routine is not fully backwards compatible with g77. In g77, the FSEEK takes a statement label instead of a STATUS variable. If FSEEK is used in old code, change

        CALL FSEEK(UNIT, OFFSET, WHENCE, *label)

to

        INTEGER :: status
        CALL FSEEK(UNIT, OFFSET, WHENCE, status)
        IF (status /= 0) GOTO label

Please note that GNU Fortran provides the Fortran 2003 Stream facility. Programmers should consider the use of new stream IO feature in new code for future portability. See also Section 1.3.2 [Fortran 2003 status], page 4.

Standard: GNU extension

Class: Subroutine

Syntax: CALL FSEEK(UNIT, OFFSET, WHENCE[, STATUS])

Arguments:

UNIT  Shall be a scalar of type INTEGER.
OFFSET  Shall be a scalar of type INTEGER.

WHENCE Shall be a scalar of type INTEGER. Its value shall be either 0, 1 or 2.

STATUS  (Optional) shall be a scalar of type INTEGER(4).

Example:

      PROGRAM test_fseek
        INTEGER, PARAMETER :: SEEK_SET = 0, SEEK_CUR = 1, SEEK_END = 2
        INTEGER :: fd, offset, ierr
        ierr   = 0
        offset = 5
        fd     = 10
        OPEN(UNIT=fd, FILE="fseek.test")
        CALL FSEEK(fd, offset, SEEK_SET, ierr)  ! move to OFFSET
        print *, FTELL(fd), ierr
        CALL FSEEK(fd, 0, SEEK_END, ierr)       ! move to end
        print *, FTELL(fd), ierr
        CALL FSEEK(fd, 0, SEEK_SET, ierr)       ! move to beginning
        print *, FTELL(fd), ierr
        CLOSE(UNIT=fd)
      END PROGRAM

See also: Section 8.118 [FTELL], page 191,

8.117 FSTAT — Get file status

Description:

FSTAT is identical to Section 8.259 [STAT], page 275, except that information about an already opened file is obtained. The elements in VALUES are the same as described by Section 8.259 [STAT], page 275. This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL FSTAT(UNIT, VALUES [, STATUS])
STATUS = FSTAT(UNIT, VALUES)

Arguments:

UNIT  An open I/O unit number of type INTEGER.

VALUES  The type shall be INTEGER(4), DIMENSION(13).

STATUS (Optional) status flag of type INTEGER(4). Returns 0 on success and a system specific error code otherwise.

Example: See Section 8.259 [STAT], page 275, for an example.

See also: To stat a link:

Section 8.182 [LSTAT], page 229, To stat a file: Section 8.259 [STAT], page 275,

8.118 FTELL — Current stream position

Description:

Retrieves the current position within an open file.

This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

          CALL FTELL(UNIT, OFFSET)
          OFFSET = FTELL(UNIT)

Arguments:

OFFSET  Shall of type INTEGER.
UNIT  Shall of type INTEGER.

Return value:

In either syntax, OFFSET is set to the current offset of unit number UNIT, or to −1 if the unit is not currently open.

Example:

                  PROGRAM test_ftell
                    INTEGER :: i
                    OPEN(10, FILE="temp.dat")
                    CALL ftell(10,i)
                    WRITE(*,*) i
                  END PROGRAM

See also: Section 8.116 [FSEEK], page 189,

8.119 GAMMA — Gamma function

Description:

GAMMA(X) computes Gamma (Γ) of X. For positive, integer values of X the Gamma function simplifies to the factorial function Γ(x) = (x − 1)!.

Γ(x) = ∫_{0}^{∞} t^{x−1}e^{−t} dt

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: X = GAMMA(X)

Arguments:

X Shall be of type REAL and neither zero nor a negative integer.

Return value:

The return value is of type REAL of the same kind as X.

Example:

                  program test_gamma
                    real :: x = 1.0
                    x = gamma(x) ! returns 1.0
                  end program test_gamma

Specific names:

Name  Argument  Return type Standard

          DGAMMA(X)      REAL(8) X      REAL(8)  GNU extension

See also: Logarithm of the Gamma function:

Section 8.179 [LOG GAMMA], page 227,

8.120 GERROR — Get last system error message

Description:

Returns the system error message corresponding to the last system error. This resembles the functionality of strerror(3) in C.

Standard: GNU extension Class: Subroutine

Syntax: CALL GERROR(RESULT)

Arguments:

RESULT  Shall be of type CHARACTER and of default kind.

Example:

                  PROGRAM test_gerror
                    CHARACTER(len=100) :: msg
                    CALL gerror(msg)
                    WRITE(*,*) msg
                  END PROGRAM

See also: Section 8.146 [IERRNO], page 209, Section 8.214 [PERROR], page 249,

8.121 GETARG — Get command line arguments

Description:

Retrieve the POS-th argument that was passed on the command line when the containing program was invoked. This intrinsic routine is provided for backwards compatibility with GNU Fortran 77. In new code, programmers should consider the use of the Section 8.123 [GET COMMAND ARGUMENT], page 194, intrinsic defined by the Fortran 2003 standard.

Standard: GNU extension Class: Subroutine

Syntax: CALL GETARG(POS, VALUE)

Arguments:

POS Shall be of type INTEGER and not wider than the default integer kind; POS ≥ 0

VALUE  Shall be of type CHARACTER and of default kind.

Return value:

After GETARG returns, the VALUE argument holds the POSth command line argument. If VALUE cannot hold the argument, it is truncated to fit the length of VALUE. If there are less than POS arguments specified at the command line, VALUE will be filled with blanks. If POS = 0, VALUE is set to the name of the program (on systems that support this feature).

Example:

      PROGRAM test_getarg
        INTEGER :: i
        CHARACTER(len=32) :: arg
        DO i = 1, iargc()
          CALL getarg(i, arg)
          WRITE (*,*) arg
        END DO
      END PROGRAM

See also: GNU Fortran 77 compatibility function:

Section 8.139 [IARGC], page 204, Fortran 2003 functions and subroutines: Section 8.122 [GET COMMAND], page 193, Section 8.123 [GET COMMAND ARGUMENT], page 194, Section 8.68 [COMMAND ARGUMENT COUNT], page 157,

8.122 GET_COMMAND — Get the entire command line

Description:

Retrieve the entire command line that was used to invoke the program.

Standard: Fortran 2003 and later

Class: Subroutine

Syntax: CALL GET_COMMAND([COMMAND, LENGTH, STATUS])

Arguments:

COMMAND (Optional) shall be of type CHARACTER and of default kind.

LENGTH (Optional) Shall be of type INTEGER and of default kind.

STATUS (Optional) Shall be of type INTEGER and of default kind.

Return value:

If COMMAND is present, stores the entire command line that was used to invoke the program in COMMAND. If LENGTH is present, it is assigned the length of the command line. If STATUS is present, it is assigned 0 upon success of the command, -1 if COMMAND is too short to store the command line, or a positive value in case of an error.

Example:

      PROGRAM test_get_command
        CHARACTER(len=255) :: cmd
        CALL get_command(cmd)
        WRITE (*,*) TRIM(cmd)
      END PROGRAM

See also: Section 8.123 [GET COMMAND ARGUMENT], page 194, Section 8.68 [COMMAND ARGUMENT COUNT], page 157,

8.123 GET_COMMAND_ARGUMENT — Get command line arguments

Description:

Retrieve the NUMBER-th argument that was passed on the command line when the containing program was invoked.

Standard: Fortran 2003 and later

Class: Subroutine

Syntax: CALL GET_COMMAND_ARGUMENT(NUMBER [, VALUE, LENGTH, STATUS])

Arguments:

NUMBER Shall be a scalar of type INTEGER and of default kind, NUMBER ≥ 0

VALUE (Optional) Shall be a scalar of type CHARACTER and of default kind.

LENGTH (Optional) Shall be a scalar of type INTEGER and of default kind.

STATUS (Optional) Shall be a scalar of type INTEGER and of default kind.

Return value:

After GET_COMMAND_ARGUMENT returns, the VALUE argument holds the NUM-BER-th command line argument. If VALUE cannot hold the argument, it is truncated to fit the length of VALUE. If there are less than NUMBER arguments specified at the command line, VALUE will be filled with blanks. If NUMBER = 0, VALUE is set to the name of the program (on systems that support this feature). The LENGTH argument contains the length of the NUM-BER-th command line argument. If the argument retrieval fails, STATUS is a positive number; if VALUE contains a truncated command line argument, STATUS is -1; and otherwise the STATUS is zero.

Example:

      PROGRAM test_get_command_argument
        INTEGER :: i
        CHARACTER(len=32) :: arg
        i = 0
        DO

          CALL get_command_argument(i, arg)
          IF (LEN_TRIM(arg) == 0) EXIT
          WRITE (*,*) TRIM(arg)
          i = i+1
        END DO
      END PROGRAM

See also: Section 8.122 [GET COMMAND], page 193, Section 8.68 [COMMAND ARGUMENT COUNT], page 157,

8.124 GETCWD — Get current working directory

Description:

Get current working directory. This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL GETCWD(C [, STATUS])
STATUS = GETCWD(C)

Arguments:

C  The type shall be CHARACTER and of default kind.

STATUS (Optional) status flag. Returns 0 on success, a system specific and nonzero error code otherwise.

Example:

      PROGRAM test_getcwd
        CHARACTER(len=255) :: cwd
        CALL getcwd(cwd)
        WRITE(*,*) TRIM(cwd)
      END PROGRAM

See also: Section 8.60 [CHDIR], page 150,

8.125 GETENV — Get an environmental variable

Description:

Get the VALUE of the environmental variable NAME. This intrinsic routine is provided for backwards compatibility with GNU Fortran 77. In new code, programmers should consider the use of the Section 8.126 [GET ENVIRONMENT VARIABLE], page 196, intrinsic defined by the Fortran 2003 standard. Note that GETENV need not be thread-safe. It is the responsibility of the user to ensure that the environment is not being updated concurrently with a call to the GETENV intrinsic.

Standard: GNU extension Class: Subroutine

Syntax: CALL GETENV(NAME, VALUE)

Arguments:

NAME  Shall be of type CHARACTER and of default kind.
VALUE  Shall be of type CHARACTER and of default kind.

Return value:

Stores the value of NAME in VALUE. If VALUE is not large enough to hold the data, it is truncated. If NAME is not set, VALUE will be filled with blanks.

Example:

      PROGRAM test_getenv
        CHARACTER(len=255) :: homedir
        CALL getenv("HOME", homedir)
        WRITE (*,*) TRIM(homedir)
      END PROGRAM

See also: Section 8.126 [GET ENVIRONMENT VARIABLE], page 196,

8.126 GET_ENVIRONMENT_VARIABLE — Get an environmental

variable

Description:

Get the VALUE of the environmental variable NAME. Note that GET_ENVIRONMENT_VARIABLE need not be thread-safe. It is the responsibility of the user to ensure that the environment is not being updated concurrently with a call to the GET_ENVIRONMENT_VARIABLE intrinsic.

Standard: Fortran 2003 and later

Class: Subroutine

Syntax: CALL GET_ENVIRONMENT_VARIABLE(NAME[, VALUE, LENGTH, STATUS,
TRIM_NAME)

Arguments:

NAME Shall be a scalar of type CHARACTER and of default kind.

VALUE (Optional) Shall be a scalar of type CHARACTER and of default kind.

LENGTH (Optional) Shall be a scalar of type INTEGER and of default kind.

STATUS (Optional) Shall be a scalar of type INTEGER and of default kind.

TRIM NAME (Optional) Shall be a scalar of type LOGICAL and of default kind.

Return value:

Stores the value of NAME in VALUE. If VALUE is not large enough to hold the data, it is truncated. If NAME is not set, VALUE will be filled with blanks. Argument LENGTH contains the length needed for storing the environment variable NAME or zero if it is not present. STATUS is -1 if VALUE is present but too short for the environment variable; it is 1 if the environment variable does not exist and 2 if the processor does not support environment variables; in all other cases STATUS is zero. If TRIM NAME is present with the value .FALSE., the trailing blanks in NAME are significant; otherwise they are not part of the environment variable name.

Example:

                  PROGRAM test_getenv
                    CHARACTER(len=255) :: homedir
                    CALL get_environment_variable("HOME", homedir)
                    WRITE (*,*) TRIM(homedir)
                  END PROGRAM

8.127 GETGID — Group ID function

Description:

Returns the numerical group ID of the current process.

Standard: GNU extension

Class: Function

Syntax: RESULT = GETGID()

Return value:

The return value of GETGID is an INTEGER of the default kind.

Example: See GETPID for an example.

See also: Section 8.129 [GETPID], page 198, Section 8.130 [GETUID], page 198,

8.128 GETLOG — Get login name

Description:

Gets the username under which the program is running.

Standard: GNU extension

Class: Subroutine

Syntax: CALL GETLOG(C)

Arguments:

C  Shall be of type CHARACTER and of default kind.

Return value:

Stores the current user name in C. (On systems where POSIX functions geteuid and getpwuid are not available, and the getlogin function is not implemented either, this will return a blank string.)

Example:

                  PROGRAM TEST_GETLOG
                    CHARACTER(32) :: login

                    CALL GETLOG(login)
                    WRITE(*,*) login
                  END PROGRAM

See also: Section 8.130 [GETUID], page 198,

8.129 GETPID — Process ID function

Description:

Returns the numerical process identifier of the current process.

Standard: GNU extension Class: Function

Syntax: RESULT = GETPID()

Return value:

The return value of GETPID is an INTEGER of the default kind.

Example:

                  program info
                    print *, "The current process ID is ", getpid()
                    print *, "Your numerical user ID is ", getuid()
                    print *, "Your numerical group ID is ", getgid()
                  end program info

See also: Section 8.127 [GETGID], page 197, Section 8.130 [GETUID], page 198,

8.130 GETUID — User ID function

Description:

Returns the numerical user ID of the current process.

Standard: GNU extension Class: Function

Syntax: RESULT = GETUID()

Return value:

The return value of GETUID is an INTEGER of the default kind.

Example: See GETPID for an example. See also: Section 8.129 [GETPID], page 198,

Section 8.128 [GETLOG], page 197,

8.131 GMTIME — Convert time to GMT info

Description:

Given a system time value TIME (as provided by the Section 8.269 [TIME], page 282, intrinsic), fills VALUES with values extracted from it appropriate to the UTC time zone (Universal Coordinated Time, also known in some countries as GMT, Greenwich Mean Time), using gmtime(3).

This intrinsic routine is provided for backwards compatibility with GNU Fortran 77. In new code, programmers should consider the use of the Section 8.82 [DATE AND TIME], page 166, intrinsic defined by the Fortran 95 standard.

Standard: GNU extension

Class: Subroutine

Syntax: CALL GMTIME(TIME, VALUES)

Arguments:

TIME An INTEGER scalar expression corresponding to a system time, with INTENT(IN).

VALUES A default INTEGER array with 9 elements, with

            INTENT(OUT).

Return value:

The elements of VALUES are assigned as follows:

1. Seconds after the minute, range 0–59 or 0–61 to allow for leap seconds 2. Minutes after the hour, range 0–59

3. Hours past midnight, range 0–23

4. Day of month, range 1–31

5. Number of months since January, range 0–11

6. Years since 1900

7. Number of days since Sunday, range 0–6

8. Days since January 1, range 0–365

9. Daylight savings indicator: positive if daylight savings is in effect, zero if not, and negative if the information is not available.

See also: Section 8.82 [DATE AND TIME], page 166, Section 8.81 [CTIME], page 165, Section 8.183 [LTIME], page 230, Section 8.269 [TIME], page 282, Section 8.270 [TIME8], page 283,

8.132 HOSTNM — Get system host name

Description:

Retrieves the host name of the system on which the program is running. This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL HOSTNM(C [, STATUS])
STATUS = HOSTNM(NAME)

Arguments:

C  Shall of type CHARACTER and of default kind.

STATUS (Optional) status flag of type INTEGER. Returns 0 on success, or a system specific error code otherwise.

Return value:

In either syntax, NAME is set to the current hostname if it can be obtained, or to a blank string otherwise.

8.133 HUGE — Largest number of a kind

Description:

HUGE(X) returns the largest number that is not an infinity in the model of the type of X.

Standard: Fortran 90 and later

Class: Inquiry function

Syntax: RESULT = HUGE(X)

Arguments:

X  Shall be of type REAL or INTEGER.

Return value:

The return value is of the same type and kind as X

Example:

                  program test_huge_tiny
                    print *, huge(0), huge(0.0), huge(0.0d0)
                    print *, tiny(0.0), tiny(0.0d0)
                  end program test_huge_tiny

8.134 HYPOT — Euclidean distance function

Description:

HYPOT(X,Y) is the Euclidean distance function. It is equal to \sqrt{X^{2} + Y^{2}}, without undue underflow or overflow.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = HYPOT(X, Y)

Arguments:

X  The type shall be REAL.

Y The type and kind type parameter shall be the same as X.

Return value:

The return value has the same type and kind type parameter as X.

Example:

                  program test_hypot
                    real(4) :: x = 1.e0_4, y = 0.5e0_4
                    x = hypot(x,y)
                  end program test_hypot

8.135 IACHAR — Code in ASCII collating sequence

Description:

IACHAR(C) returns the code for the ASCII character in the first character position of C.

Standard: Fortran 95 and later, with KIND argument Fortran 2003 and later

Class: Elemental function

Syntax: RESULT = IACHAR(C [, KIND])

Arguments:

C  Shall be a scalar CHARACTER, with INTENT(IN)

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind.

Example:

      program test_iachar
        integer i
        i = iachar(' ')
      end program test_iachar

Note: See Section 8.143 [ICHAR], page 206, for a discussion of converting between numerical values and formatted string representations.

See also: Section 8.5 [ACHAR], page 113, Section 8.59 [CHAR], page 149, Section 8.143 [ICHAR], page 206,

8.136 IALL — Bitwise AND of array elements

Description:

Reduces with bitwise AND the elements of ARRAY along dimension DIM if the corresponding element in MASK is TRUE.

Standard: Fortran 2008 and later

Class: Transformational function

Syntax:

RESULT = IALL(ARRAY[, MASK])
RESULT = IALL(ARRAY, DIM[, MASK])

Arguments:

ARRAY  Shall be an array of type INTEGER

DIM (Optional) shall be a scalar of type INTEGER with a value in the range from 1 to n, where n equals the rank of ARRAY.

MASK (Optional) shall be of type LOGICAL and either be a scalar or an array of the same shape as ARRAY.

Return value:

The result is of the same type as ARRAY. If DIM is absent, a scalar with the bitwise ALL of all elements in ARRAY is returned. Otherwise, an array of rank n-1, where n equals the rank of ARRAY, and a shape similar to that of ARRAY with dimension DIM dropped is returned.

Example:

      PROGRAM test_iall
        INTEGER(1) :: a(2)
        a(1) = b'00100100'
        a(2) = b'01101010'
        ! prints 00100000
        PRINT '(b8.8)', IALL(a)
      END PROGRAM

See also: Section 8.138 [IANY], page 203, Section 8.153 [IPARITY], page 212, Section 8.137 [IAND], page 202,

8.137 IAND — Bitwise logical and

Description:

Bitwise logical AND.

Standard: Fortran 90 and later, with boz-literal-constant Fortran 2008 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = IAND(I, J)

Arguments:

I  The type shall be INTEGER or a boz-literal-constant.

J The type shall be INTEGER with the same kind type parameter as I or a boz-literal-constant. I and J shall not both be boz-literal-constants.

Return value:

The return type is INTEGER with the kind type parameter of the arguments. A boz-literal-constant is converted to an INTEGER with the kind type parameter of the other argument as-if a call to Section 8.149 [INT], page 210, occurred.

Example:

      PROGRAM test_iand
        INTEGER :: a, b
        DATA a / Z'F' /, b / Z'3' /
        WRITE (*,*) IAND(a, b)
      END PROGRAM

Specific names:

Name  Argument  Return type Standard

IAND(A)        INTEGER A      INTEGER  Fortran 90 and later

BIAND(A)       INTEGER(1) A   INTEGER(1) GNU extension
IIAND(A)       INTEGER(2) A   INTEGER(2) GNU extension
JIAND(A)       INTEGER(4) A   INTEGER(4) GNU extension
KIAND(A)       INTEGER(8) A   INTEGER(8) GNU extension

See also: Section 8.152 [IOR], page 212, Section 8.145 [IEOR], page 208, Section 8.141 [IBITS], page 205, Section 8.142 [IBSET], page 206, Section 8.140 [IBCLR], page 204, Section 8.208 [NOT], page 245,

8.138 IANY — Bitwise OR of array elements

Description:

Reduces with bitwise OR (inclusive or) the elements of ARRAY along dimension DIM if the corresponding element in MASK is TRUE.

Standard: Fortran 2008 and later

Class: Transformational function

Syntax:

RESULT = IANY(ARRAY[, MASK])
RESULT = IANY(ARRAY, DIM[, MASK])

Arguments:

ARRAY  Shall be an array of type INTEGER

DIM (Optional) shall be a scalar of type INTEGER with a value in the range from 1 to n, where n equals the rank of ARRAY.

MASK (Optional) shall be of type LOGICAL and either be a scalar or an array of the same shape as ARRAY.

Return value:

The result is of the same type as ARRAY. If DIM is absent, a scalar with the bitwise OR of all elements in ARRAY is returned. Otherwise, an array of rank n-1, where n equals the rank of ARRAY, and a shape similar to that of ARRAY with dimension DIM dropped is returned.

Example:

      PROGRAM test_iany
        INTEGER(1) :: a(2)
        a(1) = b'00100100'
        a(2) = b'01101010'
        ! prints 01101110
        PRINT '(b8.8)', IANY(a)
      END PROGRAM

See also: Section 8.153 [IPARITY], page 212, Section 8.136 [IALL], page 201, Section 8.152 [IOR], page 212,

8.139 IARGC — Get the number of command line arguments

Description:

IARGC returns the number of arguments passed on the command line when the containing program was invoked. This intrinsic routine is provided for backwards compatibility with GNU Fortran 77. In new code, programmers should consider the use of the Section 8.68 [COMMAND ARGUMENT COUNT], page 157, intrinsic defined by the Fortran 2003 standard.

Standard: GNU extension

Class: Function

Syntax: RESULT = IARGC()

Arguments:

None

Return value:

The number of command line arguments, type INTEGER(4).

Example: See Section 8.121 [GETARG], page 192,

See also: GNU Fortran 77 compatibility subroutine:

Section 8.121 [GETARG], page 192, Fortran 2003 functions and subroutines: Section 8.122 [GET COMMAND], page 193, Section 8.123 [GET COMMAND ARGUMENT], page 194, Section 8.68 [COMMAND ARGUMENT COUNT], page 157,

8.140 IBCLR — Clear bit

Description:

IBCLR returns the value of I with the bit at position POS set to zero.

Standard: Fortran 90 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = IBCLR(I, POS)

Arguments:

I  The type shall be INTEGER.
POS  The type shall be INTEGER.

Return value:

The return value is of type INTEGER and of the same kind as I.

Specific names:

Name  Argument  Return type Standard

IBCLR(A)       INTEGER A      INTEGER  Fortran 90 and later
BBCLR(A)       INTEGER(1) A   INTEGER(1) GNU extension
IIBCLR(A)      INTEGER(2) A   INTEGER(2) GNU extension
JIBCLR(A)      INTEGER(4) A   INTEGER(4) GNU extension
KIBCLR(A)      INTEGER(8) A   INTEGER(8) GNU extension

See also: Section 8.141 [IBITS], page 205, Section 8.142 [IBSET], page 206, Section 8.137 [IAND], page 202, Section 8.152 [IOR], page 212, Section 8.145 [IEOR], page 208, Section 8.203 [MVBITS], page 242,

8.141 IBITS — Bit extraction

Description:

IBITS extracts a field of length LEN from I, starting from bit position POS and extending left for LEN bits. The result is right-justified and the remaining bits are zeroed. The value of POS+LEN must be less than or equal to the value

BIT_SIZE(I).

Standard: Fortran 90 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = IBITS(I, POS, LEN)

Arguments:

I  The type shall be INTEGER.
POS  The type shall be INTEGER.
LEN  The type shall be INTEGER.

Return value:

The return value is of type INTEGER and of the same kind as I.

Specific names:

Name  Argument  Return type Standard

IBITS(A)       INTEGER A      INTEGER  Fortran 90 and later
BBITS(A)       INTEGER(1) A   INTEGER(1) GNU extension
IIBITS(A)      INTEGER(2) A   INTEGER(2) GNU extension
JIBITS(A)      INTEGER(4) A   INTEGER(4) GNU extension
KIBITS(A)      INTEGER(8) A   INTEGER(8) GNU extension

See also: Section 8.48 [BIT SIZE], page 142, Section 8.140 [IBCLR], page 204, Section 8.142 [IBSET], page 206, Section 8.137 [IAND], page 202, Section 8.152 [IOR], page 212, Section 8.145 [IEOR], page 208,

8.142 IBSET — Set bit

Description:

IBSET returns the value of I with the bit at position POS set to one.

Standard: Fortran 90 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = IBSET(I, POS)

Arguments:

I  The type shall be INTEGER.
POS  The type shall be INTEGER.

Return value:

The return value is of type INTEGER and of the same kind as I.

Specific names:

Name  Argument  Return type Standard

IBSET(A)       INTEGER A      INTEGER  Fortran 90 and later
BBSET(A)       INTEGER(1) A   INTEGER(1) GNU extension
IIBSET(A)      INTEGER(2) A   INTEGER(2) GNU extension
JIBSET(A)      INTEGER(4) A   INTEGER(4) GNU extension
KIBSET(A)      INTEGER(8) A   INTEGER(8) GNU extension

See also: Section 8.140 [IBCLR], page 204, Section 8.141 [IBITS], page 205, Section 8.137 [IAND], page 202, Section 8.152 [IOR], page 212, Section 8.145 [IEOR], page 208, Section 8.203 [MVBITS], page 242,

8.143 ICHAR — Character-to-integer conversion function

Description:

ICHAR(C) returns the code for the character in the first character position of C in the system’s native character set. The correspondence between characters and their codes is not necessarily the same across different GNU Fortran implementations.

Standard: Fortran 77 and later, with KIND argument Fortran 2003 and later

Class: Elemental function

Syntax: RESULT = ICHAR(C [, KIND])

Arguments:

C  Shall be a scalar CHARACTER, with INTENT(IN)

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind.

Example:

      program test_ichar
        integer i
        i = ichar(' ')
      end program test_ichar

Specific names:

Name  Argument  Return type Standard

ICHAR(C)       CHARACTER C    INTEGER(4) Fortran 77 and later

Note: No intrinsic exists to convert between a numeric value and a formatted character string representation – for instance, given the CHARACTER value '154', obtaining an INTEGER or REAL value with the value 154, or vice versa. Instead, this functionality is provided by internal-file I/O, as in the following example:

      program read_val
        integer value
        character(len=10) string, string2
        string = '154'
        ! Convert a string to a numeric value
        read (string,'(I10)') value
        print *, value
        ! Convert a value to a formatted string
        write (string2,'(I10)') value
        print *, string2
      end program read_val

See also: Section 8.5 [ACHAR], page 113, Section 8.59 [CHAR], page 149, Section 8.135 [IACHAR], page 201,

8.144 IDATE — Get current local time subroutine (day/month/year)

Description:

IDATE(VALUES) Fills VALUES with the numerical values at the current local time. The day (in the range 1-31), month (in the range 1-12), and year appear in elements 1, 2, and 3 of VALUES, respectively. The year has four significant digits. This intrinsic routine is provided for backwards compatibility with GNU Fortran 77. In new code, programmers should consider the use of the Section 8.82 [DATE AND TIME], page 166, intrinsic defined by the Fortran 95 standard.

Standard: GNU extension

Class: Subroutine

Syntax: CALL IDATE(VALUES)

Arguments:

VALUES

            The type shall be INTEGER, DIMENSION(3) and the

kind shall be the default integer kind.

Return value:

Does not return anything.

Example:

      program test_idate
        integer, dimension(3) :: tarray
        call idate(tarray)
        print *, tarray(1)
        print *, tarray(2)
        print *, tarray(3)
      end program test_idate

See also: Section 8.82 [DATE AND TIME], page 166,

8.145 IEOR — Bitwise logical exclusive or

Description:

IEOR returns the bitwise Boolean exclusive-OR of I and J.

Standard: Fortran 90 and later, with boz-literal-constant Fortran 2008 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = IEOR(I, J)

Arguments:

I  The type shall be INTEGER or a boz-literal-constant.

J The type shall be INTEGER with the same kind type parameter as I or a boz-literal-constant. I and J shall not both be boz-literal-constants.

Return value:

The return type is INTEGER with the kind type parameter of the arguments. A boz-literal-constant is converted to an INTEGER with the kind type parameter of the other argument as-if a call to Section 8.149 [INT], page 210, occurred.

Specific names:

Name  Argument  Return type Standard

IEOR(A)        INTEGER A      INTEGER  Fortran 90 and later
BIEOR(A)       INTEGER(1) A   INTEGER(1) GNU extension
IIEOR(A)       INTEGER(2) A   INTEGER(2) GNU extension
JIEOR(A)       INTEGER(4) A   INTEGER(4) GNU extension
KIEOR(A)       INTEGER(8) A   INTEGER(8) GNU extension

See also: Section 8.152 [IOR], page 212, Section 8.137 [IAND], page 202, Section 8.141 [IBITS], page 205, Section 8.142 [IBSET], page 206, Section 8.140 [IBCLR], page 204, Section 8.208 [NOT], page 245,

8.146 IERRNO — Get the last system error number

Description:

Returns the last system error number, as given by the C errno variable.

Standard: GNU extension

Class: Function

Syntax: RESULT = IERRNO()

Arguments:

None

Return value:

The return value is of type INTEGER and of the default integer kind.

See also: Section 8.214 [PERROR], page 249,

8.147 IMAGE_INDEX — Function that converts a cosubscript to an image index

Description:

Returns the image index belonging to a cosubscript.

Standard: Fortran 2008 and later

Class: Inquiry function.

Syntax: RESULT = IMAGE_INDEX(COARRAY, SUB)

Arguments:

COARRAY Coarray of any type.

SUB default integer rank-1 array of a size equal to the corank of COARRAY.

Return value:

Scalar default integer with the value of the image index which corresponds to the cosubscripts. For invalid cosubscripts the result is zero.

Example:

                  INTEGER :: array[2,-1:4,8,*]
                  ! Writes  28 (or 0 if there are fewer than 28 images)
                  WRITE (*,*) IMAGE_INDEX (array, [2,0,3,1])

See also: Section 8.268 [THIS IMAGE], page 281, Section 8.210 [NUM IMAGES], page 246,

8.148 INDEX — Position of a substring within a string

Description:

Returns the position of the start of the first occurrence of string SUBSTRING as a substring in STRING, counting from one. If SUBSTRING is not present in STRING, zero is returned. If the BACK argument is present and true, the return value is the start of the last occurrence rather than the first.

Standard: Fortran 77 and later, with KIND argument Fortran 2003 and later

Class: Elemental function

Syntax: RESULT = INDEX(STRING, SUBSTRING [, BACK [, KIND]])

Arguments:

STRING  Shall be a scalar CHARACTER, with INTENT(IN)
SUBSTRING Shall be a scalar CHARACTER, with INTENT(IN)

BACK (Optional) Shall be a scalar LOGICAL, with

            INTENT(IN)

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind.

Specific names:

Name  Argument Return type  Standard

INDEX(STRING,SUBSTRING)  CHARACTER INTEGER(4) Fortran 77 and later

See also: Section 8.236 [SCAN], page 261, Section 8.282 [VERIFY], page 289,

8.149 INT — Convert to integer type

Description:

Convert to integer type

Standard: Fortran 77 and later, with boz-literal-constant Fortran 2008 and later.

Class: Elemental function

Syntax: RESULT = INT(A [, KIND))

Arguments:

A Shall be of type INTEGER, REAL, or COMPLEX or a bozliteral-constant.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

These functions return a INTEGER variable or array under the following rules:

(A)  If A is of type INTEGER, INT(A) = A

(B) If A is of type REAL and |A| < 1, INT(A) equals 0. If |A| ≥ 1, then INT(A) is the integer whose magnitude is the largest integer that does not exceed the magnitude of A and whose sign is the same as the sign of A.

(C) If A is of type COMPLEX, rule B is applied to the real part of A.

Example:

                  program test_int
                    integer :: i = 42
                    complex :: z = (-3.7, 1.0)
                    print *, int(i)
                    print *, int(z), int(z,8)
                  end program

Specific names:

Name  Argument  Return type Standard

          INT(A)         REAL(4) A      INTEGER  Fortran 77 and later
          IFIX(A)        REAL(4) A      INTEGER  Fortran 77 and later
          IDINT(A)       REAL(8) A      INTEGER  Fortran 77 and later

8.150 INT2 — Convert to 16-bit integer type

Description:

Convert to a KIND=2 integer type. This is equivalent to the standard INT intrinsic with an optional argument of KIND=2, and is only included for backwards compatibility.

Standard: GNU extension Class: Elemental function

Syntax: RESULT = INT2(A)

Arguments:

A  Shall be of type INTEGER, REAL, or COMPLEX.

Return value:

The return value is a INTEGER(2) variable.

See also: Section 8.149 [INT], page 210, Section 8.151 [INT8], page 211,

8.151 INT8 — Convert to 64-bit integer type

Description:

Convert to a KIND=8 integer type. This is equivalent to the standard INT intrinsic with an optional argument of KIND=8, and is only included for backwards compatibility.

Standard: GNU extension Class: Elemental function

Syntax: RESULT = INT8(A)

Arguments:

A  Shall be of type INTEGER, REAL, or COMPLEX.

Return value:

The return value is a INTEGER(8) variable.

See also: Section 8.149 [INT], page 210, Section 8.150 [INT2], page 211,

8.152 IOR — Bitwise logical or

Description:

IOR returns the bitwise Boolean inclusive-OR of I and J.

Standard: Fortran 90 and later, with boz-literal-constant Fortran 2008 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = IOR(I, J)

Arguments:

I  The type shall be INTEGER or a boz-literal-constant.

J The type shall be INTEGER with the same kind type parameter as I or a boz-literal-constant. I and J shall not both be boz-literal-constants.

Return value:

The return type is INTEGER with the kind type parameter of the arguments. A boz-literal-constant is converted to an INTEGER with the kind type parameter of the other argument as-if a call to Section 8.149 [INT], page 210, occurred.

Specific names:

Name  Argument  Return type Standard

IOR(A)         INTEGER A      INTEGER  Fortran 90 and later
BIOR(A)        INTEGER(1) A   INTEGER(1) GNU extension
IIOR(A)        INTEGER(2) A   INTEGER(2) GNU extension
JIOR(A)        INTEGER(4) A   INTEGER(4) GNU extension
KIOR(A)        INTEGER(8) A   INTEGER(8) GNU extension

See also: Section 8.145 [IEOR], page 208, Section 8.137 [IAND], page 202, Section 8.141 [IBITS], page 205, Section 8.142 [IBSET], page 206, Section 8.140 [IBCLR], page 204, Section 8.208 [NOT], page 245,

8.153 IPARITY — Bitwise XOR of array elements

Description:

Reduces with bitwise XOR (exclusive or) the elements of ARRAY along dimension DIM if the corresponding element in MASK is TRUE.

Standard: Fortran 2008 and later

Class: Transformational function

Syntax:

RESULT = IPARITY(ARRAY[, MASK])
RESULT = IPARITY(ARRAY, DIM[, MASK])

Arguments:

ARRAY  Shall be an array of type INTEGER

DIM (Optional) shall be a scalar of type INTEGER with a value in the range from 1 to n, where n equals the rank of ARRAY.

MASK (Optional) shall be of type LOGICAL and either be a scalar or an array of the same shape as ARRAY.

Return value:

The result is of the same type as ARRAY. If DIM is absent, a scalar with the bitwise XOR of all elements in ARRAY is returned. Otherwise, an array of rank n-1, where n equals the rank of ARRAY, and a shape similar to that of ARRAY with dimension DIM dropped is returned.

Example:

      PROGRAM test_iparity
        INTEGER(1) :: a(2)
        a(1) = int(b'00100100', 1)
        a(2) = int(b'01101010', 1)
        ! prints 01001110
        PRINT '(b8.8)', IPARITY(a)
      END PROGRAM

See also: Section 8.138 [IANY], page 203, Section 8.136 [IALL], page 201, Section 8.145 [IEOR], page 208, Section 8.213 [PARITY], page 249,

8.154 IRAND — Integer pseudo-random number

Description:

IRAND(FLAG) returns a pseudo-random number from a uniform distribution between 0 and a system-dependent limit (which is in most cases 2147483647). If FLAG is 0, the next number in the current sequence is returned; if FLAG is 1, the generator is restarted by CALL SRAND(0); if FLAG has any other value, it is used as a new seed with SRAND. This intrinsic routine is provided for backwards compatibility with GNU Fortran 77. It implements a simple modulo generator as provided by g77. For new code, one should consider the use of Section 8.224 [RANDOM NUMBER], page 255, as it implements a superior algorithm.

Standard: GNU extension

Class: Function

Syntax: RESULT = IRAND(I)

Arguments:

I  Shall be a scalar INTEGER of kind 4.

Return value:

The return value is of INTEGER(kind=4) type.

Example:

                  program test_irand
                    integer,parameter :: seed = 86456
                    call srand(seed)
                    print *, irand(), irand(), irand(), irand()
                    print *, irand(seed), irand(), irand(), irand()
                  end program test_irand

8.155 IS_CONTIGUOUS — Test whether an array is contiguous

Description:

IS_CONTIGUOUS tests whether an array is contiguous.

Standard: Fortran 2008 and later

Class: Inquiry function

Syntax: RESULT = IS_CONTIGUOUS(ARRAY)

Arguments:

ARRAY  Shall be an array of any type.

Return value:

Returns a LOGICAL of the default kind, which .TRUE. if ARRAY is contiguous and false otherwise.

Example:

                  program test
                    integer :: a(10)
                    a = [1,2,3,4,5,6,7,8,9,10]
                    call sub (a)      ! every element, is contiguous
                    call sub (a(::2)) ! every other element, is noncontiguous
                  contains
                    subroutine sub (x)
                      integer :: x(:)
                      if (is_contiguous (x)) then
                        write (*,*) 'X is contiguous'
                      else
                        write (*,*) 'X is not contiguous'
                      end if
                    end subroutine sub
                  end program test

8.156 IS_IOSTAT_END — Test for end-of-file value

Description:

IS_IOSTAT_END tests whether an variable has the value of the I/O status “end of file”. The function is equivalent to comparing the variable with the IOSTAT_END parameter of the intrinsic module ISO_FORTRAN_ENV.

Standard: Fortran 2003 and later

Class: Elemental function

Syntax: RESULT = IS_IOSTAT_END(I)

Arguments:

I  Shall be of the type INTEGER.

Return value:

Returns a LOGICAL of the default kind, which .TRUE. if I has the value which indicates an end of file condition for IOSTAT= specifiers, and is .FALSE. otherwise.

Example:

                  PROGRAM iostat
                    IMPLICIT NONE
                    INTEGER :: stat, i
                    OPEN(88, FILE='test.dat')
                    READ(88, *, IOSTAT=stat) i
                    IF(IS_IOSTAT_END(stat)) STOP 'END OF FILE'
                  END PROGRAM

8.157 IS_IOSTAT_EOR — Test for end-of-record value

Description:

IS_IOSTAT_EOR tests whether an variable has the value of the I/O status “end of record”. The function is equivalent to comparing the variable with the IOSTAT_ EOR parameter of the intrinsic module ISO_FORTRAN_ENV.

Standard: Fortran 2003 and later

Class: Elemental function

Syntax: RESULT = IS_IOSTAT_EOR(I)

Arguments:

I  Shall be of the type INTEGER.

Return value:

Returns a LOGICAL of the default kind, which .TRUE. if I has the value which indicates an end of file condition for IOSTAT= specifiers, and is .FALSE. otherwise.

Example:

                  PROGRAM iostat
                    IMPLICIT NONE
                    INTEGER :: stat, i(50)
                    OPEN(88, FILE='test.dat', FORM='UNFORMATTED')
                    READ(88, IOSTAT=stat) i
                    IF(IS_IOSTAT_EOR(stat)) STOP 'END OF RECORD'
                  END PROGRAM

8.158 ISATTY — Whether a unit is a terminal device

Description:

Determine whether a unit is connected to a terminal device.

Standard: GNU extension

Class: Function

Syntax: RESULT = ISATTY(UNIT)

Arguments:

UNIT  Shall be a scalar INTEGER.

Return value:

Returns .TRUE. if the UNIT is connected to a terminal device, .FALSE. otherwise.

Example:

      PROGRAM test_isatty
        INTEGER(kind=1) :: unit
        DO unit = 1, 10
          write(*,*) isatty(unit=unit)
        END DO
      END PROGRAM

See also: Section 8.276 [TTYNAM], page 286,

8.159 ISHFT — Shift bits

Description:

ISHFT returns a value corresponding to I with all of the bits shifted SHIFT places. A value of SHIFT greater than zero corresponds to a left shift, a value of zero corresponds to no shift, and a value less than zero corresponds to a right shift. If the absolute value of SHIFT is greater than BIT_SIZE(I), the value is undefined. Bits shifted out from the left end or right end are lost; zeros are shifted in from the opposite end.

Standard: Fortran 90 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = ISHFT(I, SHIFT)

Arguments:

I  The type shall be INTEGER.
SHIFT  The type shall be INTEGER.

Return value:

The return value is of type INTEGER and of the same kind as I.

Specific names:

Name  Argument  Return type Standard

ISHFT(A)       INTEGER A      INTEGER  Fortran 90 and later
BSHFT(A)       INTEGER(1) A   INTEGER(1) GNU extension
IISHFT(A)      INTEGER(2) A   INTEGER(2) GNU extension
JISHFT(A)      INTEGER(4) A   INTEGER(4) GNU extension
KISHFT(A)      INTEGER(8) A   INTEGER(8) GNU extension

See also: Section 8.160 [ISHFTC], page 217,

8.160 ISHFTC — Shift bits circularly

Description:

ISHFTC returns a value corresponding to I with the rightmost SIZE bits shifted circularly SHIFT places; that is, bits shifted out one end are shifted into the opposite end. A value of SHIFT greater than zero corresponds to a left shift, a value of zero corresponds to no shift, and a value less than zero corresponds to a right shift. The absolute value of SHIFT must be less than SIZE. If the SIZE argument is omitted, it is taken to be equivalent to BIT_SIZE(I).

Standard: Fortran 90 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = ISHFTC(I, SHIFT [, SIZE])

Arguments:

I  The type shall be INTEGER.
SHIFT  The type shall be INTEGER.

SIZE (Optional) The type shall be INTEGER; the value must be greater than zero and less than or equal to BIT_

            SIZE(I).

Return value:

The return value is of type INTEGER and of the same kind as I.

Specific names:

Name  Argument  Return type Standard

ISHFTC(A)      INTEGER A      INTEGER  Fortran 90 and later
BSHFTC(A)      INTEGER(1) A   INTEGER(1) GNU extension
IISHFTC(A)     INTEGER(2) A   INTEGER(2) GNU extension
JISHFTC(A)     INTEGER(4) A   INTEGER(4) GNU extension
KISHFTC(A)     INTEGER(8) A   INTEGER(8) GNU extension

See also: Section 8.159 [ISHFT], page 216,

8.161 ISNAN — Test for a NaN

Description:

ISNAN tests whether a floating-point value is an IEEE Not-a-Number (NaN).

Standard: GNU extension

Class: Elemental function

Syntax: ISNAN(X)

Arguments:

X  Variable of the type REAL.

Return value:

Returns a default-kind LOGICAL. The returned value is TRUE if X is a NaN and FALSE otherwise.

Example:

                  program test_nan
                    implicit none
                    real :: x
                    x = -1.0
                    x = sqrt(x)
                    if (isnan(x)) stop '"x" is a NaN'
                  end program test_nan

8.162 ITIME — Get current local time subroutine (hour/minutes/seconds)

Description:

ITIME(VALUES) Fills VALUES with the numerical values at the current local time. The hour (in the range 1-24), minute (in the range 1-60), and seconds (in the range 1-60) appear in elements 1, 2, and 3 of VALUES, respectively. This intrinsic routine is provided for backwards compatibility with GNU Fortran 77. In new code, programmers should consider the use of the Section 8.82 [DATE AND TIME], page 166, intrinsic defined by the Fortran 95 standard.

Standard: GNU extension

Class: Subroutine

Syntax: CALL ITIME(VALUES)

Arguments:

VALUES

                      The type shall be INTEGER, DIMENSION(3) and the

kind shall be the default integer kind.

Return value:

Does not return anything.

Example:

                  program test_itime
                    integer, dimension(3) :: tarray
                    call itime(tarray)
                    print *, tarray(1)
                    print *, tarray(2)
                    print *, tarray(3)
                  end program test_itime

See also: Section 8.82 [DATE AND TIME], page 166,

8.163 KILL — Send a signal to a process

Description:

Sends the signal specified by SIG to the process PID. See kill(2).

This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Standard: GNU extension Class: Subroutine, function

Syntax:

          CALL KILL(PID, SIG [, STATUS])
          STATUS = KILL(PID, SIG)

Arguments:

PID  Shall be a scalar INTEGER with INTENT(IN).
SIG  Shall be a scalar INTEGER with INTENT(IN).

STATUS [Subroutine](Optional) Shall be a scalar INTEGER. Returns 0 on success; otherwise a system-specific error code is returned.

STATUS [Function] The kind type parameter is that of pid. Returns 0 on success; otherwise a system-specific error code is returned.

See also: Section 8.2 [ABORT], page 111, Section 8.101 [EXIT], page 179,

8.164 KIND — Kind of an entity

Description:

KIND(X) returns the kind value of the entity X.

Standard: Fortran 95 and later

Class: Inquiry function

Syntax: K = KIND(X)

Arguments:

X

                      Shall be of type LOGICAL, INTEGER, REAL, COMPLEX

or CHARACTER. It may be scalar or array valued.

Return value:

The return value is a scalar of type INTEGER and of the default integer kind.

Example:

                  program test_kind
                    integer,parameter :: kc = kind(' ')
                    integer,parameter :: kl = kind(.true.)
                    print *, "The default character kind is ", kc
                    print *, "The default logical kind is ", kl
                  end program test_kind

8.165 LBOUND — Lower dimension bounds of an array

Description:

Returns the lower bounds of an array, or a single lower bound along the DIM dimension.

Standard: Fortran 90 and later, with KIND argument Fortran 2003 and later Class: Inquiry function

Syntax: RESULT = LBOUND(ARRAY [, DIM [, KIND]])

Arguments:

ARRAY  Shall be an array, of any type.
DIM  (Optional) Shall be a scalar INTEGER.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind. If DIM is absent, the result is an array of the lower bounds of ARRAY. If DIM is present, the result is a scalar corresponding to the lower bound of the array along that dimension. If ARRAY is an expression rather than a whole array or array structure component, or if it has a zero extent along the relevant dimension, the lower bound is taken to be 1.

See also: Section 8.277 [UBOUND], page 287, Section 8.166 [LCOBOUND], page 220,

8.166 LCOBOUND — Lower codimension bounds of an array

Description:

Returns the lower bounds of a coarray, or a single lower cobound along the DIM codimension.

Standard: Fortran 2008 and later

Class: Inquiry function

Syntax: RESULT = LCOBOUND(COARRAY [, DIM [, KIND]])

Arguments:

ARRAY  Shall be an coarray, of any type.
DIM  (Optional) Shall be a scalar INTEGER.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind. If DIM is absent, the result is an array of the lower cobounds of COARRAY. If DIM is present, the result is a scalar corresponding to the lower cobound of the array along that codimension.

See also: Section 8.278 [UCOBOUND], page 287, Section 8.165 [LBOUND], page 219,

8.167 LEADZ — Number of leading zero bits of an integer

Description:

LEADZ returns the number of leading zero bits of an integer.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = LEADZ(I)

Arguments:

I  Shall be of type INTEGER.

Return value:

The type of the return value is the default INTEGER. If all the bits of I are zero, the result value is BIT_SIZE(I).

Example:

      PROGRAM test_leadz
        WRITE (*,*) BIT_SIZE(1)  ! prints 32
        WRITE (*,*) LEADZ(1)     ! prints 31
      END PROGRAM

See also: Section 8.48 [BIT SIZE], page 142, Section 8.272 [TRAILZ], page 284, Section 8.215 [POPCNT], page 250, Section 8.216 [POPPAR], page 250,

8.168 LEN — Length of a character entity

Description:

Returns the length of a character string. If STRING is an array, the length of an element of STRING is returned. Note that STRING need not be defined when this intrinsic is invoked, since only the length, not the content, of STRING is needed.

Standard: Fortran 77 and later, with KIND argument Fortran 2003 and later

Class: Inquiry function

Syntax: L = LEN(STRING [, KIND])

Arguments:

STRING Shall be a scalar or array of type CHARACTER, with

            INTENT(IN)

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind.

Specific names:

Name  Argument  Return type Standard

LEN(STRING)    CHARACTER      INTEGER  Fortran 77 and later

See also: Section 8.169 [LEN TRIM], page 222, Section 8.9 [ADJUSTL], page 116, Section 8.10 [ADJUSTR], page 116,

8.169 LEN_TRIM — Length of a character entity without trailing blank characters

Description:

Returns the length of a character string, ignoring any trailing blanks.

Standard: Fortran 90 and later, with KIND argument Fortran 2003 and later

Class: Elemental function

Syntax: RESULT = LEN_TRIM(STRING [, KIND])

Arguments:

STRING  Shall be a scalar of type CHARACTER, with INTENT(IN)

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind.

See also: Section 8.168 [LEN], page 221, Section 8.9 [ADJUSTL], page 116, Section 8.10 [ADJUSTR], page 116,

8.170 LGE — Lexical greater than or equal

Description:

Determines whether one string is lexically greater than or equal to another string, where the two strings are interpreted as containing ASCII character codes. If the String A and String B are not the same length, the shorter is compared as if spaces were appended to it to form a value that has the same length as the longer. In general, the lexical comparison intrinsics LGE, LGT, LLE, and LLT differ from the corresponding intrinsic operators .GE., .GT., .LE., and .LT., in that the latter use the processor’s character ordering (which is not ASCII on some targets), whereas the former always use the ASCII ordering.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = LGE(STRING_A, STRING_B)

Arguments:

STRING A Shall be of default CHARACTER type.
STRING B Shall be of default CHARACTER type.

Return value:

Returns .TRUE. if STRING_A >= STRING_B, and .FALSE. otherwise, based on

the ASCII ordering.

Specific names:

Name  Argument Return type  Standard

LGE(STRING_A,STRING_B)  CHARACTER LOGICAL  Fortran 77 and later

See also: Section 8.171 [LGT], page 223, Section 8.173 [LLE], page 224, Section 8.174 [LLT], page 225,

8.171 LGT — Lexical greater than

Description:

Determines whether one string is lexically greater than another string, where the two strings are interpreted as containing ASCII character codes. If the String A and String B are not the same length, the shorter is compared as if spaces were appended to it to form a value that has the same length as the longer. In general, the lexical comparison intrinsics LGE, LGT, LLE, and LLT differ from the corresponding intrinsic operators .GE., .GT., .LE., and .LT., in that the latter use the processor’s character ordering (which is not ASCII on some targets), whereas the former always use the ASCII ordering.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = LGT(STRING_A, STRING_B)

Arguments:

STRING A Shall be of default CHARACTER type.
STRING B Shall be of default CHARACTER type.

Return value:

Returns .TRUE. if STRING_A > STRING_B, and .FALSE. otherwise, based on the ASCII ordering.

Specific names:

Name  Argument Return type  Standard

LGT(STRING_A,STRING_B)  CHARACTER LOGICAL  Fortran 77 and later

See also: Section 8.170 [LGE], page 222, Section 8.173 [LLE], page 224, Section 8.174 [LLT], page 225,

8.172 LINK — Create a hard link

Description:

Makes a (hard) link from file PATH1 to PATH2. A null character (CHAR(0)) can be used to mark the end of the names in PATH1 and PATH2; otherwise, trailing blanks in the file names are ignored. If the STATUS argument is supplied, it contains 0 on success or a nonzero error code upon return; see link(2). This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension Class: Subroutine, function

Syntax:

CALL LINK(PATH1, PATH2 [, STATUS])
STATUS = LINK(PATH1, PATH2)

Arguments:

PATH1  Shall be of default CHARACTER type.
PATH2  Shall be of default CHARACTER type.
STATUS  (Optional) Shall be of default INTEGER type.

See also: Section 8.262 [SYMLNK], page 278, Section 8.280 [UNLINK], page 288,

8.173 LLE — Lexical less than or equal

Description:

Determines whether one string is lexically less than or equal to another string, where the two strings are interpreted as containing ASCII character codes. If the String A and String B are not the same length, the shorter is compared as if spaces were appended to it to form a value that has the same length as the longer.

In general, the lexical comparison intrinsics LGE, LGT, LLE, and LLT differ from the corresponding intrinsic operators .GE., .GT., .LE., and .LT., in that the latter use the processor’s character ordering (which is not ASCII on some targets), whereas the former always use the ASCII ordering.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = LLE(STRING_A, STRING_B)

Arguments:

STRING A Shall be of default CHARACTER type.
STRING B Shall be of default CHARACTER type.

Return value:

Returns .TRUE. if STRING_A <= STRING_B, and .FALSE. otherwise, based on

the ASCII ordering.

Specific names:

Name  Argument Return type  Standard

LLE(STRING_A,STRING_B)  CHARACTER LOGICAL  Fortran 77 and later

See also: Section 8.170 [LGE], page 222, Section 8.171 [LGT], page 223, Section 8.174 [LLT], page 225,

8.174 LLT — Lexical less than

Description:

Determines whether one string is lexically less than another string, where the two strings are interpreted as containing ASCII character codes. If the String A and String B are not the same length, the shorter is compared as if spaces were appended to it to form a value that has the same length as the longer. In general, the lexical comparison intrinsics LGE, LGT, LLE, and LLT differ from the corresponding intrinsic operators .GE., .GT., .LE., and .LT., in that the latter use the processor’s character ordering (which is not ASCII on some targets), whereas the former always use the ASCII ordering.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = LLT(STRING_A, STRING_B)

Arguments:

STRING A Shall be of default CHARACTER type.
STRING B Shall be of default CHARACTER type.

Return value:

Returns .TRUE. if STRING_A < STRING_B, and .FALSE. otherwise, based on the ASCII ordering.

Specific names:

Name  Argument Return type  Standard

LLT(STRING_A,STRING_B)  CHARACTER LOGICAL  Fortran 77 and later

See also: Section 8.170 [LGE], page 222, Section 8.171 [LGT], page 223, Section 8.173 [LLE], page 224,

8.175 LNBLNK — Index of the last non-blank character in a string

Description:

Returns the length of a character string, ignoring any trailing blanks. This is identical to the standard LEN_TRIM intrinsic, and is only included for backwards compatibility.

Standard: GNU extension

Class: Elemental function

Syntax: RESULT = LNBLNK(STRING)

Arguments:

STRING  Shall be a scalar of type CHARACTER, with INTENT(IN)

Return value:

The return value is of INTEGER(kind=4) type.

See also: Section 8.148 [INDEX intrinsic], page 209, Section 8.169 [LEN TRIM], page 222,

8.176 LOC — Returns the address of a variable

Description:

LOC(X) returns the address of X as an integer.

Standard: GNU extension Class: Inquiry function

Syntax: RESULT = LOC(X)

Arguments:

X  Variable of any type.

Return value:

The return value is of type INTEGER, with a KIND corresponding to the size (in bytes) of a memory address on the target machine.

Example:

                  program test_loc
                    integer :: i
                    real :: r
                    i = loc(r)
                    print *, i
                  end program test_loc

8.177 LOG — Natural logarithm function

Description:

LOG(X) computes the natural logarithm of X, i.e. the logarithm to the base e.

Standard: Fortran 77 and later, has GNU extensions Class: Elemental function

Syntax: RESULT = LOG(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value is of type REAL or COMPLEX. The kind type parameter is the same as X. If X is COMPLEX, the imaginary part ω is in the range −π < ω ≤ π.

Example:

                  program test_log
                    real(8) :: x = 2.7182818284590451_8
                    complex :: z = (1.0, 2.0)
                    x = log(x)    ! will yield (approximately) 1
                    z = log(z)
                  end program test_log

Specific names:

Name  Argument  Return type Standard

ALOG(X)        REAL(4) X      REAL(4)  Fortran 77 or later
DLOG(X)        REAL(8) X      REAL(8)  Fortran 77 or later
CLOG(X)        COMPLEX(4) X   COMPLEX(4) Fortran 77 or later
ZLOG(X)        COMPLEX(8) X   COMPLEX(8) GNU extension
CDLOG(X)       COMPLEX(8) X   COMPLEX(8) GNU extension

8.178 LOG10 — Base 10 logarithm function

Description:

LOG10(X) computes the base 10 logarithm of X.

Standard: Fortran 77 and later Class: Elemental function

Syntax: RESULT = LOG10(X)

Arguments:

X  The type shall be REAL.

Return value:

The return value is of type REAL or COMPLEX. The kind type parameter is the same as X.

Example:

      program test_log10
        real(8) :: x = 10.0_8
        x = log10(x)
      end program test_log10

Specific names:

Name  Argument  Return type Standard

ALOG10(X)      REAL(4) X      REAL(4)  Fortran 77 and later
DLOG10(X)      REAL(8) X      REAL(8)  Fortran 77 and later

8.179 LOG_GAMMA — Logarithm of the Gamma function

Description:

LOG_GAMMA(X) computes the natural logarithm of the absolute value of the Gamma (Γ) function.

Standard: Fortran 2008 and later Class: Elemental function

Syntax: X = LOG_GAMMA(X)

Arguments:

X Shall be of type REAL and neither zero nor a negative integer.

Return value:

The return value is of type REAL of the same kind as X.

Example:

      program test_log_gamma
        real :: x = 1.0
        x = lgamma(x) ! returns 0.0
      end program test_log_gamma

Specific names:

Name  Argument  Return type Standard

LGAMMA(X)      REAL(4) X      REAL(4)  GNU extension
ALGAMA(X)      REAL(4) X      REAL(4)  GNU extension
DLGAMA(X)      REAL(8) X      REAL(8)  GNU extension

See also: Gamma function:

Section 8.119 [GAMMA], page 191,

8.180 LOGICAL — Convert to logical type

Description:

Converts one kind of LOGICAL variable to another.

Standard: Fortran 90 and later

Class: Elemental function

Syntax: RESULT = LOGICAL(L [, KIND])

Arguments:

L  The type shall be LOGICAL.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is a LOGICAL value equal to L, with a kind corresponding to KIND, or of the default logical kind if KIND is not given.

See also: Section 8.149 [INT], page 210, Section 8.228 [REAL], page 257, Section 8.62 [CMPLX], page 152,

8.181 LSHIFT — Left shift bits

Description:

LSHIFT returns a value corresponding to I with all of the bits shifted left by SHIFT places. SHIFT shall be nonnegative and less than or equal to BIT_ SIZE(I), otherwise the result value is undefined. Bits shifted out from the left end are lost; zeros are shifted in from the opposite end. This function has been superseded by the ISHFT intrinsic, which is standard in Fortran 95 and later, and the SHIFTL intrinsic, which is standard in Fortran 2008 and later.

Standard: GNU extension

Class: Elemental function

Syntax: RESULT = LSHIFT(I, SHIFT)

Arguments:

I  The type shall be INTEGER.
SHIFT  The type shall be INTEGER.

Return value:

The return value is of type INTEGER and of the same kind as I.

See also: Section 8.159 [ISHFT], page 216, Section 8.160 [ISHFTC], page 217, Section 8.233 [RSHIFT], page 260, Section 8.244 [SHIFTA], page 267, Section 8.245 [SHIFTL], page 267, Section 8.246 [SHIFTR], page 268,

8.182 LSTAT — Get file status

Description:

LSTAT is identical to Section 8.259 [STAT], page 275, except that if path is a symbolic link, then the link itself is statted, not the file that it refers to.

The elements in VALUES are the same as described by Section 8.259 [STAT], page 275.

This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL LSTAT(NAME, VALUES [, STATUS])
STATUS = LSTAT(NAME, VALUES)

Arguments:

NAME The type shall be CHARACTER of the default kind, a valid path within the file system.

VALUES  The type shall be INTEGER(4), DIMENSION(13).

STATUS (Optional) status flag of type INTEGER(4). Returns 0 on success and a system specific error code otherwise.

Example: See Section 8.259 [STAT], page 275, for an example.

See also: To stat an open file:

Section 8.117 [FSTAT], page 190, To stat a file: Section 8.259 [STAT], page 275,

8.183 LTIME — Convert time to local time info

Description:

Given a system time value TIME (as provided by the Section 8.269 [TIME], page 282, intrinsic), fills VALUES with values extracted from it appropriate to the local time zone using localtime(3). This intrinsic routine is provided for backwards compatibility with GNU Fortran 77. In new code, programmers should consider the use of the Section 8.82 [DATE AND TIME], page 166, intrinsic defined by the Fortran 95 standard.

Standard: GNU extension

Class: Subroutine

Syntax: CALL LTIME(TIME, VALUES)

Arguments:

TIME An INTEGER scalar expression corresponding to a system time, with INTENT(IN).

VALUES A default INTEGER array with 9 elements, with

            INTENT(OUT).

Return value:

The elements of VALUES are assigned as follows:

1. Seconds after the minute, range 0–59 or 0–61 to allow for leap seconds 2. Minutes after the hour, range 0–59

3. Hours past midnight, range 0–23

4. Day of month, range 1–31

5. Number of months since January, range 0–11

6. Years since 1900

7. Number of days since Sunday, range 0–6

8. Days since January 1, range 0–365

9. Daylight savings indicator: positive if daylight savings is in effect, zero if not, and negative if the information is not available.

See also: Section 8.82 [DATE AND TIME], page 166, Section 8.81 [CTIME], page 165, Section 8.131 [GMTIME], page 198, Section 8.269 [TIME], page 282, Section 8.270 [TIME8], page 283,

8.184 MALLOC — Allocate dynamic memory

Description:

MALLOC(SIZE) allocates SIZE bytes of dynamic memory and returns the address of the allocated memory. The MALLOC intrinsic is an extension intended to be used with Cray pointers, and is provided in GNU Fortran to allow the user to compile legacy code. For new code using Fortran 95 pointers, the memory allocation intrinsic is ALLOCATE.

Standard: GNU extension

Class: Function

Syntax: PTR = MALLOC(SIZE)

Arguments:

SIZE  The type shall be INTEGER.

Return value:

The return value is of type INTEGER(K), with K such that variables of type INTEGER(K) have the same size as C pointers (sizeof(void *)).

Example: The following example demonstrates the use of MALLOC and FREE with Cray pointers.

                  program test_malloc
                    implicit none
                    integer i
                    real*8 x(*), z
                    pointer(ptr_x,x)
                    ptr_x = malloc(20*8)
                    do i = 1, 20
                      x(i) = sqrt(1.0d0 / i)
                    end do
                    z = 0
                    do i = 1, 20
                      z = z + x(i)
                      print *, z
                    end do
                    call free(ptr_x)
                  end program test_malloc

See also: Section 8.115 [FREE], page 188,

8.185 MASKL — Left justified mask

Description:

MASKL(I[, KIND]) has its leftmost I bits set to 1, and the remaining bits set to 0.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = MASKL(I[, KIND])

Arguments:

I  Shall be of type INTEGER.
KIND  Shall be a scalar constant expression of type INTEGER.

Return value:

The return value is of type INTEGER. If KIND is present, it specifies the kind value of the return type; otherwise, it is of the default integer kind.

See also: Section 8.186 [MASKR], page 232,

8.186 MASKR — Right justified mask

Description:

MASKL(I[, KIND]) has its rightmost I bits set to 1, and the remaining bits set to 0.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = MASKR(I[, KIND])

Arguments:

I  Shall be of type INTEGER.
KIND  Shall be a scalar constant expression of type INTEGER.

Return value:

The return value is of type INTEGER. If KIND is present, it specifies the kind value of the return type; otherwise, it is of the default integer kind.

See also: Section 8.185 [MASKL], page 231,

8.187 MATMUL — matrix multiplication

Description:

Performs a matrix multiplication on numeric or logical arguments.

Standard: Fortran 90 and later

Class: Transformational function

Syntax: RESULT = MATMUL(MATRIX_A, MATRIX_B)

Arguments:

MATRIX A

                      An array of INTEGER, REAL, COMPLEX, or LOGICAL

type, with a rank of one or two.

MATRIX B An array of INTEGER, REAL, or COMPLEX type if MA-TRIX A is of a numeric type; otherwise, an array of LOGICAL type. The rank shall be one or two, and the first (or only) dimension of MATRIX B shall be equal to the last (or only) dimension of MATRIX A. MATRIX A and MATRIX B shall not both be rank one arrays.

Return value:

The matrix product of MATRIX A and MATRIX B. The type and kind of the result follow the usual type and kind promotion rules, as for the * or .AND. operators.

8.188 MAX — Maximum value of an argument list

Description:

Returns the argument with the largest (most positive) value.

Standard: Fortran 77 and later Class: Elemental function

Syntax: RESULT = MAX(A1, A2 [, A3 [, ...]])

Arguments:

A1  The type shall be INTEGER or REAL.

A2, A3, ...

An expression of the same type and kind as A1. (As a GNU extension, arguments of different kinds are permitted.)

Return value:

The return value corresponds to the maximum value among the arguments, and has the same type and kind as the first argument.

Specific names:

Name  Argument  Return type Standard

          MAX0(A1)       INTEGER(4) A1  INTEGER(4) Fortran 77 and later
          AMAX0(A1)      INTEGER(4) A1  REAL(MAX(X)) Fortran 77 and later
          MAX1(A1)       REAL A1        INT(MAX(X)) Fortran 77 and later
          AMAX1(A1)      REAL(4) A1     REAL(4)  Fortran 77 and later
          DMAX1(A1)      REAL(8) A1     REAL(8)  Fortran 77 and later

See also: Section 8.190 [MAXLOC], page 234, Section 8.191 [MAXVAL], page 235, Section 8.196 [MIN], page 237,

8.189 MAXEXPONENT — Maximum exponent of a real kind

Description:

MAXEXPONENT(X) returns the maximum exponent in the model of the type of X.

Standard: Fortran 90 and later Class: Inquiry function

Syntax: RESULT = MAXEXPONENT(X)

Arguments:

X  Shall be of type REAL.

Return value:

The return value is of type INTEGER and of the default integer kind.

Example:

                  program exponents
                    real(kind=4) :: x
                    real(kind=8) :: y
                    print *, minexponent(x), maxexponent(x)
                    print *, minexponent(y), maxexponent(y)
                  end program exponents

8.190 MAXLOC — Location of the maximum value within an array

Description:

Determines the location of the element in the array with the maximum value, or, if the DIM argument is supplied, determines the locations of the maximum element along each row of the array in the DIM direction. If MASK is present, only the elements for which MASK is .TRUE. are considered. If more than one element in the array has the maximum value, the location returned is that of the first such element in array element order if the BACK is not present, or is false; if BACK is true, the location returned is that of the last such element. If the array has zero size, or all of the elements of MASK are .FALSE., then the result is an array of zeroes. Similarly, if DIM is supplied and all of the elements of MASK along a given row are zero, the result value for that row is zero.

Standard: Fortran 95 and later; ARRAY of CHARACTER and the KIND argument are available in Fortran 2003 and later. The BACK argument is available in Fortran 2008 and later.

Class: Transformational function

Syntax:

RESULT = MAXLOC(ARRAY, DIM [, MASK] [,KIND] [,BACK])
RESULT = MAXLOC(ARRAY [, MASK] [,KIND] [,BACK])

Arguments:

ARRAY  Shall be an array of type INTEGER or REAL.

DIM (Optional) Shall be a scalar of type INTEGER, with a value between one and the rank of ARRAY, inclusive. It may not be an optional dummy argument.

MASK Shall be of type LOGICAL, and conformable with ARRAY.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

BACK  (Optional) A scalar of type LOGICAL.

Return value:

If DIM is absent, the result is a rank-one array with a length equal to the rank of ARRAY. If DIM is present, the result is an array with a rank one less than the rank of ARRAY, and a size corresponding to the size of ARRAY with the DIM dimension removed. If DIM is present and ARRAY has a rank of one, the result is a scalar. If the optional argument KIND is present, the result is an integer of kind KIND, otherwise it is of default kind.

See also: Section 8.108 [FINDLOC], page 183, Section 8.188 [MAX], page 232, Section 8.191 [MAXVAL], page 235,

8.191 MAXVAL — Maximum value of an array

Description:

Determines the maximum value of the elements in an array value, or, if the DIM argument is supplied, determines the maximum value along each row of the array in the DIM direction. If MASK is present, only the elements for which MASK is .TRUE. are considered. If the array has zero size, or all of the elements of MASK are .FALSE., then the result is -HUGE(ARRAY) if ARRAY is numeric, or a string of nulls if ARRAY is of character type.

Standard: Fortran 90 and later

Class: Transformational function

Syntax:

RESULT = MAXVAL(ARRAY, DIM [, MASK])
RESULT = MAXVAL(ARRAY [, MASK])

Arguments:

ARRAY  Shall be an array of type INTEGER or REAL.

DIM (Optional) Shall be a scalar of type INTEGER, with a value between one and the rank of ARRAY, inclusive. It may not be an optional dummy argument.

MASK (Optional) Shall be of type LOGICAL, and conformable with ARRAY.

Return value:

If DIM is absent, or if ARRAY has a rank of one, the result is a scalar. If DIM is present, the result is an array with a rank one less than the rank of ARRAY, and a size corresponding to the size of ARRAY with the DIM dimension removed. In all cases, the result is of the same type and kind as ARRAY.

See also: Section 8.188 [MAX], page 232, Section 8.190 [MAXLOC], page 234,

8.192 MCLOCK — Time function

Description:

Returns the number of clock ticks since the start of the process, based on the function clock(3) in the C standard library. This intrinsic is not fully portable, such as to systems with 32-bit INTEGER types but supporting times wider than 32 bits. Therefore, the values returned by this intrinsic might be, or become, negative, or numerically less than previous values, during a single run of the compiled program.

Standard: GNU extension

Class: Function

Syntax: RESULT = MCLOCK()

Return value:

The return value is a scalar of type INTEGER(4), equal to the number of clock ticks since the start of the process, or -1 if the system does not support

clock(3).

See also: Section 8.81 [CTIME], page 165, Section 8.131 [GMTIME], page 198, Section 8.183 [LTIME], page 230, Section 8.192 [MCLOCK], page 235, Section 8.269 [TIME], page 282,

8.193 MCLOCK8 — Time function (64-bit)

Description:

Returns the number of clock ticks since the start of the process, based on the function clock(3) in the C standard library. Warning: this intrinsic does not increase the range of the timing values over that returned by clock(3). On a system with a 32-bit clock(3), MCLOCK8 will return a 32-bit value, even though it is converted to a 64-bit INTEGER(8) value. That means overflows of the 32-bit value can still occur. Therefore, the values returned by this intrinsic might be or become negative or numerically less than previous values during a single run of the compiled program.

Standard: GNU extension

Class: Function

Syntax: RESULT = MCLOCK8()

Return value:

The return value is a scalar of type INTEGER(8), equal to the number of clock ticks since the start of the process, or -1 if the system does not support

clock(3).

See also: Section 8.81 [CTIME], page 165, Section 8.131 [GMTIME], page 198, Section 8.183 [LTIME], page 230, Section 8.192 [MCLOCK], page 235, Section 8.270 [TIME8], page 283,

8.194 MERGE — Merge variables

Description:

Select values from two arrays according to a logical mask. The result is equal to TSOURCE if MASK is .TRUE., or equal to FSOURCE if it is .FALSE..

Standard: Fortran 90 and later

Class: Elemental function

Syntax: RESULT = MERGE(TSOURCE, FSOURCE, MASK)

Arguments:

TSOURCE May be of any type.

FSOURCE Shall be of the same type and type parameters as TSOURCE.

MASK  Shall be of type LOGICAL.

Return value:

The result is of the same type and type parameters as TSOURCE.

8.195 MERGE_BITS — Merge of bits under mask

Description:

MERGE_BITS(I, J, MASK) merges the bits of I and J as determined by the mask. The i-th bit of the result is equal to the i-th bit of I if the i-th bit of MASK is 1; it is equal to the i-th bit of J otherwise.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = MERGE_BITS(I, J, MASK)

Arguments:

I  Shall be of type INTEGER or a boz-literal-constant.

J Shall be of type INTEGER with the same kind type parameter as I or a boz-literal-constant. I and J shall not both be boz-literal-constants.

MASK Shall be of type INTEGER or a boz-literal-constant and of the same kind as I.

Return value:

The result is of the same type and kind as I.

8.196 MIN — Minimum value of an argument list

Description:

Returns the argument with the smallest (most negative) value.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = MIN(A1, A2 [, A3, ...])

Arguments:

A1  The type shall be INTEGER or REAL.

A2, A3, ...

An expression of the same type and kind as A1. (As a GNU extension, arguments of different kinds are permitted.)

Return value:

The return value corresponds to the minimum value among the arguments, and has the same type and kind as the first argument.

Specific names:

Name  Argument  Return type Standard

MIN0(A1)       INTEGER(4) A1  INTEGER(4) Fortran 77 and later
AMIN0(A1)      INTEGER(4) A1  REAL(4)  Fortran 77 and later
MIN1(A1)       REAL A1        INTEGER(4) Fortran 77 and later
AMIN1(A1)      REAL(4) A1     REAL(4)  Fortran 77 and later
DMIN1(A1)      REAL(8) A1     REAL(8)  Fortran 77 and later

See also: Section 8.188 [MAX], page 232, Section 8.198 [MINLOC], page 238, Section 8.199 [MINVAL], page 239,

8.197 MINEXPONENT — Minimum exponent of a real kind

Description:

MINEXPONENT(X) returns the minimum exponent in the model of the type of X.

Standard: Fortran 90 and later

Class: Inquiry function

Syntax: RESULT = MINEXPONENT(X)

Arguments:

X  Shall be of type REAL.

Return value:

The return value is of type INTEGER and of the default integer kind.

Example: See MAXEXPONENT for an example.

8.198 MINLOC — Location of the minimum value within an array

Description:

Determines the location of the element in the array with the minimum value, or, if the DIM argument is supplied, determines the locations of the minimum element along each row of the array in the DIM direction. If MASK is present, only the elements for which MASK is .TRUE. are considered. If more than one element in the array has the minimum value, the location returned is that of the first such element in array element order if the BACK is not present, or is false; if BACK is true, the location returned is that of the last such element. If the array has zero size, or all of the elements of MASK are .FALSE., then the result is an array of zeroes. Similarly, if DIM is supplied and all of the elements of MASK along a given row are zero, the result value for that row is zero.

Standard: Fortran 90 and later; ARRAY of CHARACTER and the KIND argument are available in Fortran 2003 and later. The BACK argument is available in Fortran 2008 and later.

Class: Transformational function Syntax:

RESULT = MINLOC(ARRAY, DIM [, MASK] [,KIND] [,BACK])
RESULT = MINLOC(ARRAY [, MASK], [,KIND] [,BACK])

Arguments:

ARRAY Shall be an array of type INTEGER, REAL or

            CHARACTER.

DIM (Optional) Shall be a scalar of type INTEGER, with a value between one and the rank of ARRAY, inclusive. It may not be an optional dummy argument.

MASK Shall be of type LOGICAL, and conformable with ARRAY.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

BACK  (Optional) A scalar of type LOGICAL.

Return value:

If DIM is absent, the result is a rank-one array with a length equal to the rank of ARRAY. If DIM is present, the result is an array with a rank one less than the rank of ARRAY, and a size corresponding to the size of ARRAY with the DIM dimension removed. If DIM is present and ARRAY has a rank of one, the result is a scalar. If the optional argument KIND is present, the result is an integer of kind KIND, otherwise it is of default kind.

See also: Section 8.108 [FINDLOC], page 183, Section 8.196 [MIN], page 237, Section 8.199 [MINVAL], page 239,

8.199 MINVAL — Minimum value of an array

Description:

Determines the minimum value of the elements in an array value, or, if the DIM argument is supplied, determines the minimum value along each row of the array in the DIM direction. If MASK is present, only the elements for which MASK is .TRUE. are considered. If the array has zero size, or all of the elements of MASK are .FALSE., then the result is HUGE(ARRAY) if ARRAY is numeric, or a string of CHAR(255) characters if ARRAY is of character type.

Standard: Fortran 90 and later

Class: Transformational function

Syntax:

RESULT = MINVAL(ARRAY, DIM [, MASK])
RESULT = MINVAL(ARRAY [, MASK])

Arguments:

ARRAY  Shall be an array of type INTEGER or REAL.

DIM (Optional) Shall be a scalar of type INTEGER, with a value between one and the rank of ARRAY, inclusive. It may not be an optional dummy argument.

MASK Shall be of type LOGICAL, and conformable with ARRAY.

Return value:

If DIM is absent, or if ARRAY has a rank of one, the result is a scalar. If DIM is present, the result is an array with a rank one less than the rank of ARRAY, and a size corresponding to the size of ARRAY with the DIM dimension removed. In all cases, the result is of the same type and kind as ARRAY.

See also: Section 8.196 [MIN], page 237, Section 8.198 [MINLOC], page 238,

8.200 MOD — Remainder function

Description:

MOD(A,P) computes the remainder of the division of A by P.

Standard: Fortran 77 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = MOD(A, P)

Arguments:

A  Shall be a scalar of type INTEGER or REAL.

P Shall be a scalar of the same type and kind as A and not equal to zero. (As a GNU extension, arguments of different kinds are permitted.)

Return value:

The return value is the result of A - (INT(A/P) * P). The type and kind of the return value is the same as that of the arguments. The returned value has the same sign as A and a magnitude less than the magnitude of P. (As a GNU extension, kind is the largest kind of the actual arguments.)

Example:

      program test_mod
        print *, mod(17,3)
        print *, mod(17.5,5.5)
        print *, mod(17.5d0,5.5)
        print *, mod(17.5,5.5d0)
        print *, mod(-17,3)
        print *, mod(-17.5,5.5)
        print *, mod(-17.5d0,5.5)
        print *, mod(-17.5,5.5d0)
        print *, mod(17,-3)
        print *, mod(17.5,-5.5)
        print *, mod(17.5d0,-5.5)
        print *, mod(17.5,-5.5d0)
      end program test_mod

Specific names:

Name  Arguments  Return type Standard

MOD(A,P)       INTEGER A,P      INTEGER  Fortran 77 and later

AMOD(A,P)      REAL(4) A,P      REAL(4)  Fortran 77 and later
DMOD(A,P)      REAL(8) A,P      REAL(8)  Fortran 77 and later
BMOD(A,P)      INTEGER(1) A,P   INTEGER(1) GNU extension
IMOD(A,P)      INTEGER(2) A,P   INTEGER(2) GNU extension
JMOD(A,P)      INTEGER(4) A,P   INTEGER(4) GNU extension
KMOD(A,P)      INTEGER(8) A,P   INTEGER(8) GNU extension

See also: Section 8.201 [MODULO], page 241,

8.201 MODULO — Modulo function

Description:

MODULO(A,P) computes the A modulo P.

Standard: Fortran 95 and later Class: Elemental function

Syntax: RESULT = MODULO(A, P)

Arguments:

A  Shall be a scalar of type INTEGER or REAL.

P Shall be a scalar of the same type and kind as A. It shall not be zero. (As a GNU extension, arguments of different kinds are permitted.)

Return value:

The type and kind of the result are those of the arguments. (As a GNU extension, kind is the largest kind of the actual arguments.) If A and P are of type INTEGER:

MODULO(A,P) has the value R such that A=Q*P+R, where Q is an integer and R is between 0 (inclusive) and P (exclusive).

If A and P are of type REAL:

          MODULO(A,P) has the value of A - FLOOR (A / P) * P.

The returned value has the same sign as P and a magnitude less than the magnitude of P.

Example:

      program test_modulo
        print *, modulo(17,3)
        print *, modulo(17.5,5.5)
        print *, modulo(-17,3)
        print *, modulo(-17.5,5.5)
        print *, modulo(17,-3)
        print *, modulo(17.5,-5.5)
      end program

See also: Section 8.200 [MOD], page 240,

8.202 MOVE_ALLOC — Move allocation from one object to another

Description:

MOVE_ALLOC(FROM, TO) moves the allocation from FROM to TO. FROM will become deallocated in the process.

Standard: Fortran 2003 and later

Class: Pure subroutine

Syntax: CALL MOVE_ALLOC(FROM, TO)

Arguments:

FROM

            ALLOCATABLE, INTENT(INOUT), may be of any type

and kind.

TO

            ALLOCATABLE, INTENT(OUT), shall be of the same

type, kind and rank as FROM.

Return value:

None

Example:

      program test_move_alloc
          integer, allocatable :: a(:), b(:)
          allocate(a(3))
          a = [ 1, 2, 3 ]
          call move_alloc(a, b)
          print *, allocated(a), allocated(b)
          print *, b
      end program test_move_alloc

8.203 MVBITS — Move bits from one integer to another

Description:

Moves LEN bits from positions FROMPOS through FROMPOS+LEN-1 of FROM to positions TOPOS through TOPOS+LEN-1 of TO. The portion of argument TO not affected by the movement of bits is unchanged. The values of FROMPOS+LEN-

1 and TOPOS+LEN-1 must be less than BIT_SIZE(FROM).

Standard: Fortran 90 and later, has overloads that are GNU extensions

Class: Elemental subroutine

Syntax: CALL MVBITS(FROM, FROMPOS, LEN, TO, TOPOS)

Arguments:

FROM  The type shall be INTEGER.
FROMPOS The type shall be INTEGER.
LEN  The type shall be INTEGER.

TO The type shall be INTEGER, of the same kind as FROM.

TOPOS  The type shall be INTEGER.

Specific names:

Name  Argument  Return type Standard

MVBITS(A)      INTEGER A      INTEGER  Fortran 90 and later
BMVBITS(A)     INTEGER(1) A   INTEGER(1) GNU extension
IMVBITS(A)     INTEGER(2) A   INTEGER(2) GNU extension
JMVBITS(A)     INTEGER(4) A   INTEGER(4) GNU extension
KMVBITS(A)     INTEGER(8) A   INTEGER(8) GNU extension

See also: Section 8.140 [IBCLR], page 204, Section 8.142 [IBSET], page 206, Section 8.141 [IBITS], page 205, Section 8.137 [IAND], page 202, Section 8.152 [IOR], page 212, Section 8.145 [IEOR], page 208,

8.204 NEAREST — Nearest representable number

Description:

NEAREST(X, S) returns the processor-representable number nearest to X in the direction indicated by the sign of S.

Standard: Fortran 90 and later

Class: Elemental function

Syntax: RESULT = NEAREST(X, S)

Arguments:

X  Shall be of type REAL.
S  Shall be of type REAL and not equal to zero.

Return value:

The return value is of the same type as X. If S is positive, NEAREST returns the processor-representable number greater than X and nearest to it. If S is negative, NEAREST returns the processor-representable number smaller than X and nearest to it.

Example:

      program test_nearest
        real :: x, y
        x = nearest(42.0, 1.0)
        y = nearest(42.0, -1.0)
        write (*,"(3(G20.15))") x, y, x - y
      end program test_nearest

8.205 NEW_LINE — New line character

Description:

NEW_LINE(C) returns the new-line character.

Standard: Fortran 2003 and later

Class: Inquiry function

Syntax: RESULT = NEW_LINE(C)

Arguments:

C The argument shall be a scalar or array of the type

                      CHARACTER.

Return value:

Returns a CHARACTER scalar of length one with the new-line character of the same kind as parameter C.

Example:

                  program newline
                    implicit none
                    write(*,'(A)') 'This is record 1.'//NEW_LINE('A')//'This is record 2.'
                  end program newline

8.206 NINT — Nearest whole number

Description:

NINT(A) rounds its argument to the nearest whole number.

Standard: Fortran 77 and later, with KIND argument Fortran 90 and later

Class: Elemental function

Syntax: RESULT = NINT(A [, KIND])

Arguments:

A  The type of the argument shall be REAL.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

Returns A with the fractional portion of its magnitude eliminated by rounding to the nearest whole number and with its sign preserved, converted to an INTEGER of the default kind.

Example:

                  program test_nint
                    real(4) x4
                    real(8) x8
                    x4 = 1.234E0_4
                    x8 = 4.321_8
                    print *, nint(x4), idnint(x8)
                  end program test_nint

Specific names:

Name  Argument  Return Type Standard

          NINT(A)        REAL(4) A      INTEGER  Fortran 77 and later
          IDNINT(A)      REAL(8) A      INTEGER  Fortran 77 and later

See also: Section 8.58 [CEILING], page 149, Section 8.109 [FLOOR], page 184,

8.207 NORM2 — Euclidean vector norms

Description:

Calculates the Euclidean vector norm (L_{2} norm) of ARRAY along dimension DIM.

Standard: Fortran 2008 and later

Class: Transformational function

Syntax:

RESULT = NORM2(ARRAY[, DIM])

Arguments:

ARRAY  Shall be an array of type REAL

DIM (Optional) shall be a scalar of type INTEGER with a value in the range from 1 to n, where n equals the rank of ARRAY.

Return value:

The result is of the same type as ARRAY. If DIM is absent, a scalar with the square root of the sum of all elements in ARRAY squared is returned. Otherwise, an array of rank n−1, where n equals the rank of ARRAY, and a shape similar to that of ARRAY with dimension DIM dropped is returned.

Example:

      PROGRAM test_sum
        REAL :: x(5) = [ real :: 1, 2, 3, 4, 5 ]
        print *, NORM2(x)  ! = sqrt(55.) ~ 7.416
      END PROGRAM

8.208 NOT — Logical negation

Description:

NOT returns the bitwise Boolean inverse of I.

Standard: Fortran 90 and later, has overloads that are GNU extensions

Class: Elemental function

Syntax: RESULT = NOT(I)

Arguments:

I  The type shall be INTEGER.

Return value:

The return type is INTEGER, of the same kind as the argument.

Specific names:

Name  Argument  Return type Standard

NOT(A)         INTEGER A      INTEGER  Fortran 95 and later
BNOT(A)        INTEGER(1) A   INTEGER(1) GNU extension
INOT(A)        INTEGER(2) A   INTEGER(2) GNU extension
JNOT(A)        INTEGER(4) A   INTEGER(4) GNU extension
KNOT(A)        INTEGER(8) A   INTEGER(8) GNU extension

See also: Section 8.137 [IAND], page 202, Section 8.145 [IEOR], page 208, Section 8.152 [IOR], page 212, Section 8.141 [IBITS], page 205, Section 8.142 [IBSET], page 206, Section 8.140 [IBCLR], page 204,

8.209 NULL — Function that returns an disassociated pointer

Description:

Returns a disassociated pointer. If MOLD is present, a disassociated pointer of the same type is returned, otherwise the type is determined by context. In Fortran 95, MOLD is optional. Please note that Fortran 2003 includes cases where it is required.

Standard: Fortran 95 and later

Class: Transformational function

Syntax: PTR => NULL([MOLD])

Arguments:

MOLD (Optional) shall be a pointer of any association status and of any type.

Return value:

A disassociated pointer.

Example:

      REAL, POINTER, DIMENSION(:) :: VEC => NULL ()

See also: Section 8.22 [ASSOCIATED], page 124,

8.210 NUM_IMAGES — Function that returns the number of images

Description:

Returns the number of images.

Standard: Fortran 2008 and later. With DISTANCE or FAILED argument, Technical Specification (TS) 18508 or later

Class: Transformational function

Syntax: RESULT = NUM_IMAGES(DISTANCE, FAILED)

Arguments:

DISTANCE (optional, intent(in)) Nonnegative scalar integer
FAILED  (optional, intent(in)) Scalar logical expression

Return value:

Scalar default-kind integer. If DISTANCE is not present or has value 0, the number of images in the current team is returned. For values smaller or equal distance to the initial team, it returns the number of images index on the ancestor team which has a distance of DISTANCE from the invoking team. If DISTANCE is larger than the distance to the initial team, the number of images of the initial team is returned. If FAILED is not present the total number of images is returned; if it has the value .TRUE., the number of failed images is returned, otherwise, the number of images which do have not the failed status.

Example:

      INTEGER :: value[*]
      INTEGER :: i
      value = THIS_IMAGE()
      SYNC ALL
      IF (THIS_IMAGE() == 1) THEN
        DO i = 1, NUM_IMAGES()
          WRITE(*,'(2(a,i0))') 'value[', i, '] is ', value[i]
        END DO
      END IF

See also: Section 8.268 [THIS IMAGE], page 281, Section 8.147 [IMAGE INDEX], page 209,

8.211 OR — Bitwise logical OR

Description:

Bitwise logical OR. This intrinsic routine is provided for backwards compatibility with GNU Fortran 77. For integer arguments, programmers should consider the use of the Section 8.152 [IOR], page 212, intrinsic defined by the Fortran standard.

Standard: GNU extension

Class: Function

Syntax: RESULT = OR(I, J)

Arguments:

I The type shall be either a scalar INTEGER type or a scalar LOGICAL type or a boz-literal-constant.

J The type shall be the same as the type of I or a bozliteral-constant. I and J shall not both be boz-literal-constants. If either I and J is a boz-literal-constant, then the other argument must be a scalar INTEGER.

Return value:

The return type is either a scalar INTEGER or a scalar LOGICAL. If the kind type parameters differ, then the smaller kind type is implicitly converted to larger kind, and the return has the larger kind. A boz-literal-constant is converted to an INTEGER with the kind type parameter of the other argument as-if a call to Section 8.149 [INT], page 210, occurred.

Example:

      PROGRAM test_or
        LOGICAL :: T = .TRUE., F = .FALSE.

        INTEGER :: a, b
        DATA a / Z'F' /, b / Z'3' /
        WRITE (*,*) OR(T, T), OR(T, F), OR(F, T), OR(F, F)
        WRITE (*,*) OR(a, b)
      END PROGRAM

See also: Fortran 95 elemental function:

Section 8.152 [IOR], page 212,

8.212 PACK — Pack an array into an array of rank one

Description:

Stores the elements of ARRAY in an array of rank one. The beginning of the resulting array is made up of elements whose MASK equals TRUE. Afterwards, positions are filled with elements taken from VECTOR.

Standard: Fortran 90 and later

Class: Transformational function

Syntax: RESULT = PACK(ARRAY, MASK[,VECTOR])

Arguments:

ARRAY  Shall be an array of any type.

MASK Shall be an array of type LOGICAL and of the same size as ARRAY. Alternatively, it may be a LOGICAL scalar.

VECTOR (Optional) shall be an array of the same type as AR-RAY and of rank one. If present, the number of elements in VECTOR shall be equal to or greater than the number of true elements in MASK. If MASK is scalar, the number of elements in VECTOR shall be equal to or greater than the number of elements in ARRAY.

Return value:

The result is an array of rank one and the same type as that of ARRAY. If VECTOR is present, the result size is that of VECTOR, the number of TRUE values in MASK otherwise.

Example: Gathering nonzero elements from an array:

      PROGRAM test_pack_1
        INTEGER :: m(6)
        m = (/ 1, 0, 0, 0, 5, 0 /)
        WRITE(*, FMT="(6(I0, ' '))") pack(m, m /= 0)  ! "1 5"
      END PROGRAM

Gathering nonzero elements from an array and appending elements from VEC-TOR:

      PROGRAM test_pack_2
        INTEGER :: m(4)
        m = (/ 1, 0, 0, 2 /)
        ! The following results in "1 2 3 4"

                    WRITE(*, FMT="(4(I0, ' '))") pack(m, m /= 0, (/ 0, 0, 3, 4 /))
                  END PROGRAM

See also: Section 8.281 [UNPACK], page 289,

8.213 PARITY — Reduction with exclusive OR

Description:

Calculates the parity, i.e. the reduction using .XOR., of MASK along dimension DIM.

Standard: Fortran 2008 and later

Class: Transformational function

Syntax:

          RESULT = PARITY(MASK[, DIM])

Arguments:

MASK  Shall be an array of type LOGICAL

DIM (Optional) shall be a scalar of type INTEGER with a value in the range from 1 to n, where n equals the rank of MASK.

Return value:

The result is of the same type as MASK.

If DIM is absent, a scalar with the parity of all elements in MASK is returned, i.e. true if an odd number of elements is .true. and false otherwise. If DIM is present, an array of rank n − 1, where n equals the rank of ARRAY, and a shape similar to that of MASK with dimension DIM dropped is returned.

Example:

                  PROGRAM test_sum
                    LOGICAL :: x(2) = [ .true., .false. ]
                    print *, PARITY(x) ! prints "T" (true).
                  END PROGRAM

8.214 PERROR — Print system error message

Description:

Prints (on the C stderr stream) a newline-terminated error message corresponding to the last system error. This is prefixed by STRING, a colon and a space. See perror(3).

Standard: GNU extension

Class: Subroutine

Syntax: CALL PERROR(STRING)

Arguments:

STRING  A scalar of type CHARACTER and of the default kind.

See also: Section 8.146 [IERRNO], page 209,

8.215 POPCNT — Number of bits set

Description:

POPCNT(I) returns the number of bits set (’1’ bits) in the binary representation of I.

Standard: Fortran 2008 and later Class: Elemental function

Syntax: RESULT = POPCNT(I)

Arguments:

I  Shall be of type INTEGER.

Return value:

The return value is of type INTEGER and of the default integer kind.

Example:

                  program test_population
                    print *, popcnt(127),       poppar(127)
                    print *, popcnt(huge(0_4)), poppar(huge(0_4))
                    print *, popcnt(huge(0_8)), poppar(huge(0_8))
                  end program test_population

See also: Section 8.216 [POPPAR], page 250, Section 8.167 [LEADZ], page 220, Section 8.272 [TRAILZ], page 284,

8.216 POPPAR — Parity of the number of bits set

Description:

POPPAR(I) returns parity of the integer I, i.e. the parity of the number of bits set (’1’ bits) in the binary representation of I. It is equal to 0 if I has an even number of bits set, and 1 for an odd number of ’1’ bits.

Standard: Fortran 2008 and later Class: Elemental function

Syntax: RESULT = POPPAR(I)

Arguments:

I  Shall be of type INTEGER.

Return value:

The return value is of type INTEGER and of the default integer kind.

Example:

                  program test_population
                    print *, popcnt(127),       poppar(127)
                    print *, popcnt(huge(0_4)), poppar(huge(0_4))
                    print *, popcnt(huge(0_8)), poppar(huge(0_8))
                  end program test_population

See also: Section 8.215 [POPCNT], page 250, Section 8.167 [LEADZ], page 220, Section 8.272 [TRAILZ], page 284,

8.217 PRECISION — Decimal precision of a real kind

Description:

PRECISION(X) returns the decimal precision in the model of the type of X.

Standard: Fortran 90 and later

Class: Inquiry function

Syntax: RESULT = PRECISION(X)

Arguments:

X Shall be of type REAL or COMPLEX. It may be scalar or valued.

Return value:

The return value is of type INTEGER and of the default integer kind.

Example:

                  program prec_and_range
                    real(kind=4) :: x(2)
                    complex(kind=8) :: y
                    print *, precision(x), range(x)
                    print *, precision(y), range(y)
                  end program prec_and_range

See also: Section 8.241 [SELECTED REAL KIND], page 264, Section 8.226 [RANGE], page 256,

8.218 PRESENT — Determine whether an optional dummy argument is specified

Description:

Determines whether an optional dummy argument is present.

Standard: Fortran 90 and later

Class: Inquiry function

Syntax: RESULT = PRESENT(A)

Arguments:

A May be of any type and may be a pointer, scalar or array value, or a dummy procedure. It shall be the name of an optional dummy argument accessible within the current subroutine or function.

Return value:

Returns either TRUE if the optional argument A is present, or FALSE otherwise.

Example:

                  PROGRAM test_present
                    WRITE(*,*) f(), f(42)      ! "F T"
                  CONTAINS
                    LOGICAL FUNCTION f(x)

                      INTEGER, INTENT(IN), OPTIONAL :: x
                      f = PRESENT(x)
                    END FUNCTION
                  END PROGRAM

8.219 PRODUCT — Product of array elements

Description:

Multiplies the elements of ARRAY along dimension DIM if the corresponding element in MASK is TRUE.

Standard: Fortran 90 and later

Class: Transformational function

Syntax:

          RESULT = PRODUCT(ARRAY[, MASK])
          RESULT = PRODUCT(ARRAY, DIM[, MASK])

Arguments:

ARRAY  Shall be an array of type INTEGER, REAL or COMPLEX.

DIM (Optional) shall be a scalar of type INTEGER with a value in the range from 1 to n, where n equals the rank of ARRAY.

MASK (Optional) shall be of type LOGICAL and either be a scalar or an array of the same shape as ARRAY.

Return value:

The result is of the same type as ARRAY.

If DIM is absent, a scalar with the product of all elements in ARRAY is returned. Otherwise, an array of rank n-1, where n equals the rank of ARRAY, and a shape similar to that of ARRAY with dimension DIM dropped is returned.

Example:

                  PROGRAM test_product
                    INTEGER :: x(5) = (/ 1, 2, 3, 4 ,5 /)
                    print *, PRODUCT(x)                    ! all elements, product = 120
                    print *, PRODUCT(x, MASK=MOD(x, 2)==1) ! odd elements, product = 15
                  END PROGRAM

See also: Section 8.261 [SUM], page 277,

8.220 RADIX — Base of a model number

Description:

RADIX(X) returns the base of the model representing the entity X.

Standard: Fortran 90 and later

Class: Inquiry function

Syntax: RESULT = RADIX(X)

Arguments:

X  Shall be of type INTEGER or REAL

Return value:

The return value is a scalar of type INTEGER and of the default integer kind.

Example:

                  program test_radix
                    print *, "The radix for the default integer kind is", radix(0)
                    print *, "The radix for the default real kind is", radix(0.0)
                  end program test_radix

See also: Section 8.241 [SELECTED REAL KIND], page 264,

8.221 RAN — Real pseudo-random number

Description:

For compatibility with HP FORTRAN 77/iX, the RAN intrinsic is provided as an alias for RAND. See Section 8.222 [RAND], page 253, for complete documentation.

Standard: GNU extension

Class: Function

See also: Section 8.222 [RAND], page 253, Section 8.224 [RANDOM NUMBER], page 255,

8.222 RAND — Real pseudo-random number

Description:

RAND(FLAG) returns a pseudo-random number from a uniform distribution between 0 and 1. If FLAG is 0, the next number in the current sequence is returned; if FLAG is 1, the generator is restarted by CALL SRAND(0); if FLAG has any other value, it is used as a new seed with SRAND. This intrinsic routine is provided for backwards compatibility with GNU Fortran 77. It implements a simple modulo generator as provided by g77. For new code, one should consider the use of Section 8.224 [RANDOM NUMBER], page 255, as it implements a superior algorithm.

Standard: GNU extension

Class: Function

Syntax: RESULT = RAND(I)

Arguments:

I  Shall be a scalar INTEGER of kind 4.

Return value:

The return value is of REAL type and the default kind.

Example:

                  program test_rand

integer,parameter :: seed = 86456
call srand(seed)
print *, rand(), rand(), rand(), rand()
print *, rand(seed), rand(), rand(), rand()
end program test_rand

See also: Section 8.258 [SRAND], page 275, Section 8.224 [RANDOM NUMBER], page 255,

8.223 RANDOM_INIT — Initialize a pseudo-random number generator

Description:

Initializes the state of the pseudorandom number generator used by RANDOM_

NUMBER.

Standard: Fortran 2018

Class: Subroutine

Syntax: CALL RANDOM_INIT(REPEATABLE, IMAGE_DISTINCT)

Arguments:

REPEATABLE Shall be a scalar with a LOGICAL type, and it is INTENT(IN). If it is .true., the seed is set to a processor-dependent value that is the same each time RANDOM_INIT is called from the same image. The term “same image” means a single instance of program execution. The sequence of random numbers is different for repeated execution of the program. If it is .false., the seed is set to a processor-dependent value.

IMAGE DISTINCT Shall be a scalar with a LOGICAL type, and it is INTENT(IN). If it is .true., the seed is set to a processor-dependent value that is distinct from th seed set by a call to RANDOM_INIT in another image. If it is .false., the seed is set to a value that does depend which image called RANDOM_INIT.

Example:

program test_random_seed
implicit none
real x(3), y(3)
call random_init(.true., .true.)
call random_number(x)
call random_init(.true., .true.)
call random_number(y)
! x and y are the same sequence
if (any(x /= y)) call abort
end program test_random_seed

See also: Section 8.224 [RANDOM NUMBER], page 255, Section 8.225 [RANDOM SEED], page 255,

8.224 RANDOM_NUMBER — Pseudo-random number

Description:

Returns a single pseudorandom number or an array of pseudorandom numbers from the uniform distribution over the range 0 ≤ x < 1. The runtime-library implements the xoshiro256** pseudorandom number generator (PRNG). This generator has a period of 2^{256}−1, and when using multiple threads up to 2^{128} threads can each generate 2^{128} random numbers before any aliasing occurs. Note that in a multi-threaded program (e.g. using OpenMP directives), each thread will have its own random number state. For details of the seeding procedure, see the documentation for the RANDOM_SEED intrinsic.

Standard: Fortran 90 and later

Class: Subroutine

Syntax: CALL RANDOM_NUMBER(HARVEST)

Arguments:

HARVEST Shall be a scalar or an array of type REAL.

Example:

      program test_random_number
        REAL :: r(5,5)
        CALL RANDOM_NUMBER(r)
      end program

See also: Section 8.225 [RANDOM SEED], page 255, Section 8.223 [RANDOM INIT], page 254,

8.225 RANDOM_SEED — Initialize a pseudo-random number sequence

Description:

Restarts or queries the state of the pseudorandom number generator used by

RANDOM_NUMBER.

If RANDOM_SEED is called without arguments, it is seeded with random data retrieved from the operating system. As an extension to the Fortran standard, the GFortran RANDOM_NUMBER supports multiple threads. Each thread in a multi-threaded program has its own seed. When RANDOM_SEED is called either without arguments or with the PUT argument, the given seed is copied into a master seed as well as the seed of the current thread. When a new thread uses RANDOM_NUMBER for the first time, the seed is copied from the master seed, and forwarded N ∗ 2^{128} steps to guarantee that the random stream does not alias any other stream in the system, where N is the number of threads that have used RANDOM_NUMBER so far during the program execution.

Standard: Fortran 90 and later

Class: Subroutine

Syntax: CALL RANDOM_SEED([SIZE, PUT, GET])

Arguments:

SIZE (Optional) Shall be a scalar and of type default INTEGER, with INTENT(OUT). It specifies the minimum size of the arrays used with the PUT and GET arguments.

PUT (Optional) Shall be an array of type default INTEGER and rank one. It is INTENT(IN) and the size of the array must be larger than or equal to the number returned by the SIZE argument.

GET (Optional) Shall be an array of type default INTEGER and rank one. It is INTENT(OUT) and the size of the array must be larger than or equal to the number returned by the SIZE argument.

Example:

                  program test_random_seed
                    implicit none
                    integer, allocatable :: seed(:)
                    integer :: n
                    call random_seed(size = n)
                    allocate(seed(n))
                    call random_seed(get=seed)
                    write (*, *) seed
                  end program test_random_seed

See also: Section 8.224 [RANDOM NUMBER], page 255, Section 8.223 [RANDOM INIT], page 254,

8.226 RANGE — Decimal exponent range

Description:

RANGE(X) returns the decimal exponent range in the model of the type of X.

Standard: Fortran 90 and later

Class: Inquiry function

Syntax: RESULT = RANGE(X)

Arguments:

          X  Shall be of type INTEGER, REAL or COMPLEX.

Return value:

The return value is of type INTEGER and of the default integer kind.

Example: See PRECISION for an example.

See also: Section 8.241 [SELECTED REAL KIND], page 264, Section 8.217 [PRECISION], page 251,

8.227 RANK — Rank of a data object

Description:

RANK(A) returns the rank of a scalar or array data object.

Standard: Technical Specification (TS) 29113

Class: Inquiry function

Syntax: RESULT = RANK(A)

Arguments:

A  can be of any type

Return value:

The return value is of type INTEGER and of the default integer kind. For arrays, their rank is returned; for scalars zero is returned.

Example:

      program test_rank
        integer :: a
        real, allocatable :: b(:,:)
        print *, rank(a), rank(b) ! Prints:  0  2
      end program test_rank

8.228 REAL — Convert to real type

Description:

REAL(A [, KIND]) converts its argument A to a real type. The REALPART function is provided for compatibility with g77, and its use is strongly discouraged.

Standard: Fortran 77 and later, with KIND argument Fortran 90 and later, has GNU extensions

Class: Elemental function

Syntax:

RESULT = REAL(A [, KIND])
RESULT = REALPART(Z)

Arguments:

A  Shall be INTEGER, REAL, or COMPLEX.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

These functions return a REAL variable or array under the following rules:

(A) REAL(A) is converted to a default real type if A is an integer or real variable.

(B) REAL(A) is converted to a real type with the kind type parameter of A if A is a complex variable.

(C) REAL(A, KIND) is converted to a real type with kind type parameter KIND if A is a complex, integer, or real variable.

Example:

      program test_real
        complex :: x = (1.0, 2.0)
        print *, real(x), real(x,8), realpart(x)
      end program test_real

Specific names:

Name  Argument  Return type Standard

FLOAT(A)       INTEGER(4)     REAL(4)  Fortran 77 and later
DFLOAT(A)      INTEGER(4)     REAL(8)  GNU extension
FLOATI(A)      INTEGER(2)     REAL(4)  GNU extension (-fdec)
FLOATJ(A)      INTEGER(4)     REAL(4)  GNU extension (-fdec)
FLOATK(A)      INTEGER(8)     REAL(4)  GNU extension (-fdec)
SNGL(A)        REAL(8)        REAL(4)  Fortran 77 and later

See also: Section 8.83 [DBLE], page 167,

8.229 RENAME — Rename a file

Description:

Renames a file from file PATH1 to PATH2. A null character (CHAR(0)) can be used to mark the end of the names in PATH1 and PATH2; otherwise, trailing blanks in the file names are ignored. If the STATUS argument is supplied, it contains 0 on success or a nonzero error code upon return; see rename(2). This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension Class: Subroutine, function Syntax:

CALL RENAME(PATH1, PATH2 [, STATUS])
STATUS = RENAME(PATH1, PATH2)

Arguments:

PATH1  Shall be of default CHARACTER type.
PATH2  Shall be of default CHARACTER type.
STATUS  (Optional) Shall be of default INTEGER type.

See also: Section 8.172 [LINK], page 223,

8.230 REPEAT — Repeated string concatenation

Description:

Concatenates NCOPIES copies of a string.

Standard: Fortran 90 and later Class: Transformational function

Syntax: RESULT = REPEAT(STRING, NCOPIES)

Arguments:

STRING  Shall be scalar and of type CHARACTER.
NCOPIES Shall be scalar and of type INTEGER.

Return value:

A new scalar of type CHARACTER built up from NCOPIES copies of STRING.

Example:

                  program test_repeat
                    write(*,*) repeat("x", 5)   ! "xxxxx"
                  end program

8.231 RESHAPE — Function to reshape an array

Description:

Reshapes SOURCE to correspond to SHAPE. If necessary, the new array may be padded with elements from PAD or permuted as defined by ORDER.

Standard: Fortran 90 and later

Class: Transformational function

Syntax: RESULT = RESHAPE(SOURCE, SHAPE[, PAD, ORDER])

Arguments:

SOURCE  Shall be an array of any type.

SHAPE Shall be of type INTEGER and an array of rank one. Its values must be positive or zero.

PAD (Optional) shall be an array of the same type as SOURCE.

ORDER (Optional) shall be of type INTEGER and an array of the same shape as SHAPE. Its values shall be a permutation of the numbers from 1 to n, where n is the size of SHAPE. If ORDER is absent, the natural ordering shall be assumed.

Return value:

The result is an array of shape SHAPE with the same type as SOURCE.

Example:

                  PROGRAM test_reshape
                    INTEGER, DIMENSION(4) :: x
                    WRITE(*,*) SHAPE(x)                       ! prints "4"
                    WRITE(*,*) SHAPE(RESHAPE(x, (/2, 2/)))    ! prints "2 2"
                  END PROGRAM

See also: Section 8.243 [SHAPE], page 266,

8.232 RRSPACING — Reciprocal of the relative spacing

Description:

RRSPACING(X) returns the reciprocal of the relative spacing of model numbers near X.

Standard: Fortran 90 and later

Class: Elemental function

Syntax: RESULT = RRSPACING(X)

Arguments:

X  Shall be of type REAL.

Return value:

The return value is of the same type and kind as X. The value returned is equal

to ABS(FRACTION(X)) * FLOAT(RADIX(X))**DIGITS(X).

See also: Section 8.255 [SPACING], page 273,

8.233 RSHIFT — Right shift bits

Description:

RSHIFT returns a value corresponding to I with all of the bits shifted right by SHIFT places. SHIFT shall be nonnegative and less than or equal to BIT_ SIZE(I), otherwise the result value is undefined. Bits shifted out from the right end are lost. The fill is arithmetic: the bits shifted in from the left end are equal to the leftmost bit, which in two’s complement representation is the sign bit. This function has been superseded by the SHIFTA intrinsic, which is standard in Fortran 2008 and later.

Standard: GNU extension

Class: Elemental function

Syntax: RESULT = RSHIFT(I, SHIFT)

Arguments:

I  The type shall be INTEGER.
SHIFT  The type shall be INTEGER.

Return value:

The return value is of type INTEGER and of the same kind as I.

See also: Section 8.159 [ISHFT], page 216, Section 8.160 [ISHFTC], page 217, Section 8.181 [LSHIFT], page 228, Section 8.244 [SHIFTA], page 267, Section 8.246 [SHIFTR], page 268, Section 8.245 [SHIFTL], page 267,

8.234 SAME_TYPE_AS — Query dynamic types for equality

Description:

Query dynamic types for equality.

Standard: Fortran 2003 and later Class: Inquiry function

Syntax: RESULT = SAME_TYPE_AS(A, B)

Arguments:

A Shall be an object of extensible declared type or unlimited polymorphic.

B Shall be an object of extensible declared type or unlimited polymorphic.

Return value:

The return value is a scalar of type default logical. It is true if and only if the dynamic type of A is the same as the dynamic type of B.

See also: Section 8.104 [EXTENDS TYPE OF], page 180,

8.235 SCALE — Scale a real value

Description:

          SCALE(X,I) returns X * RADIX(X)**I.

Standard: Fortran 90 and later

Class: Elemental function

Syntax: RESULT = SCALE(X, I)

Arguments:

X  The type of the argument shall be a REAL.
I  The type of the argument shall be a INTEGER.

Return value:

The return value is of the same type and kind as X. Its value is X *

          RADIX(X)**I.

Example:

                  program test_scale
                    real :: x = 178.1387e-4
                    integer :: i = 5
                    print *, scale(x,i), x*radix(x)**i
                  end program test_scale

8.236 SCAN — Scan a string for the presence of a set of characters

Description:

Scans a STRING for any of the characters in a SET of characters.

If BACK is either absent or equals FALSE, this function returns the position of the leftmost character of STRING that is in SET. If BACK equals TRUE, the rightmost position is returned. If no character of SET is found in STRING, the result is zero.

Standard: Fortran 90 and later, with KIND argument Fortran 2003 and later

Class: Elemental function

Syntax: RESULT = SCAN(STRING, SET[, BACK [, KIND]])

Arguments:

STRING  Shall be of type CHARACTER.
SET  Shall be of type CHARACTER.
BACK  (Optional) shall be of type LOGICAL.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind.

Example:

                  PROGRAM test_scan
                    WRITE(*,*) SCAN("FORTRAN", "AO")          ! 2, found 'O'
                    WRITE(*,*) SCAN("FORTRAN", "AO", .TRUE.)  ! 6, found 'A'
                    WRITE(*,*) SCAN("FORTRAN", "C++")         ! 0, found none
                  END PROGRAM

See also: Section 8.148 [INDEX intrinsic], page 209, Section 8.282 [VERIFY], page 289,

8.237 SECNDS — Time function

Description:

SECNDS(X) gets the time in seconds from the real-time system clock. X is a reference time, also in seconds. If this is zero, the time in seconds from midnight is returned. This function is non-standard and its use is discouraged.

Standard: GNU extension

Class: Function

Syntax: RESULT = SECNDS (X)

Arguments:

T  Shall be of type REAL(4).
X  Shall be of type REAL(4).

Return value:

None

Example:

                  program test_secnds
                      integer :: i
                      real(4) :: t1, t2
                      print *, secnds (0.0)   ! seconds since midnight
                      t1 = secnds (0.0)       ! reference time
                      do i = 1, 10000000      ! do something
                      end do
                      t2 = secnds (t1)        ! elapsed time
                      print *, "Something took ", t2, " seconds."
                  end program test_secnds

8.238 SECOND — CPU time function

Description:

Returns a REAL(4) value representing the elapsed CPU time in seconds. This provides the same functionality as the standard CPU_TIME intrinsic, and is only included for backwards compatibility. This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL SECOND(TIME)
TIME = SECOND()

Arguments:

TIME  Shall be of type REAL(4).

Return value:

In either syntax, TIME is set to the process’s current runtime in seconds.

See also: Section 8.79 [CPU TIME], page 164,

8.239 SELECTED_CHAR_KIND — Choose character kind

Description:

SELECTED_CHAR_KIND(NAME) returns the kind value for the character set named NAME, if a character set with such a name is supported, or −1 otherwise. Currently, supported character sets include “ASCII” and “DEFAULT”, which are equivalent, and “ISO 10646” (Universal Character Set, UCS-4) which is commonly known as Unicode.

Standard: Fortran 2003 and later

Class: Transformational function

Syntax: RESULT = SELECTED_CHAR_KIND(NAME)

Arguments:

NAME  Shall be a scalar and of the default character type.

Example:

      program character_kind
        use iso_fortran_env
        implicit none
        integer, parameter :: ascii = selected_char_kind ("ascii")
        integer, parameter :: ucs4  = selected_char_kind ('ISO_10646')
        character(kind=ascii, len=26) :: alphabet
        character(kind=ucs4,  len=30) :: hello_world
        alphabet = ascii_"abcdefghijklmnopqrstuvwxyz"
        hello_world = ucs4_'Hello World and Ni Hao -- ' &

                                  // char (int (z'4F60'), ucs4)     &
                                  // char (int (z'597D'), ucs4)
                    write (*,*) alphabet
                    open (output_unit, encoding='UTF-8')
                    write (*,*) trim (hello_world)
                  end program character_kind

8.240 SELECTED_INT_KIND — Choose integer kind

Description:

SELECTED_INT_KIND(R) return the kind value of the smallest integer type that can represent all values ranging from −10^{R} (exclusive) to 10^{R} (exclusive). If there is no integer kind that accommodates this range, SELECTED_INT_KIND returns −1.

Standard: Fortran 90 and later

Class: Transformational function

Syntax: RESULT = SELECTED_INT_KIND(R)

Arguments:

R  Shall be a scalar and of type INTEGER.

Example:

                  program large_integers
                    integer,parameter :: k5 = selected_int_kind(5)
                    integer,parameter :: k15 = selected_int_kind(15)
                    integer(kind=k5) :: i5
                    integer(kind=k15) :: i15
                    print *, huge(i5), huge(i15)
                    ! The following inequalities are always true
                    print *, huge(i5) >= 10_k5**5-1
                    print *, huge(i15) >= 10_k15**15-1
                  end program large_integers

8.241 SELECTED_REAL_KIND — Choose real kind

Description:

SELECTED_REAL_KIND(P,R) returns the kind value of a real data type with decimal precision of at least P digits, exponent range of at least R, and with a radix of RADIX.

Standard: Fortran 90 and later, with RADIX Fortran 2008 or later

Class: Transformational function

Syntax: RESULT = SELECTED_REAL_KIND([P, R, RADIX])

Arguments:

P  (Optional) shall be a scalar and of type INTEGER.
R  (Optional) shall be a scalar and of type INTEGER.

RADIX  (Optional) shall be a scalar and of type INTEGER.

Before Fortran 2008, at least one of the arguments R or P shall be present; since Fortran 2008, they are assumed to be zero if absent.

Return value:

SELECTED_REAL_KIND returns the value of the kind type parameter of a real data type with decimal precision of at least P digits, a decimal exponent range of at least R, and with the requested RADIX. If the RADIX parameter is absent, real kinds with any radix can be returned. If more than one real data type meet the criteria, the kind of the data type with the smallest decimal precision is returned. If no real data type matches the criteria, the result is

-1 if the processor does not support a real data type with a precision greater than or equal to P, but the R and RADIX requirements can be fulfilled

-2 if the processor does not support a real type with an exponent range greater than or equal to R, but P and RADIX are fulfillable

-3 if RADIX but not P and R requirements are fulfillable

-4 if RADIX and either P or R requirements are fulfillable

-5 if there is no real type with the given RADIX

Example:

      program real_kinds
        integer,parameter :: p6 = selected_real_kind(6)
        integer,parameter :: p10r100 = selected_real_kind(10,100)
        integer,parameter :: r400 = selected_real_kind(r=400)
        real(kind=p6) :: x
        real(kind=p10r100) :: y
        real(kind=r400) :: z
        print *, precision(x), range(x)
        print *, precision(y), range(y)
        print *, precision(z), range(z)
      end program real_kinds

See also: Section 8.217 [PRECISION], page 251, Section 8.226 [RANGE], page 256, Section 8.220 [RADIX], page 252,

8.242 SET_EXPONENT — Set the exponent of the model

Description:

SET_EXPONENT(X, I) returns the real number whose fractional part is that of X and whose exponent part is I.

Standard: Fortran 90 and later

Class: Elemental function

Syntax: RESULT = SET_EXPONENT(X, I)

Arguments:

X  Shall be of type REAL.
I  Shall be of type INTEGER.

Return value:

The return value is of the same type and kind as X. The real number whose fractional part is that of X and whose exponent part if I is returned; it is

FRACTION(X) * RADIX(X)**I.

Example:

      PROGRAM test_setexp
        REAL :: x = 178.1387e-4
        INTEGER :: i = 17
        PRINT *, SET_EXPONENT(x, i), FRACTION(x) * RADIX(x)**i
      END PROGRAM

8.243 SHAPE — Determine the shape of an array

Description:

Determines the shape of an array.

Standard: Fortran 90 and later, with KIND argument Fortran 2003 and later

Class: Inquiry function

Syntax: RESULT = SHAPE(SOURCE [, KIND])

Arguments:

SOURCE Shall be an array or scalar of any type. If SOURCE is a pointer it must be associated and allocatable arrays must be allocated.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

An INTEGER array of rank one with as many elements as SOURCE has dimensions. The elements of the resulting array correspond to the extend of SOURCE along the respective dimensions. If SOURCE is a scalar, the result is the rank one array of size zero. If KIND is absent, the return value has the default integer kind otherwise the specified kind.

Example:

      PROGRAM test_shape
        INTEGER, DIMENSION(-1:1, -1:2) :: A
        WRITE(*,*) SHAPE(A)             ! (/ 3, 4 /)
        WRITE(*,*) SIZE(SHAPE(42))      ! (/ /)
      END PROGRAM

See also: Section 8.231 [RESHAPE], page 259, Section 8.252 [SIZE], page 271,

8.244 SHIFTA — Right shift with fill

Description:

SHIFTA returns a value corresponding to I with all of the bits shifted right by SHIFT places. SHIFT that be nonnegative and less than or equal to BIT_ SIZE(I), otherwise the result value is undefined. Bits shifted out from the right end are lost. The fill is arithmetic: the bits shifted in from the left end are equal to the leftmost bit, which in two’s complement representation is the sign bit.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = SHIFTA(I, SHIFT)

Arguments:

I  The type shall be INTEGER.
SHIFT  The type shall be INTEGER.

Return value:

The return value is of type INTEGER and of the same kind as I.

See also: Section 8.245 [SHIFTL], page 267, Section 8.246 [SHIFTR], page 268,

8.245 SHIFTL — Left shift

Description:

SHIFTL returns a value corresponding to I with all of the bits shifted left by SHIFT places. SHIFT shall be nonnegative and less than or equal to BIT_ SIZE(I), otherwise the result value is undefined. Bits shifted out from the left end are lost, and bits shifted in from the right end are set to 0.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = SHIFTL(I, SHIFT)

Arguments:

I  The type shall be INTEGER.
SHIFT  The type shall be INTEGER.

Return value:

The return value is of type INTEGER and of the same kind as I.

See also: Section 8.244 [SHIFTA], page 267, Section 8.246 [SHIFTR], page 268,

8.246 SHIFTR — Right shift

Description:

SHIFTR returns a value corresponding to I with all of the bits shifted right by SHIFT places. SHIFT shall be nonnegative and less than or equal to BIT_ SIZE(I), otherwise the result value is undefined. Bits shifted out from the right end are lost, and bits shifted in from the left end are set to 0.

Standard: Fortran 2008 and later Class: Elemental function

Syntax: RESULT = SHIFTR(I, SHIFT)

Arguments:

I  The type shall be INTEGER.
SHIFT  The type shall be INTEGER.

Return value:

The return value is of type INTEGER and of the same kind as I.

See also: Section 8.244 [SHIFTA], page 267, Section 8.245 [SHIFTL], page 267,

8.247 SIGN — Sign copying function

Description:

SIGN(A,B) returns the value of A with the sign of B.

Standard: Fortran 77 and later Class: Elemental function

Syntax: RESULT = SIGN(A, B)

Arguments:

A  Shall be of type INTEGER or REAL
B  Shall be of the same type and kind as A.

Return value:

The kind of the return value is that of A and B. If B ≥ 0 then the result is

          ABS(A), else it is -ABS(A).

Example:

                  program test_sign
                    print *, sign(-12,1)
                    print *, sign(-12,0)
                    print *, sign(-12,-1)
                    print *, sign(-12.,1.)
                    print *, sign(-12.,0.)
                    print *, sign(-12.,-1.)
                  end program test_sign

Specific names:

Name  Arguments  Return type Standard

          SIGN(A,B)      REAL(4) A, B      REAL(4)  Fortran 77 and later

ISIGN(A,B)     INTEGER(4) A, B   INTEGER(4) Fortran 77 and later
DSIGN(A,B)     REAL(8) A, B      REAL(8)  Fortran 77 and later

8.248 SIGNAL — Signal handling subroutine (or function)

Description:

SIGNAL(NUMBER, HANDLER [, STATUS]) causes external subroutine HANDLER

to be executed with a single integer argument when signal NUMBER occurs. If HANDLER is an integer, it can be used to turn off handling of signal NUMBER or revert to its default action. See signal(2). If SIGNAL is called as a subroutine and the STATUS argument is supplied, it is set to the value returned by signal(2).

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL SIGNAL(NUMBER, HANDLER [, STATUS])
STATUS = SIGNAL(NUMBER, HANDLER)

Arguments:

NUMBER Shall be a scalar integer, with INTENT(IN)

HANDLER

            Signal handler (INTEGER FUNCTION or SUBROUTINE)

or dummy/global INTEGER scalar. INTEGER. It is

            INTENT(IN).

STATUS (Optional) STATUS shall be a scalar integer. It has

            INTENT(OUT).

Return value:

The SIGNAL function returns the value returned by signal(2).

Example:

      program test_signal
        intrinsic signal
        external handler_print
        call signal (12, handler_print)
        call signal (10, 1)
        call sleep (30)
      end program test_signal

8.249 SIN — Sine function

Description:

SIN(X) computes the sine of X.

Standard: Fortran 77 and later

Class: Elemental function

Syntax: RESULT = SIN(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value has same type and kind as X.

Example:

      program test_sin
        real :: x = 0.0
        x = sin(x)
      end program test_sin

Specific names:

Name  Argument  Return type Standard

SIN(X)         REAL(4) X      REAL(4)  Fortran 77 and later
DSIN(X)        REAL(8) X      REAL(8)  Fortran 77 and later
CSIN(X)        COMPLEX(4) X   COMPLEX(4) Fortran 77 and later
ZSIN(X)        COMPLEX(8) X   COMPLEX(8) GNU extension
CDSIN(X)       COMPLEX(8) X   COMPLEX(8) GNU extension

See also: Inverse function:

Section 8.19 [ASIN], page 122, Degrees function: Section 8.250 [SIND], page 270,

8.250 SIND — Sine function, degrees

Description:

SIND(X) computes the sine of X in degrees. This function is for compatibility only and should be avoided in favor of standard constructs wherever possible.

Standard: GNU extension, enabled with -fdec-math. Class: Elemental function

Syntax: RESULT = SIND(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value has same type and kind as X, and its value is in degrees.

Example:

      program test_sind
        real :: x = 0.0
        x = sind(x)
      end program test_sind

Specific names:

Name  Argument  Return type Standard

SIND(X)        REAL(4) X      REAL(4)  GNU extension
DSIND(X)       REAL(8) X      REAL(8)  GNU extension

          CSIND(X)       COMPLEX(4) X   COMPLEX(4) GNU extension
          ZSIND(X)       COMPLEX(8) X   COMPLEX(8) GNU extension
          CDSIND(X)      COMPLEX(8) X   COMPLEX(8) GNU extension

See also: Inverse function:

Section 8.20 [ASIND], page 123,

Radians function:

Section 8.249 [SIN], page 269,

8.251 SINH — Hyperbolic sine function

Description:

SINH(X) computes the hyperbolic sine of X.

Standard: Fortran 90 and later, for a complex argument Fortran 2008 or later, has a GNU extension

Class: Elemental function

Syntax: RESULT = SINH(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value has same type and kind as X.

Example:

                  program test_sinh
                    real(8) :: x = - 1.0_8
                    x = sinh(x)
                  end program test_sinh

Specific names:

Name  Argument  Return type Standard

          DSINH(X)       REAL(8) X      REAL(8)  Fortran 90 and later

See also: Section 8.21 [ASINH], page 124,

8.252 SIZE — Determine the size of an array

Description:

Determine the extent of ARRAY along a specified dimension DIM, or the total number of elements in ARRAY if DIM is absent.

Standard: Fortran 90 and later, with KIND argument Fortran 2003 and later

Class: Inquiry function

Syntax: RESULT = SIZE(ARRAY[, DIM [, KIND]])

Arguments:

ARRAY Shall be an array of any type. If ARRAY is a pointer it must be associated and allocatable arrays must be allocated.

DIM (Optional) shall be a scalar of type INTEGER and its value shall be in the range from 1 to n, where n equals the rank of ARRAY.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind.

Example:

      PROGRAM test_size
        WRITE(*,*) SIZE((/ 1, 2 /))    ! 2
      END PROGRAM

See also: Section 8.243 [SHAPE], page 266, Section 8.231 [RESHAPE], page 259,

8.253 SIZEOF — Size in bytes of an expression

Description:

SIZEOF(X) calculates the number of bytes of storage the expression X occupies.

Standard: GNU extension

Class: Inquiry function

Syntax: N = SIZEOF(X)

Arguments:

X  The argument shall be of any type, rank or shape.

Return value:

The return value is of type integer and of the system-dependent kind C SIZE T (from the ISO C BINDING module). Its value is the number of bytes occupied by the argument. If the argument has the POINTER attribute, the number of bytes of the storage area pointed to is returned. If the argument is of a derived type with POINTER or ALLOCATABLE components, the return value does not account for the sizes of the data pointed to by these components. If the argument is polymorphic, the size according to the dynamic type is returned. The argument may not be a procedure or procedure pointer. Note that the code assumes for arrays that those are contiguous; for contiguous arrays, it returns the storage or an array element multiplied by the size of the array.

Example:

         integer :: i
         real :: r, s(5)
         print *, (sizeof(s)/sizeof(r) == 5)
         end

The example will print .TRUE. unless you are using a platform where default REAL variables are unusually padded.

See also: Section 8.57 [C SIZEOF], page 148, Section 8.260 [STORAGE SIZE], page 276,

8.254 SLEEP — Sleep for the specified number of seconds

Description:

Calling this subroutine causes the process to pause for SECONDS seconds.

Standard: GNU extension Class: Subroutine

Syntax: CALL SLEEP(SECONDS)

Arguments:

SECONDS The type shall be of default INTEGER.

Example:

                  program test_sleep
                    call sleep(5)
                  end

8.255 SPACING — Smallest distance between two numbers of

a given type

Description:

Determines the distance between the argument X and the nearest adjacent number of the same type.

Standard: Fortran 90 and later Class: Elemental function

Syntax: RESULT = SPACING(X)

Arguments:

X  Shall be of type REAL.

Return value:

The result is of the same type as the input argument X.

Example:

                  PROGRAM test_spacing
                    INTEGER, PARAMETER :: SGL = SELECTED_REAL_KIND(p=6, r=37)
                    INTEGER, PARAMETER :: DBL = SELECTED_REAL_KIND(p=13, r=200)
                    WRITE(*,*) spacing(1.0_SGL)      ! "1.1920929E-07"          on i686
                    WRITE(*,*) spacing(1.0_DBL)      ! "2.220446049250313E-016" on i686
                  END PROGRAM

See also: Section 8.232 [RRSPACING], page 259,

8.256 SPREAD — Add a dimension to an array

Description:

Replicates a SOURCE array NCOPIES times along a specified dimension DIM.

Standard: Fortran 90 and later Class: Transformational function

Syntax: RESULT = SPREAD(SOURCE, DIM, NCOPIES)

Arguments:

SOURCE Shall be a scalar or an array of any type and a rank less than seven.

DIM Shall be a scalar of type INTEGER with a value in the range from 1 to n+1, where n equals the rank of SOURCE.

NCOPIES Shall be a scalar of type INTEGER.

Return value:

The result is an array of the same type as SOURCE and has rank n+1 where n equals the rank of SOURCE.

Example:

      PROGRAM test_spread
        INTEGER :: a = 1, b(2) = (/ 1, 2 /)
        WRITE(*,*) SPREAD(A, 1, 2)            ! "1 1"
        WRITE(*,*) SPREAD(B, 1, 2)            ! "1 1 2 2"
      END PROGRAM

See also: Section 8.281 [UNPACK], page 289,

8.257 SQRT — Square-root function

Description:

SQRT(X) computes the square root of X.

Standard: Fortran 77 and later Class: Elemental function

Syntax: RESULT = SQRT(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value is of type REAL or COMPLEX. The kind type parameter is the same as X.

Example:

      program test_sqrt
        real(8) :: x = 2.0_8
        complex :: z = (1.0, 2.0)
        x = sqrt(x)
        z = sqrt(z)
      end program test_sqrt

Specific names:

Name  Argument  Return type Standard

SQRT(X)        REAL(4) X      REAL(4)  Fortran 77 and later
DSQRT(X)       REAL(8) X      REAL(8)  Fortran 77 and later
CSQRT(X)       COMPLEX(4) X   COMPLEX(4) Fortran 77 and later
ZSQRT(X)       COMPLEX(8) X   COMPLEX(8) GNU extension
CDSQRT(X)      COMPLEX(8) X   COMPLEX(8) GNU extension

8.258 SRAND — Reinitialize the random number generator

Description:

SRAND reinitializes the pseudo-random number generator called by RAND and IRAND. The new seed used by the generator is specified by the required argument SEED.

Standard: GNU extension

Class: Subroutine

Syntax: CALL SRAND(SEED)

Arguments:

SEED  Shall be a scalar INTEGER(kind=4).

Return value:

Does not return anything.

Example: See RAND and IRAND for examples.

Notes: The Fortran standard specifies the intrinsic subroutines RANDOM_SEED to initialize the pseudo-random number generator and RANDOM_NUMBER to generate pseudo-random numbers. These subroutines should be used in new codes. Please note that in GNU Fortran, these two sets of intrinsics (RAND, IRAND and SRAND on the one hand, RANDOM_NUMBER and RANDOM_SEED on the other hand) access two independent pseudo-random number generators.

See also: Section 8.222 [RAND], page 253, Section 8.225 [RANDOM SEED], page 255, Section 8.224 [RANDOM NUMBER], page 255,

8.259 STAT — Get file status

Description:

This function returns information about a file. No permissions are required on the file itself, but execute (search) permission is required on all of the directories in path that lead to the file. The elements that are obtained and stored in the array VALUES:

VALUES(1) Device ID

VALUES(2) Inode number

VALUES(3) File mode

VALUES(4) Number of links
VALUES(5) Owner’s uid
VALUES(6) Owner’s gid

VALUES(7)

ID of device containing directory entry for file (0 if not available)

VALUES(8) File size (bytes)
VALUES(9) Last access time
VALUES(10) Last modification time
VALUES(11) Last file status change time

VALUES(12) Preferred I/O block size (-1 if not available)
VALUES(13) Number of blocks allocated (-1 if not available)

Not all these elements are relevant on all systems. If an element is not relevant, it is returned as 0. This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL STAT(NAME, VALUES [, STATUS])
STATUS = STAT(NAME, VALUES)

Arguments:

NAME The type shall be CHARACTER, of the default kind and a valid path within the file system.

VALUES  The type shall be INTEGER(4), DIMENSION(13).

STATUS (Optional) status flag of type INTEGER(4). Returns 0 on success and a system specific error code otherwise.

Example:

      PROGRAM test_stat
        INTEGER, DIMENSION(13) :: buff
        INTEGER :: status
        CALL STAT("/etc/passwd", buff, status)
        IF (status == 0) THEN
          WRITE (*, FMT="('Device ID:',               T30, I19)") buff(1)
          WRITE (*, FMT="('Inode number:',            T30, I19)") buff(2)
          WRITE (*, FMT="('File mode (octal):',       T30, O19)") buff(3)
          WRITE (*, FMT="('Number of links:',         T30, I19)") buff(4)
          WRITE (*, FMT="('Owner''s uid:',            T30, I19)") buff(5)
          WRITE (*, FMT="('Owner''s gid:',            T30, I19)") buff(6)
          WRITE (*, FMT="('Device where located:',    T30, I19)") buff(7)
          WRITE (*, FMT="('File size:',               T30, I19)") buff(8)
          WRITE (*, FMT="('Last access time:',        T30, A19)") CTIME(buff(9))
          WRITE (*, FMT="('Last modification time',   T30, A19)") CTIME(buff(10))
          WRITE (*, FMT="('Last status change time:', T30, A19)") CTIME(buff(11))
          WRITE (*, FMT="('Preferred block size:',    T30, I19)") buff(12)
          WRITE (*, FMT="('No. of blocks allocated:', T30, I19)") buff(13)
        END IF
      END PROGRAM

See also: To stat an open file:

Section 8.117 [FSTAT], page 190, To stat a link: Section 8.182 [LSTAT], page 229,

8.260 STORAGE_SIZE — Storage size in bits

Description:

Returns the storage size of argument A in bits.

Standard: Fortran 2008 and later

Class: Inquiry function

Syntax: RESULT = STORAGE_SIZE(A [, KIND])

Arguments:

A  Shall be a scalar or array of any type.

KIND (Optional) shall be a scalar integer constant expression.

Return Value:

The result is a scalar integer with the kind type parameter specified by KIND (or default integer type if KIND is missing). The result value is the size expressed in bits for an element of an array that has the dynamic type and type parameters of A.

See also: Section 8.57 [C SIZEOF], page 148, Section 8.253 [SIZEOF], page 272,

8.261 SUM — Sum of array elements

Description:

Adds the elements of ARRAY along dimension DIM if the corresponding element in MASK is TRUE.

Standard: Fortran 90 and later

Class: Transformational function

Syntax:

RESULT = SUM(ARRAY[, MASK])
RESULT = SUM(ARRAY, DIM[, MASK])

Arguments:

ARRAY  Shall be an array of type INTEGER, REAL or COMPLEX.

DIM (Optional) shall be a scalar of type INTEGER with a value in the range from 1 to n, where n equals the rank of ARRAY.

MASK (Optional) shall be of type LOGICAL and either be a scalar or an array of the same shape as ARRAY.

Return value:

The result is of the same type as ARRAY. If DIM is absent, a scalar with the sum of all elements in ARRAY is returned. Otherwise, an array of rank n-1, where n equals the rank of ARRAY, and a shape similar to that of ARRAY with dimension DIM dropped is returned.

Example:

      PROGRAM test_sum
        INTEGER :: x(5) = (/ 1, 2, 3, 4 ,5 /)
        print *, SUM(x)                        ! all elements, sum = 15
        print *, SUM(x, MASK=MOD(x, 2)==1)     ! odd elements, sum = 9
      END PROGRAM

See also: Section 8.219 [PRODUCT], page 252,

8.262 SYMLNK — Create a symbolic link

Description:

Makes a symbolic link from file PATH1 to PATH2. A null character (CHAR(0)) can be used to mark the end of the names in PATH1 and PATH2; otherwise, trailing blanks in the file names are ignored. If the STATUS argument is supplied, it contains 0 on success or a nonzero error code upon return; see symlink(2). If the system does not supply symlink(2), ENOSYS is returned. This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL SYMLNK(PATH1, PATH2 [, STATUS])
STATUS = SYMLNK(PATH1, PATH2)

Arguments:

PATH1  Shall be of default CHARACTER type.
PATH2  Shall be of default CHARACTER type.
STATUS  (Optional) Shall be of default INTEGER type.

See also: Section 8.172 [LINK], page 223, Section 8.280 [UNLINK], page 288,

8.263 SYSTEM — Execute a shell command

Description:

Passes the command COMMAND to a shell (see system(3)). If argument STATUS is present, it contains the value returned by system(3), which is presumably 0 if the shell command succeeded. Note that which shell is used to invoke the command is system-dependent and environment-dependent. This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit. Note that the system function need not be thread-safe. It is the responsibility of the user to ensure that system is not called concurrently.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL SYSTEM(COMMAND [, STATUS])
STATUS = SYSTEM(COMMAND)

Arguments:

COMMAND Shall be of default CHARACTER type.
STATUS  (Optional) Shall be of default INTEGER type.

See also: Section 8.100 [EXECUTE COMMAND LINE], page 178, which is part of the Fortran 2008 standard and should considered in new code for future portability.

8.264 SYSTEM_CLOCK — Time function

Description:

Determines the COUNT of a processor clock since an unspecified time in the past modulo COUNT MAX, COUNT RATE determines the number of clock ticks per second. If the platform supports a monotonic clock, that clock is used and can, depending on the platform clock implementation, provide up to nanosecond resolution. If a monotonic clock is not available, the implementation falls back to a realtime clock. COUNT RATE is system dependent and can vary depending on the kind of the arguments. For kind=4 arguments (and smaller integer kinds), COUNT represents milliseconds, while for kind=8 arguments (and larger integer kinds), COUNT typically represents micro- or nanoseconds depending on resolution of the underlying platform clock. COUNT MAX usually equals HUGE(COUNT_ MAX). Note that the millisecond resolution of the kind=4 version implies that the COUNT will wrap around in roughly 25 days. In order to avoid issues with the wrap around and for more precise timing, please use the kind=8 version. If there is no clock, or querying the clock fails, COUNT is set to -HUGE(COUNT), and COUNT RATE and COUNT MAX are set to zero. When running on a platform using the GNU C library (glibc) version 2.16 or older, or a derivative thereof, the high resolution monotonic clock is available only when linking with the rt library. This can be done explicitly by adding the -lrt flag when linking the application, but is also done implicitly when using OpenMP. On the Windows platform, the version with kind=4 arguments uses the GetTickCount function, whereas the kind=8 version uses

QueryPerformanceCounter and QueryPerformanceCounterFrequency.

For more information, and potential caveats, please see the platform documentation.

Standard: Fortran 90 and later

Class: Subroutine

Syntax: CALL SYSTEM_CLOCK([COUNT, COUNT_RATE, COUNT_MAX])

Arguments:

COUNT (Optional) shall be a scalar of type INTEGER with

               INTENT(OUT).

COUNT RATE (Optional) shall be a scalar of type INTEGER or

               REAL, with INTENT(OUT).

COUNT MAX (Optional) shall be a scalar of type INTEGER with

               INTENT(OUT).

Example:

      PROGRAM test_system_clock
        INTEGER :: count, count_rate, count_max
        CALL SYSTEM_CLOCK(count, count_rate, count_max)
        WRITE(*,*) count, count_rate, count_max
      END PROGRAM

See also: Section 8.82 [DATE AND TIME], page 166, Section 8.79 [CPU TIME], page 164,

8.265 TAN — Tangent function

Description:

TAN(X) computes the tangent of X.

Standard: Fortran 77 and later, for a complex argument Fortran 2008 or later Class: Elemental function

Syntax: RESULT = TAN(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value has same type and kind as X, and its value is in radians.

Example:

                  program test_tan
                    real(8) :: x = 0.165_8
                    x = tan(x)
                  end program test_tan

Specific names:

Name  Argument  Return type Standard

          TAN(X)         REAL(4) X      REAL(4)  Fortran 77 and later
          DTAN(X)        REAL(8) X      REAL(8)  Fortran 77 and later

See also: Inverse function:

Section 8.23 [ATAN], page 126,

Degrees function:

Section 8.266 [TAND], page 280,

8.266 TAND — Tangent function, degrees

Description:

TAND(X) computes the tangent of X in degrees.

This function is for compatibility only and should be avoided in favor of standard constructs wherever possible.

Standard: GNU extension, enabled with -fdec-math. Class: Elemental function

Syntax: RESULT = TAND(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value has same type and kind as X, and its value is in degrees.

Example:

      program test_tand
        real(8) :: x = 0.165_8
        x = tand(x)
      end program test_tand

Specific names:

Name  Argument  Return type Standard

TAND(X)        REAL(4) X      REAL(4)  GNU extension
DTAND(X)       REAL(8) X      REAL(8)  GNU extension

See also: Inverse function:

Section 8.24 [ATAND], page 126, Radians function: Section 8.265 [TAN], page 280,

8.267 TANH — Hyperbolic tangent function

Description:

TANH(X) computes the hyperbolic tangent of X.

Standard: Fortran 77 and later, for a complex argument Fortran 2008 or later

Class: Elemental function

Syntax: X = TANH(X)

Arguments:

X  The type shall be REAL or COMPLEX.

Return value:

The return value has same type and kind as X. If X is complex, the imaginary part of the result is in radians. If X is REAL, the return value lies in the range −1 ≤ tanh(x) ≤ 1.

Example:

      program test_tanh
        real(8) :: x = 2.1_8
        x = tanh(x)
      end program test_tanh

Specific names:

Name  Argument  Return type Standard

TANH(X)        REAL(4) X      REAL(4)  Fortran 77 and later
DTANH(X)       REAL(8) X      REAL(8)  Fortran 77 and later

See also: Section 8.27 [ATANH], page 129,

8.268 THIS_IMAGE — Function that returns the cosubscript index of this image

Description:

Returns the cosubscript for this image.

Standard: Fortran 2008 and later. With DISTANCE argument, Technical Specification (TS) 18508 or later

Class: Transformational function

Syntax:

RESULT = THIS_IMAGE()
RESULT = THIS_IMAGE(DISTANCE)
RESULT = THIS_IMAGE(COARRAY [, DIM])

Arguments:

DISTANCE (optional, intent(in)) Nonnegative scalar integer (not permitted together with COARRAY ).

COARRAY Coarray of any type (optional; if DIM present, required).

DIM default integer scalar (optional). If present, DIM shall be between one and the corank of COARRAY.

Return value:

Default integer. If COARRAY is not present, it is scalar; if DISTANCE is not present or has value 0, its value is the image index on the invoking image for the current team, for values smaller or equal distance to the initial team, it returns the image index on the ancestor team which has a distance of DISTANCE from the invoking team. If DISTANCE is larger than the distance to the initial team, the image index of the initial team is returned. Otherwise when the COARRAY is present, if DIM is not present, a rank-1 array with corank elements is returned, containing the cosubscripts for COARRAY specifying the invoking image. If DIM is present, a scalar is returned, with the value of the

DIM element of THIS_IMAGE(COARRAY).

Example:

      INTEGER :: value[*]
      INTEGER :: i
      value = THIS_IMAGE()
      SYNC ALL
      IF (THIS_IMAGE() == 1) THEN
        DO i = 1, NUM_IMAGES()
          WRITE(*,'(2(a,i0))') 'value[', i, '] is ', value[i]
        END DO
      END IF
      ! Check whether the current image is the initial image
      IF (THIS_IMAGE(HUGE(1)) /= THIS_IMAGE())
        error stop "something is rotten here"

See also: Section 8.210 [NUM IMAGES], page 246, Section 8.147 [IMAGE INDEX], page 209,

8.269 TIME — Time function

Description:

Returns the current time encoded as an integer (in the manner of the function time(3) in the C standard library). This value is suitable for passing to Section 8.81 [CTIME], page 165, Section 8.131 [GMTIME], page 198, and Section 8.183 [LTIME], page 230. This intrinsic is not fully portable, such as to systems with 32-bit INTEGER types but supporting times wider than 32 bits. Therefore, the values returned by this intrinsic might be, or become, negative, or numerically less than previous values, during a single run of the compiled program. See Section 8.270 [TIME8], page 283, for information on a similar intrinsic that might be portable to more GNU Fortran implementations, though to fewer Fortran compilers.

Standard: GNU extension

Class: Function

Syntax: RESULT = TIME()

Return value:

The return value is a scalar of type INTEGER(4).

See also: Section 8.82 [DATE AND TIME], page 166, Section 8.81 [CTIME], page 165, Section 8.131 [GMTIME], page 198, Section 8.183 [LTIME], page 230, Section 8.192 [MCLOCK], page 235, Section 8.270 [TIME8], page 283,

8.270 TIME8 — Time function (64-bit)

Description:

Returns the current time encoded as an integer (in the manner of the function time(3) in the C standard library). This value is suitable for passing to Section 8.81 [CTIME], page 165, Section 8.131 [GMTIME], page 198, and Section 8.183 [LTIME], page 230. Warning: this intrinsic does not increase the range of the timing values over that returned by time(3). On a system with a 32-bit time(3), TIME8 will return a 32-bit value, even though it is converted to a 64-bit INTEGER(8) value. That means overflows of the 32-bit value can still occur. Therefore, the values returned by this intrinsic might be or become negative or numerically less than previous values during a single run of the compiled program.

Standard: GNU extension

Class: Function

Syntax: RESULT = TIME8()

Return value:

The return value is a scalar of type INTEGER(8).

See also: Section 8.82 [DATE AND TIME], page 166, Section 8.81 [CTIME], page 165, Section 8.131 [GMTIME], page 198, Section 8.183 [LTIME], page 230, Section 8.193 [MCLOCK8], page 236, Section 8.269 [TIME], page 282,

8.271 TINY — Smallest positive number of a real kind

Description:

TINY(X) returns the smallest positive (non zero) number in the model of the type of X.

Standard: Fortran 90 and later

Class: Inquiry function

Syntax: RESULT = TINY(X)

Arguments:

X  Shall be of type REAL.

Return value:

The return value is of the same type and kind as X

Example: See HUGE for an example.

8.272 TRAILZ — Number of trailing zero bits of an integer

Description:

TRAILZ returns the number of trailing zero bits of an integer.

Standard: Fortran 2008 and later

Class: Elemental function

Syntax: RESULT = TRAILZ(I)

Arguments:

I  Shall be of type INTEGER.

Return value:

The type of the return value is the default INTEGER. If all the bits of I are zero, the result value is BIT_SIZE(I).

Example:

                  PROGRAM test_trailz
                    WRITE (*,*) TRAILZ(8)  ! prints 3
                  END PROGRAM

See also: Section 8.48 [BIT SIZE], page 142, Section 8.167 [LEADZ], page 220, Section 8.216 [POPPAR], page 250, Section 8.215 [POPCNT], page 250,

8.273 TRANSFER — Transfer bit patterns

Description:

Interprets the bitwise representation of SOURCE in memory as if it is the representation of a variable or array of the same type and type parameters as MOLD. This is approximately equivalent to the C concept of casting one type to another.

Standard: Fortran 90 and later

Class: Transformational function

Syntax: RESULT = TRANSFER(SOURCE, MOLD[, SIZE])

Arguments:

SOURCE  Shall be a scalar or an array of any type.
MOLD  Shall be a scalar or an array of any type.
SIZE  (Optional) shall be a scalar of type INTEGER.

Return value:

The result has the same type as MOLD, with the bit level representation of SOURCE. If SIZE is present, the result is a one-dimensional array of length SIZE. If SIZE is absent but MOLD is an array (of any size or shape), the result is a one- dimensional array of the minimum length needed to contain the entirety of the bitwise representation of SOURCE. If SIZE is absent and MOLD is a scalar, the result is a scalar. If the bitwise representation of the result is longer than that of SOURCE, then the leading bits of the result correspond to those of SOURCE and any trailing bits are filled arbitrarily. When the resulting bit representation does not correspond to a valid representation of a variable of the same type as MOLD, the results are undefined, and subsequent operations on the result cannot be guaranteed to produce sensible behavior. For example, it is possible to create LOGICAL variables for which VAR and .NOT.VAR both appear to be true.

Example:

      PROGRAM test_transfer
        integer :: x = 2143289344
        print *, transfer(x, 1.0)    ! prints "NaN" on i686
      END PROGRAM

8.274 TRANSPOSE — Transpose an array of rank two

Description:

Transpose an array of rank two. Element (i, j) of the result has the value

MATRIX(j, i), for all i, j.

Standard: Fortran 90 and later

Class: Transformational function

Syntax: RESULT = TRANSPOSE(MATRIX)

Arguments:

MATRIX  Shall be an array of any type and have a rank of two.

Return value:

The result has the same type as MATRIX, and has shape (/ m, n /) if MA-TRIX has shape (/ n, m /).

8.275 TRIM — Remove trailing blank characters of a string

Description:

Removes trailing blank characters of a string.

Standard: Fortran 90 and later Class: Transformational function

Syntax: RESULT = TRIM(STRING)

Arguments:

STRING  Shall be a scalar of type CHARACTER.

Return value:

A scalar of type CHARACTER which length is that of STRING less the number of trailing blanks.

Example:

                  PROGRAM test_trim
                    CHARACTER(len=10), PARAMETER :: s = "GFORTRAN  "
                    WRITE(*,*) LEN(s), LEN(TRIM(s))  ! "10 8", with/without trailing blanks
                  END PROGRAM

See also: Section 8.9 [ADJUSTL], page 116, Section 8.10 [ADJUSTR], page 116,

8.276 TTYNAM — Get the name of a terminal device

Description:

Get the name of a terminal device. For more information, see ttyname(3). This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension Class: Subroutine, function Syntax:

          CALL TTYNAM(UNIT, NAME)
          NAME = TTYNAM(UNIT)

Arguments:

UNIT  Shall be a scalar INTEGER.
NAME  Shall be of type CHARACTER.

Example:

      PROGRAM test_ttynam
        INTEGER :: unit
        DO unit = 1, 10
          IF (isatty(unit=unit)) write(*,*) ttynam(unit)
        END DO
      END PROGRAM

See also: Section 8.158 [ISATTY], page 215,

8.277 UBOUND — Upper dimension bounds of an array

Description:

Returns the upper bounds of an array, or a single upper bound along the DIM dimension.

Standard: Fortran 90 and later, with KIND argument Fortran 2003 and later

Class: Inquiry function

Syntax: RESULT = UBOUND(ARRAY [, DIM [, KIND]])

Arguments:

ARRAY  Shall be an array, of any type.
DIM  (Optional) Shall be a scalar INTEGER.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind. If DIM is absent, the result is an array of the upper bounds of ARRAY. If DIM is present, the result is a scalar corresponding to the upper bound of the array along that dimension. If ARRAY is an expression rather than a whole array or array structure component, or if it has a zero extent along the relevant dimension, the upper bound is taken to be the number of elements along the relevant dimension.

See also: Section 8.165 [LBOUND], page 219, Section 8.166 [LCOBOUND], page 220,

8.278 UCOBOUND — Upper codimension bounds of an array

Description:

Returns the upper cobounds of a coarray, or a single upper cobound along the DIM codimension.

Standard: Fortran 2008 and later

Class: Inquiry function

Syntax: RESULT = UCOBOUND(COARRAY [, DIM [, KIND]])

Arguments:

ARRAY  Shall be an coarray, of any type.

DIM  (Optional) Shall be a scalar INTEGER.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind. If DIM is absent, the result is an array of the lower cobounds of COARRAY. If DIM is present, the result is a scalar corresponding to the lower cobound of the array along that codimension.

See also: Section 8.166 [LCOBOUND], page 220, Section 8.165 [LBOUND], page 219,

8.279 UMASK — Set the file creation mask

Description:

Sets the file creation mask to MASK. If called as a function, it returns the old value. If called as a subroutine and argument OLD if it is supplied, it is set to the old value. See umask(2).

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL UMASK(MASK [, OLD])
OLD = UMASK(MASK)

Arguments:

MASK  Shall be a scalar of type INTEGER.
OLD  (Optional) Shall be a scalar of type INTEGER.

8.280 UNLINK — Remove a file from the file system

Description:

Unlinks the file PATH. A null character (CHAR(0)) can be used to mark the end of the name in PATH; otherwise, trailing blanks in the file name are ignored. If the STATUS argument is supplied, it contains 0 on success or a nonzero error code upon return; see unlink(2). This intrinsic is provided in both subroutine and function forms; however, only one form can be used in any given program unit.

Standard: GNU extension

Class: Subroutine, function

Syntax:

CALL UNLINK(PATH [, STATUS])
STATUS = UNLINK(PATH)

Arguments:

PATH  Shall be of default CHARACTER type.
STATUS  (Optional) Shall be of default INTEGER type.

See also: Section 8.172 [LINK], page 223, Section 8.262 [SYMLNK], page 278,

8.281 UNPACK — Unpack an array of rank one into an array

Description:

Store the elements of VECTOR in an array of higher rank.

Standard: Fortran 90 and later

Class: Transformational function

Syntax: RESULT = UNPACK(VECTOR, MASK, FIELD)

Arguments:

VECTOR Shall be an array of any type and rank one. It shall have at least as many elements as MASK has TRUE values.

MASK  Shall be an array of type LOGICAL.

FIELD Shall be of the same type as VECTOR and have the same shape as MASK.

Return value:

The resulting array corresponds to FIELD with TRUE elements of MASK replaced by values from VECTOR in array element order.

Example:

                  PROGRAM test_unpack
                    integer :: vector(2)  = (/1,1/)
                    logical :: mask(4)  = (/ .TRUE., .FALSE., .FALSE., .TRUE. /)
                    integer :: field(2,2) = 0, unity(2,2)
                    ! result: unity matrix
                    unity = unpack(vector, reshape(mask, (/2,2/)), field)
                  END PROGRAM

See also: Section 8.212 [PACK], page 248, Section 8.256 [SPREAD], page 273,

8.282 VERIFY — Scan a string for characters not a given set

Description:

Verifies that all the characters in STRING belong to the set of characters in SET. If BACK is either absent or equals FALSE, this function returns the position of the leftmost character of STRING that is not in SET. If BACK equals TRUE, the rightmost position is returned. If all characters of STRING are found in SET, the result is zero.

Standard: Fortran 90 and later, with KIND argument Fortran 2003 and later

Class: Elemental function

Syntax: RESULT = VERIFY(STRING, SET[, BACK [, KIND]])

Arguments:

STRING  Shall be of type CHARACTER.
SET  Shall be of type CHARACTER.
BACK  (Optional) shall be of type LOGICAL.

KIND (Optional) A scalar INTEGER constant expression indicating the kind parameter of the result.

Return value:

The return value is of type INTEGER and of kind KIND. If KIND is absent, the return value is of default integer kind.

Example:

      PROGRAM test_verify
        WRITE(*,*) VERIFY("FORTRAN", "AO")           ! 1, found 'F'
        WRITE(*,*) VERIFY("FORTRAN", "FOO")          ! 3, found 'R'
        WRITE(*,*) VERIFY("FORTRAN", "C++")          ! 1, found 'F'
        WRITE(*,*) VERIFY("FORTRAN", "C++", .TRUE.)  ! 7, found 'N'
        WRITE(*,*) VERIFY("FORTRAN", "FORTRAN")      ! 0' found none
      END PROGRAM

See also: Section 8.236 [SCAN], page 261, Section 8.148 [INDEX intrinsic], page 209,

8.283 XOR — Bitwise logical exclusive OR

Description:

Bitwise logical exclusive or. This intrinsic routine is provided for backwards compatibility with GNU Fortran 77. For integer arguments, programmers should consider the use of the Section 8.145 [IEOR], page 208, intrinsic and for logical arguments the .NEQV. operator, which are both defined by the Fortran standard.

Standard: GNU extension

Class: Function

Syntax: RESULT = XOR(I, J)

Arguments:

I The type shall be either a scalar INTEGER type or a scalar LOGICAL type or a boz-literal-constant.

J The type shall be the same as the type of I or a bozliteral-constant. I and J shall not both be boz-literal-constants. If either I and J is a boz-literal-constant, then the other argument must be a scalar INTEGER.

Return value:

The return type is either a scalar INTEGER or a scalar LOGICAL. If the kind type parameters differ, then the smaller kind type is implicitly converted to larger kind, and the return has the larger kind. A boz-literal-constant is converted to an INTEGER with the kind type parameter of the other argument as-if a call to Section 8.149 [INT], page 210, occurred.

Chapter 8: Intrinsic Procedures 291

Example:

PROGRAM test_xor
LOGICAL :: T = .TRUE., F = .FALSE.
INTEGER :: a, b
DATA a / Z'F' /, b / Z'3' /
WRITE (*,*) XOR(T, T), XOR(T, F), XOR(F, T), XOR(F, F)
WRITE (*,*) XOR(a, b)
END PROGRAM

See also: Fortran 95 elemental function:

Section 8.145 [IEOR], page 208,

9 Intrinsic Modules

9.1 ISO_FORTRAN_ENV

Standard: Fortran 2003 and later, except when otherwise noted

The ISO_FORTRAN_ENV module provides the following scalar default-integer named constants:

ATOMIC_INT_KIND:

Default-kind integer constant to be used as kind parameter when defining integer variables used in atomic operations. (Fortran 2008 or later.)

ATOMIC_LOGICAL_KIND:

Default-kind integer constant to be used as kind parameter when defining logical variables used in atomic operations. (Fortran 2008 or later.)

CHARACTER_KINDS:

Default-kind integer constant array of rank one containing the supported kind parameters of the CHARACTER type. (Fortran 2008 or later.)

CHARACTER_STORAGE_SIZE:

Size in bits of the character storage unit.

ERROR_UNIT:

Identifies the preconnected unit used for error reporting.

FILE_STORAGE_SIZE:

Size in bits of the file-storage unit.

INPUT_UNIT:

Identifies the preconnected unit identified by the asterisk (*) in READ statement.

INT8, INT16, INT32, INT64:

Kind type parameters to specify an INTEGER type with a storage size of 16, 32, and 64 bits. It is negative if a target platform does not support the particular kind. (Fortran 2008 or later.)

INTEGER_KINDS:

Default-kind integer constant array of rank one containing the supported kind parameters of the INTEGER type. (Fortran 2008 or later.)

IOSTAT_END:

The value assigned to the variable passed to the IOSTAT= specifier of an input/output statement if an end-of-file condition occurred.

IOSTAT_EOR:

The value assigned to the variable passed to the IOSTAT= specifier of an input/output statement if an end-of-record condition occurred.

IOSTAT_INQUIRE_INTERNAL_UNIT:

Scalar default-integer constant, used by INQUIRE for the IOSTAT= specifier to denote an that a unit number identifies an internal unit. (Fortran 2008 or later.)

NUMERIC_STORAGE_SIZE:

The size in bits of the numeric storage unit.

LOGICAL_KINDS:

Default-kind integer constant array of rank one containing the supported kind parameters of the LOGICAL type. (Fortran 2008 or later.)

OUTPUT_UNIT:

Identifies the preconnected unit identified by the asterisk (*) in WRITE statement.

REAL32, REAL64, REAL128:

Kind type parameters to specify a REAL type with a storage size of 32, 64, and 128 bits. It is negative if a target platform does not support the particular kind. (Fortran 2008 or later.)

REAL_KINDS:

Default-kind integer constant array of rank one containing the supported kind parameters of the REAL type. (Fortran 2008 or later.)

STAT_LOCKED:

Scalar default-integer constant used as STAT= return value by LOCK to denote that the lock variable is locked by the executing image. (Fortran 2008 or later.)

STAT_LOCKED_OTHER_IMAGE:

Scalar default-integer constant used as STAT= return value by UNLOCK to denote that the lock variable is locked by another image. (Fortran 2008 or later.)

STAT_STOPPED_IMAGE:

Positive, scalar default-integer constant used as STAT= return value if the argument in the statement requires synchronisation with an image, which has initiated the termination of the execution. (Fortran 2008 or later.)

STAT_FAILED_IMAGE:

Positive, scalar default-integer constant used as STAT= return value if the argument in the statement requires communication with an image, which has is in the failed state. (TS 18508 or later.)

STAT_UNLOCKED:

Scalar default-integer constant used as STAT= return value by UNLOCK to denote that the lock variable is unlocked. (Fortran 2008 or later.)

The module provides the following derived type:

LOCK_TYPE:

Derived type with private components to be use with the LOCK and UNLOCK statement. A variable of its type has to be always declared as coarray and may not appear in a variable-definition context. (Fortran 2008 or later.)

The module also provides the following intrinsic procedures: Section 8.69 [COMPILER OPTIONS], page 157, and Section 8.70 [COMPILER VERSION], page 158.

9.2 ISO_C_BINDING

Standard: Fortran 2003 and later, GNU extensions

The following intrinsic procedures are provided by the module; their definition can be found in the section Intrinsic Procedures of this manual.

C_ASSOCIATED
C_F_POINTER
C_F_PROCPOINTER
C_FUNLOC
C_LOC
C_SIZEOF

The ISO_C_BINDING module provides the following named constants of type default integer, which can be used as KIND type parameters.

In addition to the integer named constants required by the Fortran 2003 standard and C_PTRDIFF_T of TS 29113, GNU Fortran provides as an extension named constants for the 128-bit integer types supported by the C compiler: C_INT128_T, C_INT_LEAST128_T, C_INT_FAST128_T. Furthermore, if _Float128 is supported in C, the named constants

C_FLOAT128 and C_FLOAT128_COMPLEX are defined.

Fortran Type  Named constant  C type  Extension

INTEGER         C_INT                   int
INTEGER         C_SHORT                 short int
INTEGER         C_LONG                  long int
INTEGER         C_LONG_LONG             long long int

INTEGER

                C_SIGNED_CHAR

                                         signed char/unsigned
                                         char

INTEGER         C_SIZE_T                size_t
INTEGER         C_INT8_T                int8_t
INTEGER         C_INT16_T               int16_t
INTEGER         C_INT32_T               int32_t
INTEGER         C_INT64_T               int64_t
INTEGER         C_INT128_T              int128_t  Ext.
INTEGER         C_INT_LEAST8_T          int_least8_t
INTEGER         C_INT_LEAST16_T         int_least16_t
INTEGER         C_INT_LEAST32_T         int_least32_t
INTEGER         C_INT_LEAST64_T         int_least64_t
INTEGER         C_INT_LEAST128_T        int_least128_t  Ext.
INTEGER         C_INT_FAST8_T           int_fast8_t
INTEGER         C_INT_FAST16_T          int_fast16_t
INTEGER         C_INT_FAST32_T          int_fast32_t
INTEGER         C_INT_FAST64_T          int_fast64_t
INTEGER         C_INT_FAST128_T         int_fast128_t  Ext.
INTEGER         C_INTMAX_T              intmax_t
INTEGER         C_INTPTR_T              intptr_t
INTEGER         C_PTRDIFF_T             ptrdiff_t  TS 29113

REAL            C_FLOAT                 float
REAL            C_DOUBLE                double
REAL            C_LONG_DOUBLE           long double
REAL            C_FLOAT128              _Float128  Ext.
COMPLEX         C_FLOAT_COMPLEX         float _Complex
COMPLEX         C_DOUBLE_COMPLEX        double _Complex
COMPLEX         C_LONG_DOUBLE_COMPLEX   long double _Complex
COMPLEX         C_FLOAT128_COMPLEX      _Float128 _Complex  Ext.
LOGICAL         C_BOOL                  _Bool
CHARACTER       C_CHAR                  char

Additionally, the following parameters of type CHARACTER(KIND=C_CHAR) are defined.

Name  C definition  Value

C_NULL_CHAR  null character  '\0'
C_ALERT  alert  '\a'
C_BACKSPACE  backspace  '\b'
C_FORM_FEED  form feed  '\f'
C_NEW_LINE  new line  '\n'
C_CARRIAGE_ RETURN  carriage return  '\r'

C_HORIZONTAL_
TAB

horizontal tab

                                                   '\t'

C_VERTICAL_TAB vertical tab  '\v'

Moreover, the following two named constants are defined:

Name  Type

C_NULL_PTR       C_PTR
C_NULL_FUNPTR    C_FUNPTR

Both are equivalent to the value NULL in C.

9.3 IEEE modules: IEEE_EXCEPTIONS, IEEE_ARITHMETIC, and
    IEEE_FEATURES

Standard: Fortran 2003 and later

   The IEEE_EXCEPTIONS, IEEE_ARITHMETIC, and IEEE_FEATURES intrinsic modules pro-

vide support for exceptions and IEEE arithmetic, as defined in Fortran 2003 and later standards, and the IEC 60559:1989 standard (Binary floating-point arithmetic for microprocessor systems). These modules are only provided on the following supported platforms:

• i386 and x86 64 processors

• platforms which use the GNU C Library (glibc)

• platforms with support for SysV/386 routines for floating point interface (including Solaris and BSDs)

• platforms with the AIX OS

For full compliance with the Fortran standards, code using the IEEE_EXCEPTIONS or IEEE_ARITHMETIC modules should be compiled with the following options: -fno-unsafe-

math-optimizations -frounding-math -fsignaling-nans.

9.4 OpenMP Modules OMP_LIB and OMP_LIB_KINDS

Standard: OpenMP Application Program Interface v4.5, OpenMP Application Program Interface v5.0 (partially supported) and OpenMP Application Program Interface v5.1 (partially supported).

The OpenMP Fortran runtime library routines are provided both in a form of two Fortran modules, named OMP_LIB and OMP_LIB_KINDS, and in a form of a Fortran include file named omp_lib.h. The procedures provided by OMP_LIB can be found in the Section “Introduction” in GNU Offloading and Multi Processing Runtime Library manual, the named constants defined in the modules are listed below.

For details refer to the actual OpenMP Application Program Interface v4.5

(https://www.openmp.org/wp-content/uploads/openmp-4.5.pdf) and OpenMP
Application Program Interface v5.0 (https://www.openmp.org/wp-content/uploads/
OpenMP-API-Specification-5.0.pdf).

OMP_LIB_KINDS provides the following scalar default-integer named constants:

omp_allocator_handle_kind
omp_alloctrait_key_kind
omp_alloctrait_val_kind
omp_depend_kind
omp_lock_kind
omp_lock_hint_kind
omp_nest_lock_kind
omp_pause_resource_kind
omp_memspace_handle_kind
omp_proc_bind_kind
omp_sched_kind
omp_sync_hint_kind

OMP_LIB provides the scalar default-integer named constant openmp_version with a value of the form yyyymm, where yyyy is the year and mm the month of the OpenMP version; for OpenMP v4.5 the value is 201511.

The following derived type:

omp_alloctrait

The following scalar integer named constants of the kind omp_sched_kind:

omp_sched_static
omp_sched_dynamic
omp_sched_guided
omp_sched_auto

And the following scalar integer named constants of the kind omp_proc_bind_kind:

omp_proc_bind_false
omp_proc_bind_true
omp_proc_bind_primary
omp_proc_bind_master
omp_proc_bind_close
omp_proc_bind_spread

The following scalar integer named constants are of the kind omp_lock_hint_kind:

omp_lock_hint_none
omp_lock_hint_uncontended
omp_lock_hint_contended
omp_lock_hint_nonspeculative
omp_lock_hint_speculative
omp_sync_hint_none
omp_sync_hint_uncontended
omp_sync_hint_contended
omp_sync_hint_nonspeculative
omp_sync_hint_speculative

And the following two scalar integer named constants are of the kind omp_pause_

resource_kind:

omp_pause_soft
omp_pause_hard

The following scalar integer named constants are of the kind omp_alloctrait_key_kind:

omp_atk_sync_hint
omp_atk_alignment
omp_atk_access
omp_atk_pool_size
omp_atk_fallback
omp_atk_fb_data
omp_atk_pinned
omp_atk_partition

The following scalar integer named constants are of the kind omp_alloctrait_val_kind:

          omp_alloctrait_key_kind:

omp_atv_default
omp_atv_false
omp_atv_true
omp_atv_contended
omp_atv_uncontended
omp_atv_serialized
omp_atv_sequential
omp_atv_private
omp_atv_all
omp_atv_thread
omp_atv_pteam
omp_atv_cgroup
omp_atv_default_mem_fb
omp_atv_null_fb
omp_atv_abort_fb
omp_atv_allocator_fb
omp_atv_environment
omp_atv_nearest
omp_atv_blocked

The following scalar integer named constants are of the kind omp_allocator_handle_

kind:
omp_null_allocator
omp_default_mem_alloc
omp_large_cap_mem_alloc
omp_const_mem_alloc
omp_high_bw_mem_alloc
omp_low_lat_mem_alloc
omp_cgroup_mem_alloc
omp_pteam_mem_alloc
omp_thread_mem_alloc

The following scalar integer named constants are of the kind omp_memspace_handle_

kind:
omp_default_mem_space
omp_large_cap_mem_space
omp_const_mem_space
omp_high_bw_mem_space
omp_low_lat_mem_space

9.5 OpenACC Module OPENACC

Standard: OpenACC Application Programming Interface v2.6

The OpenACC Fortran runtime library routines are provided both in a form of a Fortran 90 module, named OPENACC, and in form of a Fortran include file named openacc_lib.h. The procedures provided by OPENACC can be found in the Section “Introduction” in GNU Offloading and Multi Processing Runtime Library manual, the named constants defined in the modules are listed below.

For details refer to the actual OpenACC Application Programming Interface v2.6

(https://www.openacc.org/).

OPENACC provides the scalar default-integer named constant openacc_version with a value of the form yyyymm, where yyyy is the year and mm the month of the OpenACC version; for OpenACC v2.6 the value is 201711.

Contributing

Free software is only possible if people contribute to efforts to create it. We’re always in need of more people helping out with ideas and comments, writing documentation and contributing code.

If you want to contribute to GNU Fortran, have a look at the long lists of projects you can take on. Some of these projects are small, some of them are large; some are completely orthogonal to the rest of what is happening on GNU Fortran, but others are “mainstream” projects in need of enthusiastic hackers. All of these projects are important! We will eventually get around to the things here, but they are also things doable by someone who is willing and able.

Contributors to GNU Fortran

Most of the parser was hand-crafted by Andy Vaught, who is also the initiator of the whole project. Thanks Andy! Most of the interface with GCC was written by Paul Brook.

The following individuals have contributed code and/or ideas and significant help to the GNU Fortran project (in alphabetical order): − Janne Blomqvist − Steven Bosscher − Paul Brook − Tobias Burnus − François-Xavier Coudert − Bud Davis − Jerry DeLisle − Erik Edelmann − Bernhard Fischer − Daniel Franke − Richard Guenther − Richard Henderson − Katherine Holcomb − Jakub Jelinek − Niels Kristian Bech Jensen − Steven Johnson − Steven G. Kargl − Thomas Koenig − Asher Langton − H. J. Lu − Toon Moene − Brooks Moses − Andrew Pinski − Tim Prince − Christopher D. Rickett − Richard Sandiford − Tobias Schlüter − Roger Sayle − Paul Thomas − Andy Vaught − Feng Wang − Janus Weil − Daniel Kraft

The following people have contributed bug reports, smaller or larger patches, and much needed feedback and encouragement for the GNU Fortran project: − Bill Clodius − Dominique d’Humières − Kate Hedstrom − Erik Schnetter − Gerhard Steinmetz − Joost VandeVondele

Many other individuals have helped debug, test and improve the GNU Fortran compiler over the past few years, and we welcome you to do the same! If you already have done so, and you would like to see your name listed in the list above, please contact us.

Projects

Help build the test suite Solicit more code for donation to the test suite: the more extensive the testsuite, the smaller the risk of breaking things in the future! We can keep code private on request.

Bug hunting/squishing Find bugs and write more test cases! Test cases are especially very welcome, because it allows us to concentrate on fixing bugs instead of isolating them. Going through the bugzilla database at https://gcc.gnu.org/bugzilla/ to reduce testcases posted there and add more information (for example, for which version does the testcase work, for which versions does it fail?) is also very helpful.

Missing features For a larger project, consider working on the missing features required for Fortran language standards compliance (see Section 1.3 [Standards], page 3), or contributing to the implementation of extensions such as OpenMP (see Section 5.1.18 [OpenMP], page 54) or OpenACC (see Section 5.1.19 [OpenACC], page 55) that are under active development. Again, contributing test cases for these features is useful too!

GNU General Public License

Version 3, 29 June 2007 Copyright \begin{matrix}c \\ ◯ \end{matrix} 2007 Free Software Foundation, Inc. https://www.fsf.org

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The GNU General Public License is a free, copyleft license for software and other kinds of works.

The licenses for most software and other practical works are designed to take away your freedom to share and change the works. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change all versions of a program–to make sure it remains free software for all its users. We, the Free Software Foundation, use the GNU General Public License for most of our software; it applies also to any other work released this way by its authors. You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for them if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs, and that you know you can do these things.

To protect your rights, we need to prevent others from denying you these rights or asking you to surrender the rights. Therefore, you have certain responsibilities if you distribute copies of the software, or if you modify it: responsibilities to respect the freedom of others.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must pass on to the recipients the same freedoms that you received. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

Developers that use the GNU GPL protect your rights with two steps: (1) assert copyright on the software, and (2) offer you this License giving you legal permission to copy, distribute and/or modify it.

For the developers’ and authors’ protection, the GPL clearly explains that there is no warranty for this free software. For both users’ and authors’ sake, the GPL requires that modified versions be marked as changed, so that their problems will not be attributed erroneously to authors of previous versions.

Some devices are designed to deny users access to install or run modified versions of the software inside them, although the manufacturer can do so. This is fundamentally incompatible with the aim of protecting users’ freedom to change the software. The systematic pattern of such abuse occurs in the area of products for individuals to use, which is precisely where it is most unacceptable. Therefore, we have designed this version of the GPL to prohibit the practice for those products. If such problems arise substantially in other domains, we stand ready to extend this provision to those domains in future versions of the GPL, as needed to protect the freedom of users.

Finally, every program is threatened constantly by software patents. States should not allow patents to restrict development and use of software on general-purpose computers, but in those that do, we wish to avoid the special danger that patents applied to a free program could make it effectively proprietary. To prevent this, the GPL assures that patents cannot be used to render the program non-free.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS

0. Definitions. “This License” refers to version 3 of the GNU General Public License. “Copyright” also means copyright-like laws that apply to other kinds of works, such as semiconductor masks. “The Program” refers to any copyrightable work licensed under this License. Each licensee is addressed as “you”. “Licensees” and “recipients” may be individuals or organizations. To “modify” a work means to copy from or adapt all or part of the work in a fashion requiring copyright permission, other than the making of an exact copy. The resulting work is called a “modified version” of the earlier work or a work “based on” the earlier work. A “covered work” means either the unmodified Program or a work based on the Program. To “propagate” a work means to do anything with it that, without permission, would make you directly or secondarily liable for infringement under applicable copyright law, except executing it on a computer or modifying a private copy. Propagation includes copying, distribution (with or without modification), making available to the public, and in some countries other activities as well. To “convey” a work means any kind of propagation that enables other parties to make or receive copies. Mere interaction with a user through a computer network, with no transfer of a copy, is not conveying. An interactive user interface displays “Appropriate Legal Notices” to the extent that it includes a convenient and prominently visible feature that (1) displays an appropriate copyright notice, and (2) tells the user that there is no warranty for the work (except to the extent that warranties are provided), that licensees may convey the work under this License, and how to view a copy of this License. If the interface presents a list of user commands or options, such as a menu, a prominent item in the list meets this criterion.

1. Source Code. The “source code” for a work means the preferred form of the work for making modifications to it. “Object code” means any non-source form of a work. A “Standard Interface” means an interface that either is an official standard defined by a recognized standards body, or, in the case of interfaces specified for a particular programming language, one that is widely used among developers working in that language.

The “System Libraries” of an executable work include anything, other than the work as a whole, that (a) is included in the normal form of packaging a Major Component, but which is not part of that Major Component, and (b) serves only to enable use of the work with that Major Component, or to implement a Standard Interface for which an implementation is available to the public in source code form. A “Major Component”, in this context, means a major essential component (kernel, window system, and so on) of the specific operating system (if any) on which the executable work runs, or a compiler used to produce the work, or an object code interpreter used to run it.

The “Corresponding Source” for a work in object code form means all the source code needed to generate, install, and (for an executable work) run the object code and to modify the work, including scripts to control those activities. However, it does not include the work’s System Libraries, or general-purpose tools or generally available free programs which are used unmodified in performing those activities but which are not part of the work. For example, Corresponding Source includes interface definition files associated with source files for the work, and the source code for shared libraries and dynamically linked subprograms that the work is specifically designed to require, such as by intimate data communication or control flow between those subprograms and other parts of the work.

The Corresponding Source need not include anything that users can regenerate automatically from other parts of the Corresponding Source.

The Corresponding Source for a work in source code form is that same work.

2. Basic Permissions.

All rights granted under this License are granted for the term of copyright on the Program, and are irrevocable provided the stated conditions are met. This License explicitly affirms your unlimited permission to run the unmodified Program. The output from running a covered work is covered by this License only if the output, given its content, constitutes a covered work. This License acknowledges your rights of fair use or other equivalent, as provided by copyright law.

You may make, run and propagate covered works that you do not convey, without conditions so long as your license otherwise remains in force. You may convey covered works to others for the sole purpose of having them make modifications exclusively for you, or provide you with facilities for running those works, provided that you comply with the terms of this License in conveying all material for which you do not control copyright. Those thus making or running the covered works for you must do so exclusively on your behalf, under your direction and control, on terms that prohibit them from making any copies of your copyrighted material outside their relationship with you.

Conveying under any other circumstances is permitted solely under the conditions stated below. Sublicensing is not allowed; section 10 makes it unnecessary.

3. Protecting Users’ Legal Rights From Anti-Circumvention Law.

No covered work shall be deemed part of an effective technological measure under any applicable law fulfilling obligations under article 11 of the WIPO copyright treaty adopted on 20 December 1996, or similar laws prohibiting or restricting circumvention of such measures.

When you convey a covered work, you waive any legal power to forbid circumvention of technological measures to the extent such circumvention is effected by exercising rights under this License with respect to the covered work, and you disclaim any intention to limit operation or modification of the work as a means of enforcing, against the work’s users, your or third parties’ legal rights to forbid circumvention of technological measures.

4. Conveying Verbatim Copies.

You may convey verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice; keep intact all notices stating that this License and any non-permissive terms added in accord with section 7 apply to the code; keep intact all notices of the absence of any warranty; and give all recipients a copy of this License along with the Program. You may charge any price or no price for each copy that you convey, and you may offer support or warranty protection for a fee.

5. Conveying Modified Source Versions. You may convey a work based on the Program, or the modifications to produce it from the Program, in the form of source code under the terms of section 4, provided that you also meet all of these conditions: a. The work must carry prominent notices stating that you modified it, and giving a

relevant date. b. The work must carry prominent notices stating that it is released under this Li-

cense and any conditions added under section 7. This requirement modifies the requirement in section 4 to “keep intact all notices”.

c. You must license the entire work, as a whole, under this License to anyone who comes into possession of a copy. This License will therefore apply, along with any applicable section 7 additional terms, to the whole of the work, and all its parts, regardless of how they are packaged. This License gives no permission to license the work in any other way, but it does not invalidate such permission if you have separately received it.

d. If the work has interactive user interfaces, each must display Appropriate Legal Notices; however, if the Program has interactive interfaces that do not display Appropriate Legal Notices, your work need not make them do so.

A compilation of a covered work with other separate and independent works, which are not by their nature extensions of the covered work, and which are not combined with it such as to form a larger program, in or on a volume of a storage or distribution medium, is called an “aggregate” if the compilation and its resulting copyright are not used to limit the access or legal rights of the compilation’s users beyond what the individual works permit. Inclusion of a covered work in an aggregate does not cause this License to apply to the other parts of the aggregate.

6. Conveying Non-Source Forms.

You may convey a covered work in object code form under the terms of sections 4 and 5, provided that you also convey the machine-readable Corresponding Source under the terms of this License, in one of these ways: a. Convey the object code in, or embodied in, a physical product (including a phys-

ical distribution medium), accompanied by the Corresponding Source fixed on a durable physical medium customarily used for software interchange.

b. Convey the object code in, or embodied in, a physical product (including a physical distribution medium), accompanied by a written offer, valid for at least three years and valid for as long as you offer spare parts or customer support for that product model, to give anyone who possesses the object code either (1) a copy of the Corresponding Source for all the software in the product that is covered by this License, on a durable physical medium customarily used for software interchange, for a price no more than your reasonable cost of physically performing this conveying of source, or (2) access to copy the Corresponding Source from a network server at no charge.

c. Convey individual copies of the object code with a copy of the written offer to provide the Corresponding Source. This alternative is allowed only occasionally and noncommercially, and only if you received the object code with such an offer, in accord with subsection 6b.

d. Convey the object code by offering access from a designated place (gratis or for a charge), and offer equivalent access to the Corresponding Source in the same way through the same place at no further charge. You need not require recipients to copy the Corresponding Source along with the object code. If the place to copy the object code is a network server, the Corresponding Source may be on a different server (operated by you or a third party) that supports equivalent copying facilities, provided you maintain clear directions next to the object code saying where to find the Corresponding Source. Regardless of what server hosts the Corresponding Source, you remain obligated to ensure that it is available for as long as needed to satisfy these requirements.

e. Convey the object code using peer-to-peer transmission, provided you inform other peers where the object code and Corresponding Source of the work are being offered to the general public at no charge under subsection 6d.

A separable portion of the object code, whose source code is excluded from the Corresponding Source as a System Library, need not be included in conveying the object code work. A “User Product” is either (1) a “consumer product”, which means any tangible personal property which is normally used for personal, family, or household purposes, or (2) anything designed or sold for incorporation into a dwelling. In determining whether a product is a consumer product, doubtful cases shall be resolved in favor of coverage. For a particular product received by a particular user, “normally used” refers to a typical or common use of that class of product, regardless of the status of the particular user or of the way in which the particular user actually uses, or expects or is expected to use, the product. A product is a consumer product regardless of whether the product has substantial commercial, industrial or non-consumer uses, unless such uses represent the only significant mode of use of the product. “Installation Information” for a User Product means any methods, procedures, authorization keys, or other information required to install and execute modified versions of a covered work in that User Product from a modified version of its Corresponding Source.

The information must suffice to ensure that the continued functioning of the modified object code is in no case prevented or interfered with solely because modification has been made. If you convey an object code work under this section in, or with, or specifically for use in, a User Product, and the conveying occurs as part of a transaction in which the right of possession and use of the User Product is transferred to the recipient in perpetuity or for a fixed term (regardless of how the transaction is characterized), the Corresponding Source conveyed under this section must be accompanied by the Installation Information. But this requirement does not apply if neither you nor any third party retains the ability to install modified object code on the User Product (for example, the work has been installed in ROM). The requirement to provide Installation Information does not include a requirement to continue to provide support service, warranty, or updates for a work that has been modified or installed by the recipient, or for the User Product in which it has been modified or installed. Access to a network may be denied when the modification itself materially and adversely affects the operation of the network or violates the rules and protocols for communication across the network. Corresponding Source conveyed, and Installation Information provided, in accord with this section must be in a format that is publicly documented (and with an implementation available to the public in source code form), and must require no special password or key for unpacking, reading or copying.

7. Additional Terms. “Additional permissions” are terms that supplement the terms of this License by making exceptions from one or more of its conditions. Additional permissions that are applicable to the entire Program shall be treated as though they were included in this License, to the extent that they are valid under applicable law. If additional permissions apply only to part of the Program, that part may be used separately under those permissions, but the entire Program remains governed by this License without regard to the additional permissions. When you convey a copy of a covered work, you may at your option remove any additional permissions from that copy, or from any part of it. (Additional permissions may be written to require their own removal in certain cases when you modify the work.) You may place additional permissions on material, added by you to a covered work, for which you have or can give appropriate copyright permission. Notwithstanding any other provision of this License, for material you add to a covered work, you may (if authorized by the copyright holders of that material) supplement the terms of this License with terms: a. Disclaiming warranty or limiting liability differently from the terms of sections 15

and 16 of this License; or b. Requiring preservation of specified reasonable legal notices or author attributions

in that material or in the Appropriate Legal Notices displayed by works containing it; or

c. Prohibiting misrepresentation of the origin of that material, or requiring that modified versions of such material be marked in reasonable ways as different from the original version; or d. Limiting the use for publicity purposes of names of licensors or authors of the

material; or e. Declining to grant rights under trademark law for use of some trade names, trade-

marks, or service marks; or f. Requiring indemnification of licensors and authors of that material by anyone who

conveys the material (or modified versions of it) with contractual assumptions of liability to the recipient, for any liability that these contractual assumptions directly impose on those licensors and authors.

All other non-permissive additional terms are considered “further restrictions” within the meaning of section 10. If the Program as you received it, or any part of it, contains a notice stating that it is governed by this License along with a term that is a further restriction, you may remove that term. If a license document contains a further restriction but permits relicensing or conveying under this License, you may add to a covered work material governed by the terms of that license document, provided that the further restriction does not survive such relicensing or conveying. If you add terms to a covered work in accord with this section, you must place, in the relevant source files, a statement of the additional terms that apply to those files, or a notice indicating where to find the applicable terms. Additional terms, permissive or non-permissive, may be stated in the form of a separately written license, or stated as exceptions; the above requirements apply either way.

8. Termination.

You may not propagate or modify a covered work except as expressly provided under this License. Any attempt otherwise to propagate or modify it is void, and will automatically terminate your rights under this License (including any patent licenses granted under the third paragraph of section 11). However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, you do not qualify to receive new licenses for the same material under section 10.

9. Acceptance Not Required for Having Copies. You are not required to accept this License in order to receive or run a copy of the Program. Ancillary propagation of a covered work occurring solely as a consequence of using peer-to-peer transmission to receive a copy likewise does not require acceptance.

However, nothing other than this License grants you permission to propagate or modify any covered work. These actions infringe copyright if you do not accept this License. Therefore, by modifying or propagating a covered work, you indicate your acceptance of this License to do so.

10. Automatic Licensing of Downstream Recipients. Each time you convey a covered work, the recipient automatically receives a license from the original licensors, to run, modify and propagate that work, subject to this License. You are not responsible for enforcing compliance by third parties with this License. An “entity transaction” is a transaction transferring control of an organization, or substantially all assets of one, or subdividing an organization, or merging organizations. If propagation of a covered work results from an entity transaction, each party to that transaction who receives a copy of the work also receives whatever licenses to the work the party’s predecessor in interest had or could give under the previous paragraph, plus a right to possession of the Corresponding Source of the work from the predecessor in interest, if the predecessor has it or can get it with reasonable efforts. You may not impose any further restrictions on the exercise of the rights granted or affirmed under this License. For example, you may not impose a license fee, royalty, or other charge for exercise of rights granted under this License, and you may not initiate litigation (including a cross-claim or counterclaim in a lawsuit) alleging that any patent claim is infringed by making, using, selling, offering for sale, or importing the Program or any portion of it.

11. Patents.

A “contributor” is a copyright holder who authorizes use under this License of the Program or a work on which the Program is based. The work thus licensed is called the contributor’s “contributor version”. A contributor’s “essential patent claims” are all patent claims owned or controlled by the contributor, whether already acquired or hereafter acquired, that would be infringed by some manner, permitted by this License, of making, using, or selling its contributor version, but do not include claims that would be infringed only as a consequence of further modification of the contributor version. For purposes of this definition, “control” includes the right to grant patent sublicenses in a manner consistent with the requirements of this License. Each contributor grants you a non-exclusive, worldwide, royalty-free patent license under the contributor’s essential patent claims, to make, use, sell, offer for sale, import and otherwise run, modify and propagate the contents of its contributor version. In the following three paragraphs, a “patent license” is any express agreement or commitment, however denominated, not to enforce a patent (such as an express permission to practice a patent or covenant not to sue for patent infringement). To “grant” such a patent license to a party means to make such an agreement or commitment not to enforce a patent against the party. If you convey a covered work, knowingly relying on a patent license, and the Corresponding Source of the work is not available for anyone to copy, free of charge and under the terms of this License, through a publicly available network server or other readily accessible means, then you must either (1) cause the Corresponding Source to be so available, or (2) arrange to deprive yourself of the benefit of the patent license for this particular work, or (3) arrange, in a manner consistent with the requirements of this License, to extend the patent license to downstream recipients. “Knowingly relying” means you have actual knowledge that, but for the patent license, your conveying the covered work in a country, or your recipient’s use of the covered work in a country, would infringe one or more identifiable patents in that country that you have reason to believe are valid.

If, pursuant to or in connection with a single transaction or arrangement, you convey, or propagate by procuring conveyance of, a covered work, and grant a patent license to some of the parties receiving the covered work authorizing them to use, propagate, modify or convey a specific copy of the covered work, then the patent license you grant is automatically extended to all recipients of the covered work and works based on it.

A patent license is “discriminatory” if it does not include within the scope of its coverage, prohibits the exercise of, or is conditioned on the non-exercise of one or more of the rights that are specifically granted under this License. You may not convey a covered work if you are a party to an arrangement with a third party that is in the business of distributing software, under which you make payment to the third party based on the extent of your activity of conveying the work, and under which the third party grants, to any of the parties who would receive the covered work from you, a discriminatory patent license (a) in connection with copies of the covered work conveyed by you (or copies made from those copies), or (b) primarily for and in connection with specific products or compilations that contain the covered work, unless you entered into that arrangement, or that patent license was granted, prior to 28 March 2007.

Nothing in this License shall be construed as excluding or limiting any implied license or other defenses to infringement that may otherwise be available to you under applicable patent law.

12. No Surrender of Others’ Freedom.

If conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot convey a covered work so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not convey it at all. For example, if you agree to terms that obligate you to collect a royalty for further conveying from those to whom you convey the Program, the only way you could satisfy both those terms and this License would be to refrain entirely from conveying the Program.

13. Use with the GNU Affero General Public License.

Notwithstanding any other provision of this License, you have permission to link or combine any covered work with a work licensed under version 3 of the GNU Affero General Public License into a single combined work, and to convey the resulting work. The terms of this License will continue to apply to the part which is the covered work, but the special requirements of the GNU Affero General Public License, section 13, concerning interaction through a network will apply to the combination as such.

14. Revised Versions of this License.

The Free Software Foundation may publish revised and/or new versions of the GNU General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

Each version is given a distinguishing version number. If the Program specifies that a certain numbered version of the GNU General Public License “or any later version” applies to it, you have the option of following the terms and conditions either of that numbered version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of the GNU General Public License, you may choose any version ever published by the Free Software Foundation.

If the Program specifies that a proxy can decide which future versions of the GNU General Public License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Program.

Later license versions may give you additional or different permissions. However, no additional obligations are imposed on any author or copyright holder as a result of your choosing to follow a later version.

15. Disclaimer of Warranty.

THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EXTENT PER-MITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EX-PRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFEC-TIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

16. Limitation of Liability.

IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MODIFIES AND/OR CONVEYS THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, IN-CIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUS-TAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAM-AGES.

17. Interpretation of Sections 15 and 16.

If the disclaimer of warranty and limitation of liability provided above cannot be given local legal effect according to their terms, reviewing courts shall apply local law that most closely approximates an absolute waiver of all civil liability in connection with the Program, unless a warranty or assumption of liability accompanies a copy of the Program in return for a fee.

END OF TERMS AND CONDITIONS

How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively state the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

      one line to give the program's name and a brief idea of what it does.
      Copyright (C) year name of author
      This program is free software: you can redistribute it and/or modify
      it under the terms of the GNU General Public License as published by
      the Free Software Foundation, either version 3 of the License, or (at
      your option) any later version.
      This program is distributed in the hope that it will be useful, but
      WITHOUT ANY WARRANTY; without even the implied warranty of
      MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the GNU
      General Public License for more details.
      You should have received a copy of the GNU General Public License
      along with this program.  If not, see https://www.gnu.org/licenses/.

Also add information on how to contact you by electronic and paper mail.

If the program does terminal interaction, make it output a short notice like this when it starts in an interactive mode:

      program Copyright (C) year name of author
      This program comes with ABSOLUTELY NO WARRANTY; for details type `show w'.
      This is free software, and you are welcome to redistribute it
      under certain conditions; type `show c' for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, your program’s commands might be different; for a GUI interface, you would use an “about box”.

You should also get your employer (if you work as a programmer) or school, if any, to sign a “copyright disclaimer” for the program, if necessary. For more information on this, and how to apply and follow the GNU GPL, see https://www.gnu.org/licenses/.

The GNU General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Lesser General Public License instead of this License. But first, please read

https://www.gnu.org/licenses/why-not-lgpl.html.

GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 https://www.fsf.org

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See https://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled `  ` GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with...Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Funding Free Software

If you want to have more free software a few years from now, it makes sense for you to help encourage people to contribute funds for its development. The most effective approach known is to encourage commercial redistributors to donate.

Users of free software systems can boost the pace of development by encouraging for-a-fee distributors to donate part of their selling price to free software developers—the Free Software Foundation, and others.

The way to convince distributors to do this is to demand it and expect it from them. So when you compare distributors, judge them partly by how much they give to free software development. Show distributors they must compete to be the one who gives the most.

To make this approach work, you must insist on numbers that you can compare, such as, “We will donate ten dollars to the Frobnitz project for each disk sold.” Don’t be satisfied with a vague promise, such as “A portion of the profits are donated,” since it doesn’t give a basis for comparison.

Even a precise fraction “of the profits from this disk” is not very meaningful, since creative accounting and unrelated business decisions can greatly alter what fraction of the sales price counts as profit. If the price you pay is $50, ten percent of the profit is probably less than a dollar; it might be a few cents, or nothing at all.

Some redistributors do development work themselves. This is useful too; but to keep everyone honest, you need to inquire how much they do, and what kind. Some kinds of development make much more long-term difference than others. For example, maintaining a separate version of a program contributes very little; maintaining the standard version of a program for the whole community contributes much. Easy new ports contribute little, since someone else would surely do them; difficult ports such as adding a new CPU to the GNU Compiler Collection contribute more; major new features or packages contribute the most.

By establishing the idea that supporting further development is “the proper thing to do” when distributing free software for a fee, we can assure a steady flow of resources into making more free software.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994 Free Software Foundation, Inc.

Verbatim copying and redistribution of this section is permitted without royalty; alteration is not permitted.

Option Index

gfortran’s command line options are indexed here without any initial ‘-’ or ‘--’. Where an option has both positive and negative forms (such as -foption and -fno-option), relevant entries in the manual are indexed under the most appropriate form; it may sometimes be useful to look up both forms.

A

A-predicate=answer . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 allow-invalid-boz . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Apredicate=answer . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

B

backslash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

C

c-prototypes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 c-prototypes-external . . . . . . . . . . . . . . . . . . . . . . . 33 cpp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 CC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

D

dD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 dI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 dM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Dname. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Dname=definition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 dN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 dU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

F

faggressive-function-elimination. . . . . . . . . . . 32

falign-commons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 fall-intrinsics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

fallow-argument-mismatch . . . . . . . . . . . . . . . . . . . . . 9

fblas-matmul-limit . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 fbounds-check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 fcheck. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

fcheck-array-temporaries . . . . . . . . . . . . . . . . . . . . 28

fcoarray . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 fconvert=conversion. . . . . . . . . . . . . . . . . . . . . . . . . . 24 fcray-pointer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 fd-lines-as-code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 fd-lines-as-comments. . . . . . . . . . . . . . . . . . . . . . . . . . 9 fdebug-aux-vars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 fdec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 fdec-blank-format-item . . . . . . . . . . . . . . . . . . . . . . 10 fdec-char-conversions . . . . . . . . . . . . . . . . . . . . . . . . 9 fdec-format-defaults . . . . . . . . . . . . . . . . . . . . . . . . 10 fdec-include. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 fdec-intrinsic-ints. . . . . . . . . . . . . . . . . . . . . . . . . . 10 fdec-math . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 fdec-static . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 fdec-structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 fdefault-double-8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 fdefault-integer-8 . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 fdefault-real-10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 fdefault-real-16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 fdefault-real-8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 fdollar-ok . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 fdump-fortran-global . . . . . . . . . . . . . . . . . . . . . . . . 22

fdump-fortran-optimized . . . . . . . . . . . . . . . . . . . . . 22

fdump-fortran-original . . . . . . . . . . . . . . . . . . . . . . 22 fdump-parse-tree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 fexternal-blas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 ff2c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ffixed-form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ffixed-line-length-n . . . . . . . . . . . . . . . . . . . . . . . . 11 ffpe-summary=list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ffpe-trap=list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 ffree-form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ffree-line-length-n. . . . . . . . . . . . . . . . . . . . . . . . . . 11 fimplicit-none . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 finit-character . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 finit-derived . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 finit-integer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 finit-local-zero . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 finit-logical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 finit-real . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 finline-arg-packing. . . . . . . . . . . . . . . . . . . . . . . . . . 30 finline-matmul-limit . . . . . . . . . . . . . . . . . . . . . . . . 30 finteger-4-integer-8 . . . . . . . . . . . . . . . . . . . . . . . . 12

fintrinsic-modules-path dir . . . . . . . . . . . . . . . . . 24

fmax-array-constructor . . . . . . . . . . . . . . . . . . . . . . 28 fmax-errors=n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

fmax-identifier-length=n . . . . . . . . . . . . . . . . . . . . 11

fmax-stack-var-size. . . . . . . . . . . . . . . . . . . . . . . . . . 29

fmax-subrecord-length=length . . . . . . . . . . . . . . . 24

fmodule-private . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 fno-automatic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 fno-backtrace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 fno-protect-parens . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 fno-underscoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 fopenacc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 fopenmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 fpack-derived . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 fpad-source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 fpp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 frange-check. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 freal-4-real-10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 freal-4-real-16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 freal-4-real-8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 freal-8-real-10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 freal-8-real-16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 freal-8-real-4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 frealloc-lhs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 frecord-marker=length . . . . . . . . . . . . . . . . . . . . . . . 24 frecursive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 frepack-arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

frontend-loop-interchange. . . . . . . . . . . . . . . . . . . 32

frontend-optimize . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 fsecond-underscore . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 fshort-enums. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 fsign-zero . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 fstack-arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 fsyntax-only. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ftest-forall-temp . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 fworking-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

H

H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

I

Idir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 idirafter dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 imultilib dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 iprefix prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 iquote dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 isysroot dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 isystem dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

J

Jdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

M

Mdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

N

nostdinc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

P

pedantic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 pedantic-errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

S

static-libgfortran . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 static-libquadmath . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 std=std option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

T

tail-call-workaround . . . . . . . . . . . . . . . . . . . . . . . . 28

U

Uname. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 undef. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

W

Waliasing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Walign-commons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Wall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Wampersand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Warray-temporaries . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Wc-binding-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Wcharacter-truncation . . . . . . . . . . . . . . . . . . . . . . . 18 Wcompare-reals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Wconversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Wconversion-extra . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Wdo-subscript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Werror. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Wextra. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

Wfrontend-loop-interchange. . . . . . . . . . . . . . . . . . 19

Wfunction-elimination . . . . . . . . . . . . . . . . . . . . . . . 20 Wimplicit-interface. . . . . . . . . . . . . . . . . . . . . . . . . . 19 Wimplicit-procedure. . . . . . . . . . . . . . . . . . . . . . . . . . 19 Winteger-division . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Wintrinsic-shadow . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Wintrinsics-std . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Wline-truncation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Woverwrite-recursive . . . . . . . . . . . . . . . . . . . . . . . . 19 Wpedantic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Wreal-q-constant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Wrealloc-lhs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Wrealloc-lhs-all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Wsurprising . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Wtabs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Wtargt-lifetime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Wundefined-do-loop . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Wunderflow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Wunused-dummy-argument . . . . . . . . . . . . . . . . . . . . . . 20 Wunused-parameter . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Wuse-without-only . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Wzerotrip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

Keyword Index

$

$ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

%

%LOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 %REF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 %VAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

&

& . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

_gfortran_set_args . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 _gfortran_set_convert . . . . . . . . . . . . . . . . . . . . . . . 79 _gfortran_set_fpe . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

_gfortran_set_max_subrecord_length . . . . . . . . 80

_gfortran_set_options . . . . . . . . . . . . . . . . . . . . . . . 77

_gfortran_set_record_marker . . . . . . . . . . . . . . . . 79

A

ABORT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 absolute value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 ABS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 ACCESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 ACHAR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 ACOS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 ACOSD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 ACOSH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 adjust string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 ADJUSTL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 ADJUSTR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 AIMAG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 AINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 ALARM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 ALGAMA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 aliasing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 alignment of COMMON blocks . . . . . . . . . . . . . . . . . 20, 31 all warnings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 ALLOCATED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 allocation, moving . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 allocation, status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 ALOG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 ALOG10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 AMAX0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 AMAX1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 AMIN0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 AMIN1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 AMOD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 AND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 ANINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 ANY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 area hyperbolic cosine . . . . . . . . . . . . . . . . . . . . . . . . 115 area hyperbolic sine. . . . . . . . . . . . . . . . . . . . . . . . . . . 124 area hyperbolic tangent . . . . . . . . . . . . . . . . . . . . . . . 129 argument list functions. . . . . . . . . . . . . . . . . . . . . . . . . 55 arguments, to program. . . . . 157, 192, 193, 194, 204 array descriptor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 array, add elements . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 array, AND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 array, apply condition . . . . . . . . . . . . . . . . . . . . 119, 122 array, bounds checking . . . . . . . . . . . . . . . . . . . . . . . . . 27 array, change dimensions . . . . . . . . . . . . . . . . . . . . . . 259 array, combine arrays . . . . . . . . . . . . . . . . . . . . . . . . . 236 array, condition testing. . . . . . . . . . . . . . . . . . . 119, 122 array, conditionally add elements. . . . . . . . . . . . . . 277 array, conditionally count elements . . . . . . . . . . . . 163 array, conditionally multiply elements . . . . . . . . . 252 array, contiguity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 array, count elements . . . . . . . . . . . . . . . . . . . . . . . . . 271 array, duplicate dimensions . . . . . . . . . . . . . . . . . . . 273 array, duplicate elements . . . . . . . . . . . . . . . . . . . . . . 273 array, element counting . . . . . . . . . . . . . . . . . . . . . . . 163 array, gather elements . . . . . . . . . . . . . . . . . . . . . . . . 248 array, increase dimension . . . . . . . . . . . . . . . . . 273, 289 array, indices of type real . . . . . . . . . . . . . . . . . . . . . . 50 array, location of maximum element . . . . . . . . . . . 234 array, location of minimum element . . . . . . . . . . . 238 array, lower bound . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 array, maximum value . . . . . . . . . . . . . . . . . . . . . . . . 235 array, merge arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 array, minimum value . . . . . . . . . . . . . . . . . . . . . . . . . 239 array, multiply elements . . . . . . . . . . . . . . . . . . . . . . 252 array, number of elements . . . . . . . . . . . . . . . . 163, 271 array, OR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 array, packing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 array, parity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 array, permutation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 array, product . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 array, reduce dimension . . . . . . . . . . . . . . . . . . . . . . . 248 array, rotate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 array, scatter elements . . . . . . . . . . . . . . . . . . . . . . . . 289 array, shape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 array, shift. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 array, shift circularly. . . . . . . . . . . . . . . . . . . . . . . . . . 164 array, size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 array, sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 array, transmogrify . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 array, transpose. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 array, unpacking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 array, upper bound . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 array, XOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 ASCII collating sequence . . . . . . . . . . . . . . . . . 113, 201 ASIN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 ASIND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 ASINH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 ASSOCIATED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 association status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 association status, C pointer . . . . . . . . . . . . . . . . . . 144 assumed-rank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 assumed-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 asynchronous I/O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 ATAN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 ATAN2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 ATAN2D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 ATAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 ATANH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 Atomic subroutine, add . . . . . . . . . . . . . . . . . . . . . . . 129 Atomic subroutine, ADD with fetch . . . . . . . . . . . 132 Atomic subroutine, AND. . . . . . . . . . . . . . . . . . . . . . 130 Atomic subroutine, AND with fetch . . . . . . . . . . . 133 Atomic subroutine, compare and swap . . . . . . . . 131 Atomic subroutine, define . . . . . . . . . . . . . . . . . . . . . 132 Atomic subroutine, OR . . . . . . . . . . . . . . . . . . . . . . . 135 Atomic subroutine, OR with fetch. . . . . . . . . . . . . 134 Atomic subroutine, reference . . . . . . . . . . . . . . . . . . 136 Atomic subroutine, XOR. . . . . . . . . . . . . . . . . . . . . . 137 Atomic subroutine, XOR with fetch . . . . . . . . . . . 135 ATOMIC_ADD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 ATOMIC_AND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 ATOMIC_DEFINE. . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 132 ATOMIC_FETCH_ADD . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 ATOMIC_FETCH_AND . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 ATOMIC_FETCH_OR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 ATOMIC_FETCH_XOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 ATOMIC_OR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 ATOMIC_REF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 ATOMIC_XOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 Authors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 AUTOMATIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

B

BABS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 backslash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 BACKSPACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 backtrace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23, 138 BACKTRACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 base 10 logarithm function . . . . . . . . . . . . . . . . . . . . 227 BBCLR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 BBITS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 BBSET. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 BBTEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 BESJ0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 BESJ1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 BESJN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Bessel function, first kind . . . . . . . . . . . . . . . . 138, 139 Bessel function, second kind . . . . . . . . . . . . . . 140, 141 BESSEL_J0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 BESSEL_J1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 BESSEL_JN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 BESSEL_Y0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140

BESSEL_Y1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140
BESSEL_YN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141
BESY0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140
BESY1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140
BESYN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141
BGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142
BGT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142
BIAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202
BIEOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208
binary representation . . . . . . . . . . . . . . . . . . . . . . . . . 250
BIOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212
bit intrinsics checking . . . . . . . . . . . . . . . . . . . . . . . . . . 27
BIT_SIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142
BITEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144
bits set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250
bits, AND of array elements. . . . . . . . . . . . . . . . . . . 201
bits, clear. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204
bits, extract . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205
bits, get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205
bits, merge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237
bits, move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242, 285
bits, negate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245
bits, number of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142
bits, OR of array elements . . . . . . . . . . . . . . . . . . . . 203
bits, set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206
bits, shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216
bits, shift circular . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217
bits, shift left . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228, 267
bits, shift right . . . . . . . . . . . . . . . . . . . . . . 260, 267, 268
bits, testing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144
bits, unset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204
bits, XOR of array elements. . . . . . . . . . . . . . . . . . . 212
bitwise comparison . . . . . . . . . . . . . . . . . . . . . . . 142, 143
bitwise logical and . . . . . . . . . . . . . . . . . . . . . . . 120, 202
bitwise logical exclusive or . . . . . . . . . . . . . . . 208, 290
bitwise logical not . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245
bitwise logical or . . . . . . . . . . . . . . . . . . . . . . . . . 212, 247
BJTEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144
BKTEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144
BLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143
BLT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143
BMOD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240
BMVBITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242
BNOT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245
bounds checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27
BOZ literal constants . . . . . . . . . . . . . . . . . . . . . . . . . . 50
BSHFT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216
BSHFTC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217
BTEST. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144

Keyword Index 329

C

C derived type and struct interoperability . . . . . . 69 C interoperability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 C intrinsic type interoperability . . . . . . . . . . . . . . . . 69 C pointers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 C procedure interoperability . . . . . . . . . . . . . . . . . . . 70 C variable interoperability . . . . . . . . . . . . . . . . . . . . . 70 C_ASSOCIATED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 C_F_POINTER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 C_F_PROCPOINTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 C_FUNLOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 C_LOC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 C_SIZEOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 CABS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 calling convention . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 CARRIAGECONTROL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 CCOS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 CCOSD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 CDABS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 CDCOS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 CDCOSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 CDEXP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 CDLOG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 CDSIN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 CDSIND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 CDSQRT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 ceiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121, 149 CEILING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 CEXP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 character kind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 character set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 CHAR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 CHDIR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 checking array temporaries . . . . . . . . . . . . . . . . . . . . . 27 checking subscripts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 CHMOD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 clock ticks. . . . . . . . . . . . . . . . . . . . . . . . . . . 235, 236, 279 CLOG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 CMPLX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 CO_BROADCAST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 CO_MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 CO_MIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 CO_REDUCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 CO_SUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 Coarray, gfortran caf atomic cas . . . . . . . . . . . . . 105 Coarray, gfortran caf atomic define . . . . . . . . . . 104 Coarray, gfortran caf atomic op . . . . . . . . . . . . . 106 Coarray, gfortran caf atomic ref . . . . . . . . . . . . . 104 Coarray, gfortran caf co broadcast . . . . . . . . . . . 106 Coarray, gfortran caf co max . . . . . . . . . . . . . . . . 107 Coarray, gfortran caf co min. . . . . . . . . . . . . . . . . 107 Coarray, gfortran caf co reduce . . . . . . . . . . . . . . 108 Coarray, gfortran caf co sum . . . . . . . . . . . . . . . . 108 Coarray, gfortran caf deregister . . . . . . . . . . . . . . . 91 Coarray, gfortran caf error stop. . . . . . . . . . . . . . 103 Coarray, gfortran caf error stop str . . . . . . . . . . 104 Coarray, gfortran caf event post . . . . . . . . . . . . . 100 Coarray, gfortran caf event query . . . . . . . . . . . . 102 Coarray, gfortran caf event wait . . . . . . . . . . . . . 101 Coarray, gfortran caf fail image. . . . . . . . . . . . . . 104 Coarray, gfortran caf failed images. . . . . . . . . . . . 89 Coarray, gfortran caf finish . . . . . . . . . . . . . . . . . . . 88 Coarray, gfortran caf get. . . . . . . . . . . . . . . . . . . . . . 93 Coarray, gfortran caf get by ref. . . . . . . . . . . . . . . 97 Coarray, gfortran caf image status . . . . . . . . . . . . 89 Coarray, gfortran caf init . . . . . . . . . . . . . . . . . . . . . 88 Coarray, gfortran caf is present . . . . . . . . . . . . . . . 92 Coarray, gfortran caf lock. . . . . . . . . . . . . . . . . . . . . 99 Coarray, gfortran caf num images. . . . . . . . . . . . . 89 Coarray, gfortran caf register . . . . . . . . . . . . . . . . . 90 Coarray, gfortran caf send . . . . . . . . . . . . . . . . . . . . 92 Coarray, gfortran caf send by ref . . . . . . . . . . . . . 96 Coarray, gfortran caf sendget . . . . . . . . . . . . . . . . . 94 Coarray, gfortran caf sendget by ref . . . . . . . . . . 98 Coarray, gfortran caf stopped images . . . . . . . . . 90 Coarray, gfortran caf sync all . . . . . . . . . . . . . . . . 102 Coarray, gfortran caf sync images. . . . . . . . . . . . 102 Coarray, gfortran caf sync memory . . . . . . . . . . 103 Coarray, gfortran caf this image . . . . . . . . . . . . . . 88 Coarray, gfortran caf unlock . . . . . . . . . . . . . . . . . 100 coarray, IMAGE_INDEX . . . . . . . . . . . . . . . . . . . . . . . . . 209 coarray, lower bound . . . . . . . . . . . . . . . . . . . . . . . . . . 220 coarray, NUM_IMAGES . . . . . . . . . . . . . . . . . . . . . . . . . . 246 coarray, THIS_IMAGE . . . . . . . . . . . . . . . . . . . . . . . . . . 281 coarray, upper bound . . . . . . . . . . . . . . . . . . . . . . . . . 287 coarrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 Coarrays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 code generation, conventions . . . . . . . . . . . . . . . . . . . 25 collating sequence, ASCII . . . . . . . . . . . . . . . . . 113, 201 Collectives, generic reduction . . . . . . . . . . . . . . . . . 155 Collectives, maximal value . . . . . . . . . . . . . . . . . . . . 153 Collectives, minimal value. . . . . . . . . . . . . . . . . . . . . 154 Collectives, sum of values . . . . . . . . . . . . . . . . . . . . . 156 Collectives, value broadcasting . . . . . . . . . . . . . . . . 152 command line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 command options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 command-line arguments… 157, 192, 193, 194, 204 command-line arguments, number of . . . . . 157, 204 COMMAND_ARGUMENT_COUNT . . . . . . . . . . . . . . . . . . . . . 157 COMMON. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 compiler flags inquiry function . . . . . . . . . . . . . . . . 157 compiler, name and version . . . . . . . . . . . . . . . . . . . 158 COMPILER_OPTIONS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 COMPILER_VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 complex conjugate . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 Complex function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 complex numbers, conversion to . . . . . 152, 158, 167 complex numbers, imaginary part . . . . . . . . . . . . . 117 complex numbers, real part . . . . . . . . . . . . . . 171, 257 COMPLEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 Conditional compilation. . . . . . . . . . . . . . . . . . . . . . . . 14 conditional compilation . . . . . . . . . . . . . . . . . . . . . . . . . 3 CONJG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 consistency, durability . . . . . . . . . . . . . . . . . . . . . . . . . 43 Contributing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 Contributors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 conversion, to character . . . . . . . . . . . . . . . . . . . 51, 149 conversion, to complex . . . . . . . . . . . . . . 152, 158, 167 conversion, to integer . . . . . . . 50, 201, 206, 210, 211 conversion, to logical . . . . . . . . . . . . . . . . . . . . . . 50, 228 conversion, to real. . . . . . . . . . . . . . . . . . . . . . . . 167, 257 conversion, to string . . . . . . . . . . . . . . . . . . . . . . . . . . 165 CONVERT specifier. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 core, dump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 COS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 COSD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 COSH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 cosine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 cosine, degrees . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 cosine, hyperbolic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 cosine, hyperbolic, inverse. . . . . . . . . . . . . . . . . . . . . 115 cosine, inverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 cosine, inverse, degrees . . . . . . . . . . . . . . . . . . . . . . . . 114 COTAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 cotangent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 cotangent, degrees . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 COTAN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 COUNT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 CPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 CPU_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 CSHIFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 CSIN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 CSIND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 CSQRT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 CTIME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 current date . . . . . . . . . . . . . . . . . . . . . . . . . 166, 181, 207 current time . . . . . . . . . . . . . . . 166, 181, 218, 282, 283

D

DABS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 DACOS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 DACOSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 DACOSH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 DASIN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 DASIND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 DASINH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 DATAN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 DATAN2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 DATAN2D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 DATAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 DATANH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 date, current . . . . . . . . . . . . . . . . . . . . . . . . 166, 181, 207 DATE_AND_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 DBESJ0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 DBESJ1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 DBESJN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 DBESY0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 DBESY1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 DBESYN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 DBLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 DCMPLX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 DCONJG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 DCOS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 DCOSD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 DCOSH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 DCOTAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 DCOTAND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 DDIM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 debugging information options . . . . . . . . . . . . . . . . . 22 debugging, preprocessor . . . . . . . . . . . . . . . . . . . . 14, 15 DECODE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 delayed execution . . . . . . . . . . . . . . . . . . . . . . . . 118, 273 derived type interoperability with C . . . . . . . . . . . . 69 DEXP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 DFLOAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 DGAMMA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 dialect options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 DIGITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 DIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 DIMAG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 DINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 directive, INCLUDE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 directory, options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 directory, search paths for inclusion . . . . . . . . . . . . 23 division, modulo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 division, remainder . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 DLGAMA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 DLOG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 DLOG10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 DMAX1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 DMIN1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 DMOD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 DNINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 dope vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 dot product. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 DOT_PRODUCT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 DPROD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 DREAL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 DSHIFTL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 DSHIFTR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 DSIGN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 DSIN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 DSIND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 DSINH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 DSQRT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 DTAN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 DTAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 DTANH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 DTIME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 dummy argument, unused . . . . . . . . . . . . . . . . . . . . . . 20

E

elapsed time. . . . . . . . . . . . . . . . . . . . . . . . . 172, 262, 263 Elimination of functions with

identical argument lists . . . . . . . . . . . . . . . . . . . . . . 32 ENCODE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 environment variable. . . . . . . . . . . . . . 33, 35, 195, 196 EOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 EOSHIFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 EPSILON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 ERF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 ERFC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 ERFC_SCALED. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 error function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 error function, complementary . . . . . . . . . . . . . . . . 175 error function, complementary,

exponentially-scaled . . . . . . . . . . . . . . . . . . . . . . . . 176 errors, limiting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 escape characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ETIME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 Euclidean distance . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 Euclidean vector norm . . . . . . . . . . . . . . . . . . . . . . . . 245 EVENT_QUERY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 Events, EVENT QUERY . . . . . . . . . . . . . . . . . . . . . 177 EXECUTE_COMMAND_LINE . . . . . . . . . . . . . . . . . . . . . . . 178 EXIT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 exponent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 exponential function . . . . . . . . . . . . . . . . . . . . . . . . . . 179 exponential function, inverse . . . . . . . . . . . . . 226, 227 expression size . . . . . . . . . . . . . . . . . . . . . . . . . . . 148, 272 EXP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 EXPONENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 EXTENDS_TYPE_OF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 extensions, implemented . . . . . . . . . . . . . . . . . . . . . . . 47 extensions, not implemented . . . . . . . . . . . . . . . . . . . 65 extra warnings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

F

f2c calling convention . . . . . . . . . . . . . . . . . . . . . . 25, 26 Factorial function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 FDATE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 FDL, GNU Free Documentation License . . . . . . 315 FGET. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 FGETC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 file format, fixed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 11 file format, free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 11 file operation, file number . . . . . . . . . . . . . . . . . . . . . 186 file operation, flush . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 file operation, position . . . . . . . . . . . . . . . . . . . 189, 191 file operation, read character . . . . . . . . . . . . . 182, 183 file operation, seek . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 file operation, write character . . . . . . . . . . . . 186, 187 file system, access mode. . . . . . . . . . . . . . . . . . . . . . . 112 file system, change access mode . . . . . . . . . . . . . . . 151 file system, create link . . . . . . . . . . . . . . . . . . . 223, 278 file system, file creation mask . . . . . . . . . . . . . . . . . 288 file system, file status. . . . . . . . . . . . . . . . 190, 229, 275 file system, hard link. . . . . . . . . . . . . . . . . . . . . . . . . . 223 file system, remove file . . . . . . . . . . . . . . . . . . . . . . . . 288 file system, rename file . . . . . . . . . . . . . . . . . . . . . . . . 258 file system, soft link. . . . . . . . . . . . . . . . . . . . . . . . . . . 278 file, symbolic link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 file, unformatted sequential . . . . . . . . . . . . . . . . . . . . 44 findloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 FINDLOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 flags inquiry function . . . . . . . . . . . . . . . . . . . . . . . . . 157 floating point, exponent . . . . . . . . . . . . . . . . . . . . . . . 180 floating point, fraction . . . . . . . . . . . . . . . . . . . . . . . . 188 floating point, nearest different . . . . . . . . . . . . . . . . 243 floating point, relative spacing. . . . . . . . . . . . 259, 273 floating point, scale . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 floating point, set exponent . . . . . . . . . . . . . . . . . . . 265 FLOAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 FLOATI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 FLOATJ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 FLOATK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 floor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 184 FLOOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 FLUSH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 FNUM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 form feed whitespace . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 FORMAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 FPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 FPUT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 FPUTC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 FRACTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 FREE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Front-end optimization. . . . . . . . . . . . . . . . . . . . . . . . . 32 FSEEK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 FSTAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 FTELL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 function elimination. . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 function interoperability with C . . . . . . . . . . . . . . . . 70 Further Interoperability of Fortran with C. . . . . . 74

G

g77 calling convention . . . . . . . . . . . . . . . . . . . . . . 25, 26 Gamma function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 Gamma function, logarithm of . . . . . . . . . . . . . . . . 227 GAMMA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 GCC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Generating C prototypes from

external procedures . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Generating C prototypes from Fortran

BIND(C) enteties . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 GERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 GET_COMMAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 GET_COMMAND_ARGUMENT . . . . . . . . . . . . . . . . . . . . . . . 194

GET_ENVIRONMENT_VARIABLE. . . . . . . . . . . . . . . . . . . 196

GETARG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 GETCWD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 GETENV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 GETGID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 GETLOG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 GETPID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 GETUID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 GMTIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 GNU Compiler Collection . . . . . . . . . . . . . . . . . . . . . . . 2 GNU Fortran command options . . . . . . . . . . . . . . . . . 7

H

Hollerith constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 HOSTNM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 HUGE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 hyperbolic cosine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 hyperbolic function, cosine . . . . . . . . . . . . . . . . . . . . 161 hyperbolic function, cosine, inverse . . . . . . . . . . . . 115 hyperbolic function, sine . . . . . . . . . . . . . . . . . . . . . . 271 hyperbolic function, sine, inverse . . . . . . . . . . . . . . 124 hyperbolic function, tangent . . . . . . . . . . . . . . . . . . 281 hyperbolic function, tangent, inverse . . . . . . . . . . 129 hyperbolic sine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 hyperbolic tangent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 HYPOT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200

I

I/O item lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 I/O specifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 IABS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 IACHAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 IALL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 IAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 IANY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 IARGC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 IBCLR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 IBITS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 IBSET. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 ICHAR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 IDATE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 IDIM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 IDINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 IDNINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 IEEE, ISNAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 IEOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 IERRNO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 IFIX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 IIABS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 IIAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 IIBCLR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 IIBITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 IIBSET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 IIEOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 IIOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 IISHFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 IISHFTC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 IMAG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 IMAGE_INDEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 images, cosubscript to image

index conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 images, index of this image . . . . . . . . . . . . . . . . . . . 281 images, number of . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 IMAGPART . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 IMOD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 IMVBITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 INCLUDE directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 inclusion, directory search paths for . . . . . . . . . . . . 23 INDEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 INOT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 input/output, asynchronous . . . . . . . . . . . . . . . . . . . . 45 INT2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 INT8. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 integer kind. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 integer overflow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Interoperability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 interoperability with C . . . . . . . . . . . . . . . . . . . . . . . . . 69 interoperability, derived type and struct . . . . . . . . 69 interoperability, intrinsic type . . . . . . . . . . . . . . . . . . 69 interoperability, subroutine and function . . . . . . . 70 interoperability, variable . . . . . . . . . . . . . . . . . . . . . . . 70 intrinsic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 intrinsic Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 intrinsic procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 intrinsic type interoperability with C . . . . . . . . . . . 69 intrinsics, integer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 intrinsics, math . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 intrinsics, trigonometric functions . . . . . . . . . . . . . . 62 INT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 inverse hyperbolic cosine . . . . . . . . . . . . . . . . . . . . . . 115 inverse hyperbolic sine . . . . . . . . . . . . . . . . . . . . . . . . 124 inverse hyperbolic tangent . . . . . . . . . . . . . . . . . . . . 129 IOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 IOSTAT, end of file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 IOSTAT, end of record . . . . . . . . . . . . . . . . . . . . . . . . . 215 IPARITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 IRAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 IS_IOSTAT_END . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 IS_IOSTAT_EOR. . . . . . . . . . . . . . . . . . . . . . . . . . . 214, 215 ISATTY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 ISHFT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 ISHFTC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 ISIGN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 ISNAN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 ITIME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218

J

JIABS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 JIAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 JIBCLR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 JIBITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 JIBSET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 JIEOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 JIOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 JISHFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 JISHFTC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 JMOD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 JMVBITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 JNOT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245

K

KIABS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 KIAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 KIBCLR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 KIBITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 KIBSET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 KIEOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 KILL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 kind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 219 kind, character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 kind, integer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 kind, old-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 kind, real . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 KIND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 KIOR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 KISHFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 KISHFTC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 KMOD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 KMVBITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 KNOT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245

L

L2 vector norm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 language, dialect options . . . . . . . . . . . . . . . . . . . . . . . . 8 LBOUND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 LCOBOUND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 LEADZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 left shift, combined . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 LEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 LEN_TRIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 lexical comparison of strings …. 222, 223, 224, 225 LGAMMA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 LGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 LGT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 libf2c calling convention . . . . . . . . . . . . . . . . . . . . 25, 26 libgfortran initialization, set args. . . . . . . . . . . . . . . 77 libgfortran initialization, set convert. . . . . . . . . . . . 79 libgfortran initialization, set fpe . . . . . . . . . . . . . . . . 80 libgfortran initialization,

set max subrecord length. . . . . . . . . . . . . . . . . . . . 80 libgfortran initialization, set options. . . . . . . . . . . . 77 libgfortran initialization, set record marker. . . . . 79 limits, largest number. . . . . . . . . . . . . . . . . . . . . . . . . 200 limits, smallest number . . . . . . . . . . . . . . . . . . . . . . . 284 linking, static . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 LINK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 LLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 LLT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 LNBLNK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 location of a variable in memory . . . . . . . . . . . . . . 226 LOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 226 LOG_GAMMA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 LOG10. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 logarithm function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 logarithm function with base 10 . . . . . . . . . . . . . . . 227 logarithm function, inverse . . . . . . . . . . . . . . . . . . . . 179 logical and, bitwise . . . . . . . . . . . . . . . . . . . . . . . 120, 202 logical exclusive or, bitwise . . . . . . . . . . . . . . . 208, 290 logical not, bitwise . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 logical or, bitwise . . . . . . . . . . . . . . . . . . . . . . . . 212, 247 logical, bitwise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 logical, variable representation . . . . . . . . . . . . . . . . . 41 LOG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 LOGICAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 login name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 loop interchange, Fortran . . . . . . . . . . . . . . . . . . . . . . 32 loop interchange, warning . . . . . . . . . . . . . . . . . . . . . . 19 LSHIFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 LSTAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 LTIME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230

M

MALLOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 MAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 mask, left justified . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 mask, right justified. . . . . . . . . . . . . . . . . . . . . . . . . . . 232 MASKL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 MASKR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 MATMUL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 matrix multiplication . . . . . . . . . . . . . . . . . . . . . . . . . 232 matrix, transpose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 MAX, MIN, NaN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 MAX0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 MAX1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 MAXEXPONENT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 maximum value . . . . . . . . . . . . . . . . . . . . . . . . . . 232, 235 MAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 MAXLOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 MAXVAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 MCLOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 235 MCLOCK8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 memory checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 MERGE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 MERGE_BITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 messages, error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 messages, warning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 MIN0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 MIN1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 MINEXPONENT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 minimum value . . . . . . . . . . . . . . . . . . . . . . . . . . 237, 239 MIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 MINLOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 MINVAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 Mixed-language programming . . . . . . . . . . . . . . . . . . 69 model representation, base . . . . . . . . . . . . . . . . . . . . 252 model representation, epsilon . . . . . . . . . . . . . . . . . 174 model representation, largest number. . . . . . . . . . 200 model representation, maximum exponent . . . . . 233 model representation, minimum exponent . . . . . 238 model representation, precision. . . . . . . . . . . . . . . . 251 model representation, radix . . . . . . . . . . . . . . . . . . . 252 model representation, range . . . . . . . . . . . . . . . . . . . 256 model representation, significant digits . . . . . . . . 168 model representation, smallest number . . . . . . . . 284 module entities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 module search path. . . . . . . . . . . . . . . . . . . . . . . . . 23, 24 modulo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 MOD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 MODULO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 MOVE_ALLOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 moving allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 multiply array elements . . . . . . . . . . . . . . . . . . . . . . . 252 MVBITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242

N

Namelist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 NAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 natural logarithm function . . . . . . . . . . . . . . . . . . . . 226 NEAREST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 NEW_LINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 NINT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 norm, Euclidean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 NORM2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 NOSHARED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 NOT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 NULL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 NUM_IMAGES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246

O

open, action . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 OpenACC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 55 OpenMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 54 operators, unary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 operators, xor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 options inquiry function. . . . . . . . . . . . . . . . . . . . . . . 157 options, code generation . . . . . . . . . . . . . . . . . . . . . . . 25 options, debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 options, dialect. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 options, directory search . . . . . . . . . . . . . . . . . . . . . . . 23 options, errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 options, Fortran dialect . . . . . . . . . . . . . . . . . . . . . . . . . 9 options, gfortran command. . . . . . . . . . . . . . . . . . . . . 7 options, linking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 options, negative forms. . . . . . . . . . . . . . . . . . . . . . . . . . 7 options, preprocessor. . . . . . . . . . . . . . . . . . . . . . . . . . . 14 options, real kind type promotion . . . . . . . . . . . . . . 13 options, run-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 options, runtime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 options, warnings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 OR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 output, newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 overflow handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

P

PACK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 PARAMETER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 Parity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 parity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 PARITY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 paths, search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23, 24 PERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 pointer checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 pointer, C address of pointers . . . . . . . . . . . . . . . . . 146 pointer, C address of procedures . . . . . . . . . . . . . . 146 pointer, C association status . . . . . . . . . . . . . . . . . . 144 pointer, convert C to Fortran . . . . . . . . . . . . . . . . . 145 pointer, Cray . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 pointer, cray . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188, 230 pointer, disassociated . . . . . . . . . . . . . . . . . . . . . . . . . 246 pointer, status . . . . . . . . . . . . . . . . . . . . . . . . . . . 124, 246 pointers, C. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 POPCNT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 POPPAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 positive difference. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 PRECISION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 Preprocessing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 preprocessing, assertion . . . . . . . . . . . . . . . . . . . . . . . . 16 preprocessing, define macros . . . . . . . . . . . . . . . . . . . 16 preprocessing, include path. . . . . . . . . . . . . . . . . . . . . 15 preprocessing, keep comments . . . . . . . . . . . . . . . . . . 16 preprocessing, no linemarkers . . . . . . . . . . . . . . . . . . 17 preprocessing, undefine macros . . . . . . . . . . . . . . . . . 17 preprocessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 preprocessor, debugging . . . . . . . . . . . . . . . . . . . . 14, 15 preprocessor, disable . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 preprocessor, enable. . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 preprocessor, include file handling . . . . . . . . . . . . . . 14 preprocessor, working directory . . . . . . . . . . . . . . . . 15 PRESENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 private. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 procedure interoperability with C . . . . . . . . . . . . . . 70 procedure pointer, convert C to Fortran . . . . . . . 147 process ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 product, double-precision . . . . . . . . . . . . . . . . . . . . . 170 product, matrix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 product, vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 PRODUCT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 program termination . . . . . . . . . . . . . . . . . . . . . . . . . . 179 program termination, with core dump . . . . . . . . . 111

Q

Q edit descriptor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Q exponent-letter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

R

radix, real . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 RADIX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 random number generation . . . . . . . . . . 213, 253, 255 random number generation, initialization . . . . . . 254 random number generation, seeding . . . . . . 255, 275 range checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 rank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 RAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 RAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 RANDOM_INIT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 RANDOM_NUMBER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 RANDOM_SEED. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 RANGE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 RANK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 re-association of parenthesized expressions. . . . . . 32 read character, stream mode . . . . . . . . . . . . . 182, 183 READONLY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 real kind. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 real number, exponent . . . . . . . . . . . . . . . . . . . . . . . . 180 real number, fraction. . . . . . . . . . . . . . . . . . . . . . . . . . 188 real number, nearest different . . . . . . . . . . . . . . . . . 243 real number, relative spacing . . . . . . . . . . . . . 259, 273 real number, scale . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 real number, set exponent. . . . . . . . . . . . . . . . . . . . . 265 Reallocate the LHS in assignments . . . . . . . . . . . . . 32 Reallocate the LHS in

assignments, notification . . . . . . . . . . . . . . . . . . . . . 21 REAL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 REALPART . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 record marker. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 RECORD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Reduction, XOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 remainder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 RENAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 repacking arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 REPEAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 RESHAPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 REWIND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 right shift, combined . . . . . . . . . . . . . . . . . . . . . . . . . . 172 root . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 rounding, ceiling . . . . . . . . . . . . . . . . . . . . . . . . . 121, 149 rounding, floor . . . . . . . . . . . . . . . . . . . . . . . . . . . 117, 184 rounding, nearest whole number. . . . . . . . . . . . . . . 244 RRSPACING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 RSHIFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 run-time checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

S

SAME_TYPE_AS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 SAVE statement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 SCALE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 SCAN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 search path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 search paths, for included files. . . . . . . . . . . . . . . . . . 23 SECNDS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 SECOND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 seeding a random number generator . . . . . . 255, 275 SELECTED_CHAR_KIND. . . . . . . . . . . . . . . . . . . . . . . . . . 263 SELECTED_INT_KIND . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 SELECTED_REAL_KIND. . . . . . . . . . . . . . . . . . . . . . . . . . 264 sequential, unformatted . . . . . . . . . . . . . . . . . . . . . . . . 44 SET_EXPONENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 SHAPE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 SHARE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 SHARED. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 shift, left. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 267 shift, right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 268 shift, right with fill. . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 SHIFTA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 SHIFTL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 SHIFTR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 sign copying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 SIGN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 SIGNAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 SIND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 sine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 sine, degrees . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 sine, hyperbolic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 sine, hyperbolic, inverse . . . . . . . . . . . . . . . . . . . . . . . 124 sine, inverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 sine, inverse, degrees . . . . . . . . . . . . . . . . . . . . . . . . . . 123 SIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 SINH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 size of a variable, in bits . . . . . . . . . . . . . . . . . . . . . . 142 size of an expression . . . . . . . . . . . . . . . . . . . . . 148, 272 SIZE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 SIZEOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 SLEEP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 SNGL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 SPACING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 SPREAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 SQRT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 square-root . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 SRAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 Standards. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 statement, SAVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 STAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 STATIC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 storage size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 STORAGE_SIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 stream mode, read character . . . . . . . . . . . . . 182, 183 stream mode, write character. . . . . . . . . . . . . 186, 187 string, adjust left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 string, adjust right . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 string, comparison . . . . . . . . . . . . . . 222, 223, 224, 225 string, concatenate. . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 string, find missing set . . . . . . . . . . . . . . . . . . . . . . . . 289 string, find non-blank character . . . . . . . . . . . . . . . 225 string, find subset. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 string, find substring . . . . . . . . . . . . . . . . . . . . . . . . . . 209 string, length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 string, length, without trailing whitespace . . . . . 222 string, remove trailing whitespace . . . . . . . . . . . . . 286 string, repeat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 strings, varying length . . . . . . . . . . . . . . . . . . . . . . . . . . 3 structure packing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 STRUCTURE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 subrecord . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 subroutine interoperability with C. . . . . . . . . . . . . . 70 subscript checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 substring position. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 sum array elements . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 SUM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 suppressing warnings. . . . . . . . . . . . . . . . . . . . . . . . . . . 17 symbol names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 symbol names, transforming. . . . . . . . . . . . . . . . . . . . 26 symbol names, underscores . . . . . . . . . . . . . . . . . . . . . 26 SYMLNK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 syntax checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 system, error handling. . . . . . . . . . . . . . . 192, 209, 249 system, group ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 system, host name . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 system, login name . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 system, process ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 system, signal handling . . . . . . . . . . . . . . . . . . . . . . . 269 system, system call. . . . . . . . . . . . . . . . . . . . . . . 178, 278 system, terminal . . . . . . . . . . . . . . . . . . . . . . . . . 215, 286 system, user ID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 system, working directory . . . . . . . . . . . . . . . . 150, 195 SYSTEM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 SYSTEM_CLOCK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279

T

tabulators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 tangent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 tangent, degrees . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 tangent, hyperbolic . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 tangent, hyperbolic, inverse . . . . . . . . . . . . . . . . . . . 129 tangent, inverse . . . . . . . . . . . . . . . . . . . . . . . . . . 126, 127 tangent, inverse, degrees . . . . . . . . . . . . . . . . . 126, 128 TAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 TAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 TANH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 terminate program . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 terminate program, with core dump . . . . . . . . . . . 111 THIS_IMAGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 thread-safety, threads . . . . . . . . . . . . . . . . . . . . . . . . . . 42 time, clock ticks . . . . . . . . . . . . . . . . . . . . . 235, 236, 279 time, conversion to GMT info . . . . . . . . . . . . . . . . . 198 time, conversion to local time info. . . . . . . . . . . . . 230 time, conversion to string . . . . . . . . . . . . . . . . . . . . . 165 time, current . . . . . . . . . . . . . . 166, 181, 218, 282, 283 time, elapsed . . . . . . . . . . . . . . 164, 172, 176, 262, 263 TIME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 TIME8. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 TINY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 trace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 TRAILZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 TRANSFER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 transforming symbol names . . . . . . . . . . . . . . . . . . . . 26 transpose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 TRANSPOSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 trigonometric function, cosine . . . . . . . . . . . . . . . . . 160 trigonometric function, cosine, degrees . . . . . . . . 160 trigonometric function, cosine, inverse . . . . . . . . . 114 trigonometric function, cosine,

inverse, degrees . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 trigonometric function, cotangent . . . . . . . . . . . . . 162 trigonometric function, cotangent, degrees . . . . . 162 trigonometric function, sine . . . . . . . . . . . . . . . . . . . 269 trigonometric function, sine, degrees. . . . . . . . . . . 270 trigonometric function, sine, inverse . . . . . . . . . . . 122 trigonometric function, sine, inverse, degrees . . 123 trigonometric function, tangent . . . . . . . . . . . . . . . 280 trigonometric function, tangent, degrees . . . . . . . 280 trigonometric function, tangent, inverse… 126, 127 trigonometric function, tangent,

inverse, degrees . . . . . . . . . . . . . . . . . . . . . . . . 126, 128 TRIM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 TS 29113 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 TTYNAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 type alias print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 type cast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285

U

UBOUND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 UCOBOUND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 UMASK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 underflow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 underscore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 unformatted sequential. . . . . . . . . . . . . . . . . . . . . . . . . 44 UNION. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 UNLINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 UNPACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 unused dummy argument . . . . . . . . . . . . . . . . . . . . . . 20 unused parameter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 user id. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198

V

variable attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 variable interoperability with C . . . . . . . . . . . . . . . . 70 Varying length strings. . . . . . . . . . . . . . . . . . . . . . . . . . . 3 vector product . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 VERIFY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 version of the compiler . . . . . . . . . . . . . . . . . . . . . . . . 158 VOLATILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

W

warning, C binding type . . . . . . . . . . . . . . . . . . . . . . . 18 warnings, aliasing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 warnings, alignment of COMMON blocks. . . . . . . . . . . 20 warnings, all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 warnings, ampersand. . . . . . . . . . . . . . . . . . . . . . . . . . . 18 warnings, array temporaries . . . . . . . . . . . . . . . . . . . . 18 warnings, character truncation . . . . . . . . . . . . . . . . . 18 warnings, conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 warnings, division of integers . . . . . . . . . . . . . . . . . . . 19 warnings, extra . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 warnings, function elimination. . . . . . . . . . . . . . . . . . 20 warnings, implicit interface. . . . . . . . . . . . . . . . . . . . . 19 warnings, implicit procedure . . . . . . . . . . . . . . . . . . . 19 warnings, integer division . . . . . . . . . . . . . . . . . . . . . . 19 warnings, intrinsic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 warnings, intrinsics of other standards . . . . . . . . . . 19 warnings, line truncation . . . . . . . . . . . . . . . . . . . . . . . 18 warnings, loop interchange . . . . . . . . . . . . . . . . . . . . . 19 warnings, non-standard intrinsics . . . . . . . . . . . . . . . 19 warnings, overwrite recursive . . . . . . . . . . . . . . . . . . . 19 warnings, q exponent-letter . . . . . . . . . . . . . . . . . . . . 19 warnings, suppressing . . . . . . . . . . . . . . . . . . . . . . . . . . 17 warnings, suspicious code . . . . . . . . . . . . . . . . . . . . . . 19 warnings, tabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 warnings, to errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 warnings, undefined do loop . . . . . . . . . . . . . . . . . . . . 20 warnings, underflow . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 warnings, unused dummy argument . . . . . . . . . . . . 20 warnings, unused parameter . . . . . . . . . . . . . . . . . . . . 20 warnings, use statements . . . . . . . . . . . . . . . . . . . . . . . 20 write character, stream mode. . . . . . . . . . . . . 186, 187

X

XOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 XOR reduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249

Z

ZABS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 ZCOS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 ZCOSD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 zero bits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220, 284 ZEXP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 ZLOG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 ZSIN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 ZSIND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 ZSQRT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274